A Complete Bibliography of the Journal of Theoretical Probability

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

14 August 2024 Version 1.35

Title word cross-reference

(2J_{s} − R_{s}, s ≥ 0) [YY15]. (α, d, β) [Myt98, MX04]. (c, d) [JM09]. (c, p) [MM05]. (n − k) [Özd19]. (ξ, ψ, K) [Sch99]. (y, z) [Sun22]. (Z/pZ)^{d} [HM08]. 0 [BR08a]. 1 [AGK92, BR08a]. 1 + β [FMW11]. 1/4 < H < 1/2 [ALT20]. 1/\sqrt{n} [Huj06]. 2 [FIT14, Neu92, Neu95, SW92]. 2 < α < ∞ [Gei01]. 2 × 2 [MR97, MNR99a, MNR99b]. [k]^{N} [CL16]. α [CR13, CG03, CG06, CNX21, Den89, FMW11, JL24, RY91, RS08]. B [Arc94, CM00, Den04, LT98, Rya10, Su96]. β [Ass19, Pak20b]. C^{1} [Kim08]. G [Bou05]. D [Sch94a, CFR05, God04, Rus07]. D = 1 [BBDP98]. d × d [MN02]. ℓ_{p} [CV21]. F [BL22]. F_{ψ}(Ω) [RMVP24]. \frac{1}{2}Δ + ⟨a, ∇⟩ + b [Ger92]. \frac{4}{3} [HNS14]. G [AMR08, DH21, HL14, LS21b, LL24a, Liu21, LW19, MR02, PZ22, Sun22, WZ21a, WGYX24]. G^{1} [KT96]. h [MS01a]. H = 1/10 [HN15]. H > 1/2 [NO03]. H^{2} [MT08]. k [BGT04, BMS12, GK23, KK24]. L [Bos98]. L^{0}(R_{+}^{d}; Ω, F, P) [Bou05]. L^{1} [Fan18, Sau16, ZZ99]. L^{2} [AB21, CAMBTD96, GHLT04, MW22, Mor10]. l^{2} (S, P) [Ban16]. L^{p} [Mor21, Sam24, ZM23, KS22]. L_{1} [Cer00]. L_{2} [GMS22, Naz09]. L_{p} [Kha09, Kim08, Kim09, Kim17, Aur07]. Λ [LZ24]. n_{(−)} [YY15]. [0, ∞[ [Gru14]. L log L [LRS11]. log log [Spa04]. M [MLC22, CV15, Che97b, EL95, Lee97, MST08, VC18]. M_{1} [BK15]. F_{2} [GZ24]. GL_{N} [Kem17]. H^{3} [Cza24]. N_{0} [EP15a, EP15b]. R [Bal07a]. R^{d} [Att14, MR02]. U_{N} [Kem17]. Z [BSF15, Vo23, Vol06]. Z^{2} [CF24, CF20]. Z^{d} [AK20, Cyr21, Per97]. AX + B [MŚ21]. L [Che15b]. M_{L} [Sta11]. P([′, ∞]) [AMS12]. S [CdF00]. SLE(16/3, (16/3) − 6) [GW20]. μ = μ ∗ σ [SZ90]. N [Cyr21, FUW20, GZ98, IW99, LM16a, Lee97, ML17, Zeu94]. N_{0} [Voi93]. O [Nea92, Wat21]. ω [BGZ93]. ω^{2} [Che99a]. Ω_{p} [Myr20]. p [ARV97, BW19, BG16, DPR00, Fel15, FM17, KR21, MS06, Mus06, RR92, Sha91, Sim04, Wei22, Yas20]. Φ [HSZ22, Wu09, AR04, Kie98, Mal99, VC13]. R [BG11]. R^{d} [Ilt95, KR03, Kov02, SU14, Tug14, AMR08, IW99]. R^{K} [San07]. R^{n} [BF13, BRZ96]. R_{(+)} [YY15]. R_{2} [Zeu94]. RCD(K, ∞) [Li16]. ρ^{∗} [Mil94]. S [DMCN24, JN99]. S_{∞} [KN21]. S_{n} [Din17]. S_{nk}/(2n_{k} log log n_{k})^{1/2} [Sla89]. σ [BR08a, Kom08, MT08, Li22]. SO_{0}(p, q)/SO(p) × SO(q) [Saw01]. S ⊂ l^{2} [Ban16]. ∑(H) [EOOM17]. ∑_{s(H)} [EOOM17]. t [BL22, Sha99]. τ [DP04, Eic97]. U [AL08a, AM14, Alb00, AY94, Arc94, BU01, BW02, BS99, DPR11, EL95, Fer94, GZ96, GM07, Han18, Mik94, RS95, RU00, Su96, Tei92, TZ96, Tei98a, TM98, UB98]. U(n) [Kua18]. V [Zha01a]. Z^{2} [CHL08, Zho95]. Z^{d} [BK09, GW02, Zha01b]. Z_{d} [GP01]. Z_{p}^{k} [Asc09].

-Actions [GW02]. -Adic [BW19, Fel15, FM17, Wei22, MS06, Mus06, Sha91]. -Adic-Valued [Yas20]. -Algebras [MT08]. -asymptotically [DMCN24]. -Block [LM16a]. -bounded [GMS22]. -Brownian [LL24a, LW19, WZ21a, DH21, LS21b, Liu21, Sun22, WGYX24]. -Brownian1 [PZ22]. -Circulants [BMS12]. -conditional [Bou05]. -Contractions [ZZ99]. -Cycle [Özd19]. -Dependent [CV15, Che97b, DP04, MST08, VC18, AGK92]. -Dimensional [God04, Rus07, CFR05]. -Distance [Che15b]. -Divergence [BL22]. -Domains [Sch94a]. -Ensemble [Pak20b]. -Entropy [BL22, Wu09]. -Error [ZM23]. -Expectation [HL14]. -Exponential [MW22]. -Families [VC13]. -field [BR08a]. -Fields [Kom08]. [?]FlemingLiu:2024:EMC. -Fold [GZ98]. -Indeterminate [MLC22]. -Integrability [CG03, CG06]. -Invariant [Tug14]. -Jacobi [Ass19]. -Laguerre [Ass19]. -Mixing [Kie98, Mal99]. -Moments [AR04]. -monotone [SW92]. -Nilpotent [Neu95, Neu92]. -norm [CV21]. -Parameter [IW99]. -Processes [MS01a, Rya10, TM98, AY94, Arc94]. -Pseudostable [JM09, MM05]. -Random [GP01]. -Sequences [HKW20]. -Small [Naz09]. -Solutions [KS22]. -Space [CM00]. -Spaces [Sam24]. -Stable [CR13, DPR00, FMW11, JL24, JN99, RS08, CNX21, RY91]. -Statistic [Che99a, Sha99, GM07, Fer94]. -statistics [BGZ93, AL08a, AM14, Alb00, BU01, BW02, Bos98, BS99, DPR11, Han18, RU00, Tei98a, UB98, Zha01a, Arc94, EL95, GZ96, Mik94, RS95, Su96, Tei92].

-subexponential [Wat21]. -summable [Nea92]. -Summing [KR21]. -Superprocesses [MX04, Sch99]. -term [FUW20]. -theory [Kim17, Kim08, Kim09]. -Topology [Eic97]. -Transform [BG11]. -type [ML17, TZ96]. -Uniform [CdF00]. -Uniqueness [Sau16]. -Valued [Den04, LT98, Arc94, Su96]. -variate [EL95]. -Variation [HNS14, Sim04, HSZ22]. -Version [Kha09].

1-Meixner [SC21]. 13 [Dun01]. 1st [MV98].

3D [Sla22].

8 [Cuz01].

A. [CR22]. a.s [BD94, Den89]. abelian [Bin93, Bud94, CF95, Bin88, CF97, CC17, FG00, Fel04, Myr24, SZ90, Yas00]. above [Tyk09]. Absence [KT20a, KT22]. Absolute [Ahm05, BSC01, Bre02, Bre05, DD97, Gab11, Kul11, LW12, Wei14, YZ15, Arc95a]. absolutely [DF95]. Absorption [Buc21, MS01a]. Abstract [ABK16, BZ19, Bet23, KS06]. Accessibility [Li18]. Accuracy [Pap11, BP96]. Accurate [KN10]. Acknowledgment [Che17]. acting [Kun94b]. Action [ISY23]. Actions [GW02, Lia09]. Actual [Ver12]. Ad [CK14]. Adapted [DMV07, QV12, Ret04, dlP93]. adaptive [Kou96]. Added [GH19, Jun05]. Addendum [MNR99a]. Addition [Bon15]. Additive [BPR13, Cap13, Cas08, CL99, CF00, CP12, DMNS24, Fai98, GR97, Hu15, Kha12, KY12, KP14, KS16b, Kun22, Lac00, Lem21, LL12, MS23, MT98, PRP24, Pro21, SS18, Tak08, Wat02, Wat16, CCFR92, CC95, Kuh91, Son95, Cap13]. Additive/ [Cap13]. Additivity [MdCQ07]. Adeles [Yas10]. Adic [BW19, Fel15, FM17, FP10, Wei22, Yas20, MS06, Mus06, Sha91]. Affect [AMS22]. Affine [Bro04, BK12, CKRMT16, HM08, BS88, GL94]. Age [KS98]. Age-Dependent [KS98]. Aggregate [YLY22]. Aggregation [HJ89]. Ait [CMP24]. Algebraic [FH97, FG99, Far03, FJ05, MS22, Muk19, Pan19, Voi09]. Algebras [GHS20, MT08, Sah08, Sta07]. Algorithm [Gås06]. algorithms [Kou96]. Alignments [GHL18]. Allen [HR18, Lee18]. Allocation [MM10]. Almost [AKR99, DM07, FC07, FM23a, FM23b, GW02, HR00, Hör07, Hu15, KKH09, Lou16, MS97a, Meg02, NP24, PG99, PT17, RV04, SGQ22, Ued19, Véc23, Wan10, XZ10, BD94, Kou96, Lac91, LRJW95, Mar88, PF91]. Almost-Sure [AKR99, PG99]. Along [Cov11, NP24, DL94]. Always [Huj06, Tak01, Pin94]. Amalgamated [Zit22]. amarts [Mar88]. Amenability [JS99]. Among [LV12]. Analog [Pap90]. Analogue [Myr24, Yos20, Zho15, BHL93]. Analysis [AAO17, Att14, Att21, CP11, CMP24, EJU19, FS21, FH15, GNU17, MZ06, Qin24, Tan20, WY23, WY24, ZHBS24]. Analytic [FCU12, FK20, Has12, LMS12, SG90]. analytical [HJ96b]. Analyze [JSS23]. Ancestral [Ban19]. Andersen [MV19b]. Anderson [BC18, KMV18].

Angular [CM22]. Animals [Lee97]. Anisotropic [CF22b, CF23, CF24, GPS24, NC21, YC24]. Annealed [CX15]. annealing [HS92b]. announcement [Ano89, Ano90, Ano91, Ano93, Ano96]. Anticipating [AN98, MCR08]. Anticipative [Oss09a]. Any [Che21]. Appell [FS92, GT01]. Appetites [Pac12]. Application [AAO17, BJ02, CH21, CWM04, Che20, CNX21, CKR19, DM13, EM18, FY08, GH95a, GH95b, Gut04, Kov02, OBO23, Smi15, Sob10, Tan20, BM91, Shi96, dlP93]. Applications [Ade22, Ama12, Ast23, AG18, BP10, BJS24, CFN98, CM15, DP04, DL00, Der03, Eng20, FJ12, Fan19, FGL20, FG14, GR07, GT02b, KMX18, KO10, KU23, Lin12, Oka24, RRM07, iSU13, SJ98, Ste18, SYG22, TZ20, UM14, WLG03, Wei14, Yas18, BGZ93, HS88, Jan90, Voi90, BP13, Lac98]. Applied [Iso09]. Approach [BK00, CCET06, CdGR20, Com05b, EM00, FL14, FY08, HH10, Kor16, MR05, MV00, Oss09b, Roz13, Sta07, Tri19, Yam24, Zha01b, vD17, AZ96, Bab94, BL95a, DM91, Fen96]. Approaches [ALL^{+}07]. Approximate [BM18, Bob05, CP18, Bin88, Bin93]. Approximated [EF01]. Approximation [AL08b, BS20, BL06, BL09b, CCET06, CNZ22, CNXY24, CW10, CF22b, CF23, CAMRR03, DH20, Deh98, DP22, FUW20, Gau16, Gau20, Gry21, Huj06, JL15, KL17, KPR24, KN16, LMS24a, LMS24b, LMX22, MRS07, Mec12, Mer03, MZ24, NN07, NP18, NPY22, Nov24, Pap11, PPPA16, PZ22, Pri19, Qin24, QK23, QV12, RMVP24, Sch16b, Ste20, Sza12, Vol10, WLG03, XZ09, CCFR92, Dip91, GH95a, GH95b, Kol94b, NS23b, Pec93]. Approximations [BH22, BL19b, BPY03, BX18, BX19, BX20, Bob18, BW02, CV21, CJSS24, CV07, CDM24, DQ21, EM92, JSS23, JP22, KM17, VC18, BP96, Lal91]. Arbitrary [CWM04, Hil01, Kim22, KN10, KN16, Str95]. Arc [BHH10, RYZ02, BY96]. Arc-Sine [BHH10, RYZ02, BY96]. ARCH [JP22]. Arcsine [HJ99]. Area [HSW18, KL05]. Argument [MS00]. Arise [Pan19]. Arising [App01, Löc19, YWC18]. Arithmetic [Ald10, Pec93]. ARMA [DD97]. Array [Adl02]. Arrays [ARV97, BP10, BP13, D’A00, Dav05, GC98, HP97, HJMLX16, Kal99, Kal12, RV04, SGQ22, Bin88, Bin93, Kal92a, Mas13]. ARTICLE [BP10]. ASCLT [BH24]. Associated [BM15, BR21, Bog20, CS00, Cri19, Cri20a, Cri20b, FK20, FS21, GP11, HKU12, HL22, HN05, JL08, Kob11, Küh21, Lal99, Luo15, Man04, MRS07, NN07, OBO23, Sau16, Sha00b, SYG22, Voi17b, Wat02, XY10, YSH03, Nej94, Sif95, Voi93, Yu89]. assumptions [CG96]. Asymmetric [CNX21]. Asymptotic [Ahl15, AJGO23, BN13, BGZ93, BM14, BT12, CW23, CH05, CHL08, CF00, CH01, Dav05, DR09a, DR09b, Dip91, GGX24, GT09, GT02b, GT05, GNU17, GH06, HW92, Hua19, Jan99, JS21, KMV18, KD17, Kaw13, KMX18, Kho22, KHMCSS01, KHN21, Kon95, Kon98, Kro22, LS15a, Leh15, LCP12, LR90, Mas13, MM10, Mon19, Nag15, NZ22, Per18, RU00, Saa01, Sgi98, TW14, Uch12a, Uch12b, Wad21, Wu23, Wüb13, Xue15, Xue18, YY19, ZZ19, Zei18, Bra92, Cha95, HKW90, Hec96, Mao91a, NS95, Pel96, Wha92]. asymptotical [Ben94]. Asymptotically [AK16, BMP22, Deb10, DW24, HJ02, MMW13, VB23, ZZ99, DMCN24]. Asymptotics [AP97, AFK03, BX20, BF13, Bro03, BC22, CP21, CX15, CCF^{+}02, DY95, GKS15, Ham01, Has05, HMS21, Jeg11, JY02, Kha12, Mar04, Naz09, Spa99, Spa04, Spa16, Tel06, Von00, WW09, Ilt95]. Attachment [ARS21]. Attracted [Wan17]. attracting [SYG22, YSL08]. Attraction [BK03, DM02, FC07, Fie90, MS97b, Meg02, Sep94, Sep97, Sep99, CD91, HHV89, HS93, Mee94, MS96, Sch94a]. Autocovariance [HL20, Owa16]. automaton [Fis90]. Automorphisms [HS88, Jaw97]. Autonomous [Nas22]. Autoregressive [BP21, Ros97, BL95b, Die91]. Average [Bag98, KK08, Muk19, NSB02]. Averaged [DW06]. Averages [BH99, LPP17, BS93, CG96, PF91, Web96]. Averaging [Pah15, XL23, DL94]. Avoid [Uch20]. Axelrod [LM16b]. Axiomatization [HJ89]. Azuma [Luo22]. Azuma-Type [Luo22].

B [AT17]. Backbone [FPP21]. Backward [Ama12, ABO22, AO24, BM01, BR21, CF21, CL23, CM14, DH21, Fan18, HMW24, Kar19, KL14, KS22, LS21b, LL24a, MN15, MS02, OBO23, OA23, Sun22, WZ21a, WGYX24, Xu07, XL23, ZM23]. Badly [Häg13]. Ball [Ban16, BL02, Bro03, DFMS03, Der03, Dun00, Dun01, FGG16, GLP03, HSW18, LS99, Mar04, Naz09, Roz19, Sim04, GS96, KL93, KLmS95, Pap90, RZ90, Sha93, Shi96, Sto96]. Balls [BZ19, BEK10, GHLT04]. Banach [ARV97, Ban16, BJZ19, BZ19, BGZ93, BR90, Che97b, CG21, DPR00, DFMS03, Det89, DW06, Dob94, Gri10, HH10, Hei06, KK08, KU23, KZ08, Luo22, Mae93, PP02, RS94, RV04, ST90, Sie92, Sie90, Sla89, Wit00]. Banach-Valued [KU23, ST90]. Band [Jur18, LW11]. Bandwidth [DO13]. Banks [dHN22]. barely [Adl90]. Barrier [DFPS97]. Barriers [Xu07]. Based [CF21, ER01, FV20, IMS20, JY02, KM17, Maj99, SS09, Sza12, ZH12, Mao91a, Mao91b, ST15]. Basic [Cyr21, Kon95]. Basis [Tso20]. Bayes [RS91, Ruk93]. BC [Voi17b]. be [ADG18]. become [Xue91]. Behaved [Häg13]. Behavior [CW23, CFR07, DT24, DGG99, Gra15, GLT20, Hör07, Hu20, JX20, KMV18, KMX18, KS08, KHMCSS01, LS15a, Lin18, MWW15, NZ22, RU00, RV04, Saa01, Sav10, Sgi98, TW14, Véc23, Wad21, Xue15, YY19, YLY22, Zei18, AG93, Ben94, BR92, Eis93, HW92, Hec96, Kon95, NS95, Pri96]. Behavioral [KM18]. Behaviors [GGX24, HS92b]. Behaviour [AFK03, AB17, BM14, DM23, IW24, Kra01, Kro22, Leh15, Wat22b, ZZ19]. Belief [KM18]. Bell [Kho22]. Beltrami [JW21]. Benford [BX18, BX19]. Berezin [BM91]. Bernoulli [DLSV05, HSZ22, JMPP04, MWW15, NA11, NW19, Sch16a, Sch02b]. Bernstein [Lem21, SS15, VZ14]. Bernstein-Like [Lem21]. Berry [AL07, Alb00, Ben94, BBG96, JZ23, Kar07, ZWW04]. Besov [NP07, NO03, YWS22]. Besov- [NP07]. Bessel [CFR09a, FG14, HN05, Mat17, Pro21, Shi96, Voi09, YY15]. Best [BX18, BX19, Fan16]. Beta [JL15, JK22, Tri19, Yos20]. Beta-Wishart [JK22]. Bets [SK15]. better [Xue91]. Between [CH21, DFMS03, ES22, GGPV18, Gra15, JK17, KS21, Kom08, KL08, MPS02, Min21, Rya10, Sha00b, Ald10, BGCD10, Jan90, MT08]. Beyond [GK23]. Bi [Voi17a, GHS20]. Bi-free [Voi17a]. Bi-monotonic [GHS20]. Biangle [Mil00]. Biased [Hu17, Rev01]. Bifractional [Ma24]. Bifurcating [BP21, PD23, SWF22]. Bifurcation [BCOV23]. Big [CK99, Chu93]. Bilinear [ISC01]. Binary [AAS18, BU01, FNN15, SWF22, UB98]. Binomial [IMS20, Sch16a]. Bipartite [AE99, Zhu23]. Bipolar [Bou05, Zit02]. Biregular [Zhu23]. Birth [ALP22, DG08, DFPS97, DM09, Fil09b, GMZ12, Tob90, Wan18a, XZ09, vD17]. Birth-and-Death [Fil09b]. Bismut [Tak10]. Bispectra [LR93]. Bivariate [GT01, LPP17]. BKR [GR07]. Black [Huj06]. Blackwell [Nie22]. Bloch [DMCN24]. Bloch-Type [DMCN24]. Block [BN13, DR10, KKM13, LM16a, LLW11, Lou16, Ora07, Juh90]. Block-Hankel [Lou16]. Block-Matrix [Ora07]. Block-Transposed [BN13]. Blocks [CLM04, Mór95]. Blockwise [MST08]. Blow [BY16]. Blow-up [BY16]. Bodies [Mil09]. Body [Lia97]. Böge [RZ15]. Boltzmann [Qin24, Stu22a]. Boltzmann-Type [Qin24]. Bond [GPS24]. Bondesson [Kuz20]. Bonnet [Hsu97]. Boolean [Tan20, Ued22]. Bootstrap [BK19, Rad98, dBM00]. Borel [WWLH14]. Borodin [TV20]. Bottom [Tak18]. Bouncing [GM18]. Bound [Alb00, JZ23, Jur15, Tal98, Ued19, Vil20, ZWW04, BBG96]. Boundaries [DM07, KKR04, KM98, Dan91]. Boundary [AP10, AK16, CLY13, DS00, Fra21, God04, Has10, HMS21, IR08, KV07, KS08, Kim08, MRR04, MRS07, Naz09, Von00, CS91, Lal91, RR92]. Boundary-Crossing [AP10]. Boundary-Value [Naz09]. Bounded [Aid24, Ala22, Ben03, BDK21, CM14, Dan14, Fan16, GSS21, Lac00, Lac06, Lop20, MM22, RZ10, Uch18, Wan18a, Zha01a, GMS22, Lev89, Mao92, Zuc89]. Boundedness [MR05]. Bounding [CB04, Pes01]. Bounds [AL07, Ast23, BS05, BPY03, BW02, BK89, DFY01, DM16b, Gau16, Gau20, GS24, GOT20, GHL18, HK98, HL20, Has05, HMS21, KN10, KN16, NPY22, QK23, RS18, dAC98, Ald89b, CAMBTD96, Key90]. box [GL94]. Bracketing [NP07]. Braided [FNS99]. Branching [AM14, ABKV12, Att14, Att21, Ban19, BLP16, Bul14, Bul21, CRY17, CFZ23, CY11, Dad20, DGW01, FV05, GMPV10, GKS15, HJ02, HLX18, He22, Hu15, IM24, KS98, Kes95, Kor16, KV11, LV12, LM08, LCP12, LX12, LL21, LVZ21, LPW22, LL24b, LRS11, Mad17, ML17, Nak13, Pak20a, RY24, Rév98, Stu22b, WH17, Wat02, Zha23, MSSW93, PD91]. Brannath [Zit02]. Breiman [NP24]. Bridge [Man04]. Bridges [BL97, CKR19, EHL20, FG14, HR20, Uch20, Wei22]. Broad [BLC03]. Brownian [Dun01, WZ21a, AI01, AJ22, AP97, App93, ALT20, AV15, AMS22, AIS^{+}24, BL19a, BOLP22, BM11, BDW21, BD09, BL02, BLV24, BL97, BPY03, BH99, BX16, BL11, BGT04, BF13, BDS23, BDS24, BP14, Bro03, BFW18, Bur22, Cha95, CX15, Che15b, Chi03, CY22, CW10, CCFR96, CH01, CCFR19, CL14, CK14, Dai13, DH21, DF95, DF97, DLZ04, Dey20, DP12, DM16b, Dun21, Dun00, ES99, EF01, EK01, EMP17, Eva93, FZ01, Far11, FKX06, Föl03, Fus10, GIST20, GOT23, GM18, GK88, GK91, GK02, GV18, GS96, GHX19, GZM24, HTW24, HN15, Has10, HLZ13, HM06, HM09, Hon03, HZ07, Hu99, HNS14, IO20, IPT22, IW99, IM96, Jin17, Joh89, JV07, KMV18, KO99, KM17, KL96, KM10, KP23, KM23, KL93, KL98, KL02, Lac98, Lac00]. Brownian [Lac06, LV12, LY19, Li20, LS21b, LL24a, LM19, Liu21, LZ24, Lop20, LS11, LW19, MN20, Ma24, Mal06, Mal08, Man04, MY09, Muk19, Nag15, NN07, NS95, Nod17, NO03, NOL07, OMO22, PPPA16, PP18, Pin01, Rai97, Ran06, RY24, RSZ24, Ric17, Ryv15, Sar17, Sar19, Ser96, Shi97, Sik00, Sot04, Sun22, SYG22, Sza12, TX96, Tal96b, TV06, Tso93a, Uch12a, Uch12b, Uch18, Uem04a, Vad10, VZ23, Von91, VA96, VA01, Wan10, Wan14c, WX22, Wan24, WGYX24, Wei22, Wu99, WX10, Xia98, XSR23, Yos21, Yu21, YY24, Zei18, Zha05]. Brownian1 [PZ22]. Brox [AD11, PPR22]. BSDE [El 09, Roz13]. BSDEs [FJ12]. Bulk [KO19a, KO19b]. Burgers [Don08, HS21, LR12, Moh22]. Burkholder [Ish92].

CA [Rah22]. Cadlag [CL16, FM18]. Cahn [CWM04, HR18, Lee18]. calculation [Let93]. Calculus [BT12, FRZ18, Kob11, LT12, Oss09b]. Calibrating [KS21]. Cameron [BG16]. Can [Rad98]. Canards [Sow08]. Canonical [GR10, Ste16, Arc94, Ber93, GR17]. Cantelli [Bos98, DGZ91, Tal96a, WWLH14]. Cantor [Fal89, FG92a, FG94]. Capacities [Fit99, OW98, Kol89, O’B96, SW92]. Cards [NW19, HJ96a]. CARMA [FS13]. Carnap [RZ15]. Carrying [CZ17]. Cascades [Löc19]. Case [Alb00, Asc09, AS03, BGCD10, BK12, Di 22, EHL20, GOT23, GLT20, Kab12, LS15a, Liu21, OU14, PGS16, Stu22b, WW09, ZL18, BBG96, BX16, GM91b, Mór95, Sto91]. Cases [San14]. Catalan [BB13]. Catalyst [FKX06]. Catalytic [Bul21, LM08]. Cauchy [BPY03, BH11, Dan91, DS88, Has88, OD15]. Cauchy-type [Dan91]. causal [CH05]. cellular [Fis90]. Centered [LL12, Nam22]. Centering [Gri97, Luc10, FKT94]. Centers [Der03, PRWW19]. Central [AY94, BD93, BD06, BM18, Bjö10, Blo96, Bob18, Bra97, BT17, CZ22, CP12, FKT22, GT01, GHL18, Hör07, JH03, JL24, KN21, KL01, KV18, KZ15, LT23, Lyt18, Mas21, MM07, Mil18, NN10, NT14, OR16, Par12, PS13, PD23, RW02, Sif95, Sla22, SD22, Wan18b, Xu18, Zei13, Zha11, ZRP20, AD88, BB93, BD94, Bin88, DM95, GT05, Nej94, Neu95, NP91, PS94, Zeu89, BD06, DS14, DT14, Fra07, GT01, GW02, Gru14, JH12, LBBB11, Mal99, MS97a, Mer03, NP07, PR23a, PR23b, Per97, RS04, Saw01, Vol10, YSH03, Zha05]. Certain [CGHM06, CNW09, GR23, KV19a, Lac00, Lac06, Lou16, Mar01, Nie22, Roz19, VA03, Wei22, ZM23, Die91, Fer89, Kou96, PS89, Von96]. Cesaro [Hei90, CG03, CG06, Den89, GS10]. Cesàro- [Den89]. Chain [Bra01, Fil09b, Kor08, Nie22, Rah22, DL94]. Chains [AE99, Ban19, BL06, BL09b, BK14, Bro00, Che99b, CC11, Cop18, CP12, Den20, DFPS97, DD01, DMNS24, EK01, Fil09a, FJ95, GHLR24, GR23, HH09, HM18, Hua14, Jia19, Kha12, Kua18, LM16a, Lem21, MS22, Pan19, Pel24, PV98, PD23, RS18, Ros03, San14, San16, SWF22, Smi15, WY16, Wil03, Xu13, XZZ18, Yen04, dAC98, AGK92, AK92, Ald89b, BL95a, CC95, DSC96, DN95, Fil92, GZ24, HS92b, Las96, RX94, Voi93, dA90]. Champions [DFZ19]. Chan [BH23]. Change [BNMiS06, DGR21, FY08, HL03, LW04, Pes05, RR00, Øks90]. Change-of-Variable [Pes05]. Changed [HKU12, Kob11, KP23, MV19a, Sam21, Tak18, Von93]. Changes [AI01, Fit89]. Chaos [DT24, Hu97, MT98, MR99, Tal98, AG93, PAT94]. Chaotic [Ban16]. Characterisation [Li22]. Characteristic [FCU12, GLS07, Kös11, Juh90, Pin18, Pin93]. Characterization [FM17, HLZ08, HM09, Her20, KN23, Kol94a, Las20, Law08, Nea92, Bra91, Bra93, Bra94, Dan91, DS88, IM96, KT96, Let89, Sla89, Von91, Von96, Zen95]. Characterizations [Ejs14, GR05, MSS17, OV14, SW92, Sep94]. Characterizing [Zab95]. Charges [SK20]. Chébli [Zeu89]. Chentsov [KU23]. Chern [Hsu97]. Cherny [CR22]. Chevet [MT88]. Choosing [Sto91]. Choquet [SZ90]. chosen [BK93]. Chromosomes [Dur03]. Chung [GL99, Hil22, LM19, SGQ22, WX22]. Chung-Type [SGQ22]. Chutes [DD01]. circle [BGT10, Din95]. Circuit [Kal92b]. Circulant [BHS11, CM05, KKM13, MMS07]. Circulant-Type [BHS11]. Circulants [BMS12]. Circular [Cha10, MQ21, NNR16]. Circulation [Kal91]. clarification [Bra88]. Clark [AAO17, NYY22]. Class [BT24, BM15, BD06, BHW10, Bon15, CGHM06, Chi03, CDR17, CDM24, DLSV05, GW02, GSY24, HSZ22, HJN19, HMY21, KV07, KLR22, KS12, Lac00, Li22, LTF17, MT98, OMO22, PS05, PG99, Raj99, RY24, Sch99, SK20, Spi12, Sta11, UP03, VZ14, WY17, Wat21, Wat22a, Xu98, Yin01, Zei13, dCdCV23, App93, FG92b, GS94, Kow92, Las96, PAT94, Wha92]. Class- [Li22]. Classes [EOOM17, Far03, NP07, DGZ91, KT94, Kol94b, Tal96b, Xue91]. Classical [BGCD10, Jia10, NPPS16, Rai97, Ued22]. Classification [BLC03, DP12, HL04, BLBE^{+}94]. Clifford [Sta07]. Close [Ver12]. Closed [Bon15, OW98]. Closure [BB97, WY17]. CLT [BL95a, CNX21, CC17, Fan16, HZ07, KZ08, Lev89, MPU97, Nag21, Nag22, OV08, Pel24, Rad98, Rai04]. Cluster [Cer00, EMP17, Kal11, Mas06, Sla89]. Clustering [Ber97, FV05, dHN22, GK91]. Clusters [DG18, Kub20, Sob10, Xue18, Zha01b]. Co [DJ18]. Co- [DJ18]. Coagulation [DGG99]. Coalesce [KM23]. Coalescence [LM16a]. Coalescing [DLZ04, LL99, HLZ13]. Coarse [HM16, Vil20]. Coboundary [Bra96, Mor21]. Coding [Deb10]. Coefficient [OU14, PGS16]. Coefficients [BH22, DQ21, DP22, FG99, FJ05, Fra07, HKHY13, HMY21, HK19, IR24, MPOT19, NS23a, OA23, Qia14, Sun22, ZZ99, ZZ12, Ger92]. Coherence [SK15]. Coin [EV18]. Coinciding [CC11]. Collapse [GK03]. Collatz [ABK16]. Collector [JT24]. Collision [Myt98]. Color [FS21, For22].

Colored [BT10, Yi23, Yok21]. Coloring [Yan13]. Comb [CF20, CF22a]. Comb-Type [CF20]. Combination [TM06]. Combinations [GKM00, KC09]. Combinatorial [CP11, Rai97]. Common [San07]. Commutative [COM05a, LS23, Neu92, PPPA16]. Commutators [SW09]. Commuting [CC17, Rah22]. Compact [AB21, BSC01, CC17, FG00, Jaw97, Jia10, JS00, LT98, LM20, LMX23, MM20, Myr24, Rai97, Sep99, Sep01, Yas00, Bin88, Bin93, Bor89, HS88, Min90, Rub95, Voi96]. compactly [Dob94]. Compactness [OW98, Wan19]. Comparison [FJ12, GL06, GL07a, GL07b, KS22, Krö90, Li92, RRM07, Sar19, Sha00b, Smi15, Wan13, Yan09, RX93]. Compensators [AZ11, CCET06]. Competing [Sar19]. Competitive [Sol03]. Complete [BMP22, FJ17, GC98, KK08, LT91, LRJW95, SGQ22, Yu90]. Completely [SS15]. Complex [CW23, DM23, FG99, Glo13, Roc17, Ubø95, BD93]. Complexes [MPW18]. Components [OV08]. Composition [DO11, HO96]. compositions [CL91]. Compound [App00, CV15, CV21, HK98, JWZ15, KD17, KPR24, LS15a]. Concave [AP97, KL08, LR12]. Concentrated [Rai07]. Concentration [Ald10, BGT10, BC15, DFY01, GZ98, GSS21, Her20, Jaw97, Kho22, Ret03, Ret04, SS22]. Concept [Rao05]. Concerning [BR97]. Condition [BC18, Car02, Che99a, Cop18, FGL20, FKM18, GNT20, JL24, LRS11, RS04, WGYX24, Wu04, XZZ18]. Conditional [BCOV23, BR08a, BK19, DF88, Ejs14, KN23, LHS21, OV08, vZ18, Bou05, Pla92, RS95, Swe89, Zab93]. Conditioned [ADG18, BL97, CKR19, Cor16, MS01a, PP24, Pop21, Raf13, Uch20, Ser96]. Conditioning [ACPM24, Ber93, He17, Kal11, KM23, ST15, Pla89]. Conditioning-based [ST15]. Conditions [AV09, Ahm15, Ahm17, BH24, BR99, CP18, Chi22, Fra21, Gau16, Han18, MT98, MRR04, MPU97, RZ19, Rio09, Vol10, WW09, Xu07, YGHL08, Bud94]. Cone [Ahl15, BR23, HL04, NW02, Sar17, Ber95]. Cone-Like [Ahl15]. Cones [AW04, CKRMT16, HL01, Kol16, Kol17, MN97, PAR07, SU19, Voi09]. confidence [Dip91, Yu89]. Configurations [JW14]. Conformal [LZ19]. conjugate [GH95b]. Conjugates [iS13]. Conjugation [VB23]. Connected [LZ19, Neu08, Neu09, Riv03, Sha00a, Shi13, Neu92, Neu95, NS95, Nob91]. Connection [GH19, GPT15, MPS02]. connections [O’B96]. Connectivity [Mee92]. Consequences [Cha00]. Conservation [Roz13]. Conservative [Abr18, Bai22, Hu20, Owa16, Abr19]. Consistency [BK19, EM00, RR94]. Consistent [ZL18]. Constant [FK20, PP18, Wer05, Ben94]. Constants [Bro03, DH20, Wüb13]. Construct [OMO22]. Constructed [JX20]. Construction [BS20, HTW24, TT18]. Constructions [BdSV22, Kua18]. Constructive [Reb20]. Constructiveness [Her13]. Contact [BM14, BdSV22, Can19, DS00, Han97, Jun05, MPS02, MS00, Swa09, Xue15, Kon95]. Containing [Aid24, Wei96]. Content [KP23]. Context [RZ15]. Contingency [Sim24, Wu23]. Continuations [Has12]. Continued [BBH16, FWS18, Lyo00, Sze09, Ber95]. Continuities [LZ24]. Continuity [Bar00, Bed07, BSC01, BH88, Bre02, Bre05, Gab11, Gue09, HKHY13, JV07, JM15, Kub20, Kul11, Lee97, LMX23, MT98, MR05, OPST24, Pin01, Sol03, WX22, Wan24, Wat00, YZ15, KL96, MR92]. Continuous [Ala22, Aya13, BW19, BHY19, BLP16, BCJM13, CFZ23, Cop18, CMS03, DF97, FM18, FKZ07, HLX18, He22, JK17, Kim22, LPW22, MT98, MR97, MN02, PGS16, Pak20a, RS18, Saa01, Sen00, Sun22, Vil20, WZ21a, ZL18, vZ18, Blo96, DF95, Fil92]. Continuous-State [CFZ23, HLX18, He22, LPW22, RS18]. Continuous-Time [Sen00, Vil20, Fil92]. Continuum [Har01, Tyk09]. Contract [NSB02]. Contract-on-Average [NSB02]. contracting [HS88]. Contraction [Gar04, Gei01, Wu09]. Contractions [ZZ99]. Contractive [Jaw97, Wer05]. Contractivity [Car02]. contractors [Mao92]. Contributions [Sun97]. Control [ABK16, DFS99, KRS10]. Controllability [CP18]. Controlled [SS98]. Converge [He11]. Convergence [Ahl15, Ahm05, Als03, AV15, AIS^{+}24, BP10, BMT99, BL09a, BJL16, BHY19, BL19b, BJP01, Bin93, BGT10, BCG13, Büc15, BK19, BR97, But22, CV15, CB04, CdHH97, CW10, CNW09, CH99, CK08, CM05, Dai13, DM88b, DL00, DPR11, Den20, DM97, Fan17, FKT94, FKX06, FM23a, FM23b, FJ17, Ful08, Fus10, GW20, Gau16, GC98, Git03, Glo13, HH09, He22, HP97, HM03, JM05, JO00, JS00, Jur18, KK08, KKH09, Kol01, KL07, KV11, Kul11, LHS21, LS05, LL17, LW23, LW16, Löc20, MN20, Mad17, MW22, MM03, MZ24, Mor10, MV19b, Muk21, NN07, Nic09, OPST24, Oka24, PT08, PT17, Puh03, Ros03, RMVP24, San07, Sen13, SGQ22, Spa16, Sta22, Ste18, Stu22a, Tei92, TM98, Ued19, VZ23, WH17, Wu99, XY10, XL23, Xue91]. convergence [BB93, BS93, Bud94, D’A95, Den89, DL94, Fer94, GK88, Gri92a, Gri92b, HJ96a, JP93, LRJW95, MR96, Mar88, NP91, Pes95, RY91, RRS92, RX93, ST15, Swe89, Tom96, Yu90, BP13, Blo96]. Convergent [RV04, Rus07]. Converging [Haa07]. Converse [Kol17]. Convex [BL11, CdF00, CAMRR03, DPR00, JP22, Kle13, KL98, Lop20, LWW24, Mil09, NPY22, Sar17, TM06, Vir98, Buc90, BR90, Let93, RRMZ94, RS91, RZ90]. Convexity [DM16a, KRS10, SS15]. Convolution [AJGO23, BM92, BR97, BB98, CR01, CGW20, HM03, JGM11, Jas10, Jaw07, Kuz20, MR97, NA11, Nej94, Wat22a, Zho15, BH88, Bor89, Bud94, DM88a, Eis93, HS88, SZ90]. Convolutions [CC07, FCU12, GKM00, GZ98, HSZ22, Min90, Nic09, Rub95, Voi09, Vol92]. Coordinate [Men16]. Copulas [Res13, ST15]. Coriolis [Sau16]. Corner [Kin95]. Corral [KV03, KV19b]. Correction [Abr19, Bha17a, Cri20a, CF23, KO19a, KV19b, Nag22]. Corrections [LBBB11]. Correlated [NZ22]. Correlation [CM22, Hei21, Hu97, LP03, RW02, XZ10]. Corresponding [DM16b, Kli13, RS24, YY15, Kal92b]. cost [DY95]. Costs [LMX22]. Countable [Her13, PF16, DM88b]. Countably [SK15, KK94]. countably-valued [KK94]. counterexample [Bis90]. counterexamples [Shi93a]. Counterparts [RRM07]. Counting [GLS07, KK22]. Coupled [LWW24, QZ11]. Coupling [DP04, HA19, KT20b, Kuw07, Mur19, Pel01, Wan14b, Web96]. Couplings [BBDP98, KM17, PP18]. Coupon [JT24]. Covariance [BM15, Cap13, JWZ21, Jur18, Kös11, Kro22, LM20, Lyt18, Ma23, MS99, Min21, OS15, Ver12, Zit22]. Covariation [MN00]. Cover [Abe21, Her20, Ald91, BK89, KLNS89, Pal92]. coverage [DR93]. Covering [Mór95, Nam22, Shi13, Zuc89, Ald89a, Ald89b]. Covers [GM11]. Cox [Kal11]. Cramér [BL09c, Che99a, CAFR07, FGL20, Gra94, JPW24, JWZ15, RW05, Sha99]. Cramér-Type [JPW24, RW05, JWZ15]. Crank [Nag15]. Credibility [Pla89]. Criteria [App02, DMV07, GT18, MS22, MP06, San14, ZRP20]. Criterion [AKR99, KLR22]. Critical [ADG18, AC19, AS03, BT16, BM14, BdSV22, Bul14, CM22, He22, Hör07, IW99, Jun05, KN11, Lin18, Rév98, Ste18, Wad19, Wad21, Xue15, Xue18, YWS22, BX16, Kes95, YWC18]. Crossing [AP10, KM98, HMS21]. crossings [Has10, RR92]. Csáki [GL99]. cube [Mat89]. Cumulant [Pri15, Pri19, LR90]. Cumulants [CC07, DE13, Gri92b]. Curie [FKT22, KK18, KT20b]. Current [Kum11]. Currents [FIT14]. Curvature [FK20, PP18, Vil20, Yok21]. Curvatures [KV18]. Curve [Shi97, Shi13]. Curved [KM98, NP24]. Curves [AP97, BCJM13, Pes05]. Cut [Abb17, dT09]. Cut-Invariant [Abb17]. Cutoff [GNS24, Gho20]. Cutpoints [CFR10, TW24]. Cycle [Özd19]. Cycles [GH19, KT16]. cyclic [Fis90]. Cylinders [CLY13]. Cylindrical [Cri19, Cri20b, FM18, KR21, Cri20a].

D [FIT14]. D-Stochastic [FIT14]. Darling [DE18, EM89, KT20a]. Darmois [FG00, Fel15]. Data [BHM08, CF21, DT14, JX20, Kro22]. Davis [WZK23]. Death [ALP22, DG08, DFPS97, DM09, Fil09b, GMZ12, KM19, Wan18a, XZ09, vD17, Tob90]. Death-Process [KM19]. Decay [GP15, GZ98, MS01a, ML17, Ret03, Ret04]. Decision [Sol03]. Decks [NW19]. Decomposability [BK03, MSS17, Wat00, Sie92, WY10b]. Decomposable [Fan16, Pak20a, Rai04, Raj15, Wat22b]. Decomposed [VAÇA21]. Decomposition [AZ01, Ban16, Bur03, Che20, EM18, FG14, FPP21, HJKPS07, Lia09, MS97b, Sie06, Ber93, CLM93, Dob90, DIM94, TZ96]. Decompositions [Nik06, Slo01]. deconvolution [LR88]. Decorrelation [RY23]. Decoupled [BR21]. Decoupling [CG21, KV03, KV19b, AG93]. Decreasing [Cov11]. Deducing [Aus11]. Defect [Ros19]. Defined [Muk19]. Definetti [KS06]. Defining [Din17]. Definite [BR23, Hir03, JT19, Sat92]. Definition [BL22, Ryv15]. Deformation [Luo98]. Deformations [Cap13, RX94]. Deformed [Jur15]. Degeneracy [Hu20]. Degenerate [BK10, Fan19, LX12, Liu21, LH21, OU21, GZ96]. Degenerate/Singular [OU21]. Degenerating [Kim08]. Degree [DM16a, Ejs14, Kho22, OW19, Zuc89]. Degrees [Luc24]. Delay [LT12, LMS24b, Nas22]. Delayed [Kar19]. Delays [Liu18]. deleted [RR94].

Densities [AP97, CFN98, CH21, FCU12, HKHY13, KM11, LBBB11, MLC22, MSM07, OU21, Son06, Szt10, WY17, Wat22b, BRZ96]. Density [Aus11, BJ02, BGT04, BP14, FMW11, Gao03, GOT20, Har22, HT13, IP22, KD17, KS12, KHMCSS01, Nak19, PD23, RR08, Uch18, WW09, Wan19, Bra92, Die91, RC88]. Denumerable [Che20, GMZ12, WCL24]. Deny [SZ90]. Dependence [HL20, KS21, KK21, Lew01, CG96, Sun95]. Dependent [Aid24, BP10, BP13, BH23, BT24, BL09a, BU01, CV15, CV21, Che97b, CS20, CK08, CDM24, DP04, DT14, Gan22, GT09, GPT15, HLZ13, HKW16, Jin17, KS98, KV19a, KM17, Kli13, KKH09, Kro22, Kur07, Löc20, Luo15, Ma23, MST08, Muk21, NS23a, PG99, Pel01, Per97, RY24, RS24, Rio09, Sla22, SD22, UP03, VC18, WZZ22, AGK92, Bra92, Ho95, PS94, Pel96, Pem90]. Depths [AAS18]. Derangements [Smi15]. Derivative [Fan19, GHX19, GZM24, HNS14, JM15, Wan14a, Yu21, YY24]. Derivatives [Tho19]. Descents [CV07]. Determinacy [LS15b]. Determinantal [Cyr21]. Determinants [DT17, MN97]. Determinate [Neu08, Neu09]. Determination [BA15]. Determined [CD20, KM16, Wol10]. Deterministic [KM97, PP18, Fis90, Lew93]. Deviation [BP23, BDS23, BDS24, BC22, CHdHS22, DD97, DW22, FGL20, FWS18, FHL22, GL06, GL07a, GL07b, Gar08, HSW18, JO00, JWZ15, KL08, LL21, LWZ24, LL12, MWW15, Mar23, Puh03, RSZ24, RW05, Sha99, Tak11, Ued19, Wu04, Zha23, EL95, HR18, KC91, O’B96]. Deviations [AGK^{+}13, BC99, Bla00, BR08b, BD97, Bry03, BFW18, Cer00, Che06, Com05b, DH21, DS98, DJ18, DE13, Eic97, EL19, FZ01, FSZ21, FK14, Gao03, GH06, Hec99, Hon03, HN04, JL15, JWZ21, JY22, JPW24, KD17, KS16b, Li99, LL04, LL24b, LV21, Löc19, LM12, Lou12, LMR02, MS97a, MS01b, MP10, Nod17, Nyr09, Pac12, Pak20b, RXZ10, RWZ15, RYZ02, Sla22, Spi12, Tak08, Tra08, Wan11, WH17, Wan17, Yas20, YWC18, Zha17, Zha19, dAC98, vZ18, BM94, BM91, Ilt95, PAT94, Sch94b, dA90]. Diaconis [Hil22]. Diagonal [BS13, LL24a, BLBE^{+}94, Glo91]. Diagonally [CL23]. Diagonals [FL13]. Diameter [AKR99]. Diamond [AC19]. Dichotomy [AR23, MTU14]. Difference [CH21, Jan99, JK17, Raj15, Shi97, LP93, PV94]. Differences [Ben03, CM15, Tei98b, Lev89, Yu90]. Different [Smi15]. Differentiability [Kol98, MN89]. Differential [Ahm15, Aid24, App01, ALT20, ABO22, AO24, BM01, BOLP22, BH22, BR21, BS03, BL19b, CP18, CF21, CL23, CM14, CR22, DO11, DH21, DP22, FU19, Fan18, Fan19, GIST20, GT02a, GOT23, GOT20, HKU12, HA19, HL22, HS21, HT13, HMW24, HZ19, IP22, IR24, Kal20, Kar11, Kar19, KO19a, KO19b, KL14, KP24, Kob11, KM11, KS22, Kul23, LY19, LS21b, LMS24b, LL24a, LY16, LW19, LH21, MRR04, MCR08, MN15, MZ24, Nak19, Nas22, Ose09, Ose11, Ose14, OBO23, OA23, PZ22, RXZ10, RWZ15, RZ19, Sch98, Sun22, Tak10, TT02, VR22, VZ14, WZ21a, WGYX24, XL23, XSR23, Ye24, Yi23, ZZ12, ZM23, dCdCV23, dSEE18, Det89, Kun94a, Mao91a, Mao91b, Mao92, Sif95]. differential-functional [Mao92]. Diffusion [AD11, APP24, Bal07a, BY16, BJ02, BCOV23, CJSS24, CZ17, DW22, Far11, FL16, HJ02, JSS23, KT16, Kuw07, LMS24b, MSY16, PPR22, Pes01, QZ11, QK23, Sch02b, Sin07, Son21a, SYG22, Véc23, YSL08, Yan17, YZ15, Zha11, dCdCV23, Bab94, Ber93, Fen96, Krö90, Øks90, Von93]. Diffusions [AHK99, AP10, AJ22, BL97, BK10, BC22, CP11, Cri23, Fra07, GP98, HR20, Kli13, KLS03, Löc19, Löc20, MS01a, Pac16, SV05, SS21, Tug18, VZ14, Yam24, Yin01, dT09, PS89, Von96]. Digital [GMW17]. Dihedral [McC11]. Dilated [PT04]. Dimension [BM02, BCJM13, CWM04, CR13, CORS17, Che21, FG14, Han03, KW14, KM10, Li16, Oka21, PXY20, Sch98, Yan18, YC24, Le 19, TX96, Zho93]. Dimension-Free [Li16]. Dimensional [AIV24, Bar19, Bar04, BX20, BDK21, BEG18, BT18, Cer00, Che06, Cri23, DT24, FRZ18, God04, HJ02, HKW20, Hue00, JS21, JX20, Jur18, Lee18, LT98, Luc97, Luc10, MS01a, MBP10, Nak19, Pop21, Rus07, SS21, Sau16, Spi12, Uch16, WZZ22, Xue18, Yam24, YBH16, ZBH23, dCdCV23, Ber93, CL23, CFR05, Fis90, Hei90, LL24a, MP14, Sco90, Sie90, Sta11, Tso93b, ZHBS24]. Dimensions [Bur22, Cas97, GPS24, IW99, Kum11, LL03, Mat17, OV14, BPP96, DF95, DY95, GL94, Tel95, Voi96]. Dini [BHY19, HZ19]. Diophantine [Bob18]. direct [Car88]. Directed [AC19, CGHM06, GH19, GM11, Gra12, Led18, Mor10, TV20, VZ23, Yan13, AZ96]. Directions [CO23]. Dirichlet [AE99, BW07, CW17, CMS03, ESC14, Ger92, GP15, GT18, IW24, IMS20, Las20, LL23, SU14, Tel06, WW15]. disappearing [AI96]. Disasters [ML17]. Disc [Uch12a, Uch12b]. Disconnected [AE99, Eva89]. Discontinuous [AIS^{+}24, OA23, Fer89, Son95]. Discounted [YLY22]. discrepancies [Str95]. Discrete [AAO17, BT06, BR97, Cop18, Den20, FM17, FKZ07, Git03, HS98a, JM05, JK17, Kaw13, Key05, Key07, Leh15, LM08, MS22, MM03, MR97, NYY22, PRP24, Ret03, Ret04, RS08, YLY22, Car88, DM96, GS95]. Discrete-Time [Den20, MS22, NYY22, YLY22, GS95]. Discretization [AV15, AIS^{+}24, PR23a, PR23b]. Discrimination [Rya10]. Disintegrations [BR99]. Disorder [CGHM06]. Disordered [AC19]. Dispersion [Voi17b, CLM93]. Dissipation [Jaw07]. Distance [Bar04, Che15b, GGPV18, Kro22, LL17, LMX22, NPY22, PP18, Qin24, Sch16b, XZ09, Xu98, Jan90]. Distances [CFZ23, ES22, JY02, Kom08, MT08]. Distorted [GV18]. Distortion [BHM08, DV11]. Distributed [BP23, BJLO04, EMP17, ORS13, Ver16, Yas20, KS96]. Distribution [AT17, BH23, BJZ19, BM15, BK09, BCOV23, BGEM19, BPY03, BW07, BMS12, Buc90, DM01, Dav05, DM97, DM16b, DLL15, Fil09b, GT02b, GLM05, HM05, HLZ08, JMP12, JS21, Jur18, KD17, KN23, Kle14, KL24, KL08, Lin18, LW11, Mas13, MMS07, Mat12, MNR99a, MNR99b, MN02, NA11, PAS12, Sar17, Tso20, Uch12a, Uch12b, Uch16, Ver16, Wie03, XZ09, Yos20, Zho15, ZBH23, ZHBS24, Ber95, BR90, DM88b, FG92b, GM91b, HKW90, HP94, Has88, Kal91, NS23b, Ruk93, Sun95, Wha92]. Distribution-independent [Buc90]. distribution-valued [FG92b]. Distributional [AIR09, Döb17, GR05, TV16]. distributionally [Num94].

Distributions [All03, AW04, AM12, AMR08, AR06, BMT99, BN13, Bha17b, Bon15, Cas97, CB04, CD20, CNX21, DR09a, DR09b, ES22, FM17, FNS99, GKM00, GMZ12, HR00, IMS20, JQ19, KV19a, KL24, KK17, Kul11, Lac06, LLW11, Lin12, Lou12, MR02, MRU15, MS23, Nag15, NO00, OD15, PS05, PAR07, RY24, iS13, iSU13, SK07, Sch16a, Sie06, Tak19, Ued22, VC13, Wat00, Wat21, Wat22a, Wat22b, Yam24, Yas00, dBM00, Bet95, Bin93, Bra91, Bra93, Bra94, BM92, Bud94, CL91, Dan91, Fie90, Kun94b, Pla92, RS94, RS95, Sie90, Sie94, Von91, Von96, Wha92, Zen95, Bha17a]. Divergence [BL22, CS00, HH09, KM97, Key96, Kim09, Kim17, KLR22, Kli13, LHS21, Xu98, PF91, Shi93a]. Divergence-Distance [Xu98]. divergence-rate [Shi93a]. Divergences [Min21]. Diverges [KS16a]. Divide [FS21]. Divisibility [BNMiS06, PAS12, Ruz88]. Divisible [CdGR20, FM20, MR05, Owa16, PV19, PAR07, RZ97, iS13, iSU13, Ued22, VC18, Wat21, Yas00, BG95, Bin93, BRZ96, DM96, Sha91, Sif95]. Dobrushin [ZZ99]. Doeblin [Chu93]. Does [AMS22, Tak01]. Domain [FC07, Sep94, Sep97, Sep99, CD91, Von91]. Domains [BK03, BF13, Fie90, HMY21, Kim08, Kim09, Kim14, LS11, LWW24, MS97b, Meg02, MSY16, Sch94a, WZZ22, HHV89, HS93, Mee94, MS96, Kim17]. Domany [KKST04]. dominated [KW93]. Domination [Cyr21, Tem14]. Donsker [SW20]. Donsker-Type [SW20]. Doob [Bur03, DIM94]. Double [Kaz21, KM10, LS21b, LX17, Mon19, TT02, Tug18]. Double-Well [Tug18]. Doubly [Ama12, ABO22, AO24, BM01, Kar19, MS02, OA23]. Down [Tak01]. Drawdown [BL19a, VAÇA21, WZ21b]. Drawup [BL19a]. Drift [AHK99, BOLP22, BS03, BK12, CE24, CK00, DFPR15, EM05, HZ19, IO20, JY22, Jin17, Kur07, LS15a, Luo15, TV06, Wan14b, XZZ18]. Drift-Free [BK12]. Drifted [Far11, Sin07]. Drifts [BHY19, LH21, MMW13, iSU13, Ye24]. Driven [Ama12, ALT20, AJGO23, AIS^{+}24, BSW10, BDS23, BDS24, CW23, CK00, DH21, Don08, El 09, FU19, Fra07, GIST20, GOT20, HKU12, HTW24, KP24, KM11, KS22, Kur07, LY19, LS21b, LL24a, LG19, LW19, MS02, MBP10, MZ24, Nag15, NN07, Oka14, OA23, PS05, PZ22, Sun22, SYG22, WZ21a, WGYX24, XSR23, App93, Cho17, GOT23, Kul23]. Driving [BR21, PGS16]. Dual [FM18, FM20, FM23a, FM23b, Kua18, Sud00, Nej94]. Dualities [AJ22, KKST04]. Duality [DFPS97, FFS97, Fil09b, FL16, HM16, LS23, Rig20, Wei96, Fil92, dHN22]. Duals [GR07, SS18]. Dubins [All03, BM99, KT00, Kha09]. Dudley [OV14]. Dufresne [CL99]. Duration [VAÇA21]. Dvoretzky [Men16]. Dynamic [GP01, GPS06]. Dynamical [KO19b, Lal99, Men23, PGS16, Yam24, KO19a]. Dynamically [ZL18]. Dynamics [CO23, GSY24, IR24, Kor16, Fis90].

Edge [ARS21]. Edge-Step [ARS21]. Edges [Gan22, Jun05]. Edgeworth [MJM14, Nam22]. Effective [WB23, Tet91]. Efficient [AG91]. Eigenvalue [BN13, Bro03, CHdHS22, CO20, GKR14, Lyt18, OS15, OR16, Pak20b, XZ10, Zhu23]. Eigenvalues [Bah19, BN22, CC11, DS14, FL14, FK18, HM05, JMP12, JQ19, JWZ21, JW21, LLW11, LW11, MMS07, Wie03, Zha19]. Eigenvectors [Cap13]. Eight [GK02]. Elapsed [Pes01]. Electrical [Sol99]. elementary [DM95]. Elements [ARV97, KS21, KK08, RV04]. Elliptic [AT17, CS00, GOT23, MPOT19, OR16]. Elworthy [Tak10]. Embedding [Neu08, Neu09, OP09]. Embrechts [Wat22a]. Empirical [BGEM19, Ber97, BGT10, Bry03, Büc15, BC22, CB04, DL00, DLSV05, DW22, DT14, Eic97, EM00, ER01, Hec99, JQ19, Jin19, Jur15, Jur18, Kem17, KM17, Kro22, KZ15, KL08, LW23, LL03, NP07, Tra08, Wu99, dAC98, AY94, Arc95a, DM91, HP94, Kol94b, NP91, Str95, Sun95, dA90]. emulates [Fis90]. End [Sta22]. Endomorphisms [CC17]. Energy [KL17, RR08, TV20]. Engel [FWS18]. Enjoying [Law08]. Enlarged [Di 22]. Enlargement [AZ11, KL14]. Ensemble [HM05, HKW16, Pak20b]. Ensembles [BL13, DN12, FKK21, JL15, JQ19, KO19a, KO19b, KK18, KKM13, NNR16, Tri19, VB23, Zha19]. Entirely [AD14]. Entries [HKW16, Jia10, Nem18, ORS13, Tso20]. Entropic [BCK15]. Entropies [CST23]. Entropy [AZ11, BL22, BCG13, Deb14, JS00, Kle13, LL04, NP07, QZ11, Rao05, SJ98, SWF22, WY16, Wu09]. Entry [JX20]. Envelope [Par09]. Environment [ABKV12, AD14, Ban19, CGHM06, CY11, FK14, GMPV10, GKS15, HJ02, HS98b, KMV18, LVZ21, Mor10, Nak13, Tak01, Véc23, WH17, Wan18a, Xu18, dT09, AZ96, Zho95]. Environments [CFZ23, GR23, HLX18, KLS03]. Epidemic [Luc24, Xue18]. Equal [Fit99]. equalities [BCJM13]. Equality [TV16]. Equation [AMS12, AM22, BT10, Bal22, Bau04, BSW10, BZ14, BBDP98, CWM04, CLY13, CF21, CE24, Don08, DM23, Gau16, HA19, HS21, KO19a, KO19b, KHN21, Lee18, LWZ24, LG19, LY16, Mar04, MP24b, Moh22, RR08, Sch24, YY19, YWS22, Cai90, Fen96, HR18, Kun94a, Mao91a, Mao91b, Pap90, SZ90]. Equations [Ahm15, Ahm17, Aid24, AHK99, App01, ALT20, ABO22, AO24, Bag98, BM01, BOLP22, BY16, BH22, BR21, BS03, BL19b, CMP15, CL23, CM14, Cho17, CS00, CR22, DO11, DFPR15, DH21, DQ21, DMCN24, DGR21, DP22, FU19, Fan18, Fan19, GIST20, GOT23, GOT20, HKU12, HL22, HTW24, HT13, HMW24, HZ19, IP22, IR24, Jan99, JM09, Jeg11, KM19, Kal20, Kar11, Kar19, KL14, KP24, Kob11, KM11, KT16, KS22, Küh21, Kul23, Kur07, LT12, LY19, LS21b, LMS24b, LL24a, LTF17, Liu18, LW19, LH21, LWW24, MRR04, MCR08, MN15, Men23, MZ24, MSY16, Nak19, Nas22, NO00, OMO22, OPST24, Oka14, OD15, OT17, OBO23, OA23, PZ22, Qin24, Raj15, RXZ10, RWZ15, RZ19, Sau16, Sla22, Sun22, Tak10, TT02, VR22, VZ14, WZ21a, WZZ22, WGYX24, WY23]. Equations [WY24, Xu18, XL23, XSR23, Ye24, Yi23, ZZ12, ZM23, dCdCV23, dSEE18, Det89, FG92b, HO96, Jan90, Mao92, NS23b]. equilibria [MSSW93]. Equilibrium [DGG99, Löc20]. Equivalence [AJL17, AR23, RZ97, ST06]. Equivalence-Singularity [AR23]. Equivalences [GHLR24]. Equivalent [BPR13, Sot04, WW09, Wat22a]. Erased [Sch08]. Erdos [BCK15, CHdHS22, DK18, DE18, EM89, KS96, Mar23]. Ergodic [App02, DMNS24, Jaw04, Kal02, Kaz21, KM24, OV08, PR23a, PR23b, San16, SWF22, SS98, Sta98, Tan07, Tem23, Wad19, WY16, Xu98, Zwe07, DL94, Web96]. Ergodic-Type [Wad19]. Ergodicity [BH23, BL95b, GHLR24, GR23, LY19, Lop20, MW22, Pis04, Qia14, RZ97, Sen93, Son20, ZZ99, CNP93, FJ95]. Erratum [Cuz01, Dun01, EP15a, GR17, JH12, Kim17, Mas13, Neu09, Uch12b]. Error [AV09, AAO17, CCF^{+}02, CK08, Gau20, KN16, KM11, Nag15, NYY22, NS23b, Qin24, ZM23]. Errors [DD97]. Escape [Hua14, LV12, Wat16]. Esseen [JZ23, AL07, Alb00, Ben94, BBG96, Kar07, ZWW04]. Essential [Sol99]. Estimates [AHK99, AB17, CH21, Die91, Dun00, Far11, FGG16, GOT23, GZ98, Gri10, ISC01, KS12, Küh21, KM08, LS99, Mim13, Mus06, NYY22, Ose11, Sar17, Sim04, Son06, Son20, Szt10, Uch12a, Uch12b, Wan14a, Wan14b, Wan19, Wei14, Wei22, ZM23, KL93, LP93, LR90, Dun01]. Estimating [LPP17]. Estimation [BJ02, BESF23, BESF24, CFN98, CHL08, DD97, FF04, GIST20, Kal99, LR88, LM12, Mal19, AG91, Yu89]. Estimations [JY22]. Estimator [DJ18, PD23, ZZ19]. Estimators [EM00, Gao03, RS91]. Euclidean [DE18, Gan22, GGPV18, HSW18, KL07, Oka24]. Euler [BHY19, CJSS24, DQ21, JSS23, MZ24, NS23b, Yas10]. European [Huj06]. evaluated [HP94]. Evaluations [ZL18]. Even [Rad98]. Eventual [Mao91a]. Everywhere [MS97a]. Evolution [DFPR15, DM23, FNN15, Jeg11, Kol01, KT16, Liu18, WY23, WY24, FG92b]. Evolutions [ISY23]. Evolving [MP24a, VR22]. Exact [BM11, GHLT04, GH21, Gri92a, Hu97, ISC01, LZ24, LMX23, Mur19, Naz09, NN07, Spa04, WX22, Fer94, Fil96, HS92b]. Examples [BR21, CHNR22]. Exceedence [Far03]. Exceptional [Lac05]. Excess [Deb14]. Excessive [HN20]. Excessiveness [PS89]. Exchangeability [Res13]. Exchangeable [CL16, HR00, Hir03, JH03, JH12, Kal90, Kal99, KS06, PF16, Wu04, EM92, TY19, Zab95]. Excited [AIV24, DP17]. Exclusion [FJ17]. Excursion [CH04, FY08, KV18]. Excursions [CH01, CRS01, PPR22]. Existence [AR04, BR99, CL23, Cop18, DGR21, Fan18, FM18, IW99, Kal20, Las96, MSY16, OPST24, RZ10, Sch99, Tak19, Wan14b, WCL24, YZ14, YGHL08, dSEE18, dHN22, AHP91]. Exit [Lac00, Lac06, LS11, MMW13, PRP24, Pis04, Pro21, Son21a, Tug18, Wei22, Von96]. Exit-Problem [Tug18]. Expansion [BD09, FGL20, GH06, Hu97, Hua19, Ina10, KD17, KM11, Mal08, MM10, Red11, Sch09]. Expansions [GH21, GT05, MJM14, Nam22, Pin01, AG93, BGZ93, Kol94b, PAT94, Szu92]. Expectation [HL14, JL24, Jur15, LHS21]. Expectations [Ast23, LL17, Zab93]. Expected [BF13, FH97, LW12, NS23a, Pal92, GMS22]. Explicit [BS05, CL91, Lac06, LMS24b, MBP10, OP09, Son21a, Wan14b, Bor89, Let93]. Exploration [AD07]. Explosion [SU14, Zha18, HMY21]. Explosive [JPW24]. Exponent [Bal07b, Gei01, Key90]. Exponential [AIM15, BH23, BLC03, BL15, Can19, Chi22, CDM24, DM23, FCU12, Gei01, Han18, HL01, HL04, JZ23, Kha09, MW22, MV19b, Qia14, Sie94, Sie06, Son20, Tri24, Véc22, Wat21, Wu99, XY10, Zha01a, BLBE^{+}94, CLM93, DF88, Let89]. exponentiality [Tri24]. Exponents [DP12, KMX18, Luc97, Mön22, MV93]. expressions [Pin18]. Extended [BJ02, BM94]. Extension [AN98, BZ19, CL99, Cha00, Gar04, Gut04, Koc02, MS00, Sen00, Gol88, Has88]. Extensions [LLL15, VA01, VA03, WWLH14, VA96]. Extinction [Can19, HLZ13]. Extrema [Ran06]. Extremal [AJGO23, BBH16]. Extreme [CS20, JWZ21, Sep01]. Extremes [Per18, TZ20, VAÇA21, Voi17a]. Eyring [BD16].

F [Git03]. Faces [Voi17a]. Facilitated [MV19b]. Factor [CD20, DM88a]. Factorization [Gra99, KP14, Raj99, Yen04]. Factorizations [JT19]. Factors [LMX22, OW19]. Fails [OV08, Rad98]. Families [BLC03, BH11, HL01, HL04, Sud00, VC13, BLBE^{+}94, DF88, Let89, Lev89]. Family [AGK^{+}13, Hue00, Wil03, PD91]. Fast [PGS16]. Fastest [Fil09a]. Fat [Kar23]. Favorite [LXY19]. fBm [JM15]. Feature [LM16b]. Feller [AD07, FKZ23, GV18, Gut04, Küh21, Las96, Sch98, Wha92]. Few [Ada10]. Feynman [Bet95, Fit89, MS12, Son95, Tak11, Tho19, Tho20]. Field [FM17, KZ06, KN23, Koc02, PV98, Son20, ZM23, BR08a, Eva93]. Fields [AZ01, AL08b, AB17, AW14, BNSS17, BEK10, BT17, CR13, CS20, Che21, CHL08, Fal02, FM17, GS10, GS11, HJST97, ISY23, Jun14, Kif13, Kin00, Kom08, KL01, KV18, Lac05, LMX23, Ma17, MM20, Ma23, Mal99, MP24a, MST08, NC21, PV19, PT17, Per97, Raf13, RS08, Roz19, Spa04, TZ20, Tem23, Vid20, Yi23, YSH03, YC24, ZRP20, Bra92, KT94, KS93, Mil94]. Figure [GK02]. Filtered [Gab11, Zit02]. Filtering [DL00, KLS03]. filters [Ho95]. Filtrations [AZ11, Di 22, KL14, Men23, MT08]. Finance [FG14, JN99]. Find [AM22]. Fine [NSB02]. Finetti [Bob05, DF88, KRS10, Res94]. Finetti-type [Res94]. Finite [ALL^{+}07, Bar20, BLS12, BHM08, BX18, BX19, BS13, Cer00, Fel04, Ful08, Glu99, Gri24, Hil01, ISY23, KS06, KN23, KN11, Lin12, Luc97, Luc10, MdCQ07, Rai04, Res13, Sch08, Sob10, Uch20, Vir98, Yan13, ZWW04, Adl90, Bud94, DF88, DSC96, Din88]. Finite-Dimensional [Luc10]. Finite-State [ISY23]. Finitely [BPR13, Cas08, GR97, Kuh91, Kun22, LZ19, NW02, NS23a, Spa10]. First [Ahl15, AK16, Chi20, Cor16, Fil09b, Glu99, Lac00, Lac06, LS11, MZ06, Nas22, SV05, Son21a, Von00, vD17, BG95, BL95b, BX16]. First-Hitting [vD17]. First-Order [Nas22, BX16]. First-Passage [Ahl15, Von00]. Fitting [ZZ19]. Fixation [LM16b]. Fixed [KL24, Yu89]. fixed-width [Yu89]. FK [GW20]. FKG [Sah08]. FKPP [Kno18]. Fleming [LZ24]. Flights [OD07]. Flow [Bal07b, Che15b, CZ17, Luo15, Yok21, Zha11]. Flows [Bai22, BD09, Lia99, OPST24, Owa16, Vad10, App93, DM88b, Kun94b]. Fluctuation [AB18, Iso09, VB23]. Fluctuations [BA20, BBDP98, CO20, DGW01, Kum11, LX12, ORS13, TV20, Tri19, Zha17, Zha19]. focal [GH95b]. focal/conjugate [GH95b]. Fock [Bel19, CGW20, KW15]. Fock-Space [Bel19]. Fokker [DGR21]. Fold [GZ98]. Forcing [LM22]. Forest [VZ23]. Forests [AG18, HL03, Sol99]. Form [CS00, CAFR07, FV09, KN23, Kim09, Kim17, Kli13]. Forms [AE99, BT16, CW17, GT18, GT02b, GT05, ISC01, LL23, Myr20, SU14, Wad21, WW15, Yas18, GH95a, GH95b, TT90]. Formula [AAO17, AN98, ABK16, BJ02, BLV24, Bha17a, Bha17b, Def13, MN00, NYY22, Pes05, SV16, Uem04a, Wan14a, Bor89, CGR89, Fil96, Fit89, Joh89]. Formulae [MX04, Tak10]. Formulas [BLP16, Fan19, JM15]. Forward [XL23]. Foster [XZZ18]. Fourier [CdGR20, Has12, HK19, OU14, Pin11, Pin93, Tal92, YWS22]. Fourth [GNT20, NPPS16]. Fractal [AW14, Bar19, BCJM13, CORS17, KMX18, LL03, FG92a, FG94, Mee92, Pin94, Tel95, Zho93]. Fractal-Dimensional [Bar19]. fractality [Yi23]. Fractals [BM02, KN18, Wol10]. Fractional [Ahm15, Ahm17, ALT20, AL08b, AV15, AMS22, AIS^{+}24, BL19a, BT10, BOLP22, BM11, BL02, BLS12, BH99, BGT04, BT12, BDS23, BDS24, Bro03, Bur22, CP18, Chi03, CW10, CNW09, Dai13, Dey20, DP12, Dun00, Dun01, DM23, Fan19, GIST20, GL22, GOT23, Gue09, GHX19, GZM24, GK23, HKU12, HTW24, HN15, HM06, HM09, HN05, IPT22, JZ23, JV07, Jun14, KMV18, KO99, KV19a, KK21, KK22, KK24, KM17, KP24, KL02, LT12, LY19, LY16, LTF17, LM19, MN20, MV19a, Mal06, Mal08, MN15, MS16, Muk19, Nag15, NN07, NO03, NOL07, Oka14, OT17, PPPA16, PT04, Ric17, Ryv15, Sot04, SYG22, TX96, Vad10, Wan14c, Wan19, WX22, Wan24, WX10, XSR23, YSL08, YY19, YWS22, Yos21, Yu21, YY24, Zei18, BX16, Tal96b]. Fractional-Colored [BT10]. Fractions [BBH16, FWS18, Lyo00, Sze09, Ber95]. Fragmentation [DGG99, Haa07, Kno18]. Fragmentations [Dad20]. Frailty [Res13]. Framework [FF04, HL14]. Fredrickson [MV19b]. Free [BGCD10, Ber05a, BK12, CDM05, Cap13, Che21, Dem08, DLL15, Ejs14, Fil09a, GZ24, Gri24, Jia19, Kar07, Kar11, Kuz20, Li16, MS23, NT14, NPPS16, PAS12, TV20, Voi17a, WB23, Yos20, Zho15, Zit22, Kuh91]. Freedman [All03, BM99]. Freely [HT16, MS23, Ued22]. Frequency [FKK21, Nik06]. Friedman [KV19b, KV03]. front [Fil96]. Frontier [KM10]. Fubini [MSM07]. Full [Cyr21]. Function [AR06, AK20, BJS24, Car02, DW24, DFY01, EM00, GL99, Ham01, Jaw97, KMV18, KD17, Kub20, KL08, LP18, LM12, NP07, Pop21, ST15, SS98, Wei14, Die91, LR88, Wei96]. Functional [AMS12, AM22, AFP23, BJL16, BD06, BK15, BHH10, Bin88, CZ22, CK22, Com05b, DO13, DFZ19, DLSV05, ER01, Ein07, FC07, Fra07, GL22, GT01, GR07, GLP03, HZ19, JM09, KL02, KL05, LW04, LM19, LM12, LH21, Mal06, Nag21, Nag22, Oka14, PR23a, PR23b, RS04, RWZ15, RW96, Rus07, Tak11, Tho20, Wan05, WW15, Zha05, AG91, Fer94, Mao92, CLdLO19]. Functionals [BL15, BD97, CH21, Che99b, CH05, CNW09, CF00, CP12, DMNS24, GSS21, Gry21, HK19, KPR24, Kli13, KL07, KL01, Kul11, Lac00, Lac06, Lem21, LL12, Man04, MT98, Pro21, Sch16b, Sik00, Ste16, Tak08, Véc22, Vid20, BX16, CCFR92, CC95, Gol88, Son95]. Functions [Ala22, Als03, ARS21, Ast23, BC15, Bur03, CdF00, DM13, DM16b, Dun21, Fai98, FCU12, FK20, FKK21, Fra21, GS24, GZ98, GNU17, Gru14, HSZ22, HN20, Hir03, IR21, KKR04, KM18, KS08, Kol98, KRS10, LPP17, LMS12, LM20, Ma23, Mal19, Mat12, MM22, Nak19, OW98, ORS13, OU14, Ret03, Ret04, Sen13, SS15, Voi17b, WZK23, Xu13, Zei13, CL91, Dob90, HP94, HS91, HK95, Kol89, Kol94b, Lac89, Pin18, Pin93, PS89, RC88, ST90, Sun95, Von93]. Fundamental [Wad19, Wad21]. Fuzzy [PP02, SU19].

Galerkin [WY23, WY24]. Galois [KZ06]. Galton [ADG18, AR04, Att21, BK00, Bro16, CZ18, He17, IM24, Lin18, Ste18, YWC18]. Game [AdC24, Gut04, GML16, SVT12]. Game-Theoretic [SVT12]. Games [Ney13, Pap11, Xue91]. Gamma [CST23, EHL20, Gau20, NA11]. Gap [BMP22]. Gaps [KO19b, MOP97, KO19a]. Gasket [HJ02, Von96]. Gaskets [Nod17]. Gauss [CD04, Hsu97]. Gaussian [KO19a, Neu09, Arc95b, Arc99, AGK^{+}13, AL08b, Aya13, AW14, BC18, Bas09, BESF23, BESF24, BSC01, BP96, Ber95, BSW10, BW02, BD97, Che21, CK99, CO20, DK18, DLM20, DFMS03, Der03, DN12, GHLT04, GOT20, GGPV18, GLP03, HJN19, Has05, HMS21, Hei96, HS91, Ho95, HK95, Hue00, Jan99, Jol06, KO19b, KL01, KV18, KLmS95, KZ20, KS93, LMS12, Li92, LS99, Li99, LW12, LR93, Lou16, LMX23, Mal19, MR92, MT98, MS01b, Mil09, Min21, Mus06, Naz09, Neu95, Neu08, NC21, NS23a, NZ22, Pak20b, Pap90, PS24, RR08, Roz19, Sep99, Sha93, Sie92, Sot04, ST06, SV16, Ste20, Sto96, Tal98, TZ20, Tel06, TT90, Tri19, Vid20, Wan05, Wei14, Wu09, Xia97, Yi23, YC24, Zha17, Zha19]. GB [Web96]. Gegenbauer [GP11]. General [AHV21, Asc09, Ass19, AL08b, BLV24, Com05b, Di 22, DE18, Fil09a, HMW24, KU23, KS22, Leh15, Lem21, LWZ24, LL12, Mat12, RWZ15, Tal98, Véc22, Voi21, WZY17, Xu07, XZZ18, YY15, ZBH23, ZHBS24, CS91, FJ95, Kol94b, Pin18, WWLH14]. Generalised [AFP23, IMS20]. Generalization [Cla21, Ein07, Eng20, JN99, Man04, Sch02b]. Generalized [Adl90, Adl02, AIV24, Ama12, AKR99, BOLP22, Bob05, BJS24, CD04, Cha90, CO23, EF01, FUW20, FJ12, FY08, GT18, GHL18, Has12, HH10, IPT22, JM09, JGM11, Jas10, KK22, KK18, Kle14, Kun94a, LL17, LG19, LZ16, MS12, MS97b, MN89, MM22, MJM14, MV00, Moh22, OA23, Pap11, Sep94, Sep97, Sep99, SWF22, Sön21b, TTQ14, WY16, WX22, Yos21, GH95a, GH95b, Jan90, Mee94, Tei92]. Generated [BL11, MP10, Owa16, Sch98, BS88, CLM93, Dob94, Gan91, GS95, Kuh91]. Generating [Asc01, Asc09, Ham01, HM08, MLC22, Sch05]. Generation [Lee18]. Generations [AD14, KV11]. Generator [AD07]. Generators [AB21, CL23, Cri19, Cri20b, Fan18, Kar19, KN21, Küh21, LL24a, Uem04b, WZ21a, Cri20a]. Generic [Ste16]. Genus [Red11]. Geodesics [NP24]. Geometric [Ald10, Bro16, CZ17, CH20, FC07, GHLR24, GGX24, MS22, Meg02, RS94, Rai07, Ros03, AHP91, Pfi90]. Geometrically [DMNS24, PV94]. Geometries [FK20]. Geometry [Per97, FG92a, FG94]. Gibbs [KN23]. Gibbsian [Bet23, JK21]. Ginzburg [BBDP98, LG19]. Girko [Ora07]. Girsanov [Liu21, Oss09a, Son06]. Given [BH02, FMW11, Luc24]. Gives [Fil09b]. Glivenko [Bos98, DGZ91, Tal96a].

Global [AMS22, CY11, KL08, Tri19, Mao91b]. GOE [RR19]. Goldie [Wat22a]. Gordin [Bra88]. Gorin [Cla21]. governed [HO96]. GPV [Fen96]. Gradient [OPST24, Sla22, Son20, Wan14a]. Gradient-Dependent [Sla22]. Graham [Hil22]. Graining [HM16]. Graph [BdSV22, FJ17, Häg13, HA19, KLL04, KLLW05, LT23, Luc24, Oka21, Rai07, Sta07, Yan18, Yan13]. Graph-Theoretic [Sta07]. Graphons [OPST24]. Graphs [ARS21, Can19, CHdHS22, Cop18, EK01, GM11, Hua14, Kho22, KN11, Mar23, Nam22, RY23, Rev01, Sch08, Sta22, Sud00, SV20, WB23, Zhu23, Ald89a, Ald89b, Ben91, DS90, Ger88, KLNS89, Pal92, Sal92, Tel95, Zuc89]. Greedy [Lee97]. Green [DW24, Fra21, Pop21]. Greenberg [DS91]. Gromov [Bjö10]. Ground [RR08]. Group [App01, Bro04, BK12, CC07, CR13, Def13, Fel15, Glo13, GZ24, GLM05, HP97, HM03, Jaw07, KZ06, KR03, Lia09, Myr20, Neu08, Özd19, Zei13, Bin88, Eva89, FKT22, HS88, Sal92, Tso93a, Neu09]. group-invariant [Sal92]. Group-Theoretic [CR13]. Groups [ALL^{+}07, App00, BC99, Bar20, BSC01, Bjö10, BB98, CF97, CC17, Cun03, Cza24, DR98, DP12, FZ01, FFS97, FG00, Fel04, FNS99, Ful08, Glu99, GPS06, HS98a, Hil01, Jaw97, Jaw04, Jia10, JS00, JN99, KT20b, Lia14, McC11, Mur19, Mus06, Myr24, NP24, Rah22, Rai97, Ret03, Ret04, Rev01, Sch02a, Sha00a, Sol99, Yas00, BP96, Bin93, Bor89, BR92, CF95, Car88, DM88a, Dan91, DM96, Eis93, FS92, Glo91, HS93, Kal91, KZ06, Kuh91, Neu92, Neu95, NS95, Nob91, RX93, Rub95, Sch94a, Sha91]. Growing [BDK21]. Growth [AKR99, App02, CY11, Dad20, KS22, Nam22, OA23, Ye24, ZZ12, Ben91]. Gruschin [Wan14a]. GUE [KN21]. Gundy [Ish92].

H [GR88a]. Haar [GGPV18, Jia10, Kol94b, ZH12]. Haar-Based [ZH12]. Haar-Invariant [Jia10]. Hail [FKM18]. Hales [Aus11]. Half [CF22a, DW24, IR08, Iso09, Kha12, MSS17, MPW18, TW24, Lal95]. Half-Comb [CF22a]. Half-Line [MSS17, Iso09, Lal95]. Half-Plane [CF22a, Kha12]. Half-Space [IR08]. Hamiltonian [Pah15]. Hankel [DT17, LLW11, LW11, Lou16]. Hanson [LS06]. Hard [QK23]. Hardy [Wei96]. Harmonic [Bag98, IR21, JK17, KKR04, KS08, LL24b, Lin18, Oka24, Pop21, Sen00, Sgi98, Xu13, Kol89, PS89]. Harmonics [CM22, Ros19]. Harmonizable [AB17]. Harnack [BH23, HZ19, LP93, Li16, LH21, Mim13, MS19]. Harris [Als03, Che99b, LL12]. Hartman [Har22]. Hastings [CH20, DS91, Gås06]. Hausdorff [BM11, Cer00, GL94, Hu99, KW14, KM10, Koc02, Le 19, MdCQ07, NC21, Sch98, Yan18, YC24]. Having [AP10, Gei01, LL12]. Hawkes [KPR24, Löc19]. Head [Lew01, MWW15]. Head-Run [Lew01]. Heaps [Abb17]. Heat [BT10, BZ14, CLY13, CE24, Kem17, KP23, KHN21, KM08, LWZ24, LTF17, NO00, Wan19, Xu18, YY19, Bab94, Pap90]. Heat-Type [NO00]. Heavily [Föl03]. Heavy [BT24, BJL16, CFR07, EJU19, FKZ07, HN04, KK17, LBBB11, MS99, MPW18, Nyr09, Owa16, dBM00]. Heavy-Tailed [BJL16, EJU19, FKZ07, HN04, MPW18, Nyr09, Owa16, dBM00, LBBB11]. Height [JS21, LPW22, Pes01]. Heights [CH01, CH04, CCFR16, IR21, DKW91]. Heisenberg [GPS06, MS05]. Hemigroups [BK03]. heredity [Zem90]. Hermite [MN20, Mal19]. Hermitian [BF24, CHNR22, DD18, Dey20, GKR14, JQ17, Nem18, Roc17]. Heterogeneous [KT20b]. Heyde [Fel04]. Hidden [Deb14]. Hierarchical [DS98]. Hierarchies [KT16]. High [BEG18, BT18, DV11, DT24, GPS24, JS21, JX20, Jur18, KM98, Mim13, Pak20b, YY19, DY95]. High-Dimensional [BT18, DT24, JS21, JX20]. Higher [Cas97, DM16a, GHX19, NO00, NYY22, OD07, OD15, Res13, Yu21, DF95, Gri92b, HO96]. Higher-Degree [DM16a]. Higher-Order [GHX19, NO00, NYY22, OD15, Res13, Yu21, Gri92b, HO96]. Highly [JMP12, DS90]. Hilbert [AL08a, Ahm17, AJ22, Bur03, CAMBTD96, Dip91, Gol88, KKH09, Liu12, LZ16, Maj99, MPU97, Mer03, Min21, Pap90, SD22]. Hill [RR08]. Hilliard [CWM04]. Hincin [Bor89]. Historical [AT98]. Hitting [AP97, CH20, Fil09a, FL14, Glu99, GMZ12, Iso09, LT23, MZ06, MW22, MU15, PS15, Uch12a, Uch12b, Uch16, vD17]. Hoeffding [Bis90, Tan07]. Hoeffding-Type [Tan07]. Hoffmann [BR23]. Hölder [Aya13, BHY19, Bed07, BCJM13, FMW11, HKHY13, JM15, KLmS95, Mim13, RS04]. Holderian [Gir17]. Holding [FK14]. holomorphic [MN89]. Homogeneous [AW04, BT16, BC18, BK12, CC11, EM18, Jan99, Kno18, Las10, Lia14, LM20, LMX23, MM20, Puh99, Tyk09, FJ95, Lal91, MMW13, Res94, RS24, Voi96]. Homogeneously [EMP17]. Homogenization [APP24]. Homology [AM23a, AM23b]. Homomorphic [HM03]. Honest [Li22]. Hopf [CD20, KP14]. Hörmander [HPZ21, RZ19]. Hua [DS88]. hull [Buc90, Let93, RS91]. Hulls [Kle13]. Hun [Zem90]. Hungarian [Raj99, Zem90]. Hunt [CRY17, LRS11, BHL93]. Hurst [GIST20, JV07, Ryv15]. Huxley [Moh22]. Hydrodynamic [LM22, Fen96]. Hyper [OD15]. Hyperbolic [Bjö10, LP18, MP24a, HO96, RSS96]. hyperboloid [CLM93]. Hypercubes [Li18, Mon19]. Hyperedge [JK21]. Hypergeometric [Voi17b]. Hypergroup [RX94]. Hypergroups [Gru14, Mat17, Mil00, BH88, LL89, Lei91, Voi90, Zeu89, Zeu94]. Hyperharmonic [HN20]. Hyperoctahedral [GLM05]. Hyperspherical [Ros19]. Hypoelliptic [BJ02, HPZ21]. Hypotheses [Wat21]. hypothesis [GR88a].

i.i.d [Cuz01, KS19, MNR99a]. I.I.D. [Alb00, BBG96, DM97, KM17, KW93, KLR22, MNR99b, MN02, BHW10, Cuz95, Mae93, Sun14]. Ideal [OMO22]. Idempotent [FM17]. identical [Bud94, FJ05]. Identically [BP23, BJLO04, Yas20, KS96, ORS13]. Identification [DM01, HK19, GM91b]. Identities [AP10, Iso09]. Identity [Cla21]. if [Rad98]. II [CHNR22, AD88, AS03, BR21, CC07, CF97, CLdLO19, CG06, DFY01, GH95b, HW92, Lei91, MT88, Ruz88, Tel95]. iid [CO20, Key05, Key07, Bog06, GM98, MWW15]. Image [HM03]. Images [Bur22, PXY20]. Imbedding [CC11]. Immigration [BLP16, Hon03, HZ07, LCP12, LL21, LVZ21, LL24b, Wan18b, YWC18, Zha05]. Impatient [EV19]. Implications [YY15]. Implied [Ros97]. Important [Ano89, Ano90, Ano91, Ano93, Ano96]. Impossible [Rya10]. Improved [GS24, GH19, Vid20]. Inclusion [CP18]. Incomplete [Ney13]. Increase [Ber94]. Increasing [AKR99, DM16a, LMR02, MY09, MS13, Xu07]. Increment [Bal15]. Increments [AB17, Ber97, Chi03, CK99, DM16b, FKZ07, FW20, KL17, LS06, LM19, LL03, Mön22, Muk19, Wan05, GK88, KLmS95, Sha93]. Independence [BA15, CF00, GHS20, HL14, MZ06, BY96]. Independent [AdC24, BP23, BJLO04, Bon15, CST23, CO20, Dan14, FL13, Gås06, GHX19, Jin19, Kab12, KN10, KN16, Kum11, LT03, Myr20, Nem18, NOL07, PGS16, RV04, Sha00b, Tei98b, Wan11, Yas20, ZWW04, Beh89, BX16, Bra94, Buc90, KS96, Sch94b]. Independently [Di 22]. Indeterminate [MLC22]. Index [BF24, FMW11, GIST20, HS92a]. Indexed [BI02, Bal15, HM09, Kra01, Saa01, SS06, SWF22, Yos21, Adl92, DIM94, HM06, IM96, Kal92a, Lac89, LL89, Lei91, MY09]. Indices [ARV97, Leh15, LT98, MŚ21]. Induced [FS21, KZ06, VB23]. Inequalities [AFP23, BR23, Bar04, BHM08, BT06, CW17, DLL15, DMNS24, GR07, GSS21, GP98, Hu97, Jia19, Kha09, KO10, KY12, Kle13, Li16, LL12, Luo22, LH21, MN15, Ose09, Ose14, RS24, Rio09, SS22, Sha00b, Tho20, UP03, UM14, WW15, WZK23, XY10, Yan09, Zha01a, dlP93, Bis90, CS95, DSC96, HK95, LP93, Wei96]. Inequality [BH23, Ben03, Dan14, Eng20, Han18, HZ19, KT00, Kha09, Lem21, LW12, Mim13, MS19, MS05, Sah08, Sch16a, TT98, Tan07, Tan20, Vid20, WY10a, Wu09, Ben94]. Infected [Swa09]. Infection [Xue15]. Infima [HN20]. Infimum [Pis04]. Infinitary [Aus11]. Infinite [BNMiS06, Bar04, Bro16, CZ22, CMP00, Deb14, FCU12, FRZ18, GH21, GLT20, Hua14, ISY23, KS19, Koc02, Kor16, LS15a, Lin18, MRS07, MBP10, OV14, PS13, PAS12, RV04, Ruz88, Sad08, Sch99, Stu22b, WB23, dCdCV23, BD94, Cha90, Hei90, Hor92, LP93, Rév93, Sen93, Sie90, Tso93b]. Infinite-Dimensional [FRZ18, MBP10, dCdCV23, Hei90, Sie90, Tso93b]. Infinite-Measure [Sch99]. infinite-particle [Cha90]. Infinitely [DM96, FM20, MR05, Owa16, PV19, PAR07, RZ97, iS13, iSU13, Sha91, Ued22, VC18, Wat21, Yas00, Bin93, BRZ96, Kuh91, Sif95]. Infinitesimal [AD07]. Infinity [DM02, EM05, Haa07]. Influence [Ruk93]. Information [Che13, HH09, Ney13, Rao05, ZBH23, ZHBS24]. Information-Plus-Noise-Type [ZBH23]. Inhomogeneous [Can19, CHdHS22, DGR21, GK03, Han03, Lia14, Löc20, Mar23, Spa10, SV20, dHN22]. Initial [AZ11, BC18, BOLP22, CLY13, Haa07, HLZ13]. Initial-Boundary [CLY13]. Innovations [BJL16]. Instability [Ben91]. Instantaneous [Che20, WCL24, Ho95]. Instead [EV18]. Insurance [JN99]. Integer [Ahl15, Mat17]. Integers [God04, SK20]. Integrability [CdF00, CG03, CG06, Chi22, JL24, OV14, QSS18, GZ96]. Integrable [CdF00, MT08, CNZ22]. Integral [BZ19, BLV24, CNW09, Har22, HLZ13, Hua19, Jol06, KW15, Lac98, Las20, MRU15, Ran06, iS13, TT18, BGZ93, CGR89, Mao92, NP96, Øks90, Sat92, Sco90]. Integrals [Ala22, AV15, AIS^{+}24, Bai22, BW19, BG11, Bre02, Bre05, DK16, Gar08, HJN19, KL05, LMS24a, LW16, MM22, MV00, MS05, NP18, NN10, NO03, PT08, Pri15, Pri19, SV16, Sta07, ST90, Szu92, Ubø95]. Integrands [AIS^{+}24]. Integrated [DJ18, GLT20]. Integration [BL02, Dey20, KR21, SS09, Tso93a, BM91]. Integro [Ahm15, CP18]. Integrodifferential [Ahm17, DMCN24]. Intensity [Löc20, Mim13]. Interacting [BC22, CP11, Def13, DL00, DPR11, DS98, KW15, Löc19, Sud00, Fis90]. Interaction [AT98, Bry03, LPW22, AI96]. Interactions [AD10, LM22]. Interchange [FJ17]. Interface [BBDP98, HA19]. Interfaces [APP24, Lee18]. Interlaced [Bra01]. Interlacements [Li20]. Interlacing [FL14, Sob10]. Intermediate [CdHH97, WW09]. Intermediately [KS19]. Interpolation [BGCD10]. Interpretation [KM18]. Interpretations [CH20]. interrelation [Kou96, ST15]. Intersection [BGT04, GHX19, IW99, JM15, NOL07, WX10, Yu21, BGT15, GZM24, HNS14, Sza12, YY24]. Intertwining [ACPM24, LL23]. Intertwinings [Ass19, FL14]. Interval [Lac00, Lac06, LL99, PF91]. interval-valued [PF91]. Intervals [Bah19, CL18, Yu89]. Intolerance [HL03]. introduction [Ald89a]. Introductory [Ano88]. Invariance [BK14, CM15, D’A00, DMV07, Gir17, HL04, Jir17, Kif13, KLLW05, Kol16, Mer03, MP06, PV20, Rus07, Sep97, UP03, Wan14c, Zeu94, BL95a, Kou96, Let89, Luo15, Mik94, Pec93, Sad08]. Invariant [Abb17, Bal07a, CD04, CE24, Don08, Jia10, Kal12, Kal20, LMS24b, Mal19, MSY16, PD23, Puh03, Tug14, VA03, YZ14, Zha19, CLM93, CGR89, Hor92, Las96, Sal92]. Inverse [Kol98, MV19a, Mat12, PXY20, Ber95]. Inversion [AP10, FCU12, Law08, Pin93]. Inversions [CV07, Gra15, iS13]. Involving [Son21a, Spa16]. Irreducible [Ful08]. Irregular [AO24, BH22, Nod17]. Irrevocably [Men23]. Ising [For22, GW20, JS99]. Isometric [Lia99, MV00, Ben91]. Isomorphisms [LL23]. Isoperimetric [Bar04, Wüb13, Ger88]. Isotropic [BD09, GKM00, Jun14, LM20, LMX23, Ma17, MM20, CS91]. Itô [Bha17a, Pri15, BL19b, Bha17b, Hu97, Luo15, Nas22, PT08, RS24, SV16, YY15]. Itô-Wiener [Hu97]. Itô’s [AN98, BJ02, MN00]. Iterated [Als03, App02, Arc99, Car02, Che97b, Che99b, Che15a, CHM13, DK18, Den04, ER01, EF01, GM07, Hu99, Jia99, Kal11, KLL04, Kov02, Led18, LT98, LW97, LM19, MCR08, RWZ15, Sep99, Sep01, SW21, SS98, Spa10, TM98, Xia98, Adl90, Arc95a, Arc95b, Cha95, CCFR96, DM91, Gri92a, HP94, Ho95, HPLVS95, KL96, Lac89, Lew92, Lew93, PV94, Sla89, Su96, Voi93, ST15]. Iterates [ANS23]. Iterating [CK14]. Iterations [Pac16]. Iterative [QK23]. Ito [HJST97]. IV [Voi17a].

J [Cuz01, Dun01, Git03]. J.-F [Git03]. Jacobi [AM12, Ass19, AVW24, DD18, Dem08, Gru14]. Jewett [Aus11]. Johnson [RZ15]. Join [SRL07]. Joint [BT24, Bre05, BK19, Ran06, Yan17]. jointly [TY19]. Jørgensen [BR23]. Jump [BD16, CZ18, CJSS24, CDR17, DS98, Fra07, JSS23, LG19, Löc20, Neu08, Neu09, OBO23, SU14, SW21, Son21a, Vil20, Zha18, Cha90]. Jump-Diffusions [Fra07]. Jump-Type [SU14, Cha90]. Jumps [BY16, Del15, DJ18, DM23, JO21, Kaw13, KL14, Mim13, Oka24, Qia14, RZ19, Son20, Tak10, Wan18a].

Kabanov [LMS24a]. Kac [BCK15, Fit89, Jan90, Joh89, KT20a, MS12, Rat20, Son95, Tak11, Tho19, Tho20]. Kantorovich [RW96]. Kantorovich-Functional [RW96]. Kaplan [BS99]. Karolyi [BH23]. Kearns [Sch16a]. Kelvin [BZ06]. Kendall [AJGO23, JGM11]. Kernel [AJ22, BH11, CHL08, EM00, Gao03, Kem17, KM08, PD23, Sat92]. Kernel-Type [EM00]. Kernels [FKZ23, JT19, Zha01a]. Key [Kor08]. Khinchin [Raj99]. Khintchine [CO24]. Kiefer [CH04, Deh98, Gol88]. Killed [IR08, KP23, SV06, Uch18]. killing [Von91, BFW18, KV18]. Kind [Ade22, BESF23, BESF24]. Kinds [QSS18, ZM23]. Kinzel [KKST04]. Klebanov [Myr24]. Knight [Has88]. Known [Fan16]. Kolmogorov [CM00, Gau20, Gut04, HJMLX16, KU23, LQR11, Sau16, Sch16b]. Komlos [Kol94b]. Kramers [BD16]. Kuczek [MS00]. Kuelbs [Kle13].

Labeled [Cyr21]. Ladder [IR21, SV20]. Ladders [DD01]. Laguerre [Ass19, JL15]. Lamplighter [KN18, Rev01]. Landau [BBDP98, LG19]. Landscape [Tug14, Tug18]. Laplace [FHL22, FCU12, JW21, Maj99, NW19, OD15, Pin11, Sch02b]. Laplacian [CO24]. Large [Abb17, ADG18, AP97, AHV21, AO24, AK20, AdC24, BC99, BM15, BP21, BES12, Bla00, BS99, Bou22, BD97, Bry03, BFW18, BC22, CdF00, Cer00, CHdHS22, CG03, CG06, CMP15, CRY17, CK00, CLM04, Com05b, DH21, DLZ04, DS98, DJ18, DW22, Eic97, FZ01, FWS18, FHL22, FK14, Gao03, Gar08, GH21, GP01, Gut04, HL20, He17, Hec99, HR18, HJST97, Hon03, HN04, JO00, JL15, JQ17, KD17, KS19, Kho22, KHN21, LQR11, LL21, LL24b, LWZ24, LV21, Löc19, LM12, Lou16, Lou12, MWW15, MS97a, Mar23, MOP97, NW19, Nod17, NZ22, Nyr09, Pac12, Pak20b, PS15, PAT94, PP02, Puh03, QSS18, RXZ10, RWZ15, RW05, RR00, RYZ02, Sch94b, Sha99, Sla22, Spa99, Spi12, Ste18, SYG22, Sun14, Tak08, Tak11, TM06, Ter16]. Large [Tra08, Ued19, Wan11, Wan17, Wan18a, Wit00, Wu04, XY10, Xue18, Yas20, YBH16, YWC18, ZBH23, ZHBS24, dA90, vZ18, BM94, BM91, CG96, Dob90, Dob94, Dur92, EL95, Hei90, HM95, Ilt95, KC91, LT91, O’B96, Su96, Tom96, Voi90, Voi96]. large-deviation [KC91, O’B96]. Large-Dimensional [ZBH23, ZHBS24]. Large-Scale [DLZ04]. Larger [Adl02]. Largest [JX20, Pak20b, Per18]. Last [CL18, Gra12, Led18]. Last-Passage [Gra12]. later [Tal96a]. Lattice [Ahl15, AC19, CW23, God04, Kab12, Lee97, TW24, Vir98, Wat22a, Zha01b]. Lattices [Xue15]. Law [Abb17, AD11, AHV21, Arc99, AK20, BB13, BD16, BHH10, BR08a, BGT10, BS99, Bou22, Bre02, BB98, Cha10, Che97b, Che99b, Che15a, CRY17, Chi20, CLM04, CGW20, CHM13, Cuz01, Dai13, DK18, Den04, DM07, ER01, EM05, FKM18, FL13, GL99, GNT20, GP01, Gut04, HP94, Her13, HKW16, HJST97, HJ99, HJMLX16, Jin21, JV07, KLL04, KK18, KN21, Kov02, KS16b, Led18, LT03, LQR11, LM19, Mad17, MCR08, Mor10, Neu08, Neu09, PP02, RWZ15, RZ19, RYZ02, Sep97, Sep99, Sep01, SVT12, Sik00, Sot04, Spa99, Spa10, Sun14, Tei98a, TM06, Ter16, TM98, Wan18a, Wit00, XY10, Yan17, YBH16, Arc95b, BY96, BX16, Cha95, CG96, CD91, Cuz95, Dob90, Dob94, FKT94, Gri92a, Hei90, Ho95, HM95, Lew92, Lew93, MSSW93, PV94]. law [RRS92, RSS96, Sep94, Sla89, Su96, Voi93, BX18, BX19]. Laws [ARV97, Adl02, AdC24, BCG13, Bog06, Bos98, Bre05, Bre10, BJS24, CdF00, CL99, CG03, CG06, CAMRR03, DO13, EM18, FC07, GM07, GH21, HT16, HH08, HPLVS95, Jia99, JN99, Kab12, Kem17, KS19, KC09, Kie98, Lac89, LT98, LW97, Meg02, MST08, Pak20a, Pel01, QSS18, RR00, Roz13, San07, SW21, SGQ22, Spa04, SYG22, Tei98b, Voi90, Wan17, YZ15, Adl90, DM91, HHV89, Hei96, KS96, MS96]. Lazy [Hil01]. Least [DD97, KL17]. Lebesgue [HLZ13, Pac12, UA19]. Leffler [BJS24]. Lemma [Aus11, Che97a, WWLH14, dlP93]. Length [Deb10, EHL20, KLL04, KLLW05, Lew01, MS13]. Lent [RZ19]. LePage [BJP01]. Level [GL06, GL07a, GL07b, Hu99, KM98, KL01, MRS07, MU15, Mou07, Zha23]. Level-wise [MRS07]. levelwise [ST15]. Lévy [Cri20a, Kur07, AZ01, Ama12, App00, AB21, AK16, AL08b, BS20, BL15, BLS12, BJP01, BX20, Bla00, BSW10, CL18, CO24, CFZ23, Cho17, CKR19, CH99, Cov11, Cri19, Cri20b, Del15, DM02, Don08, DK16, El 09, EM05, Eva89, EM18, FU19, FUW20, Fan17, FM20, Fra07, FS13, GS94, Gra15, HKU12, HL22, HLX18, KV07, KW14, KMX18, KP24, KS12, KM24, KR21, KS22, KL05, Kul23, KPP14, Le 19, LXY19, Lia14, LZ16, Luc97, LX17, LX20, MTU14, MV19a, MT98, Mar01, Mas21, MBP10, Mim13, OP09, OU21, Oss09a, Oss09b, OA23, PXY20, PS05, Pis04, Rat20, Res13, Rus07, Sav10, SU19, SVT12, Ste16, Szu92, TV16, VAÇA21, Véc22, Véc23, Wan14b, WZ21b, Wol10, WZY17, Zit22, Bor89, KS12, Wan13]. Lévy- [Rus07]. Lévy-Driven [KS22, Cho17, Kul23]. Lévy-Hincin [Bor89]. Lévy-Noise [Fra07]. Lévy-Type [KM24, KS12, Wan13]. Li [Kle13, Tak10]. Li-Type [Tak10]. Liberating [VB23]. Liberation [Ued19]. Libitum [CK14]. libraries [Kow92]. Lie [Neu09, Nob91, BP96, DM88a, FS92, Glo13, HS88, HP97, JN99, Lia14, Nas22, Neu92, Neu95, NS95, Neu08, Pri15, Sch94a, Tso93a]. Life [Bog20, Kov22]. Lifetime [BF13]. Lifting [RS18]. Liftings [MSM07]. Light [MQ21]. Lightning [CDQ22]. Like [Ahl15, Ina10, Jin21, Koc02, Lem21, MP10, PV20, San14, San16, Uem04b, WW15, Yan18]. Likelihood [SW20, BB93, CGR89, Yu89]. LIL [AL08a, Ein07, GZ96, GM98, GK91, KL02, KL05, LS06, Rus07, Sav10, WX22]. Limit [ABKV12, Ahm05, ANS23, AR04, AD11, AVW24, Bai22, BT24, BJZ19, Ban19, BP23, BD06, BK15, BG11, BM18, BH99, BGK17, Bjö10, BCK15, Bob18, Bog06, Bra97, BT17, Buc21, CD04, CZ22, CK22, Che15a, CK00, COM05a, Cri20b, CCFR16, CCFR19, CP12, DO13, Del15, DT24, DS14, DLSV05, DT14, FC07, FY08, FKT22, Fra07, GL22, GT01, GW02, GT05, Gru14, HK16, HSZ22, Hör07, IM24, Jas10, JH03, JH12, JL24, Jin19, JL08, Kab12, KS98, KM92, KO19a, KO19b, KT22, Kha12, Kol98, KN21, KU23, KL01, KV18, KL98, KZ15, LBBB11, LM08, LLW11, LW04, LRR03, LW11, LT23, Lyt18, Mal99, Mal06, MS97a, Mas21, MM07, Meg02, Mer03, Mil00, Mil18, MR97, MNR99a, MNR99b, MN02, NNR16, NP07, Nob91, NN10]. Limit [NT14, OR16, Owa16, PR23a, PR23b, Par12, PP24, PS13, PD23, Per97, RS04, Raf13, RS95, RW02, Saw01, Sla22, SD22, Sze09, Ued22, UB98, Voi96, Voi17b, Vol10, Wad19, Wan05, Wan10, Wan18b, Wil03, XZ10, Xu18, Yin01, YY24, YSH03, Zei13, Zha05, Zha11, ZRP20, Zwe07, Adl92, AD88, AY94, BD93, BB93, BGZ93, BD94, BX16, Bin88, Blo96, Chu93, DM95, DF88, Fen96, Fer94, HS91, Lac91, MR96, Nej94, Neu95, NP91, PS94, Sif95, TT90, Zab93, Zeu89]. Limiting [Ala22, AR06, BM15, BMS12, Cap13, IMS20, JX20, KKM13, Kra01, MQ21, Ora07, RY24, RV04, RR97, ZBH23, ZHBS24]. Limits [Arc94, BT16, CLdLO19, CdGR20, Dad20, DLZ04, FU19, HS21, Hu15, JH03, JH12, Kar23, KZ20, Neu92, Sam24, Uch20, Zit22, Ald91, Cha90, HW92]. Lindeberg [EL19]. Line [CFR09b, CFR10, MSS17, SS06, TW24, Uch16, Iso09, Lal95]. Lineages [Ban19]. Linear [AM12, BN22, BJL16, BB97, CHL08, CO20, DW06, DS14, FU19, FS21, GKM00, GP98, Jan99, KC09, KT16, Lyt18, MPU97, MP10, Myr20, OS15, OR16, Oka14, OA23, PS13, SV05, SYG22, VZ14, YSL08, Yi23, CH05, DM96, LR93, Vol06, Zen95]. Lines [MPW18]. Link [DFMS03]. Liouville [Ben91, GH95a, Sch09]. Lipschitz [Kim17, Als03, CLY13, DQ21, HMY21, Kim09, KV18, Qia14, Von91]. Lipschitzian [DP22]. lists [Fil96]. LLN [HH10, Yas18]. Local [AZ01, AD11, AFK03, AK20, AMS22, BK09, Bla00, BGT04, BK14, Bul14, Che97a, CY11, CFR05, CCFR19, CF22a, CF24, EF01, ESC14, Eva89, Eva93, Fal02, FV09, Far94, FH97, FKT22, GOT23, GS94, GP15, GM07, GNT20, Gue09, GP11, GHX19, GZM24, HK16, HL22, HKHY13, HLZ13, He22, Hsu97, HS98b, HNS14, Hu17, IW99, Jan99, JV07, JM15, KZ06, KO99, Kaw13, KT22, Kim22, Koc02, LS15a, LMX23, MP14, Mön22, Myt98, NOL07, PP24, Pes05, Sam21, Sch24, Son21a, Ste18, Stu22a, Sza12, Véc23, WY10b, Wil03, WX10, Xia98, Ye24, Yu21, YY24, AHP91, ALP22, BGZ93, BGT15, CW17, CS95, CCFR96, DM88a, DM95, GMS22, Küh21, LM22, MR92, NS95, Shi96]. Locally [BSC01, CdF00, Jaw97, Mal19, Myr24, TZ20, Yas00, Bin88, Bin93, Bor89, RRMZ94]. Location [Gan22, Lou16]. Location-Dependent [Gan22].

Locations [Per18, Ran06]. Loewner [HS21]. Log [BH23, GSS21, SW20, BB93]. Log-Harnack [BH23]. Log-Likelihood [SW20, BB93]. Log-Sobolev-Type [GSS21]. Logarithm [Arc99, Che97b, Che99b, Che15a, CHM13, DK18, Den04, ER01, GM07, HJMLX16, Jia99, KLL04, Kov02, Led18, LT98, LT03, LW97, LM19, MCR08, RWZ15, Sep99, Sep01, SW21, Spa10, TM98, Adl90, Arc95a, Arc95b, Cha95, DM91, Gri92a, HP94, Ho95, HPLVS95, Lac89, Lew92, Lew93, PV94, Sla89, Su96, Voi93]. Logarithmic [BH99, BT06, GL06, GL07a, GL07b, Ose11]. Logical [RZ15]. Logistic [AD00, AS03]. Lognormal [Kle14]. Long [BT24, BSF15, CS20, CRS01, CK08, DLSV05, Gra15, Hu20, KK21, Lew01, Lou12, WLG03, Yas18]. Long-Range [CK08, KK21, BSF15]. Long-Range-Dependent [CS20]. Long-Tailed [Lou12]. Long-Time [Hu20]. Longest [Lew01, LMR02, MWW15, MS13]. Longstaff [BH23]. Loop [FZ01, Sch08]. Loop-Erased [Sch08]. Loss [MU15]. Losses [YLY22]. Lotka [He11]. Lovász [Che97a]. Low [BEG18, Hei21, OW19, Arc95a]. Low- [BEG18]. Low-Degree [OW19]. Lower [Ald89b, Aur07, BHW10, DM16b, GHL18, HK98, Key90, LMR02, RSZ24, Tal96b, Tal98, Vil20, Zha23, CAMBTD96, Li92]. Löwner [Bau04]. LSL [GS11]. Lukacs [Kol16]. Lumpings [Pan19]. Lyapunov [Bal07b, Key90, XZZ18].

M. [Bra88]. Macroscopic [Yi23]. Magic [BEG18]. magnitude [HKW90, Wha92]. magnitude-Winsorized [HKW90]. Major [Kol94b]. Majorant [KL08, LR12]. Malliavin [BT12, LT12, Oss09b]. Mallows [Jin19, MS13]. Mandelbrot [BJ19]. Manifolds [AB21, CZ17, GPT15, LW23, Ben91]. Mantissas [BHM08]. Many [MP14, NW02, SK15]. Many-dimensional [MP14]. Map [App02]. Mappings [iS13, BS88]. Maps [AD00, AS03, Cár17, CD04, Oka24, Ste18, Stu22a, BK93]. Marchenko [AVW24]. Marcinkiewicz [LQR11, AHV21, HH10, Su96, Sun14, Tei98a]. Marckert [Git03]. Margin [Wu23]. Marginal [KL24]. Marginals [Mil09, San07]. Mark [LP18]. Marked [Del15]. Markov [AGK92, AK92, Ald89b, AE99, AB18, AJGO23, ACPM24, AR23, BL95a, BI02, Bal15, Ban19, BL06, BL09b, BR21, BK10, Bra01, BK14, Bro00, Cai90, Che99b, CC11, Che20, Cop18, CL16, CC95, CP12, Dad20, Deb14, Den20, DL94, DSC96, DD01, DN95, DMNS24, EF01, EK01, Fil92, Fit89, Fit99, FJ95, GHLR24, GR23, GR88a, GR88b, Glo91, GZ24, GSY24, HH09, HP13, HR20, HM18, HJ89, Hua14, HS92b, Jia19, KM19, Kha12, KO10, Kin95, Kin00, KP14, Kol89, Kor08, Kua18, LM16a, Las96, Law08, Lem21, LS21a, Li99, LCP12, LL24b, Lia09, MS12, MS22, MR92, MRS07, ML17, Nie22, OBO23, PRP24, Pan19, Par09, Pel24, PV98, PD23, RS18, Ros03, RX94, San14, San16, SWF22, SW21, Smi15, Sol03, Spi12, SS18]. Markov [Tak11, Tak19, Tob90, Véc22, Voi93, Von93, VA03, WY16, Wer05, Wil03, WCL24, Xu13, XZZ18, Yas10, Yen04, Zab95, Zha18, dA90, dAC98].

Markovian [BD16, FV05, Luc24]. Marriage [Pac12]. Martin [BG16, CS91, God04, IR08, Lal91]. Martingale [BR21, BPR13, Bin88, CH99, Cri19, Cri20a, Cri20b, CM15, GC98, GLS07, HH10, Hu15, IR24, KS16a, Mer03, MP10, MZ24, QV12, SW21, Tei98b, Tri19, Vol10, WH17, WZK23, Zha01b, AZ96, BL95a, Bin93, Det89, GS95, IM96, LLL15, Lev89, PV94, Str95, Yu90]. Martingale-Like [MP10]. Martingales [Ben03, BHW10, CM14, FGL20, Gei01, GH06, GPT15, Jia99, Lac98, Luo22, MT08, MN00, Ney13, Oka24, Ose14, Saa01, Sam21, Slo01, EM89, Fie90, GMS22, GS95, Mar88, Pes95]. Maruyama [BHY19, CJSS24, DQ21, JSS23]. Mass [AdC24, CZ22, Haa07, Rig20]. Matrices [Ada10, AT17, BR23, Bah19, BN22, BMT99, BN13, BM15, BHS11, BS13, BF24, CO24, CC07, Cap13, Cha10, CR01, CHNR22, CO20, CM05, DR10, DM97, FK18, FL13, GGPV18, HM05, HL20, HH08, JMP12, Jia10, JQ17, JWZ21, Jur15, Jur18, KKM13, Kös11, LLW11, LW11, Lou16, Lyt18, MMS07, MŚ21, MQ21, Mor21, MR97, MNR99a, MNR99b, MN02, Nag21, Nag22, Nem18, Noi21, NZ22, ORS13, OS15, OR16, Pac16, Pak00, PP24, QX20, Red11, Roc17, Ste20, TW14, Wie03, XZ10, YBH16, ZBH23, ZHBS24, Zit22, Bra96, CL91, CNP93, DS88, HM95, Juh90, Sen93]. Matrix [BGCD10, BGEM19, BL13, DD18, DR10, DN12, DS14, Don23, HKW16, KK18, KN21, Kov02, LM20, Ma23, MS99, ORS13, Ora07, PW22, PPPA16, Reb20, Tso20, Ued19, VB23, Ver12, Voi09, Yas18, YGHL08, Zha17, Hec96, TY19]. Matrix-Normalized [Kov02]. Matsumoto [Kol17]. Max [BT24, BGCD10, DH20, Meg02, Ued22]. Max-Infinitely [Ued22]. Max-Semistable [Meg02]. Max-Stable [DH20]. Maxima [App02, FH97, KZ20, Nak19, Far94]. Maximal [CHdHS22, CFR05, Eng20, GP98, He17, HM03, Jia19, KY12, Kuw07, Ney13, Ose14, Pes01, RS24, UP03, CGR89, GMS22, Key90, KW93]. Maximum [Ahm05, AR06, AW14, Bul21, DM01, LS11, PT17, Pfi90, Rév98, Tak18, TW14, VAÇA21, GM91b, Yu89]. Maxtrimmed [GML16]. McKean [BH22, DQ21, LWW24, Qin24, Tug18]. McMillan [NP24]. Mean [All03, BL11, BH02, Bog20, CM14, Deb10, FG99, GH21, HMW24, KS19, KL24, Lin12, LV21, LT23, MV98, Son20, Yok21, ZM23, dBM00, Adl90, Bra91, Let93, Pfi90, PRWW19, RR94]. Mean-Field [Son20, ZM23]. Meander [IO20]. Means [Ald10, BP23, ER01, KL07, TT98, Wu99, Pec93]. Measurability [YGHL08]. Measurable [DFPR15, SGQ22]. Measure [ALL^{+}07, AI01, Bal07a, BM11, Bar00, BM02, BC22, CD04, CRY17, CK00, CW10, DMCN24, Don08, DW22, Far03, Föl03, HLZ13, JK17, Jin19, Koc02, KKM13, KL02, Kul11, LMS24b, Lin18, MR97, Myt98, NC21, Pop21, Sch99, Sgi98, Tel06, Tem14, UA19, YY15, YC24, CLM93, DF95, Din95, DS88, MV93, Pap90, RZ90, Ser96, WWLH14]. Measures [ADHP97, AR23, BA15, BSC01, BPR13, BH02, BGT10, Bry03, BR97, But22, BB98, CF97, Cas97, CE24, Cri23, Deb10, DFMS03, Der03, DN12, Eic97, EM18, FY08, FM20, Gab11, GR10, Gas19, Glo13, Gra15, Has12, HP97, Hu99, JW14, Jaw04, Kal02, Kal20, KL08, Kun22, Las10, LP03, LW23, Luc97, Luc10, MS06, MZ24, Min21, MSY16, Nic09, NSB02, Noi21, OW98, Puh03, Raj15, RRM07, Sad08, Sam21, Sam24, SS15, Sha00a, SW09, Tem14, TTQ14, Tra08, YZ14, dAC98, BD93, BK93, BR92, BRZ96, CF95, DM88a, Din88, DN95, Fal94, Hor92, Kal90, Kol94a, Min90, Nob91, Pin94, PRWW19, RRMZ94, Rub95, Sch94a, Sha91, Sie92, SZ90, dA90, GR17]. Measuring [ES99]. Mechanics [MLC22]. Mechanisms [RY24]. Media [KM08, Mor10]. Medium [CX15, DF97]. Meeting [LM16a]. Mehler [AFP23]. Meier [BS99]. Meixner [BL13, Sta11, SC21]. Membrane [ZX24]. Memory [DLSV05, WLG03, Yas18]. Merge [Meg02]. Merging [KC09, Pap11]. Mesoscopic [Bah19, BN22]. Metastable [DW22]. Method [AM22, Bis96, CB04, CNX21, CNZ22, CNXY24, DW06, EL19, ES22, Gau20, GS24, JSS23, KM19, RZ19, Fen96]. methods [SG90]. Metric [Att21, Cer00, FW20, Jaw07, JO00, Kle13, KU23, LL04, Ma23, Ma24, Maj99, NP07, Tel06, TM06, CAMBTD96, Sat92]. Metric-Dependent [Ma23]. Metrics [LU98, ST15]. Metropolis [CH20, Gås06, MZ06]. Meyer [Bur03, DIM94]. Meyer-Decomposition [Bur03]. Migrating [Bri97]. Mild [BR21, HZ19, Moh22]. Minimal [Als02, Hu15, JW14, Nak13, Yam24, Cai90]. Minimality [CMP00]. Minimizing [Gol88]. Minimum [Hue00, LS11, TW14, Pfi90]. Minkowski [LP18]. Mirakyan [ANS23]. Mirror [Har01]. Mises [EL95]. Mises-statistics [EL95]. Missing [CZ22]. misunderstanding [Bra88]. Mittag [BJS24]. Mittag-Leffler [BJS24]. Mixed [KK21, Yos21, dSEE18, HM95]. Mixing [AV09, Ahm05, Bra97, Bra01, BJ16, BT17, Büc15, CH20, Cun03, GH19, Gir17, Han18, Jaw04, Kie98, Mal99, Mer03, PV19, PS15, Rah22, RS18, RZ97, Zwe07, AY94, BS93, CP93, KT96]. mixture [HS91]. mixture-type [HS91]. Mixtures [JH03, JH12, PAS12, PF16, Bin93]. Mobile [CX15]. Möbius [HM16]. Mod [BA20]. Mod-Poisson [BA20]. Model [BC18, BA20, CDQ22, CMP24, Def13, DS91, Har01, Huj06, JPW24, JS99, KKST04, KV03, KV19b, KT20b, LM16b, Leh15, Lew01, MV19b, NW19, Par12, PP18, Sch02b, YLY22, ZX24, CLM93]. Modelling [CMP24]. Models [BH23, BP21, BSF15, BD06, BEK10, BT12, CP21, CZ22, FKT22, FS21, For22, Han03, He11, JP22, LS21a, LM08, LM22, MJM14, PD23, TV20, Tri24, dCdCV23, BL95b]. Moderate [BC99, BDS23, BDS24, Che06, DE13, EL19, FGL20, FWS18, FSZ21, Gao03, GH06, Jia19, JWZ21, JY22, JPW24, JWZ15, KS16b, LMR02, MP10, Sla22, WH17, YWC18, Zha19, dAC98]. Modified [Bah19, BT06, DS00]. Modulated [FKZ07, Men23]. Moduli [AL07, Gue09, MR92, Wan24]. modulo [HS88]. Modulus [DM16b, LZ24, LMX23, Pin01, WX22, KL96]. Mokkadem [Git03]. Moment [BLP16, Ber05b, Cár17, DT17, DMNS24, Gau16, GNT20, Gru14, GKS15, Ham01, HL20, Kle14, Leh15, LS15b, MdCQ07, NPPS16, Rio09, Sha00b, Zha01a]. Moments [All03, AR04, AIM15, BK09, BEG18, DD18, DN12, Ejs14, GR10, GR17, GHL18, GLM05, Has12, ISC01, LL24b, MMW13, MS06, Mat12, Pin11, Rai04, Tei95, VB23, YY19, Bis96, Bra94, Pin18]. Monge [Hei06]. Monoid [LS23]. Monotone [CGW20, JP22, SS15, SS18, SW92].

Monotonic [KS22, Res13, GHS20]. Monotonicity [CS14, WGYX24, Xu07]. Monro [DW06]. Most [Mar01, SRL07]. Motion [AI01, AJ22, ALT20, AL08b, AV15, AMS22, BL19a, BS20, BM11, BLV24, BPY03, BL11, BF13, BDS23, BDS24, BP14, CX15, Che15b, Chi03, CW10, CCFR19, DH21, DF97, DLZ04, Dey20, DM16b, Dun21, ES99, EF01, EK01, FZ01, FMW11, Föl03, Fus10, GIST20, GOT23, GK02, GV18, GZM24, HN15, HLZ13, HM06, HM09, Hon03, HZ07, Hu99, IPT22, IW99, Jin17, JV07, KO99, KL98, KL02, Lac98, Lac00, Lac06, LV12, LLL15, Lee18, LY19, LS21b, LL24a, Lia97, LM19, Liu21, LZ24, Lop20, LS11, LW19, MN20, Mal06, Mal08, MS16, MY09, Muk19, Nag15, NN07, Nod17, NO03, OMO22, PGS16, PPPA16, PZ22, Rai97, Ran06, RSZ24, Ric17, Ryv15, Sar17, Sot04, Sun22, SYG22, Sza12, TV06, Uch18, VA01, Wan10, WZ21a, WX22, Wan24, WGYX24]. Motion [Wu99, Xia98, XSR23, Yos21, Yu21, YY24, Zei18, Zha05, App93, Cha95, CCFR96, Eva93, GK88, GK91, GS96, IM96, Joh89, KL96, KL93, NS95, Ser96, TX96, Tal96b, Tso93a, Von91, VA96]. Motion-Type [VA01, VA96]. Motions [AIS^{+}24, BH99, Bro03, CO23, CK14, Dai13, DP12, GM18, GHX19, HTW24, KM17, KP23, Ma24, NOL07, PP18, PT04, RY24, Uem04a, WX10, BX16, DF95, RY91]. move [Fil96]. move-to-front [Fil96]. Moving [AK16, KK08]. Mu [AJ22]. Mu-Brownian [AJ22]. Multi [ADG18, AIV24, CL23, FKT22, FV05, HJ02, Kov08, LL24a, MS01b, Ste18, Stu22b, Tug14, Yi23]. Multi-Dimensional [AIV24, CL23, LL24a]. Multi-fractality [Yi23]. Multi-group [FKT22]. Multi-Parameter [MS01b]. Multi-Particle [Kov08]. Multi-scale [FV05]. Multi-Type [HJ02, ADG18, Ste18, Stu22b]. Multi-wells [Tug14]. Multicolor [Dur92]. Multidimensional [Bal07b, BOLP22, DT14, Fan18, FK14, Uem04a, Wat16, Xia98, YY24, Adl92, Pin93]. Multifaceted [GLT20]. Multifractal [Att14, Att21, Bar00, TT18, Fal94]. Multifractional [Aya13, BLV24, FV09]. Multilinear [Sch24]. Multiparameter [Mal06, Mal08, Mou07, Ric17, ST06, Bet95, Ubø95]. Multiparticle [ISY23]. Multiple [Bai22, Bre02, Bre05, GKR14, GZM24, HS21, Jol06, KV11, Kum11, LX20, MR99, NN10, PT08, ST90, Szu92]. Multiplication [ALL^{+}07, Bon15]. Multiplicative [BT10, Bra96, Cap13, FFS97, LRR03, Mor21, Nik06, PAS12, Zei13, Zho15]. Multiplicatively [Hil22]. Multiplier [CdGR20]. Multiray [Yan17]. Multiresolution [ZH12]. Multiscale [BDS23, BDS24]. Multisets [CV07]. Multislice [SS22]. Multistable [FV09, LLL15, Le 19]. Multitype [BLP16, FPP21, GKS15]. Multivalued [RXZ10, RWZ15, Ter16]. Multivariate [APP24, BK15, Büc15, CL99, CNXY24, Chi18, DM01, DM16a, Fan16, FS13, GS24, GH06, Gry21, Has05, Kal99, MP24a, NPY22, Nyr09, PV19, Rai04, Sep97, Sep99, Sep01, BLBE^{+}94, CP93, GM91b, ST15, Sep94, Zen95]. Murray [Ano97, Sun97].

Narayana [AT17]. Nash [DSC96, McG91]. Natural [BLC03, Cri23, KM16, SK07, BLBE^{+}94]. Nature [SW21]. Navier [Sau16, WZZ22, YWS22]. Nearest [CFR09b, CFR09a, CFR10, KLL04, KLLW05, Puh99, D’A95, DP17, Lal91]. Nearest-Neighbor [KLLW05, Lal91]. Nearly [HN20, MS12]. Necessary [Bud94, Chi22, RS04, Wu04]. Negative [CL18, CC11, IMS20, OP09, VAÇA21, Véc23, WZ21b, Shi93b]. Negatively [BP10, BP13, JL08, NP24, Sha00b, YSH03]. Neighbor [CFR09b, CFR09a, CFR10, DP17, KLL04, KLLW05, MS00, D’A95, Lal91]. Nested [AC19]. Net [YLY22]. Networks [Men23, Sol99, McG91, Tet91]. Neumann [Fra21]. Neutral [Ahm15, CP18, LY19]. Neutralities [BW07]. Newton [AHK99]. Nice [RR08, Eis93]. Nicolson [Nag15]. Nikodym [Mar88]. Nilpotent [BC99, HT13, JN99, Neu95, Neu08, Neu09, Sha00a, BP96, Neu92, NS95, Nob91]. Nisio [HJST97]. No [BR21, CE24, Nem18]. Noise [BT10, BC18, BOLP22, BBDP98, Bre10, BC22, CW23, DW22, Fra07, Hei21, Jan99, LWZ24, LG19, LY16, MP24b, MS02, Sau16, Sla22, YY19, Yi23, Yok21, ZBH23, ZHBS24, Bet95]. Noises [FU19, Fan19, LZ16, PGS16]. noisy [Gol88]. Non [Alb00, ALP22, BSW10, BW02, BT17, CW17, CNZ22, CH05, COM05a, CS00, CHNR22, CMS03, Cza24, DP17, DMV07, DQ21, DP22, FG99, FJ05, FV05, Fra07, GT01, GPS24, GT05, Has10, HMS21, HMY21, Jan99, JQ17, Kim14, Küh21, LM22, MMW13, MS01a, Mat17, MPOT19, Nem18, NT14, ORS13, PPPA16, Qia14, RRM07, Res94, RS24, Vol06, WY17, Wil03, Yam24, BBG96, Bud94, FKT94, FJ95, Ho95, Juh90, LR93, Mór95, Sco90]. Non-Absorption [MS01a]. Non-Adapted [DMV07]. Non-causal [CH05]. Non-Central [GT01, GT05, NT14]. Non-closure [WY17]. Non-Commutative [COM05a, PPPA16]. Non-Continuous [CMS03]. Non-crossing [HMS21]. Non-crossings [Has10]. Non-Divergence [CS00]. Non-explosion [HMY21]. non-Gaussian [LR93]. Non-Hermitian [CHNR22]. Non-homogeneous [MMW13, Res94, RS24, FJ95]. non-I.I.D. [BBG96]. Non-identical [FJ05, Bud94]. Non-identically [ORS13]. non-instantaneous [Ho95]. Non-integer [Mat17]. Non-integrable [CNZ22]. non-integral-dimensional [Sco90]. Non-linear [CH05, Vol06]. Non-Lipschitz [DQ21, HMY21]. Non-Lipschitzian [DP22]. Non-local [ALP22, CW17, Küh21, LM22]. Non-Minimal [Yam24]. Non-nearest [DP17]. Non-oriented [GPS24]. non-random [FKT94]. Non-Reversible [Wil03]. Non-smooth [HMY21, Kim14]. Non-stationary [BT17]. Non-Symmetric [RRM07, Juh90, Mór95]. Non-uniformly [MPOT19]. Non-uniqueness [Cza24]. Non-Zero [FG99]. Nonatomic [GR98]. Noncausal [HK19, OU14]. Noncentral [Cha10, Lac91, TT90]. Noncommutative [Bau04, HP97, QSS18]. Noncompact [Gra99, LW23, Gra94]. Nonconventional [HK16, Kif13]. Noncritical [IM24]. Noncrossing [BB13]. Nondegenerate [MZ24, MN00]. Nondeterminism [LMX23]. Nonexplosion [Zha18]. nonextinction [Ser96]. Nonhomogeneous [SWF22, WY16, YGHL08, Zha18]. nonidentical [BM92]. Nonlinear [CE24, DL00, DM23, KLS03, LMS24b, LMX22, LY16, Luo98, MS02, Sch24, ZL18, dCdCV23, BL95b, Die91, HS91, RSS96]. Nonlocal [Ahm15, Ahm17, CP18, HPZ21]. Nonnearest [MS00]. Nonnegative [DM97, Fus10, God04, LS15b, Nik06, PP24, HM95, Sat92]. Nonnormal [HW92]. Nonparametric [LM12, AG91]. Nonrandom [Gri97]. Nonseparable [Wit00]. nonstandard [DM91]. Nonstationary [HS98a, PT17, HP94]. Nonsymmetric [GR88a]. Nontrivial [Har01]. Nonuniform [JZ23, Mal99, Sto96]. Nonunimodular [Häg13]. Norm [Ada10, BHS11, DV11, Li99, Reb20, RMVP24, BR90, CV21, TY19]. Normal [CV07, DM01, Gas19, Gau16, GS24, Gry21, KPR24, LMS24a, NPY22, Sch16b, Sep97, Yan09, Zho15, Bra93, GM91b, HHV89, Sep94]. Normality [CH05, CHL08, DM02, Fan17, Jan99, Kaw13, Bra92, LR90, Pel96]. Normalization [Che99b, HKW90]. Normalized [AR04, Dav05, DM16b, GM98, JPW24, Kov02, LS06, Mas06, Mas13, RW05, CHM13, FSZ21, HW92, Lew92, PS94, Spa16, Tei92, Wan11]. normalizers [Lew93]. Normalizing [AHV21, Bou22]. Normals [JH03, JH12]. Norming [Mee94]. Norms [Bog06, FGG16, LS99, KLmS95, Sal92, Shi96, Sto96]. Note [AIV24, BP13, Büc15, BK19, CHL08, DR98, DLL15, GKR14, Git03, Gri97, KO99, KP14, KKH09, KN11, Luc24, Pac12, Pel01, Rig20, XZZ18, GMS22, Mao91b, Sha93]. Notion [SS06]. Novel [MSS17]. Nuclear [FM18, FM20, FM23a, FM23b]. Nuclearity [Sat92]. Null [AM23a, AM23b, PGS16]. Null-Homology [AM23a, AM23b]. Null-Recurrent [PGS16]. Number [AIM15, AD14, Bag98, Bah19, BDK21, BP14, CFR10, FH97, Key05, Key07, LM16b, Muk19, NS23a, Key96]. Numbers [Abb17, Ade22, AHV21, AK20, AdC24, BS99, Bou22, Bul14, CdF00, CG03, CG06, CRY17, CLM04, Fel15, FM17, GH21, GP01, Gut04, HJST97, KM16, KS19, LQR11, PP02, QSS18, RR00, SK07, Spa99, SYG22, Sun14, TM06, Ter16, Wan18a, Wit00, XY10, CG96, Dob90, Dob94, Hei90, HM95, Key92, Key96, Su96, Voi90]. Numerical [ALT20, ZM23].

Objects [Cyr21, Jon98]. Oblique [Fra21]. Observations [Adl02, FF04, LW04, RR00]. Observed [LW97]. Obstacles [AO24, LV12]. Obtain [SRL07]. Obtuse [FH15]. Occupancies [CZ22]. Occupants [BDK21]. Occupation [CL18, CW10, CCFR19, DGW01, DN95, Din95, JM15, KL02, KPP14, LX12, SS21, WZY17, Yan17, GS96, Ser96]. Occupation-Time [JM15]. occurring [HJ96b]. Ocone [AAO17, NYY22]. Octonionic [CY22]. offs [Huj06]. Offspring [Lin18]. OK [KV19b, KV03]. Olkin [Kol16]. One [Adl92, Adl02, Ben03, BX20, BDK21, BH11, BB98, Can19, Che06, Cri23, Fan18, FGL20, HJ02, Han03, HKW20, Kno18, Lee18, Man04, MS01a, MV19b, Nak19, Pis04, SS21, Sav10, SVT12, Sik00, Spi12, Von00, Xue15, Yam24, Yan09, Adl90, Ber93, Fis90, Hei96, Kal92b, MSSW93]. One-Dimensional [BX20, BDK21, Che06, Cri23, HJ02, HKW20, Lee18, SS21, Spi12, Yam24, Ber93, Fis90]. One-Parameter [Man04, Kal92b]. One-Sided [Adl02, BH11, Fan18, FGL20, Kno18, Pis04, Sav10, Von00, Adl92, Adl90]. One-Spin [MV19b]. Operation [TM06]. Operations [BGCD10]. Operator [Ada10, ANS23, App01, BJ16, Dai13, DP12, GKR14, HHV89, KW14, Luc97, LX17, LX20, Nie22, Raj15, Sie92, Sol14, Sön21b, Wan17, Wan24, MR96, MV93, Sie90]. Operator-decomposability [Sie92]. Operator-Selfdecomposability [BJ16]. operator-stable [MR96]. Operator-Valued [App01, Sol14]. Operators [BL02, FS21, GT02a, Gas19, HPZ21, JW21, Kli13, Kol98, KR21, LHS21, Min21, Pri15, Sau16, Sch98, SGQ22, Tak18, Wad19, Bet95, Gan91, Kal92b, Mee94, Sal92, SG90, Sif95]. Opinions [LM16b]. Oppenheim [GH21]. Optimal [AL08b, BPY03, BH24, GHL18, GP98, HK98, Mal19, Mal08, NP18, Sch09, SRL07]. Optimality [Vol10, Wüb13]. Optional [HS92a, Li22]. Order [Abe21, Ast23, BGT04, Che13, CdHH97, DK18, GHX19, GK23, HKU12, Jin21, JP22, KK24, LPP17, LU98, LL23, MV98, Nas22, NO00, NYY22, OD15, Res13, Tan20, Vid20, Wat22b, YLY22, Yu21, BL95b, BX16, Gri92b, HO96, LT91, Lin12, Tom96]. Order-Preserving [LU98]. Ordered [Hir03, Sah08]. Ordering [Bro16, DM13]. Orders [HR00]. Ordinary [Nas22]. organizing [Fil96]. Oriented [Bro04, MPS02, GPS24]. Orlicz [DV11, KK17]. Ornstein [AM14, BESF23, BESF24, Cha90, GL22, Hec99, JY22, JZ23, Kaw13, Liu12, LZ16, PS05, ZZ19]. Orthogonal [CM00, CO23, Def13, DR10, FK18, GR05, GGPV18, Mat12, NNR16, Ose09, Ste20, vD17, Den89, Voi93]. Orthogonal-Polynomial [vD17]. Orthogonality [Döb17]. Orthomartingale [PV20]. Orthomartingale-Like [PV20]. Oscillating [Löc19, Vo23]. Oscillation [CM14]. Osgood [Fan18]. Other [FV09, Rus07, Zuc89]. Outdegree [He17]. Outliers [Nem18, Roc17]. overhand [Pem89]. Overshoot [Cor16, DM07].

Packing [BM02, Hu99, TX96]. Pairs [GHS20, Voi17a]. Pairwise [LS21a, Sun14]. Palindromic [JMP12, MMS07]. Palm [Kal02, Kal11]. Papangelou [RZ15]. Parabolic [BC18, CWM04, Kim08, Kim14, KHMCSS01, LS11, Mil00, RZ10, VZ14, HO96]. Parameter [CK99, Cov11, IW99, Jia99, JY22, JV07, Man04, MS01b, NO03, RS08, Ryv15, Kal92b, Mao91a]. Parameterized [FW20]. Parameters [AM12, DM01, PS24, GM91b, Mao91b]. Part [Che15b, Pin11, Kes95, Pin18, BR21, Cyr21, Fil92]. Partial [App02, BR21, BK14, Che13, CR22, DO11, DM13, FC07, HZ19, Jir17, Kal20, KP24, KLR22, KZ20, LY19, LY16, LH21, MP24b, Meg02, Nag21, Nag22, OS15, OBO23, Riv03, VZ14, Yi23, ZZ12, CD91, HS93, Kou96, Kun94a, Lac89, Sif95]. Partially [LW97, PF16, Sah08]. Particle [AM14, BDK21, Bri97, BFW18, Def13, DL00, DPR11, JK22, KS98, Kov08, Mor10, Puh99, RZ19, Sud00, Cha90, Dur92, Fis90, MSSW93]. Particles [Bul14, KM23, LL99, MM10, Sar19]. Particular [KV07]. Partition [KMV18]. Partitions [BB13, SJ98]. parts [Tso93a]. Passage [Ahl15, AK16, CL18, Chi20, Cor16, Fil09b, Gra12, Lac98, Led18, Von00].

Past [ISY23]. Pastur [AVW24]. Pastur-Type [AVW24]. Path [Aid24, AB17, BW19, GK03, HL22, HTW24, IPT22, Ina10, KMX18, MV98, Rév93, Ric17, Sön21b, Sta98, Uem04b, WY10a, Mat89]. Path-Dependent [Aid24]. Path-Wise [HL22]. Paths [Cov11, CL16, EMP17, GOT20, Kim22, MY09, Sch98]. Pathwise [BS03, CP21, FKX06, HMY21, Liu18, MPOT19, SV16, SS18, TT18]. Pay [Huj06]. Pay-offs [Huj06]. PDE [RZ10]. PDEs [Kim17, Kim09, Kim14, MS02]. Pearson [Kov22]. Péché [TV20]. Penalised [BA20]. Penalization [DG08]. Percolation [Ahl15, BJ19, BCJM13, Cer00, CORS17, Cza24, DG18, EMP17, GW20, GPS24, Gra12, Häg13, KN11, Kub20, Led18, Li18, MPS02, Sob10, SV20, Tyk09, Zha01b, FG92a, FG94, Mee92]. Percolative [Li20]. Perimeter [BL11, Let93]. Periodic [FUW20, Fra07, PS05, WY23, WY24]. Periodicity [DMCN24]. Permanents [RW02]. Permutation [Bah19, BN22, MS13, Sch05, Tso20, Wie03, EM92]. Permutations [CV07, Jin19, JMPP04]. Perpetual [DK16]. Perpetuities [AIR09]. Persistence [HKW20, Mön22, Pro21]. Persistent [CLdLO18, CLdLO19]. Perspective [Bau04]. Perturbation [BGEM19, Hec96]. Perturbations [GKR14, PW22, Pah15, Sow08]. Perturbed [AHK99, AIM15, CS14, Hu20, PGS16, Sun95, Yin01, Yok21, YZ15, HP94, HS92b]. Petersburg [BGK17, Gut04, GML16, Pap11]. Petrovskii [CWM04]. Pettis [CdF00]. Phase [BP21, CF95, CF97, Cza24, EMP17, FNS99, Har01, JS99, Li18, Vol06, LR93]. Phenomena [MP24b]. Picard [Pac16]. Pickands [DH20]. Pieri [Def13]. Pieri-Type [Def13]. Pillow [Has10]. Pinned [HR20]. Pitman [YY15]. Planar [Abe21, BPY03, BP14, CO23, HP13, Kol01, LV21, Stu22a, Sza12, GS94]. Planck [DGR21]. Plane [BL11, CF22a, CL14, Iso09, Kha12, LP18, Voi17a, Kal90, Let93]. Plus [ZBH23, ZHBS24]. Plus-Noise-Type [ZHBS24]. Poincaré [CW17, Vid20, WY10a]. Point [AR06, BL09a, Bet23, CCET06, CKR19, CP12, FMW11, JK21, KS16a, KV19a, Kul11, LM20, LMX23, MM20, MR05, MZ24, Nas22, Riv03, WY10a, GH95b, HP94, Lal91, Tso93b, Voi96]. Points [AC19, AKR99, BL09a, CM22, Che21, CFR07, Föl03, HN04, KM10, LXY19, LX17, LX20, Vir98, dT09, Buc90]. Pointwise [Aya13, CCF^{+}02]. Poisson [Neu09, AZ11, App00, BL06, BL09b, BA20, CCET06, CV15, CV21, CK00, CD91, DM23, FKM18, GK23, HK98, IMS20, JWZ15, KD17, KK21, Küh21, Kul11, LS15a, LMS24a, LMX22, MV19a, MR05, Neu08, Nov24, Pac12, Pri15, Pri19, Riv03, Sch16a, Sch16b, Sta98, WY10a]. Polar [EM18]. Policies [SRL07]. Polish [Mur19]. Pólya [CMP15, CK22, FKK21, Pem90, Raf13]. Polygons [Par12]. Polyhedral [Sar17]. Polymer [AC19, KMV18, TV20]. Polymers [CGHM06, Mor10, AZ96]. Polynomial [BT16, FH97, FG99, FKK21, God04, Nam22, XZ09, ZZ12, vD17, Ben91, Far94, GS95, Voi90]. Polynomials [AT17, Bob05, DR10, Ejs14, Far03, FJ05, GT01, GR05, GP11, Kho22, Kös11, Muk19, NNR16, OW19, PW22, Sen00, Hei96, Voi93].

Polytope [HSW18]. Polytopes [Gry21]. Population [YWC18]. Populations [dHN22]. Porosities [BJ19]. Porosity [CORS17]. Posedness [Sch24, YWS22]. Position [Hu15, Nak13]. Positive [BR23, Bal07a, Chi20, HH08, Hir03, Jon98, JT19, Par09, Pin11, Pin18, Shi93b, Sie06, TW24, TV16]. Positive-Part [Pin11, Pin18]. Positively [XY10]. Positivity [IP22]. Postulates [RZ15]. Potential [CX15, DMS09, DM23, Far11, RR08, Sin07, SV06]. Potentials [JK21, Kol94a]. Power [AIM15, DM07, FKM18, Gan22, NS23a, Zei18]. Power-Weighted [Gan22]. Powers [Ber05b, CR01, Din17, Jaw07, LS15b, MR97, NA11, Eis93, Hec96]. Precise [GOT23, IW24, Lou12, Spa99, Spa16]. Predator [Kor16]. Predictable [ABO22]. Preface [Muk97]. Preferential [ARS21]. Prescribed [BdSV22, FV09]. Presence [DJ18]. Preserving [Cár17, LU98]. Prey [Kor16]. Price [Voi21]. Primal [ACPM24]. Prime [SK20, Yos20]. Primitive [Sla22]. Principle [BDS23, BDS24, BK14, CHdHS22, D’A00, DMV07, Gar04, Gar08, Gir17, Kif13, KLLW05, LWZ24, Mar23, Mer03, MP06, Pak20b, RYZ02, Spi12, Tak11, Tak18, UP03, Wan14c, Wu04, XL23, BL95a, EL95, Glo91, HR18]. Principles [BL22, CM15, FWS18, FHL22, Gei01, Hec99, Jir17, LM12, PV20, Sep97, Sla22, Kou96, Mik94, Pec93, Zeu94]. prior [RS91, Ruk93]. Priority [Che17]. Prob [Cuz01, Dun01]. Probabilistic [Ade22, ANS23, HTW24, Sep94, WB23, ZZ12, Bab94, Ger92]. Probabilities [Aur07, BL02, Ben03, BGK17, CR01, CAMRR03, DFMS03, Der03, Dun00, Dun01, EP15a, EP15b, GL06, GL07a, GL07b, GT09, GLP03, GH06, HMS21, KKR04, LV12, Leh15, LS05, Maj99, Ney13, Ret04, Roz19, Tug14, Wei22, Wer05, Yan09, Zha23, vZ18, DM96, HS93, KLmS95, Lal95, Shi96, Sif95, Sto96]. Probability [AHK99, All03, AJL17, BS05, Bau04, BA15, BH02, Bon15, BR97, CH21, CF97, CY11, DMNS24, Eic97, FK18, FNS99, Gab11, GR97, Gri24, Has12, HT16, HLZ08, HP97, Jaw04, KM16, Kom08, LBBB11, LU98, Luc97, Luc10, Ma23, MS06, MLC22, MR97, NT14, QSS18, RRM07, Sam24, Sen13, SVT12, SW09, Ste16, Sun97, VC13, Voi17a, Wu99, BHL93, Bra91, Bra94, CF95, Dem90, GM91a, HJ96b, Las96, Nob91, PF91, PRWW19, RRS92, RRMZ94, RS91, Sha93, SZ90]. Problem [AK16, CLY13, CC11, DFMS03, DV11, Gan22, Hei06, IR24, Jaw97, JT24, Kal99, KZS12, Kle14, KRS10, MM07, MdCQ07, Ros97, Tug18, Wat22a, Det89, Ger92, GM91b, KM91, Shi93b, Tal96a]. Problems [Cri19, Cri20a, Cri20b, DW06, DFS99, KV07, Luc10, Naz09, Ald89a, GH95a, GH95b]. Procedure [BK10, Gol88]. procedures [Dip91]. Process [AD07, AR04, BLP16, BCOV23, BM14, Ber97, BdSV22, BSW10, BW07, BGT04, Büc15, Can19, CZ22, CZ18, Chi20, CK99, Cor16, CCF^{+}02, CH04, CFR09a, DK18, DD18, Dem08, DLSV05, DT14, EM00, ER01, El 09, FV05, FKX06, GL22, GM07, GLS07, HKU12, Hec99, HM08, Hil22, Hue00, JY22, JZ23, Jun05, KM19, KD17, KK21, KK22, KK24, KK08, Koc02, Kol01, KL05, KV11, Las20, LS21a, LCP12, LPW22, LL24b, LL03, MV19a, MR05, MRS07, MPS02, Mim13, PRP24, Pis04, QZ11, Raf13, Rév98, RS24, Riv03, Sch09, Sch08, Son21a, Sta98, Ste16, Swa09, Ued19, VA03, Wan18a, WZ21b, Xu98, Yan17, YY15, YWC18, BGT15, CP93, DM91, Jan90, KK94, Kol94b, Kon95, Sha93, Shi96, Tob90]. Processes [ABKV12, AZ01, AN98, Ama12, App00, AJGO23, ALP22, Ass19, AVW24, AGK^{+}13, AK16, Aya13, AIS^{+}24, BT24, BI02, BL09a, BJL16, BM15, BJ02, BNMiS06, BR21, BSF15, Bed07, BL15, BESF23, BESF24, Bel19, BLS12, BD16, Bet23, BB97, BG16, Bla00, Bog20, Bre10, BMP22, BD97, CL18, CDM05, CCET06, CRY17, Che20, CFZ23, CJSS24, CMP00, CH05, CZ17, CDR17, CKR19, CMS03, CNW09, Cov11, CL16, CK08, CKRMT16, CP12, Cyr21, DO11, Dad20, DH20, DLM20, Deb14, DG08, Deh98, DL00, Del15, DM09, DS98, DM02, Don08, DK16, DGG99, DS00, EF01, Eng20, EM05, EOOM17, FUW20, FV09, FV20, Fan17, FFS97, FL16, Fit99, FRZ18, FM18, FM20, FM23a, FM23b, FJ17, FW20, FS13, GHLT04, GMZ12, Gra15, GLP03, GLT20]. Processes [Gue09, GK23, Haa07, Häg13, HK98, HJ02, Han97, HS98a, HL22, HJN19, HMS21, He11, HLX18, He22, Hec99, HN05, HNS14, IMS20, JK21, Jin21, JSS23, Jol06, KL17, KV07, KS98, Kal11, KV19a, Kaw13, KW14, KM17, KS08, KO10, KP24, KY12, KP14, Kno18, KS12, KM24, KR03, Kon98, KR21, Kov08, KU23, KL01, KZ15, KS16b, Kur07, Kuw07, KPP14, LR12, LS15a, Law08, Les10, LS99, Li99, LL04, LXY19, LL21, LVZ21, LW23, Lia09, Lia14, LW97, LRR03, LRS11, Liu12, LZ16, LZ24, LL12, Löc19, LX17, LX20, MS12, MiS99, MTU14, MS23, Mal19, MT98, MR99, Mar01, MR05, MV98, MS01a, MS01b, Mas21, Mat17, ML17, MPU97, MP10, MBP10, Mön22, MOP97, MS00, Mou07, Naz09, NO00, OP09, OU21, OT17, Ose09, Oss09a]. Processes [Oss09b, OBO23, Owa16, OA23, Pak20a, Par09, PXY20, PS05, Pel01, PS13, PS24, Pro21, QV12, RZ15, Rat20, RZ97, RMVP24, Rus07, Rya10, SS06, Sav10, Sch98, Sen00, SW21, Sim04, Sol03, Son06, SV06, Sot04, ST06, SV16, Spi12, SS18, SW20, Tak08, Tak11, Tak18, Tak19, Tem23, TV16, Tug14, TM98, Uem04b, VAÇA21, Véc22, Vil20, WLG03, WY10a, Wan13, Wan14b, Wat02, Wol10, WZY17, WCL24, Xia97, Xu98, Xue15, Yan18, Yas10, YSH03, YZ15, ZZ19, ZH12, Zha18, Zit22, vD17, AD88, AG93, AY94, Arc94, Arc95a, Arc95b, BB93, Ber94, Blo96, Bru91, Cha90, DR93, Die91, Din88, EM92, Eva89, Fer94, FG92b, Fit89, Gan91, GS94, GR88a, GR88b, Glo91, HO96, HJ89, HPLVS95, HS92b, JP93, Kal92a, Kal92b, Kin95, KT96]. processes [Kol89, Krö90, KC91, KLmS95, Lac89, LL89, Le 19, Lei91, LR93, MT88, MR92, MN89, Nea92, NP91, NP96, Num94, PD91, Rév93, Ros94, RC88, Sto96, Tso93b, dlP93]. Product [Bar04, JQ19, Jin19, Lyt18, MR99, Nem18, Sad08, Tem14, Beh89, Hor92, NP96]. Production [QZ11]. Products [BR97, CO20, DM97, Eic97, FK18, HH08, JMPP04, Kab12, LHS21, LW12, LS15b, MNR99a, MNR99b, MN02, MSM07, QX20, Sch02a, Wie03, Yas10, BM92, Bud94, CL91, CNP93, Neu95, Sen93]. Profile [CHNR22]. Progressive [KL14]. Projections [Fal89, Jur15, MS06, Mec12, Men16]. Projective [DMV07, MP06, Rio09, Sam24, ZRP20]. Prokhorov [CLM04]. Proof [Fil09b, DM95, Hei96]. Proofs [CM14]. Propagation [Di 22].

Properties [AIR09, AB21, AJGO23, AMS22, Bar19, Bet23, CL99, CHNR22, CH01, Cyr21, DW22, Ejs14, GZ24, HL04, IPT22, JW21, Kal02, KMX18, Kho22, Kin00, KM24, KHN21, LU98, LP03, LCP12, Li20, LTF17, ML17, Nak19, PV19, Rai97, Ric17, Ryv15, SS15, Sie90, Sön21b, Sta98, TT02, Uem04b, Véc22, Voi09, Wat00, WY17, Wu23, Xia98, Buc90, Eva89, Mat89, Rév93]. Property [AD07, AP10, BI02, Bal15, Cap13, Di 22, Eng20, GV18, HKHY13, HMY21, Jas10, Kaw13, Kol17, Kun22, Law08, LMS12, Liu12, MU15, San16, Tak19, Wan14b, Ben91, Ger88, Let89, Mar88, Str95, Web96, Zem90]. Proportion [RR97]. Proximity [DR09a, DR09b]. Pruitt [Gri10]. Pseudo [BR21, DMCN24, GKM00, HKU12, Sch98]. Pseudo- [DMCN24]. Pseudo-Differential [HKU12, Sch98]. Pseudo-Isotropic [GKM00]. Pseudo-Partial [BR21]. Pseudostable [JM09, MM05]. Publications [Ano97]. Pure [SW21, Vil20, Zha18]. Pure-Jump [SW21, Vil20]. Purely [GR98].

Quadrant [Fra21]. Quadratic [BD97, CL23, GT02b, GT05, Man04, MN00, Sch24, Slo01, Yas18, GH95a, GH95b, TT90]. quadric [Gan91, SG90]. Quantile [Deh98, KN10, KZ15, RR92]. Quantiles [Bra97, KN16, Sun95]. Quantitative [DS14, RY23]. Quantities [AIM15]. Quantization [DFMS03, Der03, DV11, GLP03, JP22]. Quantum [AD10, Bar20, FFS97, FNS99, KW15, Mat17, MLC22, AK92, App93, BHL93]. Quasi [DM09, FGG16, Hor92, KT16, Luo15, Sad08, Tak19, Yam24, Ben91]. Quasi-invariance [Luo15, Sad08]. Quasi-invariant [Hor92]. quasi-isometric [Ben91]. Quasi-Linear [KT16]. Quasi-Norms [FGG16]. Quasi-stationarity [DM09]. Quasi-Stationary [Yam24, Tak19]. Quasistable [BSF15]. Quaternion [YBH16]. Quaternionic [BDW21]. Quenched [FK14, HZ07, PV20, Pel24, Stu22a, ZRP20]. Question [Bra97, LP18]. Quotient [Kol16].

Rademacher [ARV97, UM14]. Radial [Che15b, HS21]. Radii [JQ17, MQ21, QX20]. Radon [BL09c, Mar88]. Random [Abb17, Abe21, Abr18, Abr19, ALL^{+}07, Ada10, ARV97, ABKV12, AAS18, AT17, All03, ADHP97, Als02, Als03, AIM15, AB18, ARS21, AIV24, AD14, AKR99, AR06, Arc99, Asc01, Asc09, AFK03, AK20, AD00, AS03, AD10, Att14, Att21, Aur07, AG18, AdC24, AL08b, Aya13, BP10, BP13, Bag98, BR23, BN22, BW19, BC99, BS05, Ban16, BJZ19, BZ19, Bar20, BP23, Bar00, BJLO04, BK09, Bed07, BPP96, BK00, Ber05a, BS88, BM02, Ber95, BES12, BL11, BEK10, BH02, BC15, Bon15, BR08b, BMS12, BS13, BF24, BT17, BMP22, Buc21, Bul14, Bul21, BK12, BT18, CM22, CO24, Can19, CC07, CGHM06, Car88, Cas97, CLdLO18, CLdLO19, Cha10, CR13, CHdHS22, CST23, Che97b, Che06, CS14, CX15, CS20, CFZ23, CHL08, CK00, CM00, CO23]. Random [Cla21, CC17, COM05a, CY11, CS00, CHNR22, CO20, CH99, CRS01, CFR05, CFR07, CFR09b, CFR09a, CFR10, CCFR16, CCFR19, CF20, CF22b, CF23, CF24, Cun03, CDM24, Cuz01, Dan14, DP17, DG18, Den04, DW24, DM13, Der03, DR10, DN12, DT17, DS90, DS14, Don23, EL19, Ejs14, EV19, EJU19, EM05, EHL20, EP15a, EP15b, Fal02, FV20, Fan16, FH97, FG99, FJ05, FF04, FSZ21, FK20, FK18, FKZ07, FH15, FP10, FK14, Ful08, Fus10, GKR14, GR10, GR17, GR97, GR98, GMPV10, GGX24, Gas19, GS24, Gei01, GT09, GR23, Ger88, Gho20, GM11, GM98, Glu99, God04, GHL18, Gri97, GK03, Gry21, GP01, GPS06, GP11, GKS15, GS10, GS11, Ham01, HLX18, He22, HH08, Her13, HP13, Hil01, HM08, Hil22, HR00]. Random [Hir03, HJST97, HJKPS07, HZ07, HS98b, HN04, Hu15, ISC01, IR08, IR21, JM09, JH03, JH12, JQ17, JX20, JL08, JMPP04, Jon98, Jun14, Jur15, Kab12, KS21, KKR04, Kal12, Kar07, KS19, KM97, KM98, KN23, Kho22, Kie98, Kif13, KK08, Kin00, KN10, KN16, KLR22, KKH09, Kol01, KL24, Kor16, KLS03, Kov02, KV18, Kub20, KZ20, KM08, KN18, Kum11, LZ19, Las10, Les10, LP18, LT98, LT03, LLW11, LVZ21, Lia97, LS15b, LW11, LV21, Lou16, LT23, LMX23, Luc24, Lyo00, Ma17, MM20, Ma23, MMW13, Mad17, Mal99, MP24a, Mar23, McC11, MM07, Mec12, MS99, Men23, MPW18, MP10, Mil00, MS19, MM10, MM05, Mon19, Mor10, MST08, Mor21, MS13, Muk19, Muk21, MNR99a, MNR99b, Mur19, Mus06, Nag21, Nag22]. Random [NW02, Nak13, Nam22, Nem18, NC21, Nic09, NS23a, Noi21, Nyr09, OS15, OR16, OW19, Oka21, OD07, OT17, Özd19, Pac12, PW22, Pah15, Pak00, PPPA16, Par12, PV19, PP24, PG99, PT17, Per97, PS24, Pop21, PP02, QX20, Raf13, Rai07, Rai04, RY23, Reb20, RW02, Rev01, Rio09, Riv03, Roc17, RV04, Ros19, RS08, Roz19, Sch02a, Sch05, Sch08, Sen13, Sep97, Sep99, SJ98, Sha00b, Sim24, Smi15, Sob10, Sol14, SD22, Sön21b, Sow08, Spa04, Spa10, Sta11, SC21, Sta22, Ste18, Ste20, Sun14, SS09, Sza12, TT98, Tak01, TV20, TZ20, TW24, Tei98b, Tem23, Tet91, TTQ14, TT18, TW14, Uch16, Uch20, UP03, VZ23, VB23, VC18, Ver16, Vir98, Vo23, Voi09, Voi17b, Vol06, WW09, Wan11, WH17, Wan17, Wan18a, Wat02]. Random [Wol10, Wu99, Wu23, WY23, WY24, XY10, Xu18, Yan13, Yas18, Yas20, Yi23, YBH16, YSH03, YC24, Zha17, ZRP20, Zha23, ZWW04, Zhu23, dT09, Adl90, Adl92, AZ96, Ald89a, Ald89b, Beh89, BD94, Bra92, Bra96, Bra94, Buc90, CS91, CL91, CNP93, Cuz95, D’A95, Den89, Din95, DKW91, Fal89, FG92a, FG94, Far94, FKT94, GM91b, HP94, Hei96, HM95, Juh90, KLNS89, Kal90, Kal92a, Kes95, KS96, KT94, KM91, Kou96, Kuh91, KS93, Lal91, Lal95, LP93, Let93, Lew92, Lew93, LRJW95, Mae93, Mat88, Mat89, Mil94, Min90, Pal92, PS94, Pel96, Pri96, RY91, RX93, RC88, Rub95, Sch94b, Sco90, Sto91, Tal92, TY19, Zen95, Zeu94, Zho93, Zho95, Zuc89]. Random-Time [Jun14]. Randomization [Pem89]. Randomized [Tso20, GH95a, GH95b]. Randomly [Bri97, Kra01, BK93, RR94, Tei95]. Randomness [Tak01]. Range [BT24, CS14, CS20, CK08, Ham01, HSH20, KK21, Lew01, LV21, MP14, MOP97, NC21, Oka21, SV05, TV06, Xu13, Yan18, BSF15, Dur92, LP93, Le 19, RS91, Zho95]. Ranges [Che06]. Rank [Can19, CF21]. Rank-Based [CF21]. Rank-One [Can19]. Ranked [CH01, CH04]. Rapid [BM14, GZ98, Kon95]. Rare [GNS24, Key92, Key96]. Rarely [ES99]. Rate [Als03, AKR99, AV15, AIS^{+}24, BMT99, BHY19, BL19b, Ber97, BdSV22, CdHH97, CM15, Fan16, GP15, HJ96a, Hua14, Huj06, Kub20, MN20, NN07, NP91, QZ11, Ret03, XL23, CP93, D’A95, Gri92a, Shi93a]. Rates [Ahm05, App02, CB04, CRY17, CY11, Den20, FJ17, Ful08, Gau16, GH19, GK88, GK91, KL07, LS05, LL21, LW23, MR96, MW22, MS01a, NP07, RY91, Ros03, Tom96, TM98, Wat16, XY10, Xue15, Fer94, Key96, RX93]. Ratio [Ald10, FJ95, Kha12, BB93, CGR89]. Rayleigh [Sch08]. rays [GS94]. Reactant [FKX06]. Reaction [BY16, MSY16, dCdCV23, Fen96]. Real [Bag98, FK20, FK18, GLM05, HM05, JWZ21, MMS07, Mat12, Muk19, NA11, Red11, DS88]. Real-Valued [JWZ21]. Realizations [JT19]. Realized [FHL22]. Rearrangements [CM00, CLM04]. Reciprocal [CDR17]. Records [DFZ19, GLS07, Key05, Key07, Riv03]. Recovering [Hue00]. Rectangular [BG11, JL15, QX20]. Recurrence [Abr18, Abr19, Als03, DG18, GT18, HP13, KO10, Kon98, MTU14, San14, XZZ18, CLdLO18]. Recurrent [AD14, Bal07a, Che99b, KM08, LL12, McG91, PGS16, Von93]. recurrently [KW93]. Recursion [MŚ21]. Refined [Wan11]. Refinement [LS05, LQR11]. Refinements [KT00]. Reflected [Ama12, ABO22, AO24, CF21, CL23, DP22, El 09, EP15a, EP15b, Fus10, GV18, HMW24, JY22, Kal20, Kar19, LL24a, OBO23, Pis04, Sar17, WZ21b, WGYX24, Xu07, YZ15, ZZ19, Ber93]. Reflecting [CZ17, DFPS97, MRR04, YZ14]. Reflection [Cza24, JO21, Kuw07]. Reflections [BT18, LS21b]. Refracted [KPP14]. Regenerative [SS98, KC91, Num94]. Regenerativity [LS21a]. Regime [CJSS24, DT24, JSS23]. Regime-Switching [CJSS24, JSS23]. Region [Mor10]. Regions [LZ19, Dip91]. Regression [BT12, DD97]. Regular [Jas10, KT20a, KT22, Kuz20, ORS13, SD22, Xue15, vZ18]. Regularity [Aya13, Bre10, CMP00, FRZ18, HL22, Kim22, Liu12, LZ16, MV98, Mim13, NO03, RZ19, Sch99, WX10, Xu13, YGHL08, Arc95a]. Regularization [HL22, MP24b, Reb20]. Regularized [Min21]. Regularly [Bou22, Bul21, FS13, Wat22b]. Reinforced [CZ18, FP10, Vol06]. Reinforcement [Vol06]. Reinforcements [Kov08]. Related [Als03, AIM15, AIS^{+}24, Bre10, DO11, GLS07, HNS14, KO10, KLR22, Lac98, Lac06, LM08, MT98, NNR16, Nie22, NO00, OT17, RYZ02, Xia98, BY96, Cha95, RC88]. Relates [KN21]. Relation [Cri23, Gra15]. Relations [ACPM24, DFPS97, Döb17, Les10]. Relative [BCG13, DM07, KN16, OW98]. Remark [Che99a, GM91a, JP93]. Remarks [Ada10, BMT99, Bra93, Ish92, Ros97, Yos20]. Removal [Aus11]. Removed [Jun05]. Renewal [Bro04, Chi18, GT02a, KT20a, KT22, Kor08, Sgi98, Yam24, Cai90]. Renewed [IM24]. Renormalized [CX15]. Rényi [KS96, CHdHS22, Mar23]. Repeated [AD10, Ney13]. Replacement [SS22]. Representation [AL08b, Bas09, BJP01, Bob05, Bre02, Di 22, Din17, ESC14, FS13, GSY24, JK21, KZ06, MBP10, PAR07, Roz13, SS15, WB23, ZZ12, ZL18, BHL93, MS96, Ros94, RW96]. Representations [App01, For22, Ful08, GLM05, Har22, KW15, LLL15, MRU15, Wei14, Yas10, HHV89]. Reproducible [BLC03]. Reproducing [AJ22, BA20]. Rerooting [Stu22b]. Rescaled [BEK10, FKX06, He11]. Residual [Bog20, Hu97]. Residue [SK20]. Resistance [WB23, Zho93, Tet91]. Resolution [DV11, ISY23]. Resolvent [Che20]. Respect [AV15, BLV24, Bon15, Che15b, Dey20, Gei01, HJN19, JV07, KR21, NO03, Wei96]. Result [CS14, Sha99, UB98, dSEE18, Hec96, KC91]. Results [Als03, AL08b, BU01, BK19, BR97, CM14, DGR21, IO20, KS22, KL02, KV11, LM16b, MRR04, Oka14, Oka21, PXY20, PG99, SRL07, Son21a, Yin01, BY96, Bra88, Cha95, DS91, Li92, Shi93b, Wei96]. Retarded [LY19, Liu12, Liu18]. RETRACTED [BP10]. Retraction [BP13]. Retrieval [CF97, FNS99, CF95]. Return [AV09, BS05, EP15b, Lal95, EP15a]. reversal [Tso93b]. Reversals [Sch05]. Reverse [CM15]. reversed [Pes95, Str95]. Reversible [CP12, DGG99, Puh99, Wil03, Ald89b, Zho95]. Reversiblizations [CH20]. Révész [DK18]. Revisit [Che97a]. Revisited [CF22b, CF23, Tei95]. Rho [Bra01]. Rho-Mixing [Bra01]. Ricci [Che15b, Vil20]. Riemann [Sch09]. Riemannian [AB21]. Riesz [HL01, HL04, HLZ08]. Rifle [Lal99]. Rigid [Lia97]. rigorous [DS91]. Rings [Bro00]. Risk [ABK16, Leh15, YLY22, Ruk93]. Risk-Sensitive [ABK16]. Road [Yan13]. Robbins [DW06]. Robust [JL24]. Roots [Bag98]. Rosenblatt [Ano97, BGT15, Sun97]. Rotation [VA03, Yas20]. Rotation-Symmetric [Yas20]. Rotor [HSH20]. Rough [Aid24, BC18, CMP24, GOT20, HTW24, HT13, Ina10, IP22, LWZ24, MPOT19]. Row [RV04]. Rowwise [RV04]. Rubin [Kol16]. Ruin [Leh15]. Rule [HN15, Fil96, Ruk93]. Run [Lew01, MWW15]. Russo [LS06, RY23].

S [CR22]. Saddle [Lal91]. Saddle-point [Lal91]. Sahalia [CMP24]. Sahalia-Type [CMP24]. Sakhanenko [FGL20]. Salesman [Gan22]. Sample [BM15, Cap13, CL16, ER01, JWZ21, Jur18, Kös11, Lyt18, MS99, OS15, Owa16, Pri96, Ric17, Sch98, Sön21b, Sta98, Ver12, XZ10, ZWW04, Zit22, Mat89, RR92, Str95]. Sample-Path [Sta98]. Sampled [Ros03]. Sampler [MZ06]. Samples [Bog06, Wan05]. Sampling [Kal99, Kaw13, SS22, HS92a]. Sanders [BH23]. Sandpile [JW14]. Sandpiles [CdGR20, JS21]. Sarmanov [Ver16]. Sathé [BA20]. satisfying [Dob90]. Saul [Sch16a]. Sausage [GP15]. Savage [KT00, Kha09]. Scalar [CdF00, Nas22]. Scale [BHM08, Cri23, DLZ04, EM18, JH03, JH12, KRS10, LV21, Mal19, Nod17, XL23, FV05]. Scale-Distortion [BHM08]. Scale-Homogeneous [EM18]. Scaled [Yas20]. Scaling [BC22, CLdLO19, CdGR20, Dad20, FU19, Kar23, KO19a, KO19b, Sön21b, Uch20]. scanned [HS92a]. Scenery [FSZ21, GP01, Lew92, Lew93]. Schachermayer [Zit02]. Schauder [Küh21]. Scheme [ALT20, BHY19, DP22, Nag15, Ros97, NS23b]. Schemes [NN07, Qin24, ZM23]. Schoenberg [Kal12]. Scholes [Huj06]. Schreier [Sta22]. Schrödinger [DM23, Sch24, Tak18, Wad19, Wad21]. Schwartz [Kun94b]. Score [GHL18]. SDE [AT98, CK00, Luo15]. SDEs [Ama12, BHY19, HKU12, HKHY13, HMY21, MS02, MPOT19, MBP10, Nag15, NN07, Qia14, Son20, Xu07]. Search [AAS18]. Second [Abe21, Ade22, BESF23, BESF24, Lin12, Rai04, Tan20, Vid20, Wat22b, YLY22, Zhu23]. Second-Order [Abe21, Vid20, Wat22b, YLY22]. Seed [dHN22]. Seed-Banks [dHN22]. Seen [Mad17, Mor10, Swa09]. Segment [Uch16]. Segregating [HM18]. Selberg [BA20]. Selected [Yos21]. Selection [BK10]. Selective [Mil00]. Self [Bar00, BK03, BEK10, BGT04, Che99b, CHM13, Dad20, Dav05, FV20, FSZ21, GM98, GZM24, HJN19, HM09, HNS14, IW99, JPW24, JM15, KR03, KZ15, LS06, MSS17, Mal19, Mas06, Mas13, Mas21, Mön22, Pak20a, Par09, PS94, RW05, Spa16, SYG22, Sza12, Tug14, Véc22, Wan11, Wat02, WY10b, Wat16, Wat22b, YSL08, Yu21, YY24, AI96, Fal94, Fil96, GL94, HKW90, HW92, KT94, Lew92, Tei92]. self-affine [GL94]. Self-attracting [SYG22, YSL08]. Self-Decomposability [BK03, MSS17, WY10b]. Self-Decomposable [Pak20a, Wat22b]. self-interaction [AI96]. Self-Intersection [BGT04, IW99, JM15, Yu21, GZM24, HNS14, Sza12, YY24]. self-normalization [HKW90]. Self-Normalized [Dav05, GM98, JPW24, LS06, Mas06, Mas13, RW05, CHM13, FSZ21, PS94, Spa16, Wan11, HW92, Tei92]. self-organizing [Fil96]. Self-Similar [Bar00, KR03, KZ15, Mal19, Mön22, Par09, Wat02, BEK10, Dad20, HJN19, Wat16, Fal94, KT94]. Self-Similarity [HM09, Véc22]. Self-Stabilizing [FV20, Tug14]. Self-Standardized [Mas21]. Selfdecomposability [BJ16]. Selfdecomposable [AMR08, BNSS17, HT16, MS23, Sha00a]. Selfsimilar [MiS99, MS23]. Semi [BGT10, CDM24, Fai98, FKZ23, GSY24, KZ06, LLL15, MiS99, MTU14, Wan17, Fie90, RRMZ94]. Semi-Additive [Fai98]. Semi-circle [BGT10]. Semi-Exponential [CDM24]. Semi-groups [KZ06]. Semi-Lévy [MTU14]. Semi-Markov [GSY24]. Semi-martingale [LLL15]. semi-martingales [Fie90]. Semi-Selfsimilar [MiS99]. Semi-Stable [Wan17, RRMZ94]. Semi-uniform [FKZ23]. Semicircle [BB13, FL13, GNT20, HKW16, KK18]. Semicircular [Sol14, YBH16]. Semiconservative [Abr18, Abr19]. Semicontinuous [OW98]. Semigroup [HM03, IR21, Neu08, Neu09, BM92, Kal92b, Min90, Son95]. Semigroups [AFP23, AM12, Bro00, BR97, BB98, Glo13, Kuz20, LL23, Raj99, Sie06, Tho19, Tho20, Wan14a, Wan19, Bis96, BH88, Bor89, Bud94, BR92, DM88a, HS88, HJ96b, Nej94, Sie94, SZ90]. Semilinear [Ahm15, CR22, DMCN24, RZ10, WY23, WY24]. Semimartingale [GSY24, Kob11, MU15, Son21a, Mao91a, Mao91b]. Semimartingales [Bas09, Li22, Fer89]. seminorms [Li92]. Semipermeable [APP24]. semiregular [Pla92]. semisimple [Dan91]. Semistable [BK03, FC07, KW14, KMX18, KC09, LX20, MS06, MS97b, Meg02, HS88, MS96]. Sense [BLC03, Hei90]. Sensitive [ABK16]. Sensitivity [Liu18]. Separable [CdF00, Sat92, Sie92]. Separating [KM19]. Sequence [AKR99, Arc99, Key05, Key07, MWW15, Per18, Sen13]. Sequences [AM23a, AM23b, AHV21, Aur07, Ber05b, BHW10, Bou22, Bra97, BMP22, Cár17, CK08, DP04, DMV07, DT17, Gir17, GHL18, HKW20, HM03, Kaz21, KS06, Kie98, Kra01, Mer03, MP06, Mor21, O’B96, PV20, Res13, SS98, Wat02, Wat16, Wu04, AY94, Bra96, BS93, DR93, Eis93, HS91, Ho95, Pel96, PV94, TT90, Wei96, Zab95]. Sequential [Büc15]. Series [AKR99, App02, BJP01, Ber05a, BR08b, Bre02, Bre10, GGX24, HHV89, IW24, KLR22, MRU15, Mal08, MS96, Muk21, NS23a, Pin01, Sch09, Spa16, Szu92, Tan07, Yas18, AG93, Pri96, Tal92]. Set [AR06, AO24, BI02, Bal15, BM11, Ber97, DFS99, FF04, HM06, HM09, HP13, Li20, Lop20, MY09, MM22, Saa01, SS06, Uch18, Uch20, Yos21, ZWW04, DM88b, DIM94, IM96, Kal92a, Sla89, Web96]. Set-Indexed [BI02, HM09, Saa01, SS06, Yos21, HM06, MY09, DIM94, IM96, Kal92a]. Set-Valued [MM22]. Sets [AW14, BHM08, BL11, CH99, CAMRR03, DPR00, Hei21, Her13, Hu99, KV18, KL98, Mas06, MM10, Mou07, OW98, PS15, Riv03, SU19, Vir98, Zha23, DM88a, DR93, Fal89, FG92a, FG94, GL94, HS92a, Pfi90, RR94]. Setting [ABO22, KS22, NT14]. Settings [BT06]. Several [KV07]. Seymour [RY23]. Shannon [NP24]. Shape [Ahl15, BDK21, Cer00, RR08]. Shapes [GMW17]. Sharp [AV09, AIS^{+}24, CAFR07, HMS21, Ilt95, Jia19, MPU97, Ose09]. Shearer [Tem14]. Sheet [Dun00, Dun01, FIT14, KL93]. Sheets [BL02, Sön21b, Wan14c]. Shepp [KS96, TZ20]. Shift [LH21, Wat02, Wat16]. Shifted [Don23, Eis93]. Shifts [DLSV05, KT20a]. Shkolnikov [Cla21]. Short [Gra15, Tel06]. Shot [Bre10]. Shrinking [BD09]. Shuffle [Gho20, Pem89]. Shuffles [Lal99]. Shuffling [Dur03, NW19, HJ96a]. sic [MM03]. Side [Ben03]. Sided [Adl02, BH11, Fan18, FGL20, JO21, Kno18, Pis04, Sav10, Von00, Wat22b, Yan09, Adl90, Adl92, Deb10]. Siegmund [DFPS97]. Sierpiński [HJ02, Nod17, Von96]. Sigma [ISY23]. Sigma-Fields [ISY23]. Signals [HTW24]. Signed [Nic09, Sam21, Sch02b, Sch05]. Signs [FV20]. Similar [Bar00, KR03, KZ15, Mal19, Mön22, Par09, Wat02, BEK10, Dad20, Fal94, HJN19, KT94, Wat16]. Similarity [HM09, Véc22]. Simple [Abe21, AM22, BS05, CCET06, CFR05, HS98b, Kub20, Pop21, Ros97, SS09, Hei96, PD91]. Simplex [Ma23]. Simply [Neu08, Neu09, Sha00a, Neu92, Neu95, NS95, Nob91]. Simpson [HN15]. simulated [HS92b]. Simulation [TT18]. SIN [Jaw04]. Sine [BHH10, RYZ02, BY96]. Singular [AT98, AT17, Don23, GT02a, Gri24, HKHY13, HS92b, IR24, Jin17, Lou16, Luo15, LH21, OU21, Pin94, Ric17, TW14, Ye24, Ger92]. Singularities [KS93]. Singularity [AR23, EJU19, Gab11, Lyo00, MR97, MN02, Oka14]. Singularly [Yin01]. SIR [Luc24, Xue18]. Site [BDK21, Cyr21, Swa09]. Sites [ES99, Mar01]. Skellam [KK24]. Skew [BZ14, GM18, OMO22, Pah15, Pro21, VA96]. Skewness [Kov22]. Skip [Fil09a, Jia19]. Skip-Free [Fil09a, Jia19]. Skitovich [FG00, Fel15]. Skorohod [HMY21, JP93, LMS24a, NP18, NP96]. Skorokhod [BS20, NN10, OP09]. SLE [HS21]. sleeping [KK94]. Slepian [BG16]. Slim [Kar23]. Slivnyak [Kal11]. Slivnyak-Type [Kal11]. SLLN [BJLO04, CM00]. Slow [Sin07, Tak01, ZX24]. Slowdown [Far11]. Small [AHK99, AGK^{+}13, BL02, BR08b, Bro03, BC22, Che06, DFMS03, Der03, Dun00, Dun01, DW22, Fai98, Fan17, FGG16, GHLT04, GL06, GL07a, GL07b, GLP03, Jeg11, JT24, KT16, KL93, KLmS95, LS99, Li99, LL04, Liu18, Mar04, MS01b, Mim13, Naz09, Roz19, Sav10, Shi96, Sim04, Sha93, Sto96]. Smallest [Don23, Per18, XZ10]. Smooth [AGK^{+}13, AW14, BN22, HT13, MN00, Oss09b, RR19, HMY21, Kim14]. Smoothing [AM22, CMP15, MP24b, Sch24]. Smoothness [AL07, BR90, KRS10]. Snake [Git03, MM03]. Snakes [JM05]. Sobolev [Ahm17, BT06, GSS21, Kim14, KP24, LS99, NP07]. Sobolev-Type [Ahm17, NP07]. Soft [LV12]. Sojourn [DLM20, NO00, PS24]. Solution [Bal22, GM91b, KV03, LY19, WY23, WY24, Cai90, Ger92, Mao91b, RSS96, KV19b]. Solutions [AM22, BM01, BOLP22, BR21, CL23, CK00, FU19, Fan18, Gau16, GOT20, HA19, HKHY13, HZ19, JK22, KS22, Mao92, MBP10, Moh22, Nak19, OMO22, RZ10, Roz13, TT02, Wad19, Wad21, ZZ12, Kun94a]. Solvable [ALP22]. Solving [LY16, ZM23]. Some [AFP23, Ast23, Aur07, BH24, BY96, BL11, BR97, CD04, Cha95, CM14, CCFR16, DMCN24, DS91, Ejs14, GR97, Glu99, GZ24, HJ96b, HN05, HT13, IO20, IW24, Jia99, Joh89, Jol06, Kal11, Kha09, KO10, KM24, KL02, KZ15, Lac98, Lac06, LTF17, LL03, LM12, Lyo00, MRR04, MS01b, Mat89, Mus06, Nak19, NN07, NO00, Oka21, Pro21, QSS18, Ric17, Ros97, RC88, Sep97, Son21a, Sto96, TT02, Uem04b, Vol10, Wat00, XZ10, XY10, Xu13, Zha01a, Bra88, Bra96, CS95, DM91, Pfi90, Pri96, Sik00, TZ96, Voi90]. Space [Ahm17, BC18, BS20, Bel19, Bha17a, Bha17b, Bur03, CdF00, CD20, Che97b, CM00, Cho17, CY11, CGW20, DG18, DW24, DE18, DW06, EMP17, FM18, FM20, FM23a, FM23b, Gab11, GP15, Gri10, IR08, KW15, KK08, Kim14, KP24, KKH09, Kom08, LL12, Luo98, Luo22, Maj99, MPU97, Mer03, NC21, NPY22, OT17, PP02, RMVP24, Sau16, Sen00, Sta22, TM06, WY10a, YWS22, YC24, CAMBTD96, D’A95, Det89, Fie90, Gol88, GS95, Lal91, Mae93, Pap90, ST15, Sat92]. Space-Time [CY11, NC21, OT17, Lal91]. Space-Valued [Luo22]. Spaces [AL08a, ARV97, AE99, AJL17, AJ22, App00, BZ19, Bar04, Bet23, Com05b, CG21, DPR00, DFMS03, ESC14, FW20, Git03, Gra99, HH10, Hei06, KK17, KU23, KZ08, Las10, LMS24a, Li16, Lia14, Liu12, LZ16, LM20, LMX23, Luc97, Luc10, MM20, Ma24, MS06, MP24a, MM03, Min21, Oka24, OD07, PP18, RS04, RV04, Sam24, Smi15, SD22, Ste16, Tel06, Tyk09, Wit00, XZZ18, Bab94, BGZ93, BR90, Dip91, Dob94, FJ95, GM91a, Gra94, HS93, RS94, RRMZ94, RZ90, Sie92, Sie90, Sla89, Str95, Voi96, WWLH14, Wei96]. Spacing [Kal02]. Spanning [AG18, HL03, RY23, Sol99, Sto91]. Spatial [DLZ04, FRZ18, FV05, LZ24, Mil18, Mao91a, Mao91b, MSSW93]. Spatially [RY24, dHN22]. SPDE [KHMCSS01, RSS96]. SPDEs [BM01, CWM04, Kim08, YZ14]. Special [SV06]. Spectra [BES12, Ora07, Sob10, Tel95]. Spectral [BMT99, BM15, Bas09, BGEM19, BHS11, BMS12, CB04, FS13, GR10, GR17, Gas19, Her20, JQ17, Jur18, KP23, KKM13, MS97b, MQ21, Noi21, Oka21, QX20, TTQ14, ZBH23, ZHBS24, Bra92, LR90, Ros94]. Spectrally [CL18, Chi20, OP09, Pis04, TV16, VAÇA21, Véc23, WZ21b]. Spectrum [Bar00, CO24, FJ17, Gås06, HS98a, Mal19, Naz09, Nem18, Rai07, Tak18, Tri19, Fal94]. Speed [CZ18, Cri23, HSH20, RS18, BB93]. Speedup [Far11]. Sphere [Bri97]. Spheres [Lia99, Ma17]. Spherical [Ala22, BG11, CM22, Hua19]. Spider [CCFR16, CCFR19]. Spiked [Cap13, Noi21]. Spin [MV19b]. Spine [Stu22b]. Spitzer [Spa99, Wit00]. spline [GH95a, GH95b]. Square [WZK23]. Squares [AT17, Shi13]. St. [BGK17, Gut04, GML16, Pap11]. Stability [DM02, DM07, Fan18, FG92b, FKM18, Hei21, JM09, Liu18, LT91, Mao91a, Tom96]. Stabilization [FNN15]. Stabilizing [FV20, Tug14]. Stable [AK16, AB17, BS20, BK03, BJP01, BCG13, Bre02, Bre05, BJS24, CR13, CS20, CNZ22, CNXY24, Chi20, Cor16, DPR00, DH20, DW24, EM18, FMW11, Fra07, FW20, Gue09, He11, HH08, JL24, Jin21, Jun14, JN99, KS08, Koc02, KR03, KM11, LLL15, LL04, LX17, MRU15, MS16, MOP97, Mou07, OU21, OT17, Pac12, PT08, PT04, RS08, Rus07, San14, San16, Sim04, Sin07, Son06, SV06, Szt10, Tak08, Uem04b, WW15, Wan17, Yan18, ZZ99, Ber94, Blo96, Bra93, CNX21, HHV89, KT94, MR96, MT88, MV93, RY91, RS94, RRMZ94, Ros94, ST90, Sch94a, Sie90, Zen95]. Stable-Like [Jin21, Koc02, San14, San16, Uem04b, WW15, Yan18]. Stable-Type [OU21]. Standardized [AL07, Mas21]. Star [HA19, KN21]. Star-Generators [KN21]. Started [CP12]. Starting [LT23]. State [BLP16, CC11, CFZ23, Git03, HLX18, He22, ISY23, KV19a, LM08, LPW22, LL12, Löc20, RS18, RR08, Smi15, XZZ18, FJ95]. State-Dependent [KV19a, Löc20]. statement [Ano88]. States [Che20, FMW11, MM03, WCL24, DF95]. Stationarity [HM09, PF91, DM09]. Stationary [AM23a, AM23b, Arc99, AB17, BM15, Bra01, BK93, CR13, Cop18, CP12, Cyr21, DK18, DH20, Deb10, Fil09a, Fil09b, FL16, FW20, Gir17, KL17, Kon98, KL01, Las10, LY19, MP06, Mön22, NSB02, Owa16, Pel24, Per18, RS08, Sar17, TZ20, Tem23, Xu98, Yam24, ZRP20, AY94, BT17, CL91, CP93, Den89, Die91, Fil92, KLmS95, Pin94, Sha93, Tak19, TT90, Tob90]. Statistic [Che99a, GM07, Sha99, BBG96, Fer94]. Statistical [FF04, JW21, JY02, Sen13, Tri24]. Statistically [Bar00, GL94, Fal94]. Statistics [AV09, AL08a, AM14, Alb00, Ast23, BN22, BU01, BW02, Bos98, BS99, CdHH97, CO20, DK18, DPR11, DS14, Han18, LPP17, Lyt18, OS15, OR16, RU00, Sun97, TZ20, Tei98a, UB98, Zha01a, Arc94, BGZ93, EL95, GZ96, LT91, Mik94, RS95, Su96, Tei92, TZ96, Tom96, Zen95]. Steep [Che21]. Stefan [KZS12]. Stein [CNX21, CNZ22, CNXY24, ES22, Gau16, Gau20, GS24, GR05, JSS23, LMX22, Pri19]. Step [ARS21, CJSS24, Neu92, Neu95]. Steps [DP17]. Sticky [Fus10, GV18]. Stieltjes [Ber05b, BH11, Has12, Kle14]. Stirling [Ade22]. Stirring [BM14, Kon95]. Stochastic [Ahm15, Ahm17, App01, ALT20, ABO22, AO24, Bai22, BT10, Bal07b, BM01, BOLP22, BY16, BH22, BNMiS06, BR21, BS03, BL19b, BLV24, BSW10, BZ14, Bre02, Bre05, Bro16, CWM04, CP21, CP18, CR01, CLY13, CF21, CW23, CL23, CE24, CM14, Chi22, CO23, CR22, DFPR15, DH21, DM88b, DM16a, DQ21, DMCN24, DFS99, DP22, Eng20, FU19, Fan18, Fan19, FKZ23, FFS97, Fil09b, FL14, FIT14, FM23a, FM23b, FH15, GIST20, Gar08, HA19, HS98a, HL22, HJN19, HK19, Hsu97, HMW24, HZ19, Ina10, IR24, Jeg11, Kal20, KW15, Kar11, Kar19, KO19a, KO19b, KL14, Kim09, KZS12, Kim14, Kim17, KP24, Kob11, KHN21, KM11, KR21, KS22, KL05, Kul23, Kun94b, Kur07, LR12, LL89, Lee18, Lei91, Les10, LY19, LS21b, LMS24b, LWZ24]. Stochastic [LL24a, Lia99, LW16, LG19, LY16, LTF17, Liu18, LM22, Luo15, LW19, LH21, MRU15, MRR04, MCR08, Mar04, MP24b, MN15, MS02, MZ24, MV00, MSY16, Moh22, MS05, MR97, MNR99a, MNR99b, MN02, Nak19, Nas22, NO03, OU14, OV14, OBO23, OA23, Pah15, PZ22, Pri15, Pri19, Raj15, RXZ10, RWZ15, RZ19, Roz13, RMVP24, iS13, Sau16, Sch24, Sla22, Sta07, Sun22, SS09, Tak10, Tem14, TT18, TT02, VR22, Von00, Wan13, WZ21a, WZZ22, WGYX24, WY23, WY24, Xu18, XL23, XSR23, YLY22, Ye24, Yi23, YWS22, Zha11, ZH12, ZM23, dCdCV23, dSEE18, AD88, App93, Det89, Din88, Dip91, FG92b, Fis90, Gan91, HR18, HO96, Kou96, Kun94a, Mao91a, Mao91b, Mao92, NP96, NS23b, Øks90, Pec93, RC88, SG90, Sen93, Ubø95, Vol92]. Stochastically [AHK99, Yok21, Blo96]. stockbroker [KK94]. Stokes [Sau16, WZZ22, YWS22]. Stopped [LVZ21, Tei95]. Stopping [AK92, AO24, Bel19, GP98, SS06]. strassen [DR93, Ber97, CHM13, Ein07, GK91, Gri92a, KS16b, LW97]. strassen-type [DR93, CHM13]. Strategic [GR98]. Strategies [AO24]. stratified [Sch94a]. Stratonovich [Bal22, Jol06]. Stratonovich-type [Jol06]. Strictly [Chi20, EM18, MRU15]. Strings [FNN15]. Strong [AHV21, AK20, BOLP22, Bos98, BS99, BJ16, Büc15, CGHM06, CdF00, Chi18, CLM04, Cop18, CCFR92, CF22b, CM15, CDM24, Cuz01, DFPR15, EM05, Fil92, Fil09a, Fil09b, FL16, GV18, GP01, HJKPS07, JL08, JK22, KS19, KM17, KT22, Kie98, Kin00, LQR11, LMX23, MBP10, MST08, Ose09, PP02, RR94, RYZ02, Slo01, SGQ22, Sun14, Sza12, Tei98a, Tei98b, Ter16, WLG03, Wit00, XY10, XL23, CG96, Cuz95, Dob94, Ger88, Mat88, CF23]. Strongly [Bra97, BT17, CWM04, KM08, Mer03, Ho95, MR92]. Structural [Bet23]. Structure [Fal02, Meg02, NSB02, RW02, AGK92, MV93, Nej94]. Student [BBG96, Sha99]. Studentized [JWZ15]. Study [EOOM17, SW09, Zha01b, AHP91, Sto91]. Sturm [GH95a]. Sub [Tel06, Tri24, YWC18, Str95]. Sub-critical [YWC18]. Sub-exponentiality [Tri24]. Sub-Gaussian [Tel06]. sub-martingale [Str95]. Subadditive [Kaz21]. Subclass [AMR08]. Subcollection [JT24]. Subcritical [ABKV12, KK18, LVZ21]. Subdiffusive [Hu17]. Subexponential [AFK03, GT09, Lin12, WY17, Wat21]. Subexponentiality [WY10b]. Subgaussian [Men16]. Subgeometric [Den20]. Subgraphs [Ahl15]. Subgroup [Als02, HS88]. Subject [Yan13]. Sublinear [Gra12, JL24, LL17]. Submanifolds [Oka24]. Submartingale [Eng20]. Submartingales [Bog20, MP14, Nik06, Ose11, DIM94, HS92a]. Subordinate [MS19, SV06]. Subordinated [EM05, LW23, Lin12]. Subordinates [Ose11, Ose14]. Subordination [Cap13, Den20, Ose09]. Subordinator [MV19a]. Subordinators [AL07, Bar19, BJS24, SU19, SV06]. subpramarts [Xue91]. Subquotients [Pan19]. Subsequence [MS13]. Subsequences [KM97, LMR02, Sla89]. Subsets [CFR05, CF20]. subspaces [D’A95]. Subtle [UA19]. subtree [Sto91]. subtrees [PD91]. Sufficient [BR99, Chi22, RS04, Wu04, Bud94]. Sum [BT24, Dan14, JWZ15, Raf13, Ver16, ZWW04, BG95, Beh89, Lac89, Zab93]. Sum-and-Max [BT24]. Summability [Kie98]. Summable [CH05, Nea92]. Summation [EL19, GS10]. Summing [KR21]. Sums [Ahm05, AFK03, AdC24, BP10, BP13, BL22, BGK17, BK14, CV15, CV21, CST23, CHM13, Cuz01, Den04, DM13, GT09, GM98, Gri97, JL08, Jir17, JMPP04, Kab12, KS19, KW93, KN10, KN16, KLR22, Kov02, LT03, Mas06, Rio09, RV04, SD22, Spa16, VC18, Wan17, Yas20, Adl92, Bra94, Car88, Cuz95, FKT94, HKW90, HW92, KM92, KS96, Kou96, LRJW95, Mae93, PS94, Sch94b, Tei95, Wha92, Yu90]. sums-revisited [Tei95]. Sup [BL22, Li99, Shi96]. sup-norms [Shi96]. Sup-Sums [BL22]. Super [AI01, ADHP97, DF95, DF97, FKX06, FMW11, He11, HLZ13, Hon03, HZ07, RY24, RSZ24, Wan10, Zha05, Zha11, Ser96]. Super- [FMW11]. Super-Brownian [DF95, RY24, RSZ24, Ser96]. Super-coalescing [HLZ13]. Super-Diffusion [Zha11]. Super-Stable [He11]. Super-Tree [ADHP97]. Supercritical [AR04, Kub20, LL21, LRS11, Mil18, Wan18b, Wat02]. Superlinear [CW23]. Supermartingales [Cas08]. superprocess [AI96]. Superprocesses [DLZ04, FPP21, Mil18, Myt98, MX04, Sch99, Wan18b]. supOU [GLT20]. Support [BL09c, Kul23, RSZ24, XSR23]. Supports [RRMZ94]. Suprema [DM16b, TV16]. Supremum [Bar04, DH20, RSZ24, WW09, Shi96]. Sure [AKR99, DM07, FC07, FM23a, FM23b, GW02, Hör07, Hu15, KKH09, Lou16, Meg02, PG99, PT17, Ued19, Véc23, Wan10, XZ10, BD94, Kou96, Lac91, LRJW95, Mar88, PF91]. Surely [RV04]. Surface [HSW18]. Survival [CY11, GMPV10, Han97]. Switching [CJSS24, JSS23, QK23, Yin01]. Sylvester [BF24]. Symbol [RS24]. Symmetric [AR06, App00, Beh89, BR08a, Bob05, Bry03, BB98, CC07, Cas97, CR13, CRY17, CKRMT16, ESC14, GNU17, Gra99, GLM05, HM05, HJN19, HL01, HL04, Hei21, Kol16, Kol17, KKM13, LXY19, MS12, MN97, MMS07, OU21, Özd19, Pac16, PXY20, RRM07, RS08, SU14, Son06, SS09, Tak08, Tak11, Tak19, Uem04b, Yas20, Zei13, Bab94, Ber95, DS90, DS88, Fit89, Gra94, Juh90, KT94, Kol89, MR92, Mór95, Pal92, Vol92]. Symmetries [BD16, Kal92a, MM07, Nas22, MV93]. Symmetrised [Tra08]. Symmetrization [DLL15]. Symmetrizations [GR88b]. Symmetrized [Hil22, RRM07]. Symmetry [DP12, Glo91, Luc97, UA19]. System [AM14, BDK21, Hu20, JK22, Pah15, Puh99, Wer05, Fis90, Rév93]. Systems [BDS23, BDS24, BFW18, Car02, CW23, DGW01, DL00, DPR11, DS98, IM24, KS98, Lal99, LX12, Löc19, NSB02, PGS16, SS98, Sud00, VR22, Cha90, Dur92, FS92, MSSW93, Neu92, ST15]. Szász [ANS23].

Tables [Sim24, Wu23]. Tail [Aur07, AW14, Ben03, BGK17, BR92, Den04, Gei01, GT09, IW24, LP03, LS05, LMR02, MŚ21, Sar17, YLY22, AG93, DM91, Die91, Li92]. Tailed [BJL16, EJU19, FKZ07, HN04, Lou12, MPW18, Nyr09, Owa16, dBM00, LBBB11]. Tails [BT24, Bog06, Bul21, CDM24, HK98, Has05, KK17, MS99, RRM07, dlP93]. Takagi [HSZ22]. taking [Fie90]. Talagrand [Tan20]. Tanaka [MX04, SV16, Uem04a]. Tangent [Fal02, Lac05]. Taylor [Ina10]. Taylor-Like [Ina10]. Technique [DFS99, Sob10]. Techniques [Sar19]. telegraph [Jan90]. Temperature [KK18, Pak20b]. Tempered [Bha17a, Bha17b, MS16, Szt10, Fie90]. Tempering [AMS22]. Temporally [MP24a]. ten [Tal96a]. Tensor [Din17, JX20, Lyt18]. Tensors [DT24]. Term [Abe21, CE24, DFPR15, KT16, LMS24b, FUW20]. Terms [AV09, Aid24, Fen96, KW93]. Tessellations [JY02]. Test [HLZ13, Ran06]. testing [AG91]. Tests [BHW10]. th [RR92]. Their [Ast23, BW19, BGK17, Bre02, CL99, CD20, Che20, Cri23, Gas19, HKU12, KO10, Lal99, LLL15, Lin12, Ose11, Ose14, RRM07, iSU13, Sch98, SS15, Slo01, Ahm05, BLBE^{+}94, GR07, KZ06, KW93, MS96, ST15]. Theor [Cuz01, Dun01]. Theorem [AZ01, Aus11, BA20, BK15, BDK21, BM99, BG16, BCK15, Bob18, Bos98, Bou05, BT17, CV15, CD04, Chi18, CH99, CR22, CAFR07, CP12, DE18, DLSV05, ESC14, FJ12, FC07, FG00, Fel04, Fel15, FKT22, Fus10, HK16, HSZ22, HPZ21, HJST97, Hsu97, JL24, Kal12, Kaz21, KS06, KT20a, KT22, Kha12, Kol16, Kor08, KN21, KU23, KV18, Kul23, LW04, Liu21, LT23, Lyt18, Mil18, Mor21, NP24, OR16, OV14, PP24, PS13, PD23, Raj99, Sha00b, SWF22, Sla22, SW20, Wad19, Wan10, WY16, Wil03, Xu18, XSR23, YY15, YY24, Zei13, ZRP20, Zit02, AD88, BD93, BHL93, BB93, BD94, Blo96, Bra96, Chu93, DM95, DF88, Dob90, Fer89, Gra94, Has88, HS91, Ish92, MR96, MT88, McG91, Nej94, Neu95, NP91]. theorem [PS94, RX93, Sif95, TT90, Zab93, Zeu89, DS14, DT14, Fra07, GW02, Gru14, LBBB11, Mal99, MS97a, Mer03, Myr24, PR23a, PR23b, Per97, RS04, Saw01, Voi21, Vol10, YSH03, Zha05]. Theorems [ABKV12, Ahm05, ANS23, AVW24, Bai22, Ban19, BP23, BM18, BH99, Bjö10, BL09c, Buc21, CZ22, CK22, Che20, COM05a, Cri20b, CCFR16, CCFR19, Del15, FY08, GT02a, GL22, GT05, Gra99, IM24, JH03, JL08, Kal11, Kol98, KU23, KL01, KZ15, LM08, LRR03, Mal06, Mas21, Meg02, Men16, Mil00, NNR16, NN10, NT14, NPPS16, Par12, Raf13, RW02, Rig20, SS15, SS98, SD22, Sze09, Tem23, Ued22, Voi17b, Wan05, Wan18b, Zha11, Zwe07, Adl92, AY94, BGZ93, Bin88, DF88, EM89, Fer94, KM92, Krö90, Lac91, Mil94, Nob91, Res94, Voi96, BD06, GT01, JH12, NP07]. Theoretic [CR13, SVT12, Sta07]. Theory [AB18, BBH16, BT12, Bro04, CS20, Fil92, Gri24, Hör07, Ina10, Kim08, Kim14, KP24, Owa16, Smi15, SD22, SV06, Yas18, GH95a, GH95b, HJ96b, Kim09, Kim17, O’B96, SG90]. Theta [Dun21]. Thin [BF13]. Third [Ejs14, Pri19]. Three [CC11, Mil94]. Three-State [CC11]. Threshold [Ald91, BU01, DJ18, Han97, Tyk09, Xue15, Dur92]. Thresholds [NPPS16]. Tight [Bro03]. Tightness [CW10, Fan17, Tak19, DM88a]. Time [AV09, Abe21, AAO17, AHK99, AP10, AP97, AD11, AK16, BW19, BC18, BLP16, BNMiS06, BGT04, Can19, CD20, CC11, CH20, Cho17, CY11, Cop18, Cor16, CFR05, CF22a, CF24, DGW01, Den20, DGR21, Fil09b, Fit89, FY08, FKZ07, FNN15, GMZ12, GS95, Gra15, GP11, GPT15, GHX19, GZM24, HKU12, HLZ13, Her20, HNS14, Hu20, Iso09, Jeg11, Jin17, JT24, JV07, Jun14, JM15, Kar19, KO99, KM17, KP24, Kli13, Kob11, KP23, KHN21, Kur07, Lac00, Lac06, Law08, LX12, LW04, LZ16, Löc20, LT23, LS11, Luo15, Ma17, MM20, MZ06, MV19a, MS22, MP14, MU15, Myt98, NC21, NYY22, NOL07, OT17, Pes01, Pes05, RS24, Ros03, RR00, Sam21, Sau16, Sav10, Sen00, Shi97, Son21a, Sza12, Tak18, Tan07, Tel06, Tso93b, Uch12a, Uch12b, Véc23]. Time [Vil20, Von00, WH17, WZZ22, XL23, YLY22, Yas18, Yu21, YY24, YC24, Zei18, BGT15, BK89, CS95, CCFR96, Fil92, GMS22, GS96, HS98b, KLNS89, Lal91, Mat88, Øks90, PF91, Pem89, Pem90, Von93, Xue91]. Time-Change [DGR21]. Time-Changed [HKU12, Kob11, KP23, MV19a, Sam21, Tak18, Von93]. Time-Delayed [Kar19]. Time-Dependent [HLZ13, Jin17, Kli13, Kur07, Luo15, WZZ22, Pem90]. Time-Fractional [HKU12, KP24]. Time-Inhomogeneous [DGR21, Löc20]. Time-Inversion [AP10, Law08]. Time-Sampled [Ros03]. Time-Space [Sen00, GS95]. Time-Varying [MM20]. Times [AK20, BK09, Bla00, BK14, CL18, Cop18, CCFR19, DLM20, DM09, EF01, Fan17, Fil09a, FL14, FK14, Glu99, Gue09, HL22, Hu17, IW99, Kim22, Kon98, KPP14, Lac98, LS15a, LM16a, Li22, MMW13, MW22, Mön22, PRP24, PS15, PS24, Pro21, Rah22, SV05, SS21, Son21a, WX10, WZY17, Xia98, Yan17, vD17, AK92, Ald89b, Ald91, AHP91, GS94, MR92, Pal92, Shi96, Shi93b, Zuc89]. Tip [Mad17]. Toeplitz [Ada10, HM05, JMP12, LLW11, LW11, MMS07]. Top [Gho20]. Top-2 [Gho20]. Topics [Yos21]. Topological [But22, BM92, Min90]. Topology [BK15, Eic97, JP93]. Tortrat [DR98, Eis93]. Torus [Shi13]. Tossing [EV18]. Total [Gho20, KLL04, KLLW05, Qin24, Ste20, YWC18]. Totally [AE99, Eva89]. Tour [Git03, MM03]. Tour-Convergence [Git03, MM03]. Trace [Bet95]. Traces [Kol89, Nag21, Nag22]. trajectories [DL94, Sch94b]. Trajectory [ZZ19]. Transcritical [BCOV23]. Transfer [BLV24, LR88]. Transform [AM22, BG11, BL09c, FHL22, Maj99, Pin11]. Transformation [BZ06, Cha00, MS12]. Transformations [Döb17, FP10, GR05, MN97, Oka14, Oss09a, Zwe07, FG92b]. Transforms [AJ22, Has12, Son06]. Transience [Abr18, AIV24, CLdLO18, DG18, KO10, Lac98, MTU14, MS22, San14, Abr19]. Transient [AP97, AK20, BK09, CFR09b, CFR09a, CFR10, Kor08, Sin07, YY15].

Transition [AHK99, BP21, EMP17, Har01, JS99, KKR04, KZ06, KS12, Li18, RR19, Son06, Uch18, Vol06, YGHL08, Sal92]. Translated [BL06, BL09b, RZ90]. Transportation [LMX22, Rig20, DY95]. Transpose [Gho20]. Transposed [BN13]. Transpositions [Özd19, Sim24, HJ96a, Mat88]. Trap [BSF15]. Traps [PV98]. Travelling [Gan22, Kno18]. Tree [ADHP97, Att21, Bro04, BK12, Cla21, DMS09, MRS07, Puh99, SWF22, CS91, Lal91]. Trees [ADG18, AAS18, BK00, BES12, Bro16, CZ18, GMW17, GNS24, He17, He22, Hir03, Hu17, HSH20, Kar23, Kor16, Lin18, NW02, RY23, Ste18, Stu22b, TT98, Xue15, DKW91, KM91, PD91, Zuc89]. Trend [BL19a, DLM20]. Triangle [Mil00]. Triangular [Adl02, D’A00, Dav05, HP97, Mas13, MŚ21, Pak00, Neu92]. Tridiagonal [Zha17]. trigonometric [Far94]. Trimèche [Zeu89]. Trimmed [BGK17, Fan17, KS19, Mae93, Mas21, HW92, KM92, Neu95, Wha92]. Triple [Chi20]. Trispectrum [CM22]. Truncated [FK18, KS08]. Truncation [MQ21]. Tube [BT18]. Turbulence [LR12]. Turning [EV18]. Tusnady [Kol94b]. Two [AG18, CH21, Cer00, CFZ23, CK99, Cov11, Deb10, Gab11, GHX19, Häg13, Hue00, JO21, Jia99, Jin19, KV07, KT20b, Kua18, LM16b, LLL15, LT98, LM20, LMX23, MM20, MSS17, NOL07, Oka21, Pak00, Pop21, San14, Sau16, Shi93a, Sta98, Sta11, Uch16, WZZ22, WY17, Wat21, Wat22b, Xu07, XL23, YZ14, Beh89, BX16, NP96, Voi96]. Two-Boundary [KV07]. Two-Dimensional [Cer00, Hue00, LT98, Pop21, Sau16, Uch16, WZZ22, Sta11]. Two-Feature [LM16b]. Two-Parameter [CK99, Cov11]. Two-Point [LM20, LMX23, MM20, Voi96]. Two-Sided [JO21, Wat22b, Deb10]. Two-Time-Scale [XL23]. Two-Type [CFZ23]. Type [ARV97, Ahm17, AHV21, AMR08, AVW24, Bal15, BL19b, BHS11, CP21, Che99a, CFZ23, CMP24, CH99, CF20, CHM13, DK18, Def13, DMCN24, EM00, Fan18, FNN15, GSS21, Gra99, HJ02, HJMLX16, JWZ15, Kal11, KM24, KU23, LS99, LS06, Luo22, MR02, Men16, MSM07, NP07, NO00, OU21, Qin24, Sha99, SU14, SW21, SGQ22, SW20, Sun14, Tak10, TZ20, Tan07, Voi17b, VA01, Wad19, Wan14a, ZBH23, ZHBS24, ADG18, Cha90, Dan91, DR93, DS88, Fer94, Fer89, Gra94, GH95a, GH95b, HS91, Jol06, KT96, KM91, ML17, Res94, Ste18, Stu22b, TZ96, VA96, AT17, JPW24, KS12, Wan13, RW05]. Types [NW02, RR97]. Typical [KS16a, Lin18, Swa09].

Uhlenbeck [AM14, BESF23, BESF24, Cha90, GL22, Hec99, JY22, JZ23, Kaw13, Liu12, LZ16, PS05, ZZ19]. Ultrametric [DG18, DMS09]. Unbounded [CH99, DFPR15, DFS99, EMP17, MSY16, Wan18a, Xu13]. Uncertainty [Che13]. Unconditional [BK19]. Uniform [ALL^{+}07, AHK99, Asc01, Asc09, BL95a, BD09, Ber97, CdF00, Chi22, CCF^{+}02, CAMRR03, DO13, Dem90, DL94, DGZ91, EM00, FM23a, FM23b, HR00, KM16, KN16, Lop20, LL03, MW22, Par12, PXY20, Pel01, Pes95, QSS18, RRM07, RY23, RMVP24, SK07, SK20, WX22, Wu23, YC24, FKZ23, KL96, Lev89, Mao91a, Mat88, ST15]. Uniform-in-Bandwidth [DO13]. Uniformities [RRS92]. Uniformly [KN10, Mil09, Sun22, WZ21a, ZL18, BR90, Lev89, MPOT19, RZ90]. Unimodality [BJS24, HT16]. Unimodular [Mus06]. Uniquely [KM16]. Uniqueness [AT98, BS03, CL23, DFPR15, Don08, DGR21, EMP17, Fan18, Her13, HMY21, JGM11, Jin21, Kuw07, MPOT19, MSY16, Neu08, RZ10, Sau16, Tak19, Tyk09, WCL24, YZ14, dSEE18, Cza24, Las96, Ros94, Neu09]. Unit [Ban16]. Unitarily [Kal12]. Unitary [App01, AD10, KO19a, KO19b, MQ21, Zha19]. Unitriangular [Mor21]. Univariate [BGCD10]. Universal [JT24, Jur15, DGZ91]. Universality [BT16, LMS12, Mön22, NPPS16]. Unknown [AMS12]. Until [CL18]. Upper [AB17, BHW10, Chi03, HK98, KN16, OW98, Pak00, Par09, Ued19, dAC98]. Upper-Lower [BHW10]. Upward [CD20, Jia19]. Urn [BD06, CZ22, CK22, KV03, MJM14, NW19, Pem90, KV19b]. Urns [CMP15]. Using [CCET06, CM14, FL14, JSS23, Leh15, Mal19, Gri92b].

Vacant [Li20]. Value [BM14, CLY13, CS20, Don23, Jun05, Naz09, Sol03, TW14, Wei14, Xue18, YLY22]. Valued [AMS12, App01, Bur03, Che97b, CRY17, Den04, Gei01, HJST97, JWZ21, KK08, KU23, LT98, Luo22, MM22, PP02, Sol14, Yas20, Arc94, Det89, Din88, FG92b, KK94, Mae93, PF91, ST90, Su96, Tal92, Tso93a, Ubø95]. Values [AT17, App02, BS20, BP21, CdF00, CH99, Fai98, Gri24, Lou16, Mer03, Sep01, Bra91, Fie90, Juh90, Pfi90]. Vandermonde [TW14]. Vanishing [Nik06]. Variable [CJSS24, Deb10, Jin21, LM16b, Pes05, Ryv15, GM91b]. Variable-Length [Deb10]. Variable-Step [CJSS24]. Variables [AFK03, AdC24, BP10, BP13, BP23, BJLO04, BU01, Ber05a, BC15, Bon15, BR08b, CV15, CST23, Che97b, CM00, COM05a, CDM24, Cuz01, Dan14, Den04, Ejs14, EJU19, Gei01, GT09, GM98, HJKPS07, Hue00, JM09, JH03, JH12, JL08, JMPP04, KM19, Kar07, KS19, Kie98, KN10, KN16, KLR22, KKH09, Les10, LT98, LT03, LS15b, MN97, MP10, MM05, Muk21, PG99, PP02, Rio09, Sha00b, Sol14, Sun14, Tei98b, UB98, UP03, VC18, Ver16, Wan11, XY10, Yas20, ZWW04, Adl90, Adl92, Beh89, BD94, Bra94, Cuz95, Den89, HP94, Hei96, HK95, KS96, Kou96, LRJW95, Mae93, PS94, Pel96, Sch94b]. Variance [BH02, CMP00, CHNR22, Gau20, Gra12, GLT20, HK95, Hu97, LS15a, Lin18, PS13, BD94]. Variance-Gamma [Gau20]. Variant [YY15]. variate [EL95]. Variation [Aid24, BLS12, BJP01, Gho20, HSZ22, HNS14, KT20a, KT22, Kul11, MM22, MT08, Ney13, Qin24, Sim04, Slo01, SD22, Ste20, Xia97, Din88]. Variational [MN15]. Variations [MN20, Zei18, Joh89]. Variety [GMW17]. Various [Eng20]. Varying [Ban19, Bou22, Bul21, FG14, FS13, HJ02, Ma17, MM20, Wat22b, BPP96]. Vasicek [BH23, JPW24]. Vector [DM01, DV11, Din88, DJ18, Gei01, HJST97, Luc97, Luc10, Ma17, MM20, Ma23, MS06, D’A95, HS93, Tal92, Din88]. Vector-valued [Din88, Tal92].

Vectors [Arc99, Asc01, Asc09, BJP01, CV21, DM13, Fan16, GS24, HM08, KL24, Kov02, KZ20, MM07, Mec12, MS99, Rai04, Sep97, Sep99, SD22, Sta11, SC21, Wan17, Wu09]. Version [Bou05, Hil22, Kha09, Lyt18, VZ23, Voi21, Zit02, Pem90]. Versions [BHH10, FM18, DF88]. Versus [BK19, CLdLO18]. Vertex [CZ18, Kho22, Vol06]. Vertex-Reinforced [CZ18, Vol06]. Vervaat [Cha00, CCF^{+}02, CK08]. Via [ES22, Gau16, Pel01, Pin01, AL07, BL02, Ber93, BT12, DE13, Gau20, GSS21, HKW90, HM08, IR21, JK21, JM09, KV03, KV19b, LL04, Lia09, Mao92, MR92, Mer03, Pin11, Pin18, SD22, Zen95]. Viability [LW19]. View [ANS23, HN04, IR24, KZ08]. Viot [LZ24]. Viscous [Roz13]. Visited [AD14, ES99, Föl03, Mar01]. Visits [AIM15]. Viterbi [LS21a]. Vlasov [BH22, DQ21, LWW24, Qin24, Tug18]. Volatility [BESF23, BESF24, CP21, CMP24, DJ18, FHL22]. Volterra [CP21, Cho17, HL22, HTW24, He11, NS23b]. Volume [Nam22, Ben91]. Voronoi [JY02]. Voter [Han03, ZX24].

Waiting [JT24, Kon98, Shi93b]. Wald [dlP93]. Walk [Abe21, Als02, AK20, Att14, Att21, Ban16, BJZ19, BZ19, BK09, BK00, Bul14, Bul21, CS14, CRS01, CFR05, CFR07, CFR09b, CFR09a, CFR10, CF22a, CF22b, CF23, CF24, DW24, EV19, EP15a, EP15b, FSZ21, FK14, Ful08, Fus10, GK03, HS98b, Hu15, IR08, KM97, Kor16, Kub20, KM08, LV21, LT23, Mad17, Oka21, Özd19, Pop21, RY23, Rev01, Sch08, Smi15, Tak01, TT18, Uch20, Vo23, WW09, WH17, Wan18a, Yan13, Zha23, D’A95, Din95, HM95, Kes95, Lal91, Lal95, Let93, Lew92, Lew93, Mat89, Sco90, Zho93]. Walks [Abr18, Abr19, AIM15, AB18, AIV24, AD14, AD10, BR23, BW19, BC99, BS05, Bar20, Buc21, BK12, CLdLO18, CLdLO19, Che06, CC17, CY11, CS00, CCFR16, CCFR19, CF20, Cun03, DP17, DG18, EM05, FKZ07, FH15, FP10, GR97, GR98, GMPV10, GM11, Glu99, God04, GP01, GPS06, GP11, GKS15, Ham01, HP13, Hil01, HN04, Hu17, HSH20, IR21, KKR04, KM98, KN18, Kum11, LP18, MMW13, McC11, MPW18, Mil00, MS19, Mon19, Mur19, Mus06, NW02, Nak13, Nam22, Nyr09, Pak00, Sch02a, Sob10, Spa10, Sta22, SS09, Sza12, TT98, TW24, Uch16, Vir98, Voi09, Voi17b, Vol06, Wu99, Ald89a, Ald89b, BPP96, BS88, Car88, CS91, DS90, Ger88, KLNS89, KM91, Kuh91, LP93, Pal92, RX93, Tet91, Zeu94, Zho95, Zuc89]. Wall [DFPS97]. Walls [YZ14]. Walsh [VA01]. Wasserstein [CFZ23, CAMBTD96, Gau20, LL17, LW23, LMX22, XZ09]. Wasserstein-Type [CFZ23]. Watson [ADG18, AR04, Att21, BK00, Bro16, CZ18, Har22, He17, IM24, Lin18, Ste18, YWC18]. Wave [Bal22, BSW10, Mar04, MP24b]. Wavelet [AL08b, Sch09]. Waves [Kno18]. Way [MLC22]. Weak [ARV97, BM01, BK15, Bou22, BK14, Büc15, BK19, But22, CW10, CNP93, CK08, Cun03, CM05, DMV07, Di 22, FRZ18, Fus10, Gir17, Gri92b, Gut04, JM09, Jas10, Jir17, Jur18, Kol01, KM11, KV11, LL17, Mer03, MP06, Mik94, MZ24, MR97, OV14, QK23, RZ10, RV04, RZ97, Roz13, San07, iSU13, Xia97, ZZ12, JP93, Pel96, Swe89]. Weakly [ABKV12, BL09a, CV21, CGW20, Dob94, LX12, Per97, Rus07, UP03, Bra92, PS94]. Web [VZ23]. Wedges [MMW13]. Weibull [Bog06]. weight [Kal92b]. Weighted [AAS18, AdC24, BP10, BP13, BJLO04, BHW10, BR08b, CW17, Cuz01, Gan22, JL08, Kim14, KL02, LT03, MN20, RU00, SD22, Wan19, WZK23, Zei18, dBM00, CG96, Cuz95, DKW91, KS96, LRJW95, Mik94, PV94, Shi96, Yu90]. Weights [CH05, Tem23]. Weingarten [Mat12]. Weiss [FKT22, KK18, KT20b]. Welding [LZ19]. Well [Bon15, Sch24, Tug18, YWS22]. Well-Posedness [Sch24, YWS22]. wells [Tug14]. Welsh [RY23]. Which [Bra01, KW93, KN21, NSB02, BL97, MT08, OV08, Xue91]. White [FU19, Sau16, Bet95]. Whose [AMS12, RS91]. Wick [MR99]. Widder [ESC14]. width [Yu89]. Wielandt [ABK16]. Wiener [Pri15, CH21, CD20, DR93, DT24, FIT14, GP15, HK19, HPLVS95, Hu97, Jan90, KP14, Kol01, Luo98, NPY22, PT08, Rév93, Rév98, Shi96]. Wigner [AVW24, BMT99, Cap13, CGW20, Mal19, ORS13, OS15, PAS12]. Wigner- [AVW24]. Williams [McG91]. Winding [BPY03, BP14]. Windings [BDW21, CY22]. Winsorized [HKW90]. Wise [HL22, MRS07]. Wishart [AW04, BN13, Bru91, CDM05, DT24, GLM05, JK22, Let89, MN97, Mat12, NZ22, RR19, Red11]. Without [BS03, BEG18, Döb17, Kol16, SS22, Adl90, JL24, Mar88, PRWW19]. Wittmann [Den04]. Wold [BL09c, CAFR07]. Wolfowitz [Gol88]. Wong [BL19b, Fer89, PZ22]. Words [BB13]. Wreath [Sch02a, Wie03]. Wyner [Shi93b].

years [Tal96a]. Yor [Kol17]. Young [MM22].

Zakai [BL19b, Fer89, PZ22]. Zakai-Type [BL19b, Fer89]. Zero [BM11, BB98, DM02, FG99, Hei96, LVZ21, MMW13, Sav10, SVT12, Sik00, Xu98, MSSW93, RS91]. zero-one [MSSW93]. Zeroes [FK20]. Zeros [BRZ96, FG99, Muk19, NS23a, LR88]. Ziv [Shi93b]. Zygmund [AHV21, HH10, LQR11, Sun14, Tei98a]. Zygmund-Type [AHV21].

References

Akahori:2017:DTC

[AAO17] Jirô Akahori, Takafumi Amaba, and Kaori Okuma. A discrete– time Clark–Ocone formula and its application to an error analysis. Journal of Theoretical Probability, 30(3):932–960, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Aguech:2018:WDR

[AAS18] Rafik Aguech, Anis Amri, and Henning Sulzbach. On weighted depths in random binary search trees. Journal of Theoretical Probability, 31(4):1929–1951, December 2018. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-017-0773-1.pdf.

Ayache:2017:SIH

[AB17] Antoine Ayache and Geoffrey Boutard. Stationary increments harmonizable stable fields: Upper estimates on path behaviour. Journal of Theoretical Probability, 30(4):1369–1423, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Alsmeyer:2018:FTM

[AB18] Gerold Alsmeyer and Fabian Buckmann. Fluctuation theory for Markov random walks. Journal of Theoretical Probability, 31(4): 2266–2342, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Applebaum:2021:PLG

[AB21] David Applebaum and Rosemary Shewell Brockway. L^{2} properties of Lévy generators on compact Riemannian manifolds. Journal of Theoretical Probability, 34(2):1029–1042, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-019-00980-3.

Abbes:2017:CIL

[Abb17] Samy Abbes. A cut-invariant law of large numbers for random heaps. Journal of Theoretical Probability, 30(4):1692–1725, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Abe:2021:SOT

[Abe21] Yoshihiro Abe. Second-order term of cover time for planar simple random walk. Journal of Theoretical Probability, 34(3):1689– 1747, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01011-2.

Arapostathis:2016:RSC

[ABK16] Ari Arapostathis, Vivek S. Borkar, and K. Suresh Kumar. Risksensitive control and an abstract Collatz–Wielandt formula. Journal of Theoretical Probability, 29(4):1458–1484, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-015-0616-x.

Afanasyev:2012:LTW

[ABKV12] V. I. Afanasyev, C. Böinghoff, G. Kersting, and V. A. Vatutin. Limit theorems for weakly subcritical branching processes in random environment. Journal of Theoretical Probability, 25(3): 703–732, September 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0331-6.

Arharas:2022:DRB

[ABO22] Ihsan Arharas, Siham Bouhadou, and Youssef Ouknine. Doubly reflected backward stochastic differential equations in the predictable setting. Journal of Theoretical Probability, 35(1): 115–141, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01070-5.

Abramov:2018:CSR

[Abr18] Vyacheslav M. Abramov. Conservative and semiconservative random walks: Recurrence and transience. Journal of Theoretical Probability, 31(3):1900–1922, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). See correction [Abr19].

Abramov:2019:CCS

[Abr19] Vyacheslav M. Abramov. Correction to: Conservative and Semiconservative Random Walks: Recurrence and Transience. Journal of Theoretical Probability, 32(1):541–543, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-018-0871-8.pdf. See [Abr18].

Alberts:2019:NCP

[AC19] Tom Alberts and Jeremy Clark. Nested critical points for a directed polymer on a disordered diamond lattice. Journal of Theoretical Probability, 32(1):64–89, March 2019. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Arnaudon:2024:MIR

[ACPM24] Marc Arnaudon, Koléhè Coulibaly-Pasquier, and Laurent Miclo. On Markov intertwining relations and primal conditioning. Journal of Theoretical Probability, 37(3):2425–2456, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01301-5.

Andersen:1988:CLT

[AD88] N. T. Andersen and V. Dobric. The central limit theorem for stochastic processes II. Journal of Theoretical Probability, 1 (3):287–303, July 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01246631.

Athreya:2000:RLM

[AD00] K. B. Athreya and Jack Dai. Random logistic maps. I. Journal of Theoretical Probability, 13(2):595–608, April 2000. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007828804691.

Abraham:2007:FPI

[AD07] Romain Abraham and Jean-François Delmas. Feller property and infinitesimal generator of the exploration process. Journal of Theoretical Probability, 20(2):355–370, June 2007. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0082-1.

Attal:2010:RQI

[AD10] Stéphane Attal and Ameur Dhahri. Repeated quantum interactions and unitary random walks. Journal of Theoretical Probability, 23(2):345–361, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-010-0281-z.

Andreoletti:2011:LLL

[AD11] Pierre Andreoletti and Roland Diel. Limit law of the local time for Brox’s diffusion. Journal of Theoretical Probability, 24(3):

634–656, September 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0314-7.

Andreoletti:2014:NGE

[AD14] P. Andreoletti and P. Debs. The number of generations entirely visited for recurrent random walks in a random environment. Journal of Theoretical Probability, 27(2):518–538, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-012-0449-9.

Adamczak:2010:FRO

[Ada10] Radosław Adamczak. A few remarks on the operator norm of random Toeplitz matrices. Journal of Theoretical Probability, 23 (1):85–108, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-008-0201-7.

Avena:2024:LLN

[AdC24] Luca Avena and Conrado da Costa. Laws of large numbers for weighted sums of independent random variables: a game of mass. Journal of Theoretical Probability, 37(1):81–120, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01296-z.

Adell:2022:PSN

[Ade22] José A. Adell. Probabilistic Stirling numbers of the second kind and applications. Journal of Theoretical Probability, 35 (1):636–652, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01050-9.

Abraham:2018:CMT

[ADG18] Romain Abraham, Jean-François Delmas, and Hongsong Guo. Critical multi-type Galton–Watson trees conditioned to be large. Journal of Theoretical Probability, 31(2):757–788, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Allouba:1997:STR

[ADHP97] Hassan Allouba, Rick Durrett, John Hawkes, and Edwin Perkins. Super-tree random measures. Journal of Theoretical Probability, 10(3):773–794, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022666030740.

Adler:1990:GOS

[Adl90] André Adler. Generalized one-sided laws of the iterated logarithm for random variables barely with or without finite mean. Journal of Theoretical Probability, 3(4):587–597, October 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01046098.

Adler:1992:OSL

[Adl92] André Adler. One-sided limit theorems for sums of multidimensional indexed random variables. Journal of Theoretical Probability, 5(4):707–726, October 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01058726.

Adler:2002:GOS

[Adl02] André Adler. Generalized one-sided laws for the larger observations of a triangular array. Journal of Theoretical Probability, 15 (4):939–949, October 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1020640704437.

Aldous:1999:DFT

[AE99] David Aldous and Steven N. Evans. Dirichlet forms on totally disconnected spaces and bipartite Markov chains. Journal of Theoretical Probability, 12(3):839–857, July 1999. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021640218459.

Asmussen:2003:ASR

[AFK03] Søren Asmussen, Serguei Foss, and Dmitry Korshunov. Asymptotics for sums of random variables with local subexponential behaviour. Journal of Theoretical Probability, 16(2):489–518, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023535030388.

Angiuli:2023:FIS

[AFP23] Luciana Angiuli, Simone Ferrari, and Diego Pallara. Functional inequalities for some generalised Mehler semigroups. Journal of Theoretical Probability, 36(3):1762–1796, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01215-8.

Abramson:1991:ENT

[AG91] Ian Abramson and Larry Goldstein. Efficient nonparametric testing by functional estimation. Journal of Theoretical Probability, 4(1):137–159, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01046998.

Arcones:1993:DSE

[AG93] Miguel A. Arcones and Evarist Giné. On decoupling, series expansions, and tail behavior of chaos processes. Journal of Theoretical Probability, 6(1):101–122, January 1993. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01046771.

Avena:2018:TAR

[AG18] L. Avena and A. Gaudillière. Two applications of random spanning forests. Journal of Theoretical Probability, 31(4):1975– 2004, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-017-0771-3.pdf.

Aaronson:1992:SDM

[AGK92] Jon Aaronson, David Gilat, and Michael Keane. On the structure of 1-dependent Markov chains. Journal of Theoretical Probability, 5(3):545–561, July 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01060435.

Aurzada:2013:SDF

[AGK^{+}13] Frank Aurzada, Fuchang Gao, Thomas Kühn, Wenbo V. Li, and Qi-Man Shao. Small deviations for a family of smooth Gaussian processes. Journal of Theoretical Probability, 26(1):153–168, March 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-011-0380-5.

Albeverio:1999:EUT

[AHK99] Sergio Albeverio, Astrid Hilbert, and Vassili Kolokoltsov. Estimates uniform in time for the transition probability of diffusions with small drift and for stochastically perturbed Newton equations. Journal of Theoretical Probability, 12(2):293–300, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021665708716.

Ahlberg:2015:CTA

[Ahl15] Daniel Ahlberg. Convergence towards an asymptotic shape in first-passage percolation on cone-like subgraphs of the integer lattice. Journal of Theoretical Probability, 28(1):198–222, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-013-0521-0.

Ahmad:2005:MLT

[Ahm05] Ibrahim A. Ahmad. Mixing limit theorems of maximum of absolute sums and their convergence rates. Journal of Theoretical Probability, 18(4):813–826, October 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-005-7526-2.

Ahmed:2015:SNF

[Ahm15] Hamdy M. Ahmed. Semilinear neutral fractional stochastic integro–differential equations with nonlocal conditions. Journal of Theoretical Probability, 28(2):667–680, June 2015. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
013-0520-1.

Ahmed:2017:STF

[Ahm17] Hamdy M. Ahmed. Sobolev-type fractional stochastic integrodifferential equations with nonlocal conditions in Hilbert space. Journal of Theoretical Probability, 30(3):771–783, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Anderson:1991:ELT

[AHP91] J. M. Anderson, Joseph Horowitz, and L. D. Pitt. On the existence of local times: A geometric study. Journal of Theoretical Probability, 4(3):563–603, July 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01210325.

Anh:2021:MZT

[AHV21] Vu T. N. Anh, Nguyen T. T. Hien, and Vo T. H. Van. The Marcinkiewicz–Zygmund-type strong law of large numbers with general normalizing sequences. Journal of Theoretical Probability, 34(1):331–348, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-019-00973-2.

Adler:1996:SDS

[AI96] Robert J. Adler and Lidia Ivanitskaya. A superprocess with a disappearing self-interaction. Journal of Theoretical Probability, 9(1):245–261, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213743.

Adler:2001:CMS

[AI01] Robert J. Adler and Srikanth K. Iyer. On changes of measure for super Brownian motion. Journal of Theoretical Probability, 14(2):527–557, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1011124000185.

Aida:2024:RDE

[Aid24] Shigeki Aida. Rough differential equations containing pathdependent bounded variation terms. Journal of Theoretical Probability, 37(3):2130–2183, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-024-01319-3.

Alsmeyer:2015:PEM

[AIM15] Gerold Alsmeyer, Alexander Iksanov, and Matthias Meiners. Power and exponential moments of the number of visits and related quantities for perturbed random walks. Journal of Theoretical Probability, 28(1):1–40, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-012-0475-7.

Alsmeyer:2009:DPP

[AIR09] Gerold Alsmeyer, Alex Iksanov, and Uwe Rösler. On distributional properties of perpetuities. Journal of Theoretical Probability, 22(3):666–682, September 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0156-8.

Azmoodeh:2024:SRC

[AIS^{+}24] Ehsan Azmoodeh, Pauliina Ilmonen, Nourhan Shafik, Tommi Sottinen, and Lauri Viitasaari. On sharp rate of convergence for discretization of integrals driven by fractional Brownian motions and related processes with discontinuous integrands. Journal of Theoretical Probability, 37(1):721–743, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01272-7.

Alves:2024:NTG

[AIV24] Rodrigo B. Alves, Giulio Iacobelli, and Glauco Valle. A note on transience of generalized multi-dimensional excited random walks. Journal of Theoretical Probability, 37(2):1927–1943, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01311-3.

Alpay:2022:MBM

[AJ22] Daniel Alpay and Palle Jorgensen. Mu-Brownian motion, dualities, diffusions, transforms, and reproducing kernel Hilbert spaces. Journal of Theoretical Probability, 35(4):2757–2783, December 2022. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-021-01146-w.

Arendarczyk:2023:APE

[AJGO23] Marek Arendarczyk, Barbara Jasiulis-Gołdyn, and Edward Omey. Asymptotic properties of extremal Markov processes driven by Kendall convolution. Journal of Theoretical Probability, 36(4):2040–2065, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-023-01285-2.

Alpay:2017:EPS

[AJL17] Daniel Alpay, Palle Jorgensen, and David Levanony. On the equivalence of probability spaces. Journal of Theoretical Probability, 30(3):813–841, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Accardi:1992:STQ

[AK92] L. Accardi and D. Koroliuk. Stopping times for quantum Markov chains. Journal of Theoretical Probability, 5(3):521–535, July 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01060433.

Aurzada:2016:FPT

[AK16] Frank Aurzada and Tanja Kramm. The first passage time problem over a moving boundary for asymptotically stable Lévy processes. Journal of Theoretical Probability, 29(3):737– 760, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0596-x.

Asymont:2020:SLL

[AK20] Inna M. Asymont and Dmitry Korshunov. Strong law of large numbers for a function of the local times of a transient random walk in Z^{d}. Journal of Theoretical Probability, 33(4):2315– 2336, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-019-00937-6.

Appel:1999:SCA

[AKR99] Martin J. B. Appel, Michael J. Klass, and Ralph P. Russo. A series criterion for the almost-sure growth rate of the generalized diameter of an increasing sequence of random points. Journal of Theoretical Probability, 12(1):27–47, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021788309026.

Adell:2007:BEB

[AL07] José Antonio Adell and Alberto Lekuona. Berry–Esséen bounds for standardized subordinators via moduli of smoothness. Journal of Theoretical Probability, 20(2):221–235, June 2007. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0062-5.

Adamczak:2008:LSH

[AL08a] Radosław Adamczak and Rafał Latała. The LIL for U-statistics in Hilbert spaces. Journal of Theoretical Probability, 21(3): 704–744, September 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0134-6.

Ayache:2008:AGR

[AL08b] Antoine Ayache and Werner Linde. Approximation of Gaussian random fields: General results and optimal wavelet representation of the Lévy fractional motion. Journal of Theoretical Probability, 21(1):69–96, March 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0101-2.

Alam:2022:LSI

[Ala22] Irfan Alam. Limiting spherical integrals of bounded continuous functions. Journal of Theoretical Probability, 35(4):2304– 2342, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01147-9.

Alberink:2000:BEB

[Alb00] I. B. Alberink. A Berry–Esséen bound for U-statistics in the non– I.I.D. case. Journal of Theoretical Probability, 13(2):519–533, April 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007889323347.

Aldous:1989:ICP

[Ald89a] David Aldous. An introduction to covering problems for random walks on graphs. Journal of Theoretical Probability, 2(1):87–89, January 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01048271.

Aldous:1989:LBC

[Ald89b] David J. Aldous. Lower bounds for covering times for reversible Markov chains and random walks on graphs. Journal of Theoretical Probability, 2(1):91–100, January 1989. CODEN JTPREO.

ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01048272.

Aldous:1991:TLC

[Ald91] David J. Aldous. Threshold limits for cover times. Journal of Theoretical Probability, 4(1):197–211, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01047002.

Aldaz:2010:CRB

[Ald10] J. M. Aldaz. Concentration of the ratio between the geometric and arithmetic means. Journal of Theoretical Probability, 23 (2):498–508, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0215-9.

Allaart:2003:MMD

[All03] Pieter Allaart. Moments of the mean of Dubins–Freedman random probability distributions. Journal of Theoretical Probability, 16(2):471–488, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023582913550.

Abrams:2007:RMA

[ALL^{+}07] A. Abrams, H. Landau, Z. Landau, J. Pommersheim, and E. Zaslow. Random multiplication approaches uniform measure in finite groups. Journal of Theoretical Probability, 20(1):107–118, March 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-006-0051-0.

Ascione:2022:NLS

[ALP22] Giacomo Ascione, Nikolai Leonenko, and Enrica Pirozzi. Nonlocal solvable birth–death processes. Journal of Theoretical Probability, 35(2):1284–1323, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01087-4.

Alsmeyer:2002:MSR

[Als02] Gerold Alsmeyer. The minimal subgroup of a random walk. Journal of Theoretical Probability, 15(2):259–283, April 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014884725213.

Alsmeyer:2003:HRI

[Als03] Gerold Alsmeyer. On the Harris recurrence of iterated random Lipschitz functions and related convergence rate results. Journal of Theoretical Probability, 16(1):217–247, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022290807360.

Araya:2020:NSS

[ALT20] Héctor Araya, Jorge A. León, and Soledad Torres. Numerical scheme for stochastic differential equations driven by fractional Brownian motion with 1/4 < H < 1/2. Journal of Theoretical Probability, 33(3):1211–1237, September 2020. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Anshelevich:2012:SDL

[AM12] Michael Anshelevich and Wojciech Mlotkowski. Semigroups of distributions with linear Jacobi parameters. Journal of Theoretical Probability, 25(4):1173–1206, December 2012. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
012-0403-x.

Adamczak:2014:SOU

[AM14] Radoslaw Adamczak and Piotr Miloś. U-statistics of Ornstein– Uhlenbeck branching particle system. Journal of Theoretical Probability, 27(4):1071–1111, December 2014. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s10959-
013-0503-2.pdf.

Alsmeyer:2022:SMF

[AM22] Gerold Alsmeyer and Bastien Mallein. A simple method to find all solutions to the functional equation of the smoothing transform. Journal of Theoretical Probability, 35(4):2569– 2599, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01151-z.

Alsmeyer:2023:NHSa

[AM23a] Gerold Alsmeyer and Chiranjib Mukherjee. On null-homology and stationary sequences. Journal of Theoretical Probability, 36(1):1–25, March 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01249-6.

Alsmeyer:2023:NHSb

[AM23b] Gerold Alsmeyer and Chiranjib Mukherjee. On null-homology and stationary sequences. Journal of Theoretical Probability, 36(4):2476–2500, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-023-01249-6.

Aman:2012:RGB

[Ama12] Auguste Aman. Reflected generalized backward doubly SDEs driven by Lévy processes and applications. Journal of Theoretical Probability, 25(4):1153–1172, December 2012. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0328-1.

Aoyama:2008:STS

[AMR08] Takahiro Aoyama, Makoto Maejima, and Jan Rosiński. A subclass of type G selfdecomposable distributions on R^{d}. Journal of Theoretical Probability, 21(1):14–34, March 2008. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0129-3.

Aletti:2012:FEW

[AMS12] Giacomo Aletti, Caterina May, and Piercesare Secchi. A functional equation whose unknown is P([′, ∞]) valued. Journal of Theoretical Probability, 25(4):1207–1232, December 2012. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
011-0399-7.

Azmoodeh:2022:HDT

[AMS22] Ehsan Azmoodeh, Yuliya Mishura, and Farzad Sabzikar. How does tempering affect the local and global properties of fractional Brownian motion? Journal of Theoretical Probability, 35 (1):484–527, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01068-z.

Alos:1998:EIF

[AN98] Elisa Alòs and David Nualart. An extension of Itô’s formula for anticipating processes. Journal of Theoretical Probability, 11 (2):493–514, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022692024364.

Anonymous:1988:IS

[Ano88] Anonymous. Introductory statement. Journal of Theoretical Probability, 1(1):1, January 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01076285.

Anonymous:1989:IA

[Ano89] Anonymous. Important announcement. Journal of Theoretical Probability, 2(1):1, January 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01048265.

Anonymous:1990:IA

[Ano90] Anonymous. Important announcement. Journal of Theoretical Probability, 3(1):i, January 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01063324.

Anonymous:1991:IA

[Ano91] Anonymous. Important announcement. Journal of Theoretical Probability, 4(1):1, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01046991.

Anonymous:1993:IA

[Ano93] Anonymous. Important announcement. Journal of Theoretical Probability, 6(4):i, October 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01049166.

Anonymous:1996:IA

[Ano96] Anonymous. Important announcement. Journal of Theoretical Probability, 9(1):1, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02213731.

Anonymous:1997:PMR

[Ano97] Anonymous. Publications of Murray Rosenblatt. Journal of Theoretical Probability, 10(2):287–293, April 1997. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022652230974.

Akahori:2023:LTI

[ANS23] Jirô Akahori, Ryuya Namba, and Shunsuke Semba. Limit theorems for iterates of the Szász–Mirakyan operator in probabilistic view. Journal of Theoretical Probability, 36(2):1321–1338, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01199-5.

Arharas:2024:RDR

[AO24] Ihsan Arharas and Youssef Ouknine. Reflected and doubly reflected backward stochastic differential equations with irregular obstacles and a large set of stopping strategies. Journal of Theoretical Probability, 37(2):1001–1038, June 2024. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-024-
01331-7.

Anderson:1997:LTA

[AP97] J. M. Anderson and Loren D. Pitt. Large time asymptotics for Brownian hitting densities of transient concave curves. Journal of Theoretical Probability, 10(4):921–934, October 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022662616608.

Alili:2010:BCI

[AP10] L. Alili and P. Patie. Boundary-crossing identities for diffusions having the time-inversion property. Journal of Theoretical Probability, 23(1):65–84, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-009-0245-3.

Applebaum:1993:CSF

[App93] David Applebaum. On a class of stochastic flows driven by quantum Brownian motion. Journal of Theoretical Probability, 6 (1):17–32, January 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046766.

Applebaum:2000:CPP

[App00] David Applebaum. Compound Poisson processes and Lévy processes in groups and symmetric spaces. Journal of Theoretical Probability, 13(2):383–425, April 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1007845508326.

Applebaum:2001:OVS

[App01] David Applebaum. Operator-valued stochastic differential equations arising from unitary group representations. Journal of Theoretical Probability, 14(1):61–76, January 2001. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007816930696.

Appel:2002:SCG

[App02] M. J. Appel. Series criteria for growth rates of partial maxima of iterated ergodic map values. Journal of Theoretical Probability, 15(1):153–159, January 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1013843518677.

Aryasova:2024:HMD

[APP24] Olga Aryasova, Ilya Pavlyukevich, and Andrey Pilipenko. Homogenization of a multivariate diffusion with semipermeable interfaces. Journal of Theoretical Probability, 37(2):1787–1823, June 2024. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-024-01317-5.

Alsmeyer:2004:EML

[AR04] G. Alsmeyer and U. Rösler. On the existence of ϕ-moments of the limit of a normalized supercritical Galton–Watson process. Journal of Theoretical Probability, 17(4):905–928, October 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-004-0582-1.

Appel:2006:LDM

[AR06] M. J. Appel and R. P. Russo. Limiting distributions for the maximum of a symmetric function on a random point set. Journal of Theoretical Probability, 19(2):365–375, June 2006. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
006-0013-6.

Avraham-Reem:2023:ESD

[AR23] Nachi Avraham-Re’em. Equivalence-singularity dichotomy in Markov measures. Journal of Theoretical Probability, 36(3): 1437–1453, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01220-x.

Arcones:1994:LCP

[Arc94] M. A. Arcones. Limits of canonical U-processes and B-valued U-statistics. Journal of Theoretical Probability, 7(2):339–349, April 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02214273.

Arcones:1995:LILa

[Arc95a] Miguel A. Arcones. The low of the iterated logarithm for empirical processes under absolute regularity. Journal of Theoretical Probability, 8(2):433–451, April 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02212887.

Arcones:1995:LILb

[Arc95b] Miguel A. Arcones. On the law of the iterated logarithm for Gaussian processes. Journal of Theoretical Probability, 8(4): 877–903, October 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02410116.

Arcones:1999:LIL

[Arc99] Miguel A. Arcones. The law of the iterated logarithm over a stationary Gaussian sequence of random vectors. Journal of Theoretical Probability, 12(3):615–641, July 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1021667529846.

Alves:2021:PAR

[ARS21] Caio Alves, Rodrigo Ribeiro, and Rémy Sanchis. Preferential attachment random graphs with edge-step functions. Journal of Theoretical Probability, 34(1):438–476, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-019-
00959-0.

Adler:1997:WLR

[ARV97] André Adler, Andrew Rosalsky, and Andrej I. Volodin. Weak laws with random indices for arrays of random elements in Rademacher type p Banach spaces. Journal of Theoretical Probability, 10(3):605–623, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022645526197.

Athreya:2003:RLM

[AS03] K. B. Athreya and H.-J. Schuh. Random logistic maps II. The critical case. Journal of Theoretical Probability, 16(4):813–830, October 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOTP.0000011994.90898.81.

Asci:2001:GUR

[Asc01] Claudio Asci. Generating uniform random vectors. Journal of Theoretical Probability, 14(2):333–356, April 2001. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1011155412481.

Asci:2009:GUR

[Asc09] Claudio Asci. Generating uniform random vectors in Z_{p}^{k}: The general case. Journal of Theoretical Probability, 22(3):791– 809, September 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-008-0172-8.

Assiotis:2019:IGL

[Ass19] Theodoros Assiotis. Intertwinings for general β-Laguerre and β-Jacobi processes. Journal of Theoretical Probability, 32(4): 1880–1891, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-018-0842-0.pdf.

Astrauskas:2023:SBE

[Ast23] Arvydas Astrauskas. Some bounds for the expectations of functions on order statistics and their applications. Journal of Theoretical Probability, 36(2):1116–1147, June 2023. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-022-
01179-9.

Adler:1998:UHS

[AT98] Robert Adler and Roger Tribe. Uniqueness for a historical SDE with a singular interaction. Journal of Theoretical Probability, 11(2):515–533, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022644108434.

Alexeev:2017:SVD

[AT17] Nikita Alexeev and Alexander Tikhomirov. Singular values distribution of squares of elliptic random matrices and Type B Narayana polynomials. Journal of Theoretical Probability, 30 (3):1170–1190, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Attia:2014:MAB

[Att14] Najmeddine Attia. On the multifractal analysis of the branching random walk in R^{d}. Journal of Theoretical Probability, 27(4): 1329–1349, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0488-x.

Attia:2021:MAB

[Att21] Najmeddine Attia. On the multifractal analysis of branching random walk on Galton–Watson tree with random metric. Journal of Theoretical Probability, 34(1):90–102, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-019-00984-z.

Aurzada:2007:LTP

[Aur07] Frank Aurzada. On the lower tail probabilities of some random sequences in l_{p}. Journal of Theoretical Probability, 20(4): 843–858, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0095-9.

Austin:2011:DDH

[Aus11] Tim Austin. Deducing the density Hales–Jewett theorem from an infinitary removal lemma. Journal of Theoretical Probability, 24(3):615–633, September 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-011-0373-4.

Abadi:2009:SET

[AV09] Miguel Abadi and Nicolas Vergne. Sharp error terms for return time statistics under mixing conditions. Journal of Theoretical Probability, 22(1):18–37, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0199-x.

Azmoodeh:2015:RCD

[AV15] Ehsan Azmoodeh and Lauri Viitasaari. Rate of convergence for discretization of integrals with respect to fractional Brownian motion. Journal of Theoretical Probability, 28(1):396–422, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-013-0495-y.

Auer:2024:WMP

[AVW24] Martin Auer, Michael Voit, and Jeannette H. C. Woerner. Wigner- and Marchenko–Pastur-type limit theorems for Jacobi processes. Journal of Theoretical Probability, 37(2):1674– 1709, June 2024. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-024-01332-6.

Andersson:2004:WDH

[AW04] S. A. Andersson and G. G. Wojnar. Wishart distributions on homogeneous cones. Journal of Theoretical Probability, 17(4): 781–818, October 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-004-0576-z.

Azais:2014:TMS

[AW14] Jean-Marc Azaïs and Mario Wschebor. The tail of the maximum of smooth Gaussian fields on fractal sets. Journal of Theoretical Probability, 27(3):932–944, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-012-0429-0.

Arcones:1994:CLT

[AY94] M. A. Arcones and B. Yu. Central limit theorems for empirical and U-processes of stationary mixing sequences. Journal of Theoretical Probability, 7(1):47–71, January 1994. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02213360.

Ayache:2013:CGM

[Aya13] Antoine Ayache. Continuous Gaussian multifractional processes with random pointwise Hölder regularity. Journal of Theoretical Probability, 26(1):72–93, March 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-012-0418-3.

Albeverio:1996:MAD

[AZ96] Sergio Albeverio and Xian Yin Zhou. A martingale approach to directed polymers in a random environment. Journal of Theoretical Probability, 9(1):171–189, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02213739.

Albeverio:2001:DTL

[AZ01] Sergio Albeverio and Xuelei Zhao. A decomposition theorem for Lévy processes on local fields. Journal of Theoretical Probability, 14(1):1–19, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007878412949.

Ankirchner:2011:IEF

[AZ11] Stefan Ankirchner and Jakub Zwierz. Initial enlargement of filtrations and entropy of Poisson compensators. Journal of Theoretical Probability, 24(1):93–117, March 2011. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-010-0292-9.pdf.

Ben-Ari:2015:IDP

[BA15] Iddo Ben-Ari. On independence and determination of probability measures. Journal of Theoretical Probability, 28(3):968– 975, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0513-0.

Barhoumi-Andreani:2020:PMR

[BA20] Yacine Barhoumi-Andréani. A penalised model reproducing the mod-Poisson fluctuations in the Sathé–Selberg theorem. Journal of Theoretical Probability, 33(2):567–589, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Babillot:1994:PAH

[Bab94] M. Babillot. A probabilistic approach to heat diffusion on symmetric spaces. Journal of Theoretical Probability, 7(3):599–607, July 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02213570.

Bagh:1998:ANR

[Bag98] S. Bagh. Average number of real roots of random harmonic equations. Journal of Theoretical Probability, 11(4):857–868, October 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022616313073.

Bahier:2019:NEM

[Bah19] Valentin Bahier. On the number of eigenvalues of modified permutation matrices in mesoscopic intervals. Journal of Theoretical Probability, 32(2):974–1022, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Bai:2022:LTC

[Bai22] Shuyang Bai. Limit theorems for conservative flows on multiple stochastic integrals. Journal of Theoretical Probability, 35 (2):917–948, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01090-9.

Baldini:2007:IMP

[Bal07a] Michele Baldini. On the invariant measure of a positive recurrent diffusion in R. Journal of Theoretical Probability, 20(1):65–86, March 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-006-0046-x.

Baldini:2007:LEM

[Bal07b] Michele Baldini. On the Lyapunov exponent of a multidimensional stochastic flow. Journal of Theoretical Probability, 20 (2):327–337, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0071-4.

Balanca:2015:ITS

[Bal15] Paul Balança. An increment–type set–indexed Markov property. Journal of Theoretical Probability, 28(4):1271–1310, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-014-0555-y.

Balan:2022:SSW

[Bal22] Raluca M. Balan. Stratonovich solution for the wave equation. Journal of Theoretical Probability, 35(4):2643–2689, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01144-y.

Banek:2016:BRW

[Ban16] Tadeusz Banek. Banach random walk in the unit ball S ⊂ l^{2} and chaotic decomposition of l^{2} (S, P). Journal of Theoretical Probability, 29(4):1728–1735, December 2016. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-015-0620-1.pdf.

Bansaye:2019:ALL

[Ban19] Vincent Bansaye. Ancestral lineages and limit theorems for branching Markov chains in varying environment. Journal of Theoretical Probability, 32(1):249–281, March 2019. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Barral:2000:CMS

[Bar00] Julien Barral. Continuity of the multifractal spectrum of a random statistically self-similar measure. Journal of Theoretical Probability, 13(4):1027–1060, October 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1007866024819.

Barthe:2004:IDI

[Bar04] F. Barthe. Infinite dimensional isoperimetric inequalities in product spaces with the supremum distance. Journal of Theoretical Probability, 17(2):293–308, April 2004. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000020695.25095.c1.

Barker:2019:FDP

[Bar19] Adam Barker. Fractal-dimensional properties of subordinators. Journal of Theoretical Probability, 32(3):1202–1219, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10959-018-0813-5.pdf.

Baraquin:2020:RWF

[Bar20] Isabelle Baraquin. Random walks on finite quantum groups. Journal of Theoretical Probability, 33(3):1715–1736, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Basse:2009:SRG

[Bas09] Andreas Basse. Spectral representation of Gaussian semimartingales. Journal of Theoretical Probability, 22(4):811–826, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-009-0246-2.

Bauer:2004:LEN

[Bau04] R. O. Bauer. Löwner’s equation from a noncommutative probability perspective. Journal of Theoretical Probability, 17(2): 435–457, April 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000020702.23996.8f.

Basu:1993:SCC

[BB93] A. K. Basu and D. Bhattacharya. On the speed of convergence in the central limit theorem of log-likelihood ratio processes. Journal of Theoretical Probability, 6(4):619–634, October 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049167.

Bickel:1997:CLP

[BB97] Peter J. Bickel and Peter Bühlmann. Closure of linear processes. Journal of Theoretical Probability, 10(2):445–479, April 1997.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022616601841.

Byczkowska:1998:ZOL

[BB98] H. Byczkowska and T. Byczkowski. Zero–one law for symmetric convolution semigroups of measures on groups. Journal of Theoretical Probability, 11(3):633–643, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022698413824.

Banerjee:2013:NPC

[BB13] Sayan Banerjee and Arup Bose. Noncrossing partitions, Catalan words, and the semicircle law. Journal of Theoretical Probability, 26(2):386–409, June 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0365-4.

Brassesco:1998:IFC

[BBDP98] S. Brassesco, P. Buttà, A. De Masi, and E. Presutti. Interface fluctuations and couplings in the D = 1 Ginzburg–Landau equation with noise. Journal of Theoretical Probability, 11(1):25–80, January 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021642824394.

Bentkus:1996:BEB

[BBG96] V. Bentkus, M. Bloznelis, and F. Götze. A Berry–Esséen bound for Student’s statistic in the non-I.I.D. case. Journal of Theoretical Probability, 9(3):765–796, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02214086.

Bazarova:2016:ETC

[BBH16] Alina Bazarova, István Berkes, and Lajos Horváth. On the extremal theory of continued fractions. Journal of Theoretical Probability, 29(1):248–266, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-014-0577-5.

Baldi:1999:LMD

[BC99] Paolo Baldi and Lucia Caramellino. Large and moderate deviations for random walks on nilpotent groups. Journal of Theoretical Probability, 12(3):779–809, July 1999. CODEN JTPREO.

ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1021684000713.

Bobkov:2015:CFR

[BC15] Sergey G. Bobkov and Gennadiy P. Chistyakov. On concentration functions of random variables. Journal of Theoretical Probability, 28(3):976–988, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-013-0504-1.

Balan:2018:PAM

[BC18] Raluca M. Balan and Le Chen. Parabolic Anderson model with space–time homogeneous Gaussian noise and rough initial condition. Journal of Theoretical Probability, 31(4):2216–2265, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Budhiraja:2022:EMS

[BC22] Amarjit Budhiraja and Michael Conroy. Empirical measure and small noise asymptotics under large deviation scaling for interacting diffusions. Journal of Theoretical Probability, 35(1): 295–349, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01071-4.

Bobkov:2013:CSL

[BCG13] S. G. Bobkov, G. P. Chistyakov, and F. Götze. Convergence to stable laws in relative entropy. Journal of Theoretical Probability, 26(3):803–818, September 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-011-0377-0.

Broman:2013:DEH

[BCJM13] Erik I. Broman, Federico Camia, Matthijs Joosten, and Ronald Meester. Dimension (in)equalities and Hölder continuous curves in fractal percolation. Journal of Theoretical Probability, 26(3): 836–854, September 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-012-0413-8.pdf.

Bobkov:2015:EEK

[BCK15] S. G. Bobkov, G. P. Chistyakov, and H. Kösters. The entropic Erdős–Kac limit theorem. Journal of Theoretical Probability, 28(4):1520–1555, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-014-0550-3.

Benaim:2023:TBC

[BCOV23] Michel Benaïm, Nicolas Champagnat, William Oçafrain, and Denis Villemonais. Transcritical bifurcation for the conditional distribution of a diffusion process. Journal of Theoretical Probability, 36(3):1555–1571, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-022-01216-7.

Baez-Duarte:1993:CLT

[BD93] L. Báez-Duarte. Central limit theorem for complex measures. Journal of Theoretical Probability, 6(1):33–56, January 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01046767.

Berkes:1994:ASC

[BD94] I. Berkes and H. Dehling. On the almost sure (a.s.) central limit theorem for random variables with infinite variance. Journal of Theoretical Probability, 7(3):667–680, July 1994. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02213575.

Bryc:1997:LDQ

[BD97] Włodzimierz Bryc and Amir Dembo. Large deviations for quadratic functionals of Gaussian processes. Journal of Theoretical Probability, 10(2):307–332, April 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022656331883.

Basak:2006:CFC

[BD06] Gopal K. Basak and Amites Dasgupta. Central and functional Central Limit Theorems for a class of urn models. Journal of Theoretical Probability, 19(3):741–756, December 2006. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
006-0048-8.

Baxendale:2009:USE

[BD09] Peter Baxendale and Georgi Dimitroff. Uniform shrinking and expansion under isotropic Brownian flows. Journal of Theoretical Probability, 22(3):620–639, September 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-008-0193-3.

Berglund:2016:EKL

[BD16] Nils Berglund and Sébastien Dutercq. The Eyring–Kramers law for Markovian jump processes with symmetries. Journal of Theoretical Probability, 29(4):1240–1279, December 2016. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
015-0617-9.

Bezborodov:2021:STO

[BDK21] Viktor Bezborodov, Luca Di Persio, and Tyll Krueger. A shape theorem for a one-dimensional growing particle system with a bounded number of occupants per site. Journal of Theoretical Probability, 34(4):2265–2284, December 2021. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
01032-x.

Bourguin:2023:MDP

[BDS23] Solesne Bourguin, Thanh Dang, and Konstantinos Spiliopoulos. Moderate deviation principle for multiscale systems driven by fractional Brownian motion. Journal of Theoretical Probability, 36(1):??, March 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01235-y.

Bourguin:2024:MDP

[BDS24] Solesne Bourguin, Thanh Dang, and Konstantinos Spiliopoulos. Moderate deviation principle for multiscale systems driven by fractional Brownian motion. Journal of Theoretical Probability, 37(1):352–408, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01235-y.

Bethuelsen:2022:GCC

[BdSV22] Stein Andreas Bethuelsen, Gabriel Baptista da Silva, and Daniel Valesin. Graph constructions for the contact process with a prescribed critical rate. Journal of Theoretical Probability, 35 (2):863–893, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01063-4.

Baudoin:2021:QBW

[BDW21] Fabrice Baudoin, Nizar Demni, and Jing Wang. Quaternionic Brownian windings. Journal of Theoretical Probability, 34(4): 2368–2385, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01034-9.

Bednorz:2007:HCR

[Bed07] Witold Bednorz. Hölder continuity of random processes. Journal of Theoretical Probability, 20(4):917–934, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-007-0094-x.

Bodmann:2018:LHD

[BEG18] Bernhard G. Bodmann, Martin Ehler, and Manuel Gräf. From low- to high–dimensional moments without magic. Journal of Theoretical Probability, 31(4):2167–2193, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-017-0785-x.pdf.

Behboodian:1989:SSS

[Beh89] J. Behboodian. Symmetric sum and symmetric product of two independent random variables. Journal of Theoretical Probability, 2(2):267–270, April 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01053413.

Bierme:2010:SSR

[BEK10] Hermine Biermé, Anne Estrade, and Ingemar Kaj. Self-similar random fields and rescaled random balls models. Journal of Theoretical Probability, 23(4):1110–1141, December 2010. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
009-0259-x.

Belton:2019:SFS

[Bel19] Alexander C. R. Belton. On stopping Fock-space processes. Journal of Theoretical Probability, 32(1):484–526, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Benjamini:1991:ILP

[Ben91] Itai Benjamini. Instability of the Liouville property for quasiisometric graphs and manifolds of polynomial volume growth. Journal of Theoretical Probability, 4(3):631–637, July 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01210328.

Bentkus:1994:ABC

[Ben94] Vidmantas Bentkus. On the asymptotical behavior of the constant in the Berry–Esséen inequality. Journal of Theoretical Probability, 7(2):211–224, April 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214268.

Bentkus:2003:ITP

[Ben03] V. Bentkus. An inequality for tail probabilities of martingales with differences bounded from one side. Journal of Theoretical Probability, 16(1):161–173, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022234622381.

Bertoin:1993:CRO

[Ber93] Jean Bertoin. Conditioning a reflected one-dimensional diffusion via its canonical decomposition. Journal of Theoretical Probability, 6(2):303–321, April 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01047576.

Bertoin:1994:ISP

[Ber94] Jean Bertoin. Increase of stable processes. Journal of Theoretical Probability, 7(3):551–563, July 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02213568.

Bernadac:1995:RCF

[Ber95] Evelyne Bernadac. Random continued fractions and inverse Gaussian distribution on a symmetric cone. Journal of Theoretical Probability, 8(2):221–259, April 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02212879.

Berthet:1997:RCS

[Ber97] Philippe Berthet. On the rate of clustering to the Strassen set for increments of the uniform empirical process. Journal of Theoretical Probability, 10(3):557–579, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022632825288.

Bercovici:2005:SFR

[Ber05a] Hari Bercovici. Series of free random variables. Journal of Theoretical Probability, 18(4):957–965, October 2005. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
005-7540-4.

Berg:2005:PSM

[Ber05b] Christian Berg. On powers of Stieltjes moment sequences, I. Journal of Theoretical Probability, 18(4):871–889, October 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-005-7530-6.

Bhamidi:2012:SLR

[BES12] Shankar Bhamidi, Steven N. Evans, and Arnab Sen. Spectra of large random trees. Journal of Theoretical Probability, 25(3): 613–654, September 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0360-9.

Belfadli:2023:VEG

[BESF23] Rachid Belfadli, Khalifa Es-Sebaiy, and Fatima-Ezzahra Farah. Volatility estimation of Gaussian Ornstein–Uhlenbeck processes of the second kind. Journal of Theoretical Probability, 36(1): ??, March 2023. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-023-01238-9.

Belfadli:2024:VEG

[BESF24] Rachid Belfadli, Khalifa Es-Sebaiy, and Fatima-Ezzahra Farah. Volatility estimation of Gaussian Ornstein–Uhlenbeck processes of the second kind. Journal of Theoretical Probability, 37(1): 860–876, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01238-9.

Betounes:1995:TOF

[Bet95] David Betounes. Trace operators, Feynman distributions, and multiparameter white noise. Journal of Theoretical Probability, 8 (1):119–138, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213457.

Betsch:2023:SPG

[Bet23] Steffen Betsch. Structural properties of Gibbsian point processes in abstract spaces. Journal of Theoretical Probability, 36(4): 2501–2563, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01262-9.

Borisov:2013:AEL

[BF13] Denis Borisov and Pedro Freitas. Asymptotics for the expected lifetime of Brownian motion on thin domains in R^{n}. Journal of Theoretical Probability, 26(1):284–309, March 2013. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
011-0363-6.

Bouali:2024:SIR

[BF24] Mohamed Bouali and Jacques Faraut. Sylvester index of random Hermitian matrices. Journal of Theoretical Probability, 37 (1):768–813, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01232-7.

Budhiraja:2018:LDB

[BFW18] Amarjit Budhiraja, Wai-Tong (Louis) Fan, and Ruoyu Wu. Large deviations for Brownian particle systems with Killing. Journal of Theoretical Probability, 31(3):1779–1818, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Barbour:1995:FDS

[BG95] A. D. Barbour and R. Grübel. The first divisible sum. Journal of Theoretical Probability, 8(1):39–47, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02213453.

Benaych-Georges:2011:RTL

[BG11] Florent Benaych-Georges. Rectangular R-transform as the limit of rectangular spherical integrals. Journal of Theoretical Probability, 24(4):969–987, December 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-011-0362-7.

Bischoff:2016:CMT

[BG16] Wolfgang Bischoff and Andreas Gegg. The Cameron–Martin theorem for (p-)Slepian processes. Journal of Theoretical Probability, 29(2):707–715, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-014-0591-7.

Benaych-Georges:2010:MIB

[BGCD10] Florent Benaych-Georges and Thierry Cabanal-Duvillard. A matrix interpolation between classical and free max operations. I. The univariate case. Journal of Theoretical Probability, 23 (2):447–465, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0210-1.

Benaych-Georges:2019:ESD

[BGEM19] Florent Benaych-Georges, Nathanaël Enriquez, and Alkéos Michaïl. Empirical spectral distribution of a matrix under perturbation. Journal of Theoretical Probability, 32(3):1220–1251, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Berkes:2017:TPS

[BGK17] István Berkes, László Györfi, and Péter Kevei. Tail probabilities of St. Petersburg sums, trimmed sums, and their limit. Journal of Theoretical Probability, 30(3):1104–1129, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Bojdecki:2004:FBD

[BGT04] T. Bojdecki, L. G. Gorostiza, and A. Talarczyk. Fractional Brownian density process and its self-intersection local time of order k. Journal of Theoretical Probability, 17(3):717–739, July 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000040296.95910.e1.

Bobkov:2010:CEM

[BGT10] S. G. Bobkov, F. Götze, and A. N. Tikhomirov. On concentration of empirical measures and convergence to the semi-circle law. Journal of Theoretical Probability, 23(3):792–823, September 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-010-0286-7.

Bojdecki:2015:ILT

[BGT15] Tomasz Bojdecki, Luis G. Gorostiza, and Anna Talarczyk. From intersection local time to the Rosenblatt process. Journal of Theoretical Probability, 28(3):1227–1249, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-013-0535-7.pdf.

Bentkus:1993:AEI

[BGZ93] Vidmantas Bentkus, Friedrich Götze, and Ricardas Zitikis. Asymptotic expansions in the integral and local limit theorems in Banach spaces with applications to ω-statistics. Journal of Theoretical Probability, 6(4):727–780, October 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01049174.

Bloom:1988:CCS

[BH88] Walter R. Bloom and Herbert Heyer. Continuity of convolution semigroups on hypergroups. Journal of Theoretical Probability, 1(3):271–286, July 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01246630.

Berkes:1999:LTL

[BH99] István Berkes and Lajos Horváth. Limit theorems for logarithmic averages of fractional Brownian motions. Journal of Theoretical Probability, 12(4):985–1009, October 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1021641020103.

Bloomer:2002:RPM

[BH02] Lisa Bloomer and Theodore P. Hill. Random probability measures with given mean and variance. Journal of Theoretical Probability, 15(4):919–937, October 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1020688620366.

Bryc:2011:OSC

[BH11] Włodzimierz Bryc and Abdelhamid Hassairi. One-sided Cauchy– Stieltjes kernel families. Journal of Theoretical Probability, 24 (2):577–594, June 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0303-x.

Bao:2022:AMV

[BH22] Jianhai Bao and Xing Huang. Approximations of McKean– Vlasov stochastic differential equations with irregular coefficients. Journal of Theoretical Probability, 35(2):1187–1215, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01082-9.

Bai:2023:LHI

[BH23] Yifan Bai and Xing Huang. Log-Harnack inequality and exponential ergodicity for distribution dependent Chan–Karolyi– Longstaff–Sanders and Vasicek models. Journal of Theoretical Probability, 36(3):1902–1921, September 2023. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-022-
01210-z.

Berkes:2024:SOC

[BH24] István Berkes and Siegfried Hörmann. Some optimal conditions for the ASCLT. Journal of Theoretical Probability, 37 (1):209–227, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01245-w.

Bhar:2017:CIF

[Bha17a] Suprio Bhar. Correction to: An Itô Formula in the Space of Tempered Distributions. Journal of Theoretical Probability, 30(4): 1786–1787, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-017-0792-y.pdf.  See

[Bha17b].

Bhar:2017:IFS

[Bha17b] Suprio Bhar. An Itô formula in the space of tempered distributions. Journal of Theoretical Probability, 30(2):510–528, June 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). See correction [Bha17a].

Berkes:2010:FVA

[BHH10] István Berkes, Siegfried Hörmann, and Lajos Horváth. On functional versions of the arc-sine law. Journal of Theoretical Probability, 23(1):109–126, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0181-7.

Barchielli:1993:AHR

[BHL93] A. Barchielli, A. S. Holevo, and G. Lupieri. An analogue of Hunt’s representation theorem in quantum probability. Journal of Theoretical Probability, 6(2):231–265, April 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01047573.

Berger:2008:SDI

[BHM08] Arno Berger, Theodore P. Hill, and Kent E. Morrison. Scaledistortion inequalities for mantissas of finite data sets. Journal of Theoretical Probability, 21(1):97–117, March 2008. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0112-z.

Bose:2011:SNC

[BHS11] Arup Bose, Rajat Subhra Hazra, and Koushik Saha. Spectral norm of circulant-type matrices. Journal of Theoretical Probability, 24(2):479–516, June 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-009-0257-z.

Berkes:2010:ULC

[BHW10] István Berkes, Siegfried Hörmann, and Michel Weber. Upperlower class tests for weighted i.i.d. sequences and martingales. Journal of Theoretical Probability, 23(2):428–446, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-009-0219-5.

Bao:2019:CRE

[BHY19] Jianhai Bao, Xing Huang, and Chenggui Yuan. Convergence rate of Euler–Maruyama scheme for SDEs with Hölder–Dini continuous drifts. Journal of Theoretical Probability, 32(2):848–871, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10959-018-0854-9.pdf.

Balan:2002:MPS

[BI02] R. M. Balan and B. G. Ivanoff. A Markov property for setindexed processes. Journal of Theoretical Probability, 15(3): 553–588, July 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1016296330187.

Bingham:1988:FCL

[Bin88] Michael S. Bingham. Functional central limit theorems for approximate martingale arrays in a locally compact Abelian group. Journal of Theoretical Probability, 1(1):3–26, January 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01076286.

Bingham:1993:CAM

[Bin93] Michael S. Bingham. Convergence of approximate martingale arrays to mixtures of infinitely divisible distributions in locally compact abelian groups. Journal of Theoretical Probability, 6 (1):123–152, January 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046772.

Bisgaard:1990:HIC

[Bis90] Torben Maack Bisgaard. Hoeffding’s inequalities: A counterexample. Journal of Theoretical Probability, 3(1):71–80, January 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01063329.

Bisgaard:1996:MMS

[Bis96] Torben Maack Bisgaard. Method of moments on semigroups. Journal of Theoretical Probability, 9(3):631–645, July 1996.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02214079.

Bardina:2002:EDH

[BJ02] Xavier Bardina and Maria Jolis. Estimation of the density of hypoelliptic diffusion processes with application to an extended Itô’s formula. Journal of Theoretical Probability, 15(1):223–247, January 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1013899603656.

Bradley:2016:SMO

[BJ16] Richard C. Bradley and Zbigniew J. Jurek. Strong mixing and operator-selfdecomposability. Journal of Theoretical Probability, 29(1):292–306, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0575-7.

Berlinkov:2019:PMP

[BJ19] Artemi Berlinkov and Esa Järvenpää. Porosities of Mandelbrot percolation. Journal of Theoretical Probability, 32(2):608–632, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Balan:2016:FCL

[BJL16] Raluca Balan, Adam Jakubowski, and Sana Louhichi. Functional convergence of linear processes with heavy-tailed innovations. Journal of Theoretical Probability, 29(2):491–526, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10959-014-0581-9.pdf.

Baxter:2004:SWI

[BJLO04] John Baxter, Roger Jones, Michael Lin, and James Olsen. SLLN for weighted independent identically distributed random variables. Journal of Theoretical Probability, 17(1):165–181, January 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOTP.0000020480.84425.8d.

Bjorklund:2010:CLT

[Bjö10] Michael Björklund. Central limit theorems for Gromov hyperbolic groups. Journal of Theoretical Probability, 23(3):871– 887, September 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0230-x.

Bentkus:2001:LLS

[BJP01] V. Bentkus, A. Juozulynas, and V. Paulauskas. Lévy–LePage series representation of stable vectors: Convergence in variation. Journal of Theoretical Probability, 14(4):949–978, October 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1017520702943.

Bridaa:2024:GUS

[BJS24] Safa Bridaa, Wissem Jedidi, and Hristo Sendov. Generalized unimodality and subordinators, with applications to stable laws and to the Mittag-Leffler function. Journal of Theoretical Probability, 37(1):1–42, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01242-z.

Banek:2019:LDB

[BJZ19] Tadeusz Banek, Patrycja Jedrzejewska, and August M. Zapala. Limit distribution of the Banach random walk. Journal of Theoretical Probability, 32(1):47–63, March 2019. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-018-0858-5.pdf.

Broder:1989:BCT

[BK89] Andrei Z. Broder and Anna R. Karlin. Bounds on the cover time. Journal of Theoretical Probability, 2(1):101–120, January 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048273.

Burton:1993:SMR

[BK93] Robert M. Burton and Gerhard Keller. Stationary measures for randomly chosen maps. Journal of Theoretical Probability, 6(1):1–16, January 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046765.

Bennies:2000:RWA

[BK00] Jürgen Bennies and Götz Kersting. A random walk approach to Galton–Watson trees. Journal of Theoretical Probability, 13 (3):777–803, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007862612753.

Becker-Kern:2003:SSH

[BK03] Peter Becker-Kern. Stable and semistable hemigroups: Domains of attraction and self-decomposability. Journal of Theoretical Probability, 16(3):573–598, July 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025664314657.

Becker:2009:MDL

[BK09] Mathias Becker and Wolfgang König. Moments and distribution of the local times of a transient random walk on Z^{d}. Journal of Theoretical Probability, 22(2):365–374, June 2009. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
008-0168-4.

Borkar:2010:NMS

[BK10] V. S. Borkar and K. Suresh Kumar. A new Markov selection procedure for degenerate diffusions. Journal of Theoretical Probability, 23(3):729–747, September 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-009-0242-6.

Buraczewski:2012:RWA

[BK12] Dariusz Buraczewski and Konrad Kolesko. Random walks on the affine group of a homogeneous tree in the drift-free case. Journal of Theoretical Probability, 25(1):189–204, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-010-0323-6.pdf.

Bromberg:2014:WIP

[BK14] Michael Bromberg and Zemer Kosloff. Weak invariance principle for the local times of partial sums of Markov chains. Journal of Theoretical Probability, 27(2):493–517, June 2014. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
012-0438-z.

Basrak:2015:MFL

[BK15] Bojan Basrak and Danijel Krizmanić. A multivariate functional limit theorem in weak M_{1} topology. Journal of Theoretical Probability, 28(1):119–136, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-013-0510-3.

Bucher:2019:NCV

[BK19] Axel Bücher and Ivan Kojadinovic. A note on conditional versus joint unconditional weak convergence in bootstrap consistency results. Journal of Theoretical Probability, 32(3):1145– 1165, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Bae:1995:UCM

[BL95a] Jongsig Bae and Shlomo Levental. Uniform CLT for Markov chains and its invariance principle: A martingale approach. Journal of Theoretical Probability, 8(3):549–570, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02218044.

Bhattacharya:1995:ENF

[BL95b] Rabi N. Bhattacharya and Chanho Lee. Ergodicity of nonlinear first order autoregressive models. Journal of Theoretical Probability, 8(1):207–219, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02213462.

Benjamini:1997:CDW

[BL97] Itai Benjamini and Susan Lee. Conditioned diffusions which are Brownian bridges. Journal of Theoretical Probability, 10 (3):733–736, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022657828923.

Belinsky:2002:SBP

[BL02] Eduard Belinsky and Werner Linde. Small ball probabilities of fractional Brownian sheets via fractional integration operators. Journal of Theoretical Probability, 15(3):589–612, July 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1016263614257.

Barbour:2006:TPA

[BL06] A. D. Barbour and Torgny Lindvall. Translated Poisson approximation for Markov chains. Journal of Theoretical Probability, 19(3):609–630, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-006-0047-9.

Balan:2009:CPP

[BL09a] Raluca M. Balan and Sana Louhichi. Convergence of point processes with weakly dependent points. Journal of Theoretical Probability, 22(4):955–982, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-008-0176-4.

Barbour:2009:TPA

[BL09b] A. D. Barbour and Torgny Lindvall. Translated Poisson approximation for Markov chains. Journal of Theoretical Probability, 22(1):279–280, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-008-0190-6.pdf.

Boman:2009:STR

[BL09c] Jan Boman and Filip Lindskog. Support theorems for the Radon transform and Cramér–Wold theorems. Journal of Theoretical Probability, 22(3):683–710, September 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-008-0151-0.

Biane:2011:MPS

[BL11] Philippe Biane and Gérard Letac. The mean perimeter of some random plane convex sets generated by a Brownian motion. Journal of Theoretical Probability, 24(2):330–341, June 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-009-0272-0.

Bryc:2013:MME

[BL13] Włodzimierz Bryc and Gérard Letac. Meixner matrix ensembles. Journal of Theoretical Probability, 26(1):107–152, March 2013.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-011-0368-1.

Behme:2015:EFL

[BL15] Anita Behme and Alexander Lindner. On exponential functionals of Lévy processes. Journal of Theoretical Probability, 28 (2):681–720, June 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0507-y.

Bai:2019:DDF

[BL19a] Long Bai and Peng Liu. Drawdown and drawup for fractional Brownian motion with trend. Journal of Theoretical Probability, 32(3):1581–1612, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

BenAmmou:2019:RCW

[BL19b] Bilel Kacem Ben Ammou and Alberto Lanconelli. Rate of convergence for Wong–Zakai-type approximations of Itô stochastic differential equations. Journal of Theoretical Probability, 32(4): 1780–1803, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Bakhtin:2022:SSP

[BL22] V. I. Bakhtin and A. V. Lebedev. Sup-sums principles for F - divergence and a new definition for t-entropy. Journal of Theoretical Probability, 35(1):350–369, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-020-01046-5.

Blackburn:2000:LDL

[Bla00] Robert Blackburn. Large deviations of local times of Lévy processes. Journal of Theoretical Probability, 13(3):825–842, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1007866713661.

Bar-Lev:1994:DMN

[BLBE^{+}94] Shaul K. Bar-Lev, Daoud Bshouty, Peter Enis, Gérard Letac, I-Li Lu, and Donald Richards. The diagonal multivariate natural exponential families and their classification. Journal of Theoretical Probability, 7(4):883–929, October 1994. CODEN JTPREO.

ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02214378.

Bar-Lev:2003:CRN

[BLC03] Shaul K. Bar-Lev and Muriel Casalis. A classification of reproducible natural exponential families in the broad sense. Journal of Theoretical Probability, 16(1):175–196, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022286606451.

Bloznelis:1996:CLT

[Blo96] M. Bloznelis. Central limit theorem for stochastically continuous processes. Convergence to stable limit. Journal of Theoretical Probability, 9(3):541–560, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214074.

Barczy:2016:MFM

[BLP16] Mátyás Barczy, Zenghu Li, and Gyula Pap. Moment formulas for multitype continuous state and continuous time branching process with immigration. Journal of Theoretical Probability, 29(3): 958–995, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0605-0.

Bender:2012:FVF

[BLS12] Christian Bender, Alexander Lindner, and Markus Schicks. Finite variation of fractional Lévy processes. Journal of Theoretical Probability, 25(2):594–612, June 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-010-0339-y.

Bender:2024:GTF

[BLV24] Christian Bender, Joachim Lebovits, and Jacques Lévy Véhel. General transfer formula for stochastic integral with respect to multifractional Brownian motion. Journal of Theoretical Probability, 37(1):905–932, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-023-01258-5.

Brydges:1991:ABI

[BM91] David C. Brydges and Ismael Muñoz Maya. An application of Berezin integration to large deviations. Journal of Theoretical Probability, 4(2):371–389, April 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01258743.

Budzban:1992:CPN

[BM92] Greg Budzban and Arunava Mukherjea. Convolution products of nonidentical distributions on a topological semigroup. Journal of Theoretical Probability, 5(2):283–307, April 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01046736.

Broniatowski:1994:ELD

[BM94] Michel Broniatowski and David M. Mason. Extended large deviations. Journal of Theoretical Probability, 7(3):647–666, July 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF02213574.

Bhattacharya:1999:TDF

[BM99] Rabi Bhattacharya and Mukul Majumdar. On a theorem of Dubins and Freedman. Journal of Theoretical Probability, 12(4): 1067–1087, October 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021601421920.

Bally:2001:WSS

[BM01] V. Bally and A. Matoussi. Weak solutions for SPDEs and backward doubly stochastic differential equations. Journal of Theoretical Probability, 14(1):125–164, January 2001. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007825232513.

Berlinkov:2002:PMD

[BM02] Artemi Berlinkov and R. Daniel Mauldin. Packing measure and dimension of random fractals. Journal of Theoretical Probability, 15(3):695–713, July 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1016271916074.

Baraka:2011:EHM

[BM11] D. Baraka and T. S. Mountford. The exact Hausdorff measure of the zero set of fractional Brownian motion. Journal of Theoretical Probability, 24(1):271–293, March 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-009-0271-1.

Berezin:2014:ABC

[BM14] Roman Berezin and Leonid Mytnik. Asymptotic behaviour of the critical value for the contact process with rapid stirring. Journal of Theoretical Probability, 27(3):1045–1057, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-012-0470-z.

Banna:2015:LSD

[BM15] Marwa Banna and Florence Merlevède. Limiting spectral distribution of large sample covariance matrices associated with a class of stationary processes. Journal of Theoretical Probability, 28(2):745–783, June 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0508-x.

Berckmoes:2018:ACL

[BM18] Ben Berckmoes and Geert Molenberghs. Approximate central limit theorems. Journal of Theoretical Probability, 31(3):1590– 1605, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Brown:2022:RGP

[BMP22] Erin Crossen Brown, Sevak Mkrtchyan, and Jonathan Pakianathan. Random gap processes and asymptotically complete sequences. Journal of Theoretical Probability, 35(2):801–818, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s10959-021-01091-8.

Bose:2012:LSD

[BMS12] Arup Bose, Joydip Mitra, and Arnab Sen. Limiting spectral distribution of random k-circulants. Journal of Theoretical Probability, 25(3):771–797, September 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-010-0312-9.

Bai:1999:RCR

[BMT99] Z. D. Bai, Baiqi Miao, and Jhishen Tsay. Remarks on the convergence rate of the spectral distributions of Wigner matrices. Journal of Theoretical Probability, 12(2):301–311, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021617825555.

Banica:2013:AED

[BN13] Teodor Banica and Ion Nechita. Asymptotic eigenvalue distributions of block-transposed Wishart matrices. Journal of Theoretical Probability, 26(3):855–869, September 2013. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
012-0409-4.

Bahier:2022:SML

[BN22] Valentin Bahier and Joseph Najnudel. On smooth mesoscopic linear statistics of the eigenvalues of random permutation matrices. Journal of Theoretical Probability, 35(3):1640–1661, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01106-4.

Barndorff-Nielsen:2006:IDS

[BNMiS06] Ole E. Barndorff-Nielsen, Makoto Maejima, and Ken iti Sato. Infinite divisibility for stochastic processes and time change. Journal of Theoretical Probability, 19(2):411–446, June 2006. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
006-0020-7.

Barndorff-Nielsen:2017:SF

[BNSS17] Ole E. Barndorff-Nielsen, Orimar Sauri, and Benedykt Szozda. Selfdecomposable fields. Journal of Theoretical Probability, 30 (1):233–267, March 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0630-z.

Bobkov:2005:GSP

[Bob05] Sergey G. Bobkov. Generalized symmetric polynomials and an approximate de Finetti representation. Journal of Theoretical Probability, 18(2):399–412, April 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-005-3509-6.

Bobkov:2018:CLT

[Bob18] Sergey G. Bobkov. Central limit theorem and Diophantine approximations. Journal of Theoretical Probability, 31(4):2390– 2411, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Bogachev:2006:LLN

[Bog06] Leonid Bogachev. Limit laws for norms of IID samples with Weibull tails. Journal of Theoretical Probability, 19(4):849– 873, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-006-0036-z.

Bogso:2020:MRL

[Bog20] Antoine-Marie Bogso. Mean residual life processes and associated submartingales. Journal of Theoretical Probability, 33(1): 36–64, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Banos:2022:SSS

[BOLP22] David Baños, Salvador Ortiz-Latorre, and Frank Proske. Strong solutions of stochastic differential equations with generalized drift and multidimensional fractional Brownian initial noise. Journal of Theoretical Probability, 35(2):714–771, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01084-7.

Bondesson:2015:CPD

[Bon15] Lennart Bondesson. A class of probability distributions that is closed with respect to addition as well as multiplication of independent random variables. Journal of Theoretical Probability, 28(3):1063–1081, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-013-0523-y.

Born:1989:ELH

[Bor89] Eike Born. An explicit Lévy-Hincin formula for convolution semigroups on locally compact groups. Journal of Theoretical Probability, 2(3):325–342, July 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01054020.

Bose:1998:GCT

[Bos98] Arup Bose. A Glivenko–Cantelli theorem and strong laws for L-statistics. Journal of Theoretical Probability, 11(4):921–933, October 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022660730820.

Bouchard:2005:VCB

[Bou05] B. Bouchard. A version of the G-conditional bipolar theorem in L^{0}(R_{+}^{d}; Ω, F, P). Journal of Theoretical Probability, 18 (2):439–467, April 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-005-3512-y.

Boukhari:2022:WLL

[Bou22] Fakhreddine Boukhari. On a weak law of large numbers with regularly varying normalizing sequences. Journal of Theoretical Probability, 35(3):2068–2079, September 2022. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-021-
01120-6.

Bentkus:1996:AGA

[BP96] Vidmantas Bentkus and Gyula Pap. The accuracy of Gaussian approximations in nilpotent Lie groups. Journal of Theoretical Probability, 9(4):995–1017, October 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02214261.

Baek:2010:RAC

[BP10] Jong-Il Baek and Sung-Tae Park. [RETRACTED ARTICLE]: Convergence of weighted sums for arrays of negatively dependent random variables and its applications. Journal of Theoretical Probability, 23(2):362–377, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0198-y. See re-

traction [BP13].

Baek:2013:RNC

[BP13] Jong-Il Baek and Sung-Tae Park. Retraction note to: Convergence of Weighted Sums for Arrays of Negatively Dependent Random Variables and Its Applications. Journal of Theoretical Probability, 26(3):899–900, September 2013. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-013-0493-0.pdf. See [BP10].

Brassesco:2014:DWN

[BP14] Stella Brassesco and Silvana C. García Pire. On the density of the winding number of planar Brownian motion. Journal of Theoretical Probability, 27(3):899–914, September 2014. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
012-0462-z.

Bansaye:2021:PTL

[BP21] Vincent Bansaye and S. Valère Bitseki Penda. A phase transition for large values of bifurcating autoregressive models. Journal of Theoretical Probability, 34(4):2081–2116, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01033-w.

Barczy:2023:LTD

[BP23] Mátyás Barczy and Zsolt Páles. Limit theorems for deviation means of independent and identically distributed random variables. Journal of Theoretical Probability, 36(3):1626–1666, September 2023. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-022-01225-6.

Benjamini:1996:RWV

[BPP96] Itai Benjamini, Robin Pemantle, and Yuval Peres. Random walks in varying dimensions. Journal of Theoretical Probability, 9 (1):231–244, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213742.

Berti:2013:FAE

[BPR13] Patrizia Berti, Luca Pratelli, and Pietro Rigo. Finitely additive equivalent martingale measures. Journal of Theoretical Probability, 26(1):46–57, March 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-010-0337-0.

Bentkus:2003:OBC

[BPY03] V. Bentkus, G. Pap, and M. Yor. Optimal bounds for Cauchy approximations for the winding distribution of planar Brownian motion. Journal of Theoretical Probability, 16(2):345–361, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023566409916.

Byczkowski:1990:SDN

[BR90] Tomasz Byczkowski and Michał Ryznar. Smoothness of the distribution of the norm in uniformly convex Banach spaces. Journal of Theoretical Probability, 3(3):433–448, July 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01061261.

Byczkowski:1992:TBS

[BR92] Tomasz Byczkowski and Balram S. Rajput. Tail behavior for semigroups of measures on groups. Journal of Theoretical Probability, 5(3):537–544, July 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01060434.

Budzban:1997:SRC

[BR97] Greg Budzban and Imre Ruzsa. Some results concerning convergence of convolution products of probability measures on discrete semigroups. Journal of Theoretical Probability, 10(1):185–200, January 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022602700716.

Berti:1999:SCE

[BR99] Patrizia Berti and Pietro Rigo. Sufficient conditions for the existence of disintegrations. Journal of Theoretical Probability, 12 (1):75–86, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021792409934.

Berti:2008:CLS

[BR08a] Patrizia Berti and Pietro Rigo. A conditional 0–1 law for the symmetric σ-field. Journal of Theoretical Probability, 21(3): 517–526, September 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-008-0174-6.

Borovkov:2008:SDS

[BR08b] A. A. Borovkov and P. S. Ruzankin. On small deviations of series of weighted random variables. Journal of Theoretical Probability, 21(3):628–649, September 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0130-x.

Barrasso:2021:BSD

[BR21] Adrien Barrasso and Francesco Russo. Backward stochastic differential equations with no driving martingale, Markov processes and associated pseudo-partial differential equations: Part II — decoupled mild solutions and examples. Journal of Theoretical Probability, 34(3):1110–1148, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01092-7.

Bagyan:2023:HJI

[BR23] Armine Bagyan and Donald Richards. Hoffmann–Jørgensen inequalities for random walks on the cone of positive definite matrices. Journal of Theoretical Probability, 36(2):1181–1202, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01189-7.

Bradley:1988:SRM

[Bra88] Richard C. Bradley. On some results of M. I. Gordin: A clarification of a misunderstanding. Journal of Theoretical Probability, 1(2):115–119, April 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046930.

Braverman:1991:CPD

[Bra91] Michael Braverman. A characterization of probability distributions by mean values. Journal of Theoretical Probability, 4(4): 691–700, October 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01259550.

Bradley:1992:SDA

[Bra92] Richard C. Bradley. On the spectral density and asymptotic normality of weakly dependent random fields. Journal of Theoretical Probability, 5(2):355–373, April 1992. CODEN JTPREO.

ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01046741.

Braverman:1993:RCN

[Bra93] M. Sh. Braverman. Remarks on characterization of normal and stable distributions. Journal of Theoretical Probability, 6 (2):407–415, April 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01047582.

Braverman:1994:CPD

[Bra94] Michael Braverman. A characterization of probability distributions by moments of sums of independent random variables. Journal of Theoretical Probability, 7(1):187–198, January 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02213367.

Bradley:1996:MCT

[Bra96] Richard C. Bradley. A “multiplicative coboundary” theorem for some sequences of random matrices. Journal of Theoretical Probability, 9(3):659–678, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214081.

Bradley:1997:QCL

[Bra97] Richard C. Bradley. On quantiles and the central limit question for strongly mixing sequences. Journal of Theoretical Probability, 10(2):507–555, April 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022624919588.

Bradley:2001:SRM

[Bra01] Richard C. Bradley. A stationary rho-mixing Markov chain which is not “interlaced” rho-mixing. Journal of Theoretical Probability, 14(3):717–727, July 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017545123473.

Breton:2002:MSS

[Bre02] Jean-Christophe Breton. Multiple stable stochastic integrals: Series representation and absolute continuity of their law. Journal of Theoretical Probability, 15(4):877–901, October 2002.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020684519458.

Breton:2005:ACJ

[Bre05] Jean-Christophe Breton. Absolute continuity of joint laws of multiple stable stochastic integrals. Journal of Theoretical Probability, 18(1):43–77, January 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-004-2575-5.

Breton:2010:RLS

[Bre10] Jean-Christophe Breton. Regularity of the laws of shot noise series and of related processes. Journal of Theoretical Probability, 23(1):21–38, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0255-1.

Brillinger:1997:PMR

[Bri97] David R. Brillinger. A particle migrating randomly on a sphere. Journal of Theoretical Probability, 10(2):429–443, April 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022869817770.

Brown:2000:SRM

[Bro00] Kenneth S. Brown. Semigroups, rings, and Markov chains. Journal of Theoretical Probability, 13(3):871–938, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007822931408.

Bronski:2003:SBC

[Bro03] Jared C. Bronski. Small ball constants and tight eigenvalue asymptotics for fractional Brownian motions. Journal of Theoretical Probability, 16(1):87–100, January 2003. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022226420564.

Brofferio:2004:RTA

[Bro04] S. Brofferio. Renewal theory on the affine group of an oriented tree. Journal of Theoretical Probability, 17(4):819–859, October 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-004-0577-y.

Broman:2016:SOI

[Bro16] Erik I. Broman. Stochastic ordering of infinite geometric Galton– Watson trees. Journal of Theoretical Probability, 29(3):1069– 1082, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0608-x.

Bru:1991:WP

[Bru91] Marie-France Bru. Wishart processes. Journal of Theoretical Probability, 4(4):725–751, October 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01259552.

Bryc:2003:LDE

[Bry03] Włodzimierz Bryc. Large deviations of empirical measures under symmetric interaction. Journal of Theoretical Probability, 16(4): 935–955, October 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000012001.36404.18.

Byczkowski:1996:ZDI

[BRZ96] Tomasz Byczkowski, Balram S. Rajput, and Tomasz Zak. Zeros of the densities of infinitely divisible measures on R^{n}. Journal of Theoretical Probability, 9(2):533–540, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02214662.

Berger:1988:RWG

[BS88] Marc A. Berger and H. Mete Soner. Random walks generated by affine mappings. Journal of Theoretical Probability, 1 (3):239–254, July 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01246627.

Bryc:1993:CAM

[BS93] W. Bryc and W. Smoleński. On the convergence of averages of mixing sequences. Journal of Theoretical Probability, 6(3):473–483, July 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01066713.

Bose:1999:SLL

[BS99] Arup Bose and Arusharka Sen. The strong law of large numbers for Kaplan–Meier U-statistics. Journal of Theoretical Probability, 12(1):181–200, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1021752828590.

Beghdadi-Sakrani:2003:PUS

[BS03] Samia Beghdadi-Sakrani. On pathwise uniqueness of stochastic differential equations without drift. Journal of Theoretical Probability, 16(4):789–812, October 2003. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000011993.52094.c7.

Ball:2005:EBR

[BS05] Karen Ball and Jacob Sterbenz. Explicit bounds for the return probability of simple random walks. Journal of Theoretical Probability, 18(2):317–326, April 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-003-2606-7.

Bose:2013:FDR

[BS13] Arup Bose and Sanchayan Sen. Finite diagonal random matrices. Journal of Theoretical Probability, 26(3):819–835, September 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-011-0378-z.

Balan:2020:SLM

[BS20] Raluca M. Balan and Becem Saidani. Stable Lévy motion with values in the Skorokhod space: Construction and approximation. Journal of Theoretical Probability, 33(2):1061–1110, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Bendikov:2001:ACG

[BSC01] A. Bendikov and L. Saloff-Coste. On the absolute continuity of Gaussian measures on locally compact groups. Journal of Theoretical Probability, 14(3):887–898, July 2001. CODEN JTPREO.

ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1012261627108.

Barreto-Souza:2015:LRT

[BSF15] W. Barreto-Souza and L. R. G. Fontes. Long-range trap models on Z and quasistable processes. Journal of Theoretical Probability, 28(4):1500–1519, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-014-0548-x.

Bo:2010:SWE

[BSW10] Lijun Bo, Kehua Shi, and Yongjin Wang. On a stochastic wave equation driven by a non–Gaussian Lévy process. Journal of Theoretical Probability, 23(1):328–343, March 2010. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
009-0228-4.

Bobkov:2006:MLS

[BT06] Sergey G. Bobkov and Prasad Tetali. Modified logarithmic Sobolev inequalities in discrete settings. Journal of Theoretical Probability, 19(2):289–336, June 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-006-0016-3.

Balan:2010:SHE

[BT10] Raluca M. Balan and Ciprian A. Tudor. Stochastic heat equation with multiplicative fractional-colored noise. Journal of Theoretical Probability, 23(3):834–870, September 2010. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
009-0237-3.

Bourguin:2012:ATF

[BT12] Solesne Bourguin and Ciprian A. Tudor. Asymptotic theory for fractional regression models via Malliavin calculus. Journal of Theoretical Probability, 25(2):536–564, June 2012. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0302-y.

Bai:2016:UHP

[BT16] Shuyang Bai and Murad S. Taqqu. The universality of homogeneous polynomial forms and critical limits. Journal of Theoretical Probability, 29(4):1710–1727, December 2016. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
015-0613-0.

Bradley:2017:CLT

[BT17] Richard C. Bradley and Cristina Tone. A central limit theorem for non-stationary strongly mixing random fields. Journal of Theoretical Probability, 30(2):655–674, June 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Burdzy:2018:RRH

[BT18] Krzysztof Burdzy and Tvrtko Tadić. Random reflections in a high-dimensional tube. Journal of Theoretical Probability, 31 (1):466–493, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Bai:2024:JSM

[BT24] Shuyang Bai and He Tang. Joint sum-and-max limit for a class of long–range dependent processes with heavy tails. Journal of Theoretical Probability, 37(3):1958–1987, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01289-y.

Becker:2001:TRS

[BU01] Niels G. Becker and Sergey Utev. Threshold results for Ustatistics of dependent binary variables. Journal of Theoretical Probability, 14(1):97–114, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1007821131604.

Buchta:1990:DIP

[Buc90] Christian Buchta. Distribution-independent properties of the convex hull of random points. Journal of Theoretical Probability, 3(3):387–393, July 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01061259.

Bucher:2015:NWC

[Büc15] Axel Bücher. A note on weak convergence of the sequential multivariate empirical process under strong mixing. Journal of Theoretical Probability, 28(3):1028–1037, September 2015. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
013-0529-5.

Buck:2021:LTR

[Buc21] Micha Buck. Limit theorems for random walks with absorption. Journal of Theoretical Probability, 34(1):241–263, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-019-00970-5.

Budzban:1994:NSC

[Bud94] Greg Budzban. Necessary and sufficient conditions for the convergence of convolution products of non-identical distributions on finite abelian semigroups. Journal of Theoretical Probability, 7(3):635–646, July 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213573.

Bulinskaya:2014:LPN

[Bul14] Ekaterina Vladimirovna Bulinskaya. Local particles numbers in critical branching random walk. Journal of Theoretical Probability, 27(3):878–898, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-012-0441-4.

Bulinskaya:2021:MCB

[Bul21] Ekaterina Vl. Bulinskaya. Maximum of catalytic branching random walk with regularly varying tails. Journal of Theoretical Probability, 34(1):141–161, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-020-01009-w.

Burgstaller:2003:DMD

[Bur03] B. Burgstaller. Doob–Meyer-decomposition of Hilbert space valued functions. Journal of Theoretical Probability, 16(4):877–898, October 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOTP.0000011998.44302.3b.

Burrell:2022:DFB

[Bur22] Stuart A. Burrell. Dimensions of fractional Brownian images. Journal of Theoretical Probability, 35(4):2217–2238, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01131-3.

Butler:2022:WCT

[But22] Svetlana V. Butler. Weak convergence of topological measures. Journal of Theoretical Probability, 35(3):1614–1639, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01095-4.

Borovskikh:2002:BNG

[BW02] Yuri V. Borovskikh and Neville C. Weber. Bounds for non– Gaussian approximations of U-statistics. Journal of Theoretical Probability, 15(2):343–382, April 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1015062810192.

Bobecka:2007:DDP

[BW07] Konstancja Bobecka and Jacek Wesolowski. The Dirichlet distribution and process through neutralities. Journal of Theoretical Probability, 20(2):295–308, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0069-y.

Bakken:2019:CTA

[BW19] Erik Bakken and David Weisbart. Continuous time p-adic random walks and their path integrals. Journal of Theoretical Probability, 32(2):781–805, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Bi:2016:FOL

[BX16] Junna Bi and Fangjun Xu. A first-order limit law for functionals of two independent fractional Brownian motions in the critical case. Journal of Theoretical Probability, 29(3):941–957, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-015-0604-1.

Berger:2018:BFA

[BX18] Arno Berger and Chuang Xu. Best finite approximations of Benford’s Law. Journal of Theoretical Probability, ??(??):??, April 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Berger:2019:BFA

[BX19] Arno Berger and Chuang Xu. Best finite approximations of Benford’s Law. Journal of Theoretical Probability, 32(3):1525–1553, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Berger:2020:AOD

[BX20] Arno Berger and Chuang Xu. Asymptotics of one-dimensional Lévy approximations. Journal of Theoretical Probability, 33(2): 1164–1195, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Bertoin:1996:SIR

[BY96] Jean Bertoin and Marc Yor. Some independence results related to the arc-sine law. Journal of Theoretical Probability, 9 (2):447–458, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214659.

Bao:2016:BSR

[BY16] Jianhai Bao and Chenggui Yuan. Blow-up for stochastic reaction–diffusion equations with jumps. Journal of Theoretical Probability, 29(2):617–631, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-014-0589-1.

Bogdan:2006:KT

[BZ06] K. Bogdan and T. Zak. On Kelvin transformation. Journal of Theoretical Probability, 19(1):89–120, January 2006. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
006-0003-8.

Bounebache:2014:SSH

[BZ14] Said Karim Bounebache and Lorenzo Zambotti. A skew stochastic heat equation. Journal of Theoretical Probability, 27(1): 168–201, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0421-8.

Banek:2019:RWB

[BZ19] Tadeusz Banek and August M. Zapała. Random walk in balls and an extension of the Banach integral in abstract spaces. Journal of Theoretical Probability, 32(2):586–607, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-019-00890-4.pdf.

Cuesta-Albertos:2007:SFC

[CAFR07] Juan Antonio Cuesta-Albertos, Ricardo Fraiman, and Thomas Ransford. A sharp form of the Cramér–Wold theorem. Journal of Theoretical Probability, 20(2):201–209, June 2007. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0060-7.

Cai:1990:MSM

[Cai90] Haiyan Cai. The minimal solution of a Markov renewal equation. Journal of Theoretical Probability, 3(4):579–585, October 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01046097.

Cuesta-Albertos:1996:LBW

[CAMBTD96] J. A. Cuesta-Albertos, C. Matrán-Bea, and A. Tuero-Diaz. On lower bounds for the L^{2}–Wasserstein metric in a Hilbert space. Journal of Theoretical Probability, 9(2):263–283, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02214649.

Cuesta-Albertos:2003:APT

[CAMRR03] J. A. Cuesta-Albertos, C. Matrán, and J. Rodríguez-Rodríguez. Approximation to probabilities through uniform laws on convex sets. Journal of Theoretical Probability, 16(2):363–376, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023518526754.

Can:2019:EET

[Can19] Van Hao Can. Exponential extinction time of the contact process on rank-one inhomogeneous random graphs. Journal of Theoretical Probability, 32(1):106–130, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Capitaine:2013:AMF

[Cap13] M. Capitaine. Additive/ multiplicative free subordination property and limiting eigenvectors of spiked additive deformations of Wigner matrices and spiked sample covariance matrices. Journal of Theoretical Probability, 26(3):595–648, September 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-012-0416-5.

Cartwright:1988:RWD

[Car88] Donald I. Cartwright. Random walks on direct sums of discrete groups. Journal of Theoretical Probability, 1(4):341–356, October 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01048724.

Carlsson:2002:CCI

[Car02] Niclas Carlsson. A contractivity condition for iterated function systems. Journal of Theoretical Probability, 15(3):613–630, July 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1016215831096.

Carcamo:2017:MPM

[Cár17] Javier Cárcamo. Maps preserving moment sequences. Journal of Theoretical Probability, 30(1):212–232, March 2017. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
015-0647-3.

Casukhela:1997:SDR

[Cas97] Kameswarrao S. Casukhela. Symmetric distributions of random measures in higher dimensions. Journal of Theoretical Probability, 10(3):759–771, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022614013902.

Cassese:2008:FAS

[Cas08] Gianluca Cassese. Finitely additive supermartingales. Journal of Theoretical Probability, 21(3):586–603, September 2008.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-008-0164-8.

Chatterjee:2004:NMB

[CB04] S. Chatterjee and A. Bose. A new method for bounding rates of convergence of empirical spectral distributions. Journal of Theoretical Probability, 17(4):1003–1019, October 2004. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
004-0587-9.

Csaki:1995:AFM

[CC95] Endre Csáki and Miklós Csörgő. On additive functionals of Markov chains. Journal of Theoretical Probability, 8(4):905–919, October 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02410117.

Capitaine:2007:CRM

[CC07] M. Capitaine and M. Casalis. Cumulants for random matrices as convolutions on the symmetric group, II. Journal of Theoretical Probability, 20(3):505–533, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-007-0077-y.

Chen:2011:IPT

[CC11] Yong Chen and Jianmin Chen. On the imbedding problem for three-state time homogeneous Markov chains with coinciding negative eigenvalues. Journal of Theoretical Probability, 24(4): 928–938, December 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0316-5.

Cohen:2017:CRW

[CC17] Guy Cohen and Jean-Pierre Conze. CLT for random walks of commuting endomorphisms on compact Abelian groups. Journal of Theoretical Probability, 30(1):143–195, March 2017. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
015-0631-y.

Casukhela:2006:NAP

[CCET06] Jaya Casukhela, Kameswarrao S. Casukhela, and Hussain Elalaoui-Talibi. A new approach to Poisson approximation of simple point processes using compensators. Journal of Theoretical Probability, 19(4):899–909, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
006-0037-y.

Csaki:2002:PUA

[CCF^{+}02] Endre Csáki, Miklós Csörgő, Antónia Földes, Zhan Shi, and Ricardas Zitikis. Pointwise and uniform asymptotics of the Vervaat error process. Journal of Theoretical Probability, 15(4):845–875, October 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020650502619.

Csaki:1992:SAA

[CCFR92] E. Csáki, M. Csörgő, A. Földes, and P. Révész. Strong approximation of additive functionals. Journal of Theoretical Probability, 5(4):679–706, October 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01058725.

Csaki:1996:LTI

[CCFR96] E. Csáki, M. Csörgó, A. Földes, and P. Révész. The local time of iterated Brownian motion. Journal of Theoretical Probability, 9(3):717–743, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214084.

Csaki:2016:SLT

[CCFR16] Endre Csáki, Miklós Csörgő, Antónia Földes, and Pál Révész. Some limit theorems for heights of random walks on a spider. Journal of Theoretical Probability, 29(4):1685–1709, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-015-0626-8.

Csaki:2019:LTL

[CCFR19] Endre Csáki, Miklós Csörgő, Antónia Földes, and Pál Révész. Limit theorems for local and occupation times of random walks and Brownian motion on a spider. Journal of Theoretical Probability, 32(1):330–352, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Csorgo:1991:DPA

[CD91] Sándor Csörgő and Rossitza Dodunekova. The domain of partial attraction of a Poisson law. Journal of Theoretical Probability, 4 (1):169–190, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01047000.

Chakraborty:2004:IML

[CD04] P. S. Chakraborty and A. Dasgupta. Invariant measure and a limit theorem for some generalized Gauss maps. Journal of Theoretical Probability, 17(2):387–401, April 2004. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000020700.45630.5c.

Chaumont:2020:DDT

[CD20] Loïc Chaumont and Ron Doney. On distributions determined by their upward, space–time Wiener–Hopf factor. Journal of Theoretical Probability, 33(2):1011–1033, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Castaing:2000:USI

[CdF00] Charles Castaing and Paul Raynaud de Fitte. S-uniform scalar integrability and strong laws of large numbers for Pettis integrable functions with values in a separable locally convex space. Journal of Theoretical Probability, 13(1):93–134, January 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007782825974.

Cipriani:2020:SLD

[CdGR20] Alessandra Cipriani, Jan de Graaff, and Wioletta M. Ruszel. Scaling limits in divisible sandpiles: A Fourier multiplier approach. Journal of Theoretical Probability, 33(4):2061–2088, December 2020. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-019-00952-7.

Cheng:1997:RCI

[CdHH97] Shihong Cheng, Laurens de Haan, and Xin Huang. Rate of convergence of intermediate order statistics. Journal of Theoretical Probability, 10(1):1–23, January 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022613513012.

Capitaine:2005:FWP

[CDM05] M. Capitaine and C. Donati-Martin. Free Wishart processes. Journal of Theoretical Probability, 18(2):413–438, April 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-005-3511-z.

Cuny:2024:SAC

[CDM24] Christophe Cuny, Jérôme Dedecker, and Florence Merlevède. Strong approximations for a class of dependent random variables with semi-exponential tails. Journal of Theoretical Probability, 37(3):2234–2252, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-023-01306-0.

Campbell:2022:LM

[CDQ22] James T. Campbell, A. Deane, and A. Quas. The lightning model. Journal of Theoretical Probability, 35(4):2738–2756, December 2022. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-021-01135-z.

Conforti:2017:RCJ

[CDR17] Giovanni Conforti, Paolo Dai Pra, and Sylvie Roelly. Reciprocal class of jump processes. Journal of Theoretical Probability, 30 (2):551–580, June 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Chen:2024:IMN

[CE24] Le Chen and Nicholas Eisenberg. Invariant measures for the nonlinear stochastic heat equation with no drift term. Journal of Theoretical Probability, 37(2):1357–1396, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01302-4.

Cerf:2000:LDF

[Cer00] Raphaël Cerf. Large deviations of the finite cluster shape for two-dimensional percolation in the Hausdorff and L_{1} metric. Journal of Theoretical Probability, 13(2):491–517, April 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007841407417.

Carnal:1995:PRP

[CF95] H. Carnal and G. M. Fel’dman. Phase retrieval for probability measures on abelian groups. Journal of Theoretical Probability, 8(3):717–725, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02218053.

Carnal:1997:PRP

[CF97] H. Carnal and G. M. Fel’dman. Phase retrieval for probability measures on Abelian groups II. Journal of Theoretical Probability, 10(4):1065–1074, October 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022674919334.

Csaki:2000:AIA

[CF00] Endre Csáki and Antónia Földes. Asymptotic independence and additive functionals. Journal of Theoretical Probability, 13(4): 1123–1144, October 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007826310706.

Csaki:2020:RWC

[CF20] Endre Csáki and Antónia Földes. Random walks on combtype subsets of Z^{2}. Journal of Theoretical Probability, 33(4): 2233–2257, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-019-00938-5.

Chen:2021:RBS

[CF21] Zhen-Qing Chen and Xinwei Feng. Reflected backward stochastic differential equation with rank-based data. Journal of Theoretical Probability, 34(3):1213–1247, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
01026-9.

Csaki:2022:LTH

[CF22a] Endre Csáki and Antónia Földes. On the local time of the halfplane half-comb walk. Journal of Theoretical Probability, 35 (2):1247–1261, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01065-2.

Csaki:2022:SAA

[CF22b] Endre Csáki and Antónia Földes. Strong approximation of the anisotropic random walk revisited. Journal of Theoretical Probability, 35(4):2879–2895, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-021-01142-0.

See correction [CF23].

Csaki:2023:CSA

[CF23] Endre Csáki and Antónia Földes. Correction to: Strong approximation of the anisotropic random walk revisited. Journal of Theoretical Probability, 36(2):1339, June 2023. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-022-

01213-w. See [CF22b].

Csaki:2024:LTA

[CF24] Endre Csáki and Antónia Földes. On the local time of anisotropic random walk on Z^{2}. Journal of Theoretical Probability, 37(2): 1654–1673, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01297-y.

Caballero:1998:EDA

[CFN98] María Emilia Caballero, Begoña Fernández, and David Nualart. Estimation of densities and applications. Journal of Theoretical Probability, 11(3):831–851, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022614917458.

Csaki:2005:MLT

[CFR05] Endre Csáki, Antönia Földes, and Pál Révész. Maximal local time of a d-dimensional simple random walk on subsets. Journal of Theoretical Probability, 18(3):687–717, July 2005. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
005-7256-5.

Csaki:2007:BRW

[CFR07] Endre Csáki, Antónia Földes, and Pál Révész. On the behavior of random walk around heavy points. Journal of Theoretical Probability, 20(4):1041–1057, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-007-0092-z.

Csaki:2009:TNNb

[CFR09a] Endre Csáki, Antónia Földes, and Pál Révész. Transient nearest neighbor random walk and Bessel process. Journal of Theoretical Probability, 22(4):992–1009, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-008-0165-7.

Csaki:2009:TNNa

[CFR09b] Endre Csáki, Antónia Földes, and Pál Révész. Transient nearest neighbor random walk on the line. Journal of Theoretical Probability, 22(1):100–122, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0137-3.

Csaki:2010:NCT

[CFR10] Endre Csáki, Antónia Földes, and Pál Révész. On the number of cutpoints of the transient nearest neighbor random walk on the line. Journal of Theoretical Probability, 23(2):624–638, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0204-4.

Chen:2023:WTD

[CFZ23] Shukai Chen, Rongjuan Fang, and Xiangqi Zheng. Wassersteintype distances of two-type continuous-state branching processes in Lévy random environments. Journal of Theoretical Probability, 36(3):1572–1590, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-022-01211-y.

Chandra:1996:SLL

[CG96] Tapas K. Chandra and Subhashis Ghosal. The strong law of large numbers for weighted averages under dependence assumptions. Journal of Theoretical Probability, 9(3):797–809, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF02214087.

Chandra:2003:CIL

[CG03] T. K. Chandra and A. Goswami. Cesàro α-integrability and laws of large numbers I. Journal of Theoretical Probability, 16 (3):655–669, July 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1025620532404.

Chandra:2006:CIL

[CG06] T. K. Chandra and A. Goswami. Cesàro α-integrability and laws of large numbers — II. Journal of Theoretical Probability, 19(4): 789–816, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-006-0038-x.

Cox:2021:DBS

[CG21] Sonja Cox and Stefan Geiss. On decoupling in Banach spaces. Journal of Theoretical Probability, 34(3):1179–1212, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01085-6.

Carmona:2006:SDC

[CGHM06] Philippe Carmona, Francesco Guerra, Yueyun Hu, and Olivier Mejane. Strong disorder for a certain class of directed polymers in a random environment. Journal of Theoretical Probability, 19 (1):134–151, January 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-006-0010-9.

Chang:1989:RIF

[CGR89] Ted Chang, Fred Galvin, and Andrew L. Rukhin. On the ratio integral formula for the likelihood ratio of a maximal invariant. Journal of Theoretical Probability, 2(2):261–265, April 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01053412.

Crismale:2020:WMF

[CGW20] Vitonofrio Crismale, Maria Elena Griseta, and Janusz Wysoczański. Weakly monotone Fock space and monotone convolution of the Wigner law. Journal of Theoretical Probability, 33(1):268–294, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Couvreux:1999:LTM

[CH99] Jérôme Couvreux and Christian Hess. A Lévy type martingale convergence theorem for random sets with unbounded values. Journal of Theoretical Probability, 12(4):933–969, October 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021688919194.

Csaki:2001:APR

[CH01] Endre Csáki and Yueyun Hu. Asymptotic properties of ranked heights in Brownian excursions. Journal of Theoretical Probability, 14(1):77–96, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1007868914766.

Csaki:2004:REH

[CH04] Endre Csáki and Yueyun Hu. On the ranked excursion heights of a Kiefer process. Journal of Theoretical Probability, 17(1): 145–163, January 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000020479.46788.c9.

Cheng:2005:ANN

[CH05] Tsung-Lin Cheng and Hwai-Chung Ho. Asymptotic normality for non-linear functionals of non-causal linear processes with summable weights. Journal of Theoretical Probability, 18(2): 345–358, April 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-005-3506-9.

Choi:2020:HTM

[CH20] Michael C. H. Choi and Lu-Jing Huang. On hitting time, mixing time and geometric interpretations of Metropolis–Hastings reversiblizations. Journal of Theoretical Probability, 33(2):1144– 1163, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Cao:2021:EDB

[CH21] Guilan Cao and Kai He. Estimates of the difference between two probability densities of Wiener functionals and its application. Journal of Theoretical Probability, 34(2):553–579, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-00986-2.

Chari:1990:GJT

[Cha90] Ravi Chari. Generalized jump-type Ornstein–Uhlenbeck processes as limits of infinite-particle systems. Journal of Theoretical Probability, 3(1):9–29, January 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01063326.

Chan:1995:SAR

[Cha95] Terence Chan. Some asymptotic results related to the law of iterated logarithm for Brownian motion. Journal of Theoretical Probability, 8(3):643–667, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02218048.

Chaumont:2000:EVT

[Cha00] L. Chaumont. An extension of Vervaat’s transformation and its consequences. Journal of Theoretical Probability, 13(1):259–277, January 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007795228700.

Chafai:2010:CLN

[Cha10] Djalil Chafaï. Circular law for noncentral random matrices. Journal of Theoretical Probability, 23(4):945–950, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-010-0285-8.

Chakrabarty:2022:LDP

[CHdHS22] Arijit Chakrabarty, Rajat Subhra Hazra, Frank den Hollander, and Matteo Sfragara. Large deviation principle for the maximal eigenvalue of inhomogeneous Erdős–Rényi random graphs. Journal of Theoretical Probability, 35(4):2413–2441, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01138-w.

Chen:1997:RLL

[Che97a] Jeesen Chen. Revisit the Lovász local lemma. Journal of Theoretical Probability, 10(3):747–758, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022661929831.

Chen:1997:LIL

[Che97b] Xia Chen. The law of the iterated logarithm for m-dependent Banach space valued random variables. Journal of Theoretical Probability, 10(3):695–732, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022605812085.

Chebotarev:1999:RCT

[Che99a] Vladimir I. Chebotarev. Remark on the Cramér type condition for ω^{2}-statistic. Journal of Theoretical Probability, 12(1):13–25, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021736324955.

Chen:1999:LIL

[Che99b] Xia Chen. The law of the iterated logarithm for functionals of Harris recurrent Markov chains: Self normalization. Journal of Theoretical Probability, 12(2):421–445, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021630228280.

Chen:2006:MSD

[Che06] Xia Chen. Moderate and small deviations for the ranges of one-dimensional random walks. Journal of Theoretical Probability, 19(3):721–739, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-006-0032-3.

Chen:2013:POU

[Che13] Jiahua Chen. A partial order on uncertainty and information. Journal of Theoretical Probability, 26(2):349–359, June 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-011-0375-2.

Chen:2015:LLI

[Che15a] Xia Chen. The limit law of the iterated logarithm. Journal of Theoretical Probability, 28(2):721–725, June 2015. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
013-0481-4.

Cheng:2015:RPB

[Che15b] Li-Juan Cheng. The radial part of Brownian motion with respect to L-distance under Ricci flow. Journal of Theoretical Probability, 28(2):449–466, June 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-013-0512-1.

Chen:2017:AP

[Che17] Xia Chen. Acknowledgment of priority. Journal of Theoretical Probability, 30(2):700, June 2017. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/content/pdf/10.1007/s10959-
015-0649-1.pdf.

Chen:2020:RDT

[Che20] Anyue Chen. Resolvent decomposition theorems and their application in denumerable Markov processes with instantaneous states. Journal of Theoretical Probability, 33(4):2089–2118, December 2020. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-019-00941-w.

Chen:2021:SPG

[Che21] Linan Chen. Steep points of Gaussian free fields in any dimension. Journal of Theoretical Probability, 34(4):1959–2004, December 2021. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-020-01028-7.

Chiu:2003:UCI

[Chi03] Wang Chiu. On upper class of increments of fractional Brownian motion. Journal of Theoretical Probability, 16(1):117–145, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022230521472.

Chi:2018:MSR

[Chi18] Zhiyi Chi. On a multivariate strong renewal theorem. Journal of Theoretical Probability, 31(3):1235–1272, September 2018. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Chi:2020:LFP

[Chi20] Zhiyi Chi. Law of the first passage triple of a spectrally positive strictly stable process. Journal of Theoretical Probability, 33(2): 715–748, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Chikvinidze:2022:NSC

[Chi22] B. Chikvinidze. Necessary and sufficient conditions for the uniform integrability of the stochastic exponential. Journal of Theoretical Probability, 35(1):282–294, March 2022. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
01047-4.

Cheng:2008:NAN

[CHL08] Tsung-Lin Cheng, Hwai-Chung Ho, and Xuewen Lu. A note on asymptotic normality of kernel estimation for linear random fields on Z^{2}. Journal of Theoretical Probability, 21(2):267–286, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0146-x.

Csorgo:2013:STL

[CHM13] Miklós Csörgő, Zhishui Hu, and Hongwei Mei. Strassen-type law of the iterated logarithm for self-normalized sums. Journal of Theoretical Probability, 26(2):311–328, June 2013. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
011-0353-8.

Cook:2022:NHR

[CHNR22] Nicholas Cook, Walid Hachem, Jamal Najim, and David Renfrew. Non-Hermitian random matrices with a variance profile (II): Properties and examples. Journal of Theoretical Probability, 35(4):2343–2382, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01140-2.

Chong:2017:LDV

[Cho17] Carsten Chong. Lévy-driven Volterra equations in space and time. Journal of Theoretical Probability, 30(3):1014–1058, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Chung:1993:DBL

[Chu93] Kai Lai Chung. Doeblin’s big limit theorem. Journal of Theoretical Probability, 6(3):417–426, July 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01066709.

Chen:2024:VSE

[CJSS24] Peng Chen, Xinghu Jin, Tian Shen, and Zhonggen Su. Variablestep Euler–Maruyama approximations of regime-switching jump diffusion processes. Journal of Theoretical Probability, 37(2): 1597–1626, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01253-w.

Choi:1999:HBI

[CK99] Yong-Kab Choi and Norio Kôno. How big are the increments of a two-parameter Gaussian process? Journal of Theoretical Probability, 12(1):105–129, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1021796610843.

Cho:2000:LSS

[CK00] Nhansook Cho and Youngmee Kwon. Limit of solutions of a SDE with a large drift driven by a Poisson random measure. Journal of Theoretical Probability, 13(2):311–325, April 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007853810143.

Csorgo:2008:WCV

[CK08] Miklós Csörgő and Rafal Kulik. Weak convergence of Vervaat and Vervaat error processes of long-range dependent sequences. Journal of Theoretical Probability, 21(3):672–686, September 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-007-0124-8.

Curien:2014:IBM

[CK14] Nicolas Curien and Takis Konstantopoulos. Iterating Brownian motions, ad libitum. Journal of Theoretical Probability, 27 (2):433–448, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0434-3.

Cheliotis:2022:FLT

[CK22] Dimitris Cheliotis and Dimitra Kouloumpou. Functional limit theorems for the Pólya urn. Journal of Theoretical Probability, 35(3):2038–2051, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01123-3.

Conforti:2019:CPP

[CKR19] Giovanni Conforti, Tetiana Kosenkova, and Sylvie Roelly. Conditioned point processes with application to Lévy bridges. Journal of Theoretical Probability, 32(4):2111–2134, December 2019. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Cuchiero:2016:APS

[CKRMT16] Christa Cuchiero, Martin Keller-Ressel, Eberhard Mayerhofer, and Josef Teichmann. Affine processes on symmetric cones. Journal of Theoretical Probability, 29(2):359–422, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-014-0580-x.

Chamayou:1991:ESD

[CL91] Jean-François Chamayou and Gérard Letac. Explicit stationary distributions for compositions of random functions and products of random matrices. Journal of Theoretical Probability, 4 (1):3–36, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046992.

Chamayou:1999:APD

[CL99] Jean-François Chamayou and Gérard Letac. Additive properties of the Dufresne laws and their multivariate extension. Journal of Theoretical Probability, 12(4):1045–1066, October 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021649305082.

Curien:2014:BP

[CL14] Nicolas Curien and Jean-François Le Gall. The Brownian plane. Journal of Theoretical Probability, 27(4):1249–1291, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-013-0485-0.

Crane:2016:EMP

[CL16] Harry Crane and Steven P. Lalley. Exchangeable Markov processes on [k]^{N} with cadlag sample paths. Journal of Theoretical Probability, 29(1):206–230, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-014-0566-8.

Cai:2018:OTI

[CL18] Chunhao Cai and Bo Li. Occupation times of intervals until last passage times for spectrally negative Lévy processes. Journal of Theoretical Probability, 31(4):2194–2215, December 2018. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Chen:2023:EUS

[CL23] Yuyang Chen and Peng Luo. Existence and uniqueness of solutions for multi-dimensional reflected backward stochastic differential equations with diagonally quadratic generators. Journal of Theoretical Probability, 36(3):1698–1719, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01224-7.

Clancy:2021:GSI

[Cla21] David Clancy, Jr. The Gorin–Shkolnikov identity and its random tree generalization. Journal of Theoretical Probability, 34(4): 2386–2420, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01128-y.

Cenac:2018:PRW

[CLdLO18] Peggy Cénac, Arnaud Le Ny, Basile de Loynes, and Yoann Offret. Persistent random walks. I. Recurrence versus transience. Journal of Theoretical Probability, 31(1):232–243, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Cenac:2019:PRW

[CLdLO19] Peggy Cénac, Arnaud Le Ny, Basile de Loynes, and Yoann Offret. Persistent random walks. II. Functional scaling limits. Journal of Theoretical Probability, 32(2):633–658, June 2019. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Casalis:1993:DED

[CLM93] M. Casalis, G. Letac, and H. Massam. A decomposition for the exponential dispersion model generated by the invariant measure on the hyperboloid. Journal of Theoretical Probability, 6 (4):803–816, October 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01049176.

Chobanyan:2004:PBS

[CLM04] S. Chobanyan, S. Levental, and V. Mandrekar. Prokhorov blocks and strong law of large numbers under rearrangements. Journal of Theoretical Probability, 17(3):647–672, July 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000040292.52298.fd.

Chang:2013:IBV

[CLY13] Tongkeun Chang, Kijung Lee, and Minsuk Yang. On initialboundary value problem of the stochastic heat equation in Lipschitz cylinders. Journal of Theoretical Probability, 26(4):1135– 1164, December 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0444-1.

Chobanyan:2000:KSU

[CM00] Sergei Chobanyan and V. Mandrekar. On Kolmogorov SLLN under rearrangements for “orthogonal” random variables in a B-space. Journal of Theoretical Probability, 13(1):135–139, January 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007734910044.

Cureg:2005:WCC

[CM05] E. Cureg and A. Mukherjea. Weak convergence in circulant matrices. Journal of Theoretical Probability, 18(4):983–1007, October 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-005-7542-2.

Chikvinidze:2014:NPS

[CM14] B. Chikvinidze and M. Mania. New proofs of some results on bounded mean oscillation martingales using backward stochastic differential equations. Journal of Theoretical Probability, 27(4): 1213–1228, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0524-x.

Cuny:2015:SIP

[CM15] Christophe Cuny and Florence Merlevède. Strong invariance principles with rate for “reverse” martingale differences and applications. Journal of Theoretical Probability, 28(1):137–183, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-013-0506-z.

Cammarota:2022:CCP

[CM22] Valentina Cammarota and Domenico Marinucci. On the correlation of critical points and angular trispectrum for random spherical harmonics. Journal of Theoretical Probability, 35(4): 2269–2303, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01136-y.

Cheng:2000:RMI

[CMP00] R. Cheng, A. G. Miamee, and M. Pourahmadi. Regularity and minimality of infinite variance processes. Journal of Theoretical Probability, 13(4):1115–1122, October 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1007874226636.

Chauvin:2015:SEL

[CMP15] Brigitte Chauvin, Cécile Mailler, and Nicolas Pouyanne. Smoothing equations for large Pólya urns. Journal of Theoretical Probability, 28(3):923–957, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-013-0530-z.

Coffie:2024:AAS

[CMP24] Emmanuel Coffie, Xuerong Mao, and Frank Proske. On the analysis of Ait–Sahalia-type model for rough volatility modelling. Journal of Theoretical Probability, 37(1):744–767, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01269-2.

Coquet:2003:NCD

[CMS03] François Coquet, Jean Mémin, and Leszek Slomiński. On noncontinuous Dirichlet processes. Journal of Theoretical Probability, 16(1):197–216, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022238723289.

Cohn:1993:WEP

[CNP93] Harry Cohn, Olle Nerman, and Magda Peligrad. Weak ergodicity and products of random matrices. Journal of Theoretical Probability, 6(2):389–405, April 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01047581.

Corcuera:2009:CCF

[CNW09] José Manuel Corcuera, David Nualart, and Jeannette H. C. Woerner. Convergence of certain functionals of integral fractional processes. Journal of Theoretical Probability, 22(4):856–870, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0158-6.

Chen:2021:SMA

[CNX21] Peng Chen, Ivan Nourdin, and Lihu Xu. Stein’s method for asymmetric α -stable distributions, with application to the stable CLT. Journal of Theoretical Probability, 34(3):1382–1407, September 2021. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-020-01004-1.

Chen:2024:MSA

[CNXY24] Peng Chen, Ivan Nourdin, Lihu Xu, and Xiaochuan Yang. Multivariate stable approximation by Stein’s method. Journal of Theoretical Probability, 37(1):446–488, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01244-x.

Chen:2022:NIS

[CNZ22] Peng Chen, Ivan Nourdin, and Rui Zhang. Non-integrable stable approximation by Stein’s method. Journal of Theoretical Probability, 35(2):1137–1186, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01094-5.

Coston:2020:GFL

[CO20] Natalie Coston and Sean O’Rourke. Gaussian fluctuations for linear eigenvalue statistics of products of independent iid random matrices. Journal of Theoretical Probability, 33(3):1541–1612, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Cinque:2023:SDG

[CO23] Fabrizio Cinque and Enzo Orsingher. Stochastic dynamics of generalized planar random motions with orthogonal directions. Journal of Theoretical Probability, 36(4):2229–2261, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01229-2.

Campbell:2024:SLK

[CO24] Andrew Campbell and Sean O’Rourke. Spectrum of Lévy– Khintchine random Laplacian matrices. Journal of Theoretical Probability, 37(1):933–973, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-023-01275-4.

Coja-Oghlan:2005:LTN

[COM05a] A. Coja-Oghlan and J. Michałiček. Limit theorems for noncommutative random variables. Journal of Theoretical Probability, 18(3):595–614, July 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-005-7251-x.

Comman:2005:FAL

[Com05b] Henri Comman. Functional approach of large deviations in general spaces. Journal of Theoretical Probability, 18(1):187–207, January 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-004-2594-2.

Copros:2018:ECS

[Cop18] Guillaume Copros. Existence condition of strong stationary times for continuous time Markov chains on discrete graphs. Journal of Theoretical Probability, 31(3):1679–1728, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Cordero:2016:FPT

[Cor16] Fernando Cordero. The first passage time of a stable process conditioned to not overshoot. Journal of Theoretical Probability, 29(3):776–796, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-014-0592-6.

Chen:2017:FPP

[CORS17] Changhao Chen, Tuomo Ojala, Eino Rossi, and Ville Suomala. Fractal percolation, porosity, and dimension. Journal of Theoretical Probability, 30(4):1471–1498, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Covo:2011:TPL

[Cov11] Shai Covo. Two-parameter Lévy processes along decreasing paths. Journal of Theoretical Probability, 24(1):150–169, March 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-010-0277-8.

Cheng:1993:MRS

[CP93] R. Cheng and M. Pourahmadi. The mixing rate of a stationary multivariate process. Journal of Theoretical Probability, 6 (3):603–617, July 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01066720.

Chatterjee:2011:CAI

[CP11] Sourav Chatterjee and Soumik Pal. A combinatorial analysis of interacting diffusions. Journal of Theoretical Probability, 24(4): 939–968, December 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-009-0269-8.pdf.

Cuny:2012:CLT

[CP12] Christophe Cuny and Magda Peligrad. Central limit theorem started at a point for stationary processes and additive functionals of reversible Markov chains. Journal of Theoretical Probability, 25(1):171–188, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-010-0321-8.

Chadha:2018:ACN

[CP18] Alka Chadha and Dwijendra N. Pandey. Approximate controllability of a neutral stochastic fractional integro–differential inclusion with nonlocal conditions. Journal of Theoretical Probability, 31(2):705–740, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Cellupica:2021:PAV

[CP21] Miriana Cellupica and Barbara Pacchiarotti. Pathwise asymptotics for Volterra type stochastic volatility models. Journal of Theoretical Probability, 34(2):682–727, June 2021. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
00992-4.

Chakraborty:2001:CPP

[CR01] Santanu Chakraborty and B. V. Rao. Convolution powers of probabilities on stochastic matrices. Journal of Theoretical Probability, 14(2):599–603, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1011180117932.

Chakrabarty:2013:GTD

[CR13] Arijit Chakrabarty and Parthanil Roy. Group-theoretic dimension of stationary symmetric α-stable random fields. Journal of Theoretical Probability, 26(1):240–258, March 2013. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
011-0371-6.

Criens:2022:TCS

[CR22] David Criens and Moritz Ritter. On a theorem by A. S. Cherny for semilinear stochastic partial differential equations. Journal of Theoretical Probability, 35(3):2052–2067, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01107-3.

Criens:2019:CMP

[Cri19] David Criens. Cylindrical martingale problems associated with Lévy generators. Journal of Theoretical Probability, 32(3):1306– 1359, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). See correction [Cri20a].

Criens:2020:CCM

[Cri20a] David Criens. Correction to: Cylindrical Martingale Problems Associated with Lévy Generators. Journal of Theoretical Probability, 33(3):1791–1800, September 2020. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-020-01012-1.pdf. See [Cri19].

Criens:2020:LTC

[Cri20b] David Criens. Limit theorems for cylindrical martingale problems associated with Lévy generators. Journal of Theoretical Probability, 33(2):866–905, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Criens:2023:ROD

[Cri23] David Criens. On the relation of one-dimensional diffusions on natural scale and their speed measures. Journal of Theoretical Probability, 36(4):2339–2358, December 2023. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01247-8.

Csaki:2001:LER

[CRS01] Endre Csáki, Pál Révész, and Zhan Shi. Long excursions of a random walk. Journal of Theoretical Probability, 14(3):821–844, July 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017553425291.

Chen:2017:LLN

[CRY17] Zhen-Qing Chen, Yan-Xia Ren, and Ting Yang. Law of large numbers for branching symmetric hunt processes with measure– valued branching rates. Journal of Theoretical Probability, 30(3): 898–931, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Cartwright:1991:MBG

[CS91] Donald I. Cartwright and Stanley Sawyer. The Martin boundary for general isotropic random walks in a tree. Journal of Theoretical Probability, 4(1):111–136, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01046997.

Ching-Sung:1995:SIL

[CS95] Chou Ching-Sung. On some inequalities of local time. Journal of Theoretical Probability, 8(1):17–22, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02213450.

Conlon:2000:RWA

[CS00] Joseph G. Conlon and Renming Song. Random walks associated with non-divergence form elliptic equations. Journal of Theoretical Probability, 13(2):427–489, April 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1007893424255.

Chen:2014:MRR

[CS14] Lung-Chi Chen and Rongfeng Sun. A monotonicity result for the range of a perturbed random walk. Journal of Theoretical Probability, 27(3):997–1010, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http:/

/link.springer.com/article/10.1007/s10959-012-0472-x.

Chen:2020:EVT

[CS20] Zaoli Chen and Gennady Samorodnitsky. Extreme value theory for long-range-dependent stable random fields. Journal of Theoretical Probability, 33(4):1894–1918, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-019-
00951-8.

Chasapis:2023:ESI

[CST23] Giorgos Chasapis, Salil Singh, and Tomasz Tkocz. Entropies of sums of independent gamma random variables. Journal of Theoretical Probability, 36(2):1227–1242, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-022-
01192-y.

Cuny:2003:WMR

[Cun03] Christophe Cuny. Weak mixing of random walks on groups. Journal of Theoretical Probability, 16(4):923–933, October 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000012000.54810.d2.

Cuzick:1995:SLW

[Cuz95] Jack Cuzick. A strong law for weighted sums of i.i.d. random variables. Journal of Theoretical Probability, 8(3):625–641, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.

1007/BF02218047. See erratum [Cuz01].

Cuzick:2001:ESL

[Cuz01] Jack Cuzick. Erratum: A Strong Law for Weighted Sums of i.i.d. Random Variables, J. Theor. Prob. 8, 625–636 (1995). Journal of Theoretical Probability, 14(2):605, April 2001. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1011132202003. See [Cuz95].

Conger:2007:NAD

[CV07] Mark Conger and D. Viswanath. Normal approximations for descents and inversions of permutations of multisets. Journal of Theoretical Probability, 20(2):309–325, June 2007. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0070-5.

Cekanavicius:2015:CPC

[CV15] V. Cekanavicius and P. Vellaisamy. A compound Poisson convergence theorem for sums of m-dependent variables. Journal of Theoretical Probability, 28(3):1145–1164, September 2015. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
014-0540-5.

Cekanavicius:2021:CPA

[CV21] V. Cekanavicius and P. Vellaisamy. Compound Poisson approximations in ℓ_{p}-norm for sums of weakly dependent vectors. Journal of Theoretical Probability, 34(4):2241–2264, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01042-9.

Cohen:2010:TWC

[CW10] Serge Cohen and Mario Wschebor. On tightness and weak convergence in the approximation of the occupation measure of fractional Brownian motion. Journal of Theoretical Probability, 23(4):1204–1226, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-009-0249-z.

Chen:2017:WPI

[CW17] Xin Chen and Jian Wang. Weighted Poincaré inequalities for non-local Dirichlet forms. Journal of Theoretical Probability, 30 (2):452–489, June 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Chen:2023:ABS

[CW23] Zhang Chen and Bixiang Wang. Asymptotic behavior of stochastic complex lattice systems driven by superlinear noise. Journal of Theoretical Probability, 36(3):1487–1519, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01206-9.

Cardon-Weber:2004:SPP

[CWM04] C. Cardon-Weber and A. Millet. On strongly Petrovskii’s parabolic SPDEs in arbitrary dimension and application to the stochastic Cahn–Hilliard equation. Journal of Theoretical Probability, 17(1):1–49, January 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/B%3AJOTP.0000020474.79479.
fa.

Chen:2015:AAB

[CX15] Xia Chen and Jie Xiong. Annealed asymptotics for Brownian motion of renormalized potential in mobile random medium. Journal of Theoretical Probability, 28(4):1601–1650, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-014-0558-8.

Comets:2011:BRW

[CY11] Francis Comets and Nobuo Yoshida. Branching random walks in space-time random environment: Survival probability, global and local growth rates. Journal of Theoretical Probability, 24(3): 657–687, September 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0267-x.

Cho:2022:OBW

[CY22] Gunhee Cho and Guang Yang. Octonionic Brownian windings. Journal of Theoretical Probability, 35(3):1956–1973, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01127-z.

Cyr:2021:SDP

[Cyr21] Justin Cyr. Stationary determinantal processes on Z^{d} with N labeled objects per site, Part I: Basic properties and full domination. Journal of Theoretical Probability, 34(3):1321– 1365, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01062-5.

Cheng:2017:RDP

[CZ17] Li-Juan Cheng and Kun Zhang. Reflecting diffusion processes on manifolds carrying geometric flow. Journal of Theoretical Probability, 30(4):1334–1368, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Chen:2018:SVR

[CZ18] Xinxin Chen and Xiaolin Zeng. Speed of vertex-reinforced jump process on Galton–Watson trees. Journal of Theoretical Probability, 31(2):1166–1211, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Chebunin:2022:FCL

[CZ22] Mikhail Chebunin and Sergei Zuyev. Functional central limit theorems for occupancies and missing mass process in infinite urn models. Journal of Theoretical Probability, 35(1):1–19, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01053-6.

Czajkowski:2024:NUP

[Cza24] Jan Czajkowski. Non-uniqueness phase of percolation on reflection groups in H^{3}. Journal of Theoretical Probability, 37(3): 2534–2575, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-024-01313-9.

deAcosta:1990:LDE

[dA90] A. de Acosta. Large deviations for empirical measures of Markov chains. Journal of Theoretical Probability, 3(3):395–431, July 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01061260.

DAristotile:1995:NNR

[D’A95] Anthony J. D’Aristotile. The nearest neighbor random walk on subspaces of a vector space and rate of convergence. Journal of Theoretical Probability, 8(2):321–346, April 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02212882.

DAristotile:2000:IPT

[D’A00] Anthony D’Aristotile. An invariance principle for triangular arrays. Journal of Theoretical Probability, 13(2):327–341, April 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1007801726073.

deAcosta:1998:MDE

[dAC98] A. de Acosta and Xia Chen. Moderate deviations for empirical measures of Markov chains: Upper bounds. Journal of Theoretical Probability, 11(4):1075–1110, October 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022673000778.

Dadoun:2020:SSG

[Dad20] Benjamin Dadoun. Self-similar growth fragmentations as scaling limits of Markov branching processes. Journal of Theoretical Probability, 33(2):590–610, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Dai:2013:CLO

[Dai13] Hongshuai Dai. Convergence in law to operator fractional Brownian motions. Journal of Theoretical Probability, 26(3):676– 696, September 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0401-4.

Dani:1991:CCT

[Dan91] S. G. Dani. A characterization of the Cauchy-type distributions on boundaries of semisimple groups. Journal of Theoretical Probability, 4(3):625–629, July 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01210327.

Dance:2014:ISI

[Dan14] Christopher R. Dance. An inequality for the sum of independent bounded random variables. Journal of Theoretical Probability, 27(2):358–369, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0460-1.

David:2005:ADS

[Dav05] M. David. The asymptotic distribution of self-normalized triangular arrays. Journal of Theoretical Probability, 18(4):853–870, October 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-005-7529-z. See erratum [Mas13].

delBarrio:2000:WBM

[dBM00] E. del Barrio and C. Matrán. The weighted bootstrap mean for heavy-tailed distributions. Journal of Theoretical Probability, 13 (2):547–569, April 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007885222438.

daCosta:2023:RDM

[dCdCV23] Conrado da Costa, Bernardo Freitas Paulo da Costa, and Daniel Valesin. Reaction–diffusion models for a class of infinitedimensional nonlinear stochastic differential equations. Journal of Theoretical Probability, 36(2):1059–1087, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-022-
01187-9.

Davis:1997:LAD

[DD97] Richard A. Davis and William T. M. Dunsmuir. Least absolute deviation estimation for regression with ARMA errors. Journal of Theoretical Probability, 10(2):481–497, April 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022620818679.

Diaconis:2001:CLM

[DD01] Persi Diaconis and Rick Durrett. Chutes and ladders in Markov chains. Journal of Theoretical Probability, 14(3):899–926, July 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017509611178.

Deleaval:2018:MHM

[DD18] Luc Deleaval and Nizar Demni. Moments of the Hermitian matrix Jacobi process. Journal of Theoretical Probability, 31(3): 1759–1778, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Doring:2013:MDC

[DE13] Hanna Döring and Peter Eichelsbacher. Moderate deviations via cumulants. Journal of Theoretical Probability, 26(2):360–385, June 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-012-0437-0.

Dierickx:2018:GDE

[DE18] Gauthier Dierickx and Uwe Einmahl. A general Darling–Erdős theorem in Euclidean space. Journal of Theoretical Probability, 31(2):1142–1165, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Debowski:2010:VLC

[Deb10] Lukasz Debowski. Variable-length coding of two-sided asymptotically mean stationary measures. Journal of Theoretical Probability, 23(1):237–256, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-009-0264-0.

Debowski:2014:HMP

[Deb14] Lukasz Debowski. On hidden Markov processes with infinite excess entropy. Journal of Theoretical Probability, 27(2):539–551, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10959-012-0468-6.pdf.

Defosseux:2013:IPM

[Def13] Manon Defosseux. An interacting particle model and a Pieritype formula for the orthogonal group. Journal of Theoretical Probability, 26(2):568–588, June 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-012-0407-6.

Deheuvels:1998:AQP

[Deh98] Paul Deheuvels. On the approximation of quantile processes by Kiefer processes. Journal of Theoretical Probability, 11(4): 997–1018, October 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022668932637.

Delaporte:2015:LPM

[Del15] Cécile Delaporte. Lévy processes with marked jumps I: Limit theorems. Journal of Theoretical Probability, 28(4):1468–1499, December 2015. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL http://link.springer.com/
article/10.1007/s10959-014-0549-9.

Dembski:1990:UP

[Dem90] William A. Dembski. Uniform probability. Journal of Theoretical Probability, 3(4):611–626, October 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01046100.

Demni:2008:FJP

[Dem08] N. Demni. Free Jacobi process. Journal of Theoretical Probability, 21(1):118–143, March 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0110-1.

Deniel:1989:CCS

[Den89] Yves Deniel. On the a.s. Cesàro-α convergence for stationary or orthogonal random variables. Journal of Theoretical Probability, 2(4):475–485, October 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01051879.

Deng:2004:WLI

[Den04] D. Deng. Wittmann’s law of iterated logarithm for tail sums of B-valued random variables. Journal of Theoretical Probability, 17(2):367–385, April 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000020699.72759.c2.

Deng:2020:SRC

[Den20] Chang-Song Deng. Subgeometric rates of convergence for discrete-time Markov chains under discrete-time subordination. Journal of Theoretical Probability, 33(1):522–532, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Dereich:2003:SBP

[Der03] Steffen Dereich. Small ball probabilities around random centers of Gaussian measures and applications to quantization. Journal of Theoretical Probability, 16(2):427–449, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023578812641.

Dettweiler:1989:MPB

[Det89] E. Dettweiler. On the martingale problem for Banach space valued stochastic differential equations. Journal of Theoretical Probability, 2(2):159–191, April 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01053408.

Deya:2020:IRH

[Dey20] Aurélien Deya. Integration with respect to the Hermitian fractional Brownian motion. Journal of Theoretical Probability, 33 (1):295–318, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Diaconis:1988:CLT

[DF88] P. Diaconis and D. A. Freedman. Conditional limit theorems for exponential families and finite versions of de Finetti’s theorem. Journal of Theoretical Probability, 1(4):381–410, October 1988.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048727.

Dawson:1995:SBM

[DF95] Donald A. Dawson and Klaus Fleischmann. Super-Brownian motions in higher dimensions with absolutely continuous measure states. Journal of Theoretical Probability, 8(1):179–206, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02213461.

Dawson:1997:CSB

[DF97] Donald A. Dawson and Klaus Fleischmann. A continuous super–Brownian motion in a super–Brownian medium. Journal of Theoretical Probability, 10(1):213–276, January 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022606801625.

Dereich:2003:LBS

[DFMS03] S. Dereich, F. Fehringer, A. Matoussi, and M. Scheutzow. On the link between small ball probabilities and the quantization problem for Gaussian measures on Banach spaces. Journal of Theoretical Probability, 16(1):249–265, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022242924198.

DaPrato:2015:SUS

[DFPR15] G. Da Prato, F. Flandoli, E. Priola, and M. Röckner. Strong uniqueness for stochastic evolution equations with unbounded measurable drift term. Journal of Theoretical Probability, 28(4): 1571–1600, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0545-0.

Dette:1997:WSD

[DFPS97] Holger Dette, James Allen Fill, Jim Pitman, and William J. Studden. Wall and Siegmund duality relations for birth and death chains with reflecting barrier. Journal of Theoretical Probability, 10(2):349–374, April 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022660400024.

Dorroh:1999:TSC

[DFS99] J. R. Dorroh, G. Ferreyra, and P. Sundar. A technique for stochastic control problems with unbounded control set. Journal of Theoretical Probability, 12(1):255–270, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021761030407.

Deshouillers:2001:BCF

[DFY01] Jean-Marc Deshouillers, Gregory A. Freiman, and Alexander A. Yudin. On bounds for the concentration function II. Journal of Theoretical Probability, 14(3):813–820, July 2001. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1017501308452.

Dombry:2019:FRC

[DFZ19] Clément Dombry, Michael Falk, and Maximilian Zott. On functional records and champions. Journal of Theoretical Probability, 32(3):1252–1277, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Debs:2008:PBD

[DG08] Pierre Debs and Mihai Gradinaru. Penalization for birth and death processes. Journal of Theoretical Probability, 21(3):745– 771, September 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0123-9.

Dawson:2018:TRR

[DG18] D. A. Dawson and L. G. Gorostiza. Transience and recurrence of random walks on percolation clusters in an ultrametric space. Journal of Theoretical Probability, 31(1):494–526, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Durrett:1999:EBR

[DGG99] Richard Durrett, Boris L. Granovsky, and Shay Gueron. The equilibrium behavior of reversible coagulation–fragmentation processes. Journal of Theoretical Probability, 12(2):447–474, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021682212351.

Doring:2021:EUR

[DGR21] Leif Döring, Lukas Gonon, and Oleg Reichmann. Existence and uniqueness results for time-inhomogeneous time-change equations and Fokker–Planck equations. Journal of Theoretical Probability, 34(1):173–205, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-019-00969-y.

Dawson:2001:OTF

[DGW01] D. A. Dawson, L. G. Gorostiza, and A. Wakolbinger. Occupation time fluctuations in branching systems. Journal of Theoretical Probability, 14(3):729–796, July 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017597107544.

Dudley:1991:UUG

[DGZ91] R. M. Dudley, E. Giné, and J. Zinn. Uniform and universal Glivenko–Cantelli classes. Journal of Theoretical Probability, 4 (3):485–510, July 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01210321.

Debicki:2020:ASM

[DH20] Krzysztof Debicki and Enkelejd Hashorva. Approximation of supremum of max-stable stationary processes and Pickands constants. Journal of Theoretical Probability, 33(1):444–464, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Dakaou:2021:LDB

[DH21] Ibrahim Dakaou and Abdoulaye Soumana Hima. Large deviations for backward stochastic differential equations driven by G-Brownian motion. Journal of Theoretical Probability, 34 (2):499–521, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01005-0.

denHollander:2022:SIP

[dHN22] Frank den Hollander and Shubhamoy Nandan. Spatially inhomogeneous populations with seed-banks: I. Duality, existence and clustering. Journal of Theoretical Probability, 35(3):1795– 1841, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01119-z.

DiTella:2022:PWR

[Di 22] Paolo Di Tella. On the propagation of the weak representation property in independently enlarged filtrations: The general case. Journal of Theoretical Probability, 35(4):2194–2216, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01145-x.

Diebolt:1991:ETS

[Die91] Jean Diebolt. Estimates of the tail of the stationary density function of certain nonlinear autoregressive processes. Journal of Theoretical Probability, 4(4):655–667, October 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01259548.

Dozzi:1994:DMD

[DIM94] Marco Dozzi, B. Gail Ivanoff, and Ely Merzbach. Doob–Meyer decomposition for set-indexed submartingales. Journal of Theoretical Probability, 7(3):499–525, July 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02213566.

Dinculeanu:1988:VVS

[Din88] Nicolae Dinculeanu. Vector-valued stochastic processes. I. Vector measures and vector-valued stochastic processes with finite variation. Journal of Theoretical Probability, 1(2):149–169, April 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01046932.

Dinwoodie:1995:OMR

[Din95] I. H. Dinwoodie. Occupation measure for random walk on the circle. Journal of Theoretical Probability, 8(3):669–677, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02218049.

Ding:2017:TPD

[Din17] Shanshan Ding. Tensor powers of the defining representation of S_{n}. Journal of Theoretical Probability, 30(3):1191–1199, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Dippon:1991:ACR

[Dip91] Jürgen Dippon. Asymptotic confidence regions of stochastic approximation procedures in Hilbert spaces. Journal of Theoretical Probability, 4(1):37–52, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01046993.

Djellout:2018:LDT

[DJ18] Hacène Djellout and Hui Jiang. Large deviations of the threshold estimator of integrated (co-)volatility vector in the presence of jumps. Journal of Theoretical Probability, 31(3):1606–1624, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Doring:2016:PIL

[DK16] Leif Döring and Andreas E. Kyprianou. Perpetual integrals for Lévy processes. Journal of Theoretical Probability, 29(3):1192– 1198, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0607-y.

Debicki:2018:ERT

[DK18] K. Debicki and K. M. Kosiński. An Erdős–Révész type law of the iterated logarithm for order statistics of a stationary Gaussian process. Journal of Theoretical Probability, 31(1):579–597, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10959-016-0710-8.pdf.

Durrett:1991:WHR

[DKW91] Rick Durrett, Harry Kesten, and Ed Waymire. On weighted heights of random trees. Journal of Theoretical Probability, 4 (1):223–237, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01047004.

Derriennic:1994:UEC

[DL94] Yves Derriennic and Michael Lin. Uniform ergodic convergence and averaging along Markov chain trajectories. Journal of Theoretical Probability, 7(3):483–497, July 1994. CODEN JTPREO.

ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02213565.

DelMoral:2000:CEP

[DL00] P. Del Moral and M. Ledoux. Convergence of empirical processes for interacting particle systems with applications to nonlinear filtering. Journal of Theoretical Probability, 13(1):225–257, January 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007743111861.

Dong:2015:NDF

[DLL15] Zhao Dong, Jiange Li, and Wenbo V. Li. A note on distribution– free symmetrization inequalities. Journal of Theoretical Probability, 28(3):958–967, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-014-0538-z.

Debicki:2020:STG

[DLM20] Krzysztof Debicki, Peng Liu, and Zbigniew Michna. Sojourn times of Gaussian processes with trend. Journal of Theoretical Probability, 33(4):2119–2166, December 2020. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-019-
00934-9.

delaPena:1993:ITA

[dlP93] Victor H. de la Peña. Inequalities for tails of adapted processes with an application to Wald’s lemma. Journal of Theoretical Probability, 6(2):285–302, April 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01047575.

Doukhan:2005:FLT

[DLSV05] Paul Doukhan, Gabriel Lang, Donatas Surgailis, and Marie-Claude Viano. Functional limit theorem for the empirical process of a class of Bernoulli shifts with long memory. Journal of Theoretical Probability, 18(1):161–186, January 2005. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
004-2593-3.

Dawson:2004:SCB

[DLZ04] D. A. Dawson, Z. Li, and X. Zhou. Superprocesses with coalescing Brownian spatial motion as large-scale limits. Journal of Theoretical Probability, 17(3):673–692, July 2004. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000040293.67957.e3.

Dani:1988:FSM

[DM88a] S. G. Dani and M. McCrudden. On the factor sets of measures and local tightness of convolution semigroups over Lie groups. Journal of Theoretical Probability, 1(4):357–370, October 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048725.

Darling:1988:SFC

[DM88b] R. W. R. Darling and Arunava Mukherjea. Stochastic flows on a countable set: Convergence in distribution. Journal of Theoretical Probability, 1(2):121–147, April 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01046931.

Deheuvels:1991:TEP

[DM91] Paul Deheuvels and David M. Mason. A tail empirical process approach to some nonstandard laws of the iterated logarithm. Journal of Theoretical Probability, 4(1):53–85, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01046994.

Davis:1995:EPL

[DM95] Burgess Davis and David McDonald. An elementary proof of the local central limit theorem. Journal of Theoretical Probability, 8(3):693–701, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02218051.

Dani:1996:IDP

[DM96] S. G. Dani and M. McCrudden. Infinitely divisible probabilities on discrete linear groups. Journal of Theoretical Probability, 9 (1):215–229, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213741.

Dhar:1997:CDP

[DM97] S. Dhar and A. Mukherjea. Convergence in distribution of products of I.I.D. nonnegative matrices. Journal of Theoretical Probability, 10(2):375–393, April 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022612516862.

Dai:2001:IPM

[DM01] Ming Dai and Arunava Mukherjea. Identification of the parameters of a multivariate normal vector by the distribution of the maximum. Journal of Theoretical Probability, 14(1):267–298, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007889519309.

Doney:2002:SAN

[DM02] R. A. Doney and R. A. Maller. Stability and attraction to normality for Lévy processes at zero and at infinity. Journal of Theoretical Probability, 15(3):751–792, July 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1016228101053.

Doney:2007:ASR

[DM07] R. A. Doney and R. A. Maller. Almost sure relative stability of the overshoot of power law boundaries. Journal of Theoretical Probability, 20(1):47–63, March 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-006-0040-3.

Diaconis:2009:TQS

[DM09] Persi Diaconis and Laurent Miclo. On times to quasi-stationarity for birth and death processes. Journal of Theoretical Probability, 22(3):558–586, September 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-009-0234-6.

Denuit:2013:OFR

[DM13] Michel M. Denuit and Mhamed Mesfioui. Ordering functions of random vectors, with application to partial sums. Journal of Theoretical Probability, 26(2):474–479, June 2013. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
012-0402-y.

Denuit:2016:MHD

[DM16a] Michel M. Denuit and Mhamed Mesfioui. Multivariate higherdegree stochastic increasing convexity. Journal of Theoretical Probability, 29(4):1599–1623, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
015-0628-6.

Dobric:2016:LBD

[DM16b] Vladimir Dobric and Lisa Marano. Lower bounds for the distribution of suprema of Brownian increments and Brownian motion normalized by the corresponding modulus functions. Journal of Theoretical Probability, 29(1):1–31, March 2016. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
014-0570-z.

Durga:2023:EBN

[DM23] N. Durga and P. Muthukumar. Exponential behaviour of nonlinear fractional Schrödinger evolution equation with complex potential and Poisson jumps. Journal of Theoretical Probability, 36(4):1939–1955, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-023-01266-5.

Diop:2024:MPA

[DMCN24] Amadou Diop, Mamadou Moustapha Mbaye, Yong-Kui Chang, and Gaston Mandata N’Guérékata. Measure pseudo-S-asymptotically Bloch-type periodicity of some semilinear stochastic integrodifferential equations. Journal of Theoretical Probability, 37(3):2253–2276, September 2024. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-024-
01335-3.

Durmus:2024:PMI

[DMNS24] Alain Durmus, Eric Moulines, Alexey Naumov, and Sergey Samsonov. Probability and moment inequalities for additive functionals of geometrically ergodic Markov chains. Journal of Theoretical Probability, 37(3):2184–2233, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-024-01315-7.

Dellacherie:2009:UTP

[DMS09] Claude Dellacherie, Servet Martinez, and Jaime San Martin. Ultrametric and tree potential. Journal of Theoretical Probability, 22(2):311–347, June 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0209-7.

Dedecker:2007:WIP

[DMV07] Jérôme Dedecker, Florence Merlevède, and Dalibor Volný. On the weak invariance principle for non-adapted sequences under projective criteria. Journal of Theoretical Probability, 20(4): 971–1004, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0090-1.

Dinwoodie:1995:OMM

[DN95] I. H. Dinwoodie and Peter Ney. Occupation measures for Markov chains. Journal of Theoretical Probability, 8(3):679–691, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF02218050.

Dette:2012:MMR

[DN12] Holger Dette and Jan Nagel. Matrix measures, random moments, and Gaussian ensembles. Journal of Theoretical Probability, 25 (1):25–49, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0370-7.

DOvidio:2011:CPR

[DO11] Mirko D’Ovidio and Enzo Orsingher. Composition of processes and related partial differential equations. Journal of Theoretical Probability, 24(2):342–375, June 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-010-0284-9.

Deheuvels:2013:UBF

[DO13] Paul Deheuvels and Sarah Ouadah. Uniform-in-bandwidth functional limit laws. Journal of Theoretical Probability, 26(3):697– 721, September 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0376-1.

Dobric:1990:DTF

[Dob90] V. Dobric. The decomposition theorem for functions satisfying the law of large numbers. Journal of Theoretical Probability, 3 (4):489–496, October 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046091.

Dobric:1994:WCG

[Dob94] Vladimir Dobrić. Weakly compactly generated Banach spaces and the strong law of large numbers. Journal of Theoretical Probability, 7(1):129–134, January 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02213363.

Dobler:2017:DTO

[Döb17] Christian Döbler. Distributional transformations without orthogonality relations. Journal of Theoretical Probability, 30 (1):85–116, March 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0646-4.

Dong:2008:UIM

[Don08] Z. Dong. On the uniqueness of invariant measure of the Burgers equation driven by Lévy processes. Journal of Theoretical Probability, 21(2):322–335, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0143-0.

Dong:2023:SSV

[Don23] Xiaoyu Dong. The smallest singular value of a shifted random matrix. Journal of Theoretical Probability, 36(4):2448– 2475, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01255-8.

Dedecker:2004:CDS

[DP04] J. Dedecker and C. Prieur. Coupling for τ-dependent sequences and applications. Journal of Theoretical Probability, 17(4): 861–885, October 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-004-0578-x.

Didier:2012:ESG

[DP12] Gustavo Didier and Vladas Pipiras. Exponents, symmetry groups and classification of operator fractional Brownian motions. Journal of Theoretical Probability, 25(2):353–395, June 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-011-0348-5.

Davis:2017:ERW

[DP17] Burgess Davis and Jonathon Peterson. Excited random walks with non-nearest neighbor steps. Journal of Theoretical Probability, 30(4):1255–1284, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Duan:2022:ASR

[DP22] Junxia Duan and Jun Peng. An approximation scheme for reflected stochastic differential equations with non-Lipschitzian coefficients. Journal of Theoretical Probability, 35(1):575–602, March 2022. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-020-01052-7.

Davydov:2000:MSC

[DPR00] Yu. Davydov, V. Paulauskas, and A. Rackauskas. More on pstable convex sets in Banach spaces. Journal of Theoretical Probability, 13(1):39–64, January 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1007726708227.

DelMoral:2011:CSI

[DPR11] P. Del Moral, F. Patras, and S. Rubenthaler. Convergence of U-statistics for interacting particle systems. Journal of Theoretical Probability, 24(4):1002–1027, December 2011. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
011-0355-6.

Ding:2021:EMA

[DQ21] Xiaojie Ding and Huijie Qiao. Euler–Maruyama approximations for stochastic McKean–Vlasov equations with non-Lipschitz coefficients. Journal of Theoretical Probability, 34(3):1408–1425, September 2021. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-020-01041-w.

Deheuvels:1993:CST

[DR93] Paul Deheuvels and Pál Révész. On the coverage of strassen-type sets by sequences of Wiener processes. Journal of Theoretical Probability, 6(3):427–449, July 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01066710.

Dani:1998:NTG

[DR98] S. G. Dani and C. R. E. Raja. A note on Tortrat groups. Journal of Theoretical Probability, 11(2):571–576, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022600326181.

Davydov:2009:APDa

[DR09a] Youri Davydov and Vladimir Rotar. On asymptotic proximity of distributions. Journal of Theoretical Probability, 22(1):82–98, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0178-2.

Davydov:2009:APDb

[DR09b] Youri Davydov and Vladimir Rotar. On asymptotic proximity of distributions. Journal of Theoretical Probability, 22(1):99, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10959-008-0182-6.pdf.

Dette:2010:RBM

[DR10] Holger Dette and Bettina Reuther. Random block matrices and matrix orthogonal polynomials. Journal of Theoretical Probability, 23(2):378–400, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0189-z.

Dunau:1988:CTC

[DS88] Jean-Louis Dunau and Henri Senateur. A characterization of the type of the Cauchy–Hua measure on real symmetric matrices. Journal of Theoretical Probability, 1(3):263–270, July 1988.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01246629.

Devroye:1990:RWH

[DS90] Luc Devroye and Amine Sbihi. Random walks on highly symmetric graphs. Journal of Theoretical Probability, 3(4):497–514, October 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01046092.

Durrett:1991:SRR

[DS91] Richard Durrett and Jeffrey E. Steif. Some rigorous results for the Greenberg–Hastings model. Journal of Theoretical Probability, 4(4):669–690, October 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01259549.

Djehiche:1998:LDH

[DS98] Boualem Djehiche and Alexander Schied. Large deviations for hierarchical systems of interacting jump processes. Journal of Theoretical Probability, 11(1):1–24, January 1998. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021690707556.

Durrett:2000:BMC

[DS00] Rick Durrett and Rinaldo B. Schinazi. Boundary modified contact processes. Journal of Theoretical Probability, 13(2):575–594, April 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007881121529.

Dobler:2014:QCL

[DS14] Christian Döbler and Michael Stolz. A quantitative Central Limit Theorem for linear statistics of random matrix eigenvalues. Journal of Theoretical Probability, 27(3):945–953, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-012-0451-2.

Diaconis:1996:NIF

[DSC96] P. Diaconis and L. Saloff-Coste. Nash inequalities for finite Markov chains. Journal of Theoretical Probability, 9(2):459–510, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02214660.

daSilva:2018:MSD

[dSEE18] José Luís da Silva, Mohamed Erraoui, and El Hassan Essaky. Mixed stochastic differential equations: Existence and uniqueness result. Journal of Theoretical Probability, 31(2):1119–1141, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

delTenno:2009:CPD

[dT09] Ivan del Tenno. Cut points and diffusions in random environment. Journal of Theoretical Probability, 22(4):891–933, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0169-3.

Durieu:2014:EPC

[DT14] Olivier Durieu and Marco Tusche. An empirical process Central Limit Theorem for multidimensional dependent data. Journal of Theoretical Probability, 27(1):249–277, March 2014. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
012-0450-3.

Dette:2017:HDR

[DT17] Holger Dette and Dominik Tomecki. Hankel determinants of random moment sequences. Journal of Theoretical Probability, 30 (4):1539–1564, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Dhoyer:2024:LBH

[DT24] Rémy Dhoyer and Ciprian A. Tudor. Limit behavior in high-dimensional regime for the Wishart tensors in Wiener chaos. Journal of Theoretical Probability, 37(2):1445–1468, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-024-01328-2.

Dunker:2000:ESB

[Dun00] Thomas Dunker. Estimates for the small ball probabilities of the fractional Brownian sheet. Journal of Theoretical Probability, 13(2):357–382, April 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1007897525164. See er-

ratum [Dun01].

Dunker:2001:EES

[Dun01] Thomas Dunker. Erratum: Estimates for the Small Ball Probabilities of the Fractional Brownian Sheet, J. Theor. Prob. 13, 357–382 (2000). Journal of Theoretical Probability, 14(2): 607–608, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1011184218841. See [Dun00].

Duncan:2021:TFB

[Dun21] Tyrone E. Duncan. Theta functions and Brownian motion. Journal of Theoretical Probability, 34(1):81–89, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-019-00977-y.

Durrett:1992:MPS

[Dur92] R. Durrett. Multicolor particle systems with large threshold and range. Journal of Theoretical Probability, 5(1):127–152, January 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01046781.

Durrett:2003:SC

[Dur03] Rick Durrett. Shuffling chromosomes. Journal of Theoretical Probability, 16(3):725–750, July 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025676617383.

Dereich:2011:HRV

[DV11] S. Dereich and C. Vormoor. The high resolution vector quantization problem with Orlicz norm distortion. Journal of Theoretical Probability, 24(2):517–544, June 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-010-0327-2.

Dippon:2006:ARM

[DW06] Jürgen Dippon and Harro Walk. The averaged Robbins– Monro method for linear problems in a Banach space. Journal of Theoretical Probability, 19(1):166–189, January 2006. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
006-0007-4.

Dupuis:2022:LDP

[DW22] Paul Dupuis and Guo-Jhen Wu. Large deviation properties of the empirical measure of a metastable small noise diffusion. Journal of Theoretical Probability, 35(2):1049–1136, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01072-3.

Denisov:2024:GFA

[DW24] Denis Denisov and Vitali Wachtel. Green function for an asymptotically stable random walk in a half space. Journal of Theoretical Probability, 37(2):1745–1786, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-023-01283-4.

Dobric:1995:ATC

[DY95] V. Dobrić and J. E. Yukich. Asymptotics for transportation cost in high dimensions. Journal of Theoretical Probability, 8 (1):97–118, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213456.

Eisenbaum:2001:LTM

[EF01] Nathalie Eisenbaum and Antónia Földes. Local times of Markov processes approximated by a generalized iterated Brownian motion. Journal of Theoretical Probability, 14(2):559–576, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1011176017024.

Erraoui:2020:BRL

[EHL20] Mohamed Erraoui, Astrid Hilbert, and Mohammed Louriki. Bridges with random length: Gamma case. Journal of Theoretical Probability, 33(2):931–953, June 2020. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s10959-
019-00955-4.pdf.

Eichelsbacher:1997:LDP

[Eic97] Peter Eichelsbacher. Large deviations for products of empirical probability measures in the τ-topology. Journal of Theoretical Probability, 10(4):903–920, October 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022610532538.

Einmahl:2007:GSF

[Ein07] Uwe Einmahl. A generalization of Strassen’s functional LIL. Journal of Theoretical Probability, 20(4):901–915, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-007-0091-0.

Eisele:1993:TGG

[Eis93] Peter Eisele. Tortrat groups — groups with a nice behavior for sequences of shifted convolution powers. Journal of Theoretical Probability, 6(4):671–691, October 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01049170.

Ejsmont:2014:CSF

[Ejs14] Wiktor Ejsmont. Characterizations of some free random variables by properties of conditional moments of third degree polynomials. Journal of Theoretical Probability, 27(3):915– 931, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-012-0467-7.pdf.

Ercolani:2019:SAH

[EJU19] Nicholas M. Ercolani, Sabine Jansen, and Daniel Ueltschi. Singularity analysis for heavy-tailed random variables. Journal of Theoretical Probability, 32(1):1–46, March 2019. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Enriquez:2001:MCG

[EK01] Nathanaël Enriquez and Yuri Kifer. Markov chains on graphs and Brownian motion. Journal of Theoretical Probability, 14 (2):495–510, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1011119932045.

ElOtmani:2009:RBD

[El 09] Mohamed El Otmani. Reflected BSDE driven by a Lévy process. Journal of Theoretical Probability, 22(3):601–619, September 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-009-0229-3.

Eichelsbacher:1995:LDP

[EL95] Peter Eichelsbacher and Matthias Löwe. A large deviation principle for m-variate von Mises-statistics and U-statistics. Journal of Theoretical Probability, 8(4):807–824, October 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02410113.

Eichelsbacher:2019:LMM

[EL19] Peter Eichelsbacher and Matthias Löwe. Lindeberg’s method for moderate deviations and random summation. Journal of Theoretical Probability, 32(2):872–897, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Einmahl:1989:DET

[EM89] Uwe Einmahl and David M. Mason. Darling–Erdős theorems for martingales. Journal of Theoretical Probability, 2(4):437–460, October 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01051877.

Einmahl:1992:APE

[EM92] Uwe Einmahl and David M. Mason. Approximations to permutation and exchangeable processes. Journal of Theoretical Probability, 5(1):101–126, January 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01046780.

Einmahl:2000:EPA

[EM00] Uwe Einmahl and David M. Mason. An empirical process approach to the uniform consistency of kernel-type function estimators. Journal of Theoretical Probability, 13(1):1–37, January 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007769924157.

Erickson:2005:DIS

[EM05] K. B. Erickson and Ross A. Maller. Drift to infinity and the strong law for subordinated random walks and Lévy processes. Journal of Theoretical Probability, 18(2):359–375, April 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-005-3507-8.

Evans:2018:PDS

[EM18] Steven N. Evans and Ilya Molchanov. Polar decomposition of scale-homogeneous measures with application to Lévy measures of strictly stable laws. Journal of Theoretical Probability, 31(3): 1303–1321, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Erhard:2017:BPH

[EMP17] Dirk Erhard, Julián Martínez, and Julien Poisat. Brownian paths homogeneously distributed in space: Percolation phase transition and uniqueness of the unbounded cluster. Journal of Theoretical Probability, 30(3):784–812, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Englander:2020:GSP

[Eng20] János Engländer. A generalization of the submartingale property: Maximal inequality and applications to various stochastic processes. Journal of Theoretical Probability, 33(1):506–521, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Eyi-Obiang:2017:SPC

[EOOM17] Fulgence Eyi-Obiang, Youssef Ouknine, and Octave Moutsinga. On the study of processes of ∑(H) and ∑_{s(H)} classes. Journal of Theoretical Probability, 30(1):117–142, March 2017. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
015-0640-x.

Essifi:2015:ERP

[EP15a] Rim Essifi and Marc Peigné. Erratum to: Return Probabilities for the Reflected Random Walk on N_{0}. Journal of Theoretical Probability, 28(3):1250–1251, September 2015. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-014-0568-6.pdf.

Essifi:2015:RPR

[EP15b] Rim Essifi and Marc Peigné. Return probabilities for the reflected random walk on N_{0}. Journal of Theoretical Probability, 28(1):231–258, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0490-3.

Einmahl:2001:FLI

[ER01] John H. J. Einmahl and Andrew Rosalsky. The functional law of the iterated logarithm for the empirical process based on sample means. Journal of Theoretical Probability, 14(2):577–597, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1011128101094.

Eisenbaum:1999:MRV

[ES99] Nathalie Eisenbaum and Zhan Shi. Measuring the rarely visited sites of Brownian motion. Journal of Theoretical Probability, 12(3):595–613, July 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021615513008.

Ernst:2022:DBD

[ES22] Marie Ernst and Yvik Swan. Distances between distributions via Stein’s method. Journal of Theoretical Probability, 35(2): 949–987, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01075-8.

Eldredge:2014:WRT

[ESC14] Nathaniel Eldredge and Laurent Saloff-Coste. Widder’s representation theorem for symmetric local Dirichlet spaces. Journal of Theoretical Probability, 27(4):1178–1212, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-013-0484-1.

Englander:2018:TCI

[EV18] János Engländer and Stanislav Volkov. Turning a coin over instead of tossing it. Journal of Theoretical Probability, 31(2):

1097–1118, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-016-0725-1.pdf.

Englander:2019:IRW

[EV19] János Engländer and Stanislav Volkov. Impatient random walk. Journal of Theoretical Probability, 32(4):2020–2043, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10959-019-00901-4.pdf.

Evans:1989:LPL

[Eva89] Steven N. Evans. Local properties of Lévy processes on a totally disconnected group. Journal of Theoretical Probability, 2 (2):209–259, April 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01053411.

Evans:1993:LFB

[Eva93] Steven N. Evans. Local field Brownian motion. Journal of Theoretical Probability, 6(4):817–850, October 1993. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01049177.

Fainleib:1998:SVS

[Fai98] A. S. Fainleib. On small values of semi-additive functions. Journal of Theoretical Probability, 11(3):609–619, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022694312915.

Falconer:1989:PRC

[Fal89] K. J. Falconer. Projections of random Cantor sets. Journal of Theoretical Probability, 2(1):65–70, January 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01048269.

Falconer:1994:MSS

[Fal94] K. J. Falconer. The multifractal spectrum of statistically self-similar measures. Journal of Theoretical Probability, 7 (3):681–702, July 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213576.

Falconer:2002:TFL

[Fal02] Kenneth J. Falconer. Tangent fields and the local structure of random fields. Journal of Theoretical Probability, 15(3):731–750, July 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1016276016983.

Fang:2016:MCB

[Fan16] Xiao Fang. A multivariate CLT for bounded decomposable random vectors with the best known rate. Journal of Theoretical Probability, 29(4):1510–1523, December 2016. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
015-0619-7.

Fan:2017:TCT

[Fan17] Yuguang Fan. Tightness and convergence of trimmed Lévy processes to normality at small times. Journal of Theoretical Probability, 30(2):675–699, June 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Fan:2018:EUS

[Fan18] Sheng Jun Fan. Existence, uniqueness and stability of L^{1} solutions for multidimensional backward stochastic differential equations with generators of one-sided Osgood type. Journal of Theoretical Probability, 31(3):1860–1899, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Fan:2019:DFA

[Fan19] Xiliang Fan. Derivative formulas and applications for degenerate stochastic differential equations with fractional noises. Journal of Theoretical Probability, 32(3):1360–1381, September 2019. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Farahmand:1994:LMR

[Far94] K. Farahmand. Local maxima of a random trigonometric polynomial. Journal of Theoretical Probability, 7(1):175–185, January 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02213366.

Farahmand:2003:EMC

[Far03] K. Farahmand. Exceedence measure of classes of algebraic polynomials. Journal of Theoretical Probability, 16(2):419–426, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023526828571.

Faraud:2011:ESS

[Far11] Gabriel Faraud. Estimates on the speedup and slowdown for a diffusion in a drifted Brownian potential. Journal of Theoretical Probability, 24(1):194–239, March 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-009-0251-5.

Fazekas:2007:ASF

[FC07] István Fazekas and Alexey Chuprunov. An almost sure functional limit theorem for the domain of geometric partial attraction of semistable laws. Journal of Theoretical Probability, 20 (2):339–353, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0074-1.

Ferreiro-Castilla:2012:IAC

[FCU12] Albert Ferreiro-Castilla and Frederic Utzet. Inversion of analytic characteristic functions and infinite convolutions of exponential and Laplace densities. Journal of Theoretical Probability, 25(1):205–230, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0313-8.

Feldman:2004:HTF

[Fel04] G. M. Fel’dman. On the Heyde theorem for finite Abelian groups. Journal of Theoretical Probability, 17(4):929–941, October 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-004-0583-0.

Feldman:2015:SDT

[Fel15] Gennadiy Feldman. On the Skitovich–Darmois theorem for the group of p-adic numbers. Journal of Theoretical Probability, 28 (2):539–549, June 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0525-9.

Feng:1996:HLR

[Fen96] Jianfeng Feng. The hydrodynamic limit for the reaction diffusion equation — an approach in terms of the GPV method. Journal of Theoretical Probability, 9(2):285–299, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02214650.

Ferreyra:1989:WZT

[Fer89] Guillermo Ferreyra. A Wong–Zakai-type theorem for certain discontinuous semimartingales. Journal of Theoretical Probability, 2(3):313–323, July 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01054019.

Ferger:1994:ERC

[Fer94] Dietmar Ferger. On exact rates of convergence in functional limit theorems for U-statistic type processes. Journal of Theoretical Probability, 7(4):709–723, October 1994. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02214367.

Feng:2004:SFE

[FF04] De-Jun Feng and Ding Feng. On a statistical framework for estimation from random set observations. Journal of Theoretical Probability, 17(1):85–110, January 2004. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000020476.12997.c2.

Feinsilver:1997:DMS

[FFS97] Philip Feinsilver, Uwe Franz, and René Schott. Duality and multiplicative stochastic processes on quantum groups. Journal of Theoretical Probability, 10(3):795–818, July 1997. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022618114810.

Falconer:1992:GRC

[FG92a] K. J. Falconer and G. R. Grimmett. On the geometry of random Cantor sets and fractal percolation. Journal of Theoretical Probability, 5(3):465–485, July 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01060430.

Fernandez:1992:SCT

[FG92b] Begoña Fernández and Luis G. Gorostiza. Stability of a class of transformations of distribution-valued processes and stochastic evolution equations. Journal of Theoretical Probability, 5(4): 661–678, October 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01058724.

Falconer:1994:GRC

[FG94] K. J. Falconer and G. R. Grimmett. On the geometry of random Cantor sets and fractal percolation. Journal of Theoretical Probability, 7(1):209–210, January 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02213369.

Farahmand:1999:CZA

[FG99] K. Farahmand and A. Grigorash. Complex zeros of algebraic polynomial with non-zero mean random coefficients. Journal of Theoretical Probability, 12(4):1037–1044, October 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021697221011.

Feldman:2000:SDT

[FG00] G. M. Fel’dman and P. Graczyk. On the Skitovich–Darmois theorem for compact Abelian groups. Journal of Theoretical Probability, 13(3):859–869, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1007870814570.

Faraud:2014:BBD

[FG14] Gabriel Faraud and Stéphane Goutte. Bessel bridges decomposition with varying dimension: Applications to finance. Journal of Theoretical Probability, 27(4):1375–1403, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-013-0496-x.

Friedland:2016:SBE

[FGG16] Omer Friedland, Ohad Giladi, and Olivier Guédon. Small ball estimates for quasi-norms. Journal of Theoretical Probability, 29(4):1624–1643, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-015-0622-z.

Fan:2020:CMD

[FGL20] Xiequan Fan, Ion Grama, and Quansheng Liu. Cramér moderate deviation expansion for martingales with one-sided Sakhanenko’s condition and its applications. Journal of Theoretical Probability, 33(2):749–787, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Farahmand:1997:ENL

[FH97] K. Farahmand and P. Hannigan. The expected number of local maxima of a random algebraic polynomial. Journal of Theoretical Probability, 10(4):991–1002, October 1997. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022618801587.

Franz:2015:SAO

[FH15] Uwe Franz and Tarek Hamdi. Stochastic analysis for obtuse random walks. Journal of Theoretical Probability, 28(2):619–649, June 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-013-0522-z.

Feng:2022:LDP

[FHL22] Xinwei Feng, Lidan He, and Zhi Liu. Large deviation principles of realized Laplace transform of volatility. Journal of Theoretical Probability, 35(1):186–208, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-020-01055-4.

Fierro:1990:DAS

[Fie90] Raul Fierro. Domains of attraction for semi-martingales taking values in the tempered distributions space. Journal of Theoretical Probability, 3(1):31–49, January 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01063327.

Fill:1992:SSD

[Fil92] James Allen Fill. Strong stationary duality for continuous-time Markov chains. Part I: Theory. Journal of Theoretical Probability, 5(1):45–70, January 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01046778.

Fill:1996:EFM

[Fil96] James Allen Fill. An exact formula for the move-to-front rule for self-organizing lists. Journal of Theoretical Probability, 9(1): 113–160, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213737.

Fill:2009:HTF

[Fil09a] James Allen Fill. On hitting times and fastest strong stationary times for skip-free and more general chains. Journal of Theoretical Probability, 22(3):587–600, September 2009. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
009-0233-7.

Fill:2009:PTD

[Fil09b] James Allen Fill. The passage time distribution for a birth-and-death chain: Strong stationary duality gives a first stochastic proof. Journal of Theoretical Probability, 22(3):543–557, September 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-009-0235-5.

Fisch:1990:ODC

[Fis90] Robert Fisch. The one-dimensional cyclic cellular automaton: A system with deterministic dynamics that emulates an interacting particle system with stochastic dynamics. Journal of Theoretical Probability, 3(2):311–338, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01045164.

Fitzsimmons:1989:TCS

[Fit89] P. J. Fitzsimmons. Time changes of symmetric Markov processes and a Feynman–Kac formula. Journal of Theoretical Probability, 2(4):487–501, October 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01051880.

Fitzsimmons:1999:MPE

[Fit99] P. J. Fitzsimmons. Markov processes with equal capacities. Journal of Theoretical Probability, 12(1):271–292, January 1999.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021713114477.

Flandoli:2014:SCW

[FIT14] Franco Flandoli, Peter Imkeller, and Ciprian A. Tudor. 2Dstochastic currents over the Wiener sheet. Journal of Theoretical Probability, 27(2):552–575, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-012-0453-0.

Fleischer:1995:REN

[FJ95] I. Fleischer and A. Joffe. Ratio ergodicity for non-homogeneous Markov chains in general state spaces. Journal of Theoretical Probability, 8(1):31–37, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02213452.

Farahmand:2005:APN

[FJ05] K. Farahmand and Jay Jahangiri. Algebraic polynomials with non-identical random coefficients. Journal of Theoretical Probability, 18(4):827–835, October 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-005-7527-1.

Fan:2012:GCT

[FJ12] Sheng-Jun Fan and Long Jiang. A generalized comparison theorem for BSDEs and its applications. Journal of Theoretical Probability, 25(1):50–61, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-010-0293-8.

Forsstrom:2017:SCR

[FJ17] Malin P. Forsström and Johan Jonasson. The spectrum and convergence rates of exclusion and interchange processes on the complete graph. Journal of Theoretical Probability, 30(2):639– 654, June 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Fukushima:2014:QLD

[FK14] Ryoki Fukushima and Naoki Kubota. Quenched large deviations for multidimensional random walk in random environment with holding times. Journal of Theoretical Probability, 27(4):

1140–1166, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0514-z.

Forrester:2018:PAE

[FK18] Peter J. Forrester and Santosh Kumar. The probability that all eigenvalues are real for products of truncated real orthogonal random matrices. Journal of Theoretical Probability, 31(4): 2056–2071, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Flasche:2020:RZR

[FK20] Hendrik Flasche and Zakhar Kabluchko. Real zeroes of random analytic functions associated with geometries of constant curvature. Journal of Theoretical Probability, 33(1):103–133, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Forster:2021:PEP

[FKK21] Yanik-Pascal Förster, Mario Kieburg, and Holger Kösters. Polynomial ensembles and Pólya frequency functions. Journal of Theoretical Probability, 34(4):1917–1950, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
01030-z.

Foss:2018:PLC

[FKM18] Sergey Foss, Takis Konstantopoulos, and Thomas Mountford. Power law condition for stability of Poisson hail. Journal of Theoretical Probability, 31(2):684–704, June 2018. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-016-0723-3.pdf.

Finkelstein:1994:CLR

[FKT94] Mark Finkelstein, Victor M. Kruglov, and Howard G. Tucker. Convergence in law of random sums with non-random centering. Journal of Theoretical Probability, 7(3):565–598, July 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02213569.

Fleermann:2022:LCL

[FKT22] Michael Fleermann, Werner Kirsch, and Gabor Toth. Local central limit theorem for multi-group Curie–Weiss models. Journal of Theoretical Probability, 35(3):2009–2019, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01122-4.

Fleischmann:2006:PCR

[FKX06] Klaus Fleischmann, Achim Klenke, and Jie Xiong. Pathwise convergence of a rescaled super–Brownian catalyst reactant process. Journal of Theoretical Probability, 19(3):557–588, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-006-0025-2.

Foss:2007:DCT

[FKZ07] Serguei Foss, Takis Konstantopoulos, and Stan Zachary. Discrete and continuous time modulated random walks with heavytailed increments. Journal of Theoretical Probability, 20(3):581– 612, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0081-2.

Feinberg:2023:SUF

[FKZ23] Eugene A. Feinberg, Pavlo O. Kasyanov, and Michael Z. Zgurovsky. Semi-uniform Feller stochastic kernels. Journal of Theoretical Probability, 36(4):2262–2283, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01230-9.

Friesen:2013:SLM

[FL13] Olga Friesen and Matthias Löwe. The semicircle law for matrices with independent diagonals. Journal of Theoretical Probability, 26(4):1084–1096, December 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-011-0383-2.

Fill:2014:HTI

[FL14] James Allen Fill and Vince Lyzinski. Hitting times and interlacing eigenvalues: A stochastic approach using intertwinings. Journal of Theoretical Probability, 27(3):954–981, September 2014.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-012-0457-9.

Fill:2016:SSD

[FL16] James Allen Fill and Vince Lyzinski. Strong stationary duality for diffusion processes. Journal of Theoretical Probability, 29(4): 1298–1338, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0612-1.

Feldman:2017:CID

[FM17] Gennadiy Feldman and Margaryta Myronyuk. On a characterization of idempotent distributions on discrete fields and on the field of p-adic numbers. Journal of Theoretical Probability, 30 (2):608–623, June 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Fonseca-Mora:2018:ECC

[FM18] C. A. Fonseca-Mora. Existence of continuous and Càdlàg versions for cylindrical processes in the dual of a nuclear space. Journal of Theoretical Probability, 31(2):867–894, June 2018. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Fonseca-Mora:2020:LPI

[FM20] C. A. Fonseca-Mora. Lévy processes and infinitely divisible measures in the dual of a nuclear space. Journal of Theoretical Probability, 33(2):649–691, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Fonseca-Mora:2023:ASUa

[FM23a] C. A. Fonseca-Mora. Almost sure uniform convergence of stochastic processes in the dual of a nuclear space. Journal of Theoretical Probability, 36(1):1–26, March 2023. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01243-y.

Fonseca-Mora:2023:ASUb

[FM23b] C. A. Fonseca-Mora. Almost sure uniform convergence of stochastic processes in the dual of a nuclear space. Journal of Theoretical Probability, 36(4):2564–2589, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01243-y.

Fleischmann:2011:HIG

[FMW11] Klaus Fleischmann, Leonid Mytnik, and Vitali Wachtel. Hölder index at a given point for density states of super-α-stable motion of index 1 + β. Journal of Theoretical Probability, 24(1):66–92, March 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-010-0334-3.

Funk:2015:STT

[FNN15] Jacob Funk, Mihai Nica, and Michael Noyes. Stabilization time for a type of evolution on binary strings. Journal of Theoretical Probability, 28(3):848–865, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-013-0515-y.

Franz:1999:PRP

[FNS99] Uwe Franz, Daniel Neuenschwander, and René Schott. Phase retrieval for probability distributions on quantum groups and braided groups. Journal of Theoretical Probability, 12(2): 585–594, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021646631915.

Foldes:2003:AMH

[Föl03] Antónia Földes. About the measure of heavily visited points of the Brownian motion. Journal of Theoretical Probability, 16 (1):21–45, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022270202817.

Forsstrom:2022:CRI

[For22] Malin P. Forsström. Color representations of Ising models. Journal of Theoretical Probability, 35(1):603–635, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01051-8.

Frick:2010:RRW

[FP10] Sarah Bailey Frick and Karl Petersen. Reinforced random walks and adic transformations. Journal of Theoretical Probability, 23(3):920–943, September 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-010-0282-y.

Fekete:2021:BDM

[FPP21] Dorottya Fekete, Sandra Palau, and Jose Luis Pérez. Backbone decomposition of multitype superprocesses. Journal of Theoretical Probability, 34(3):1149–1178, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-021-
01076-7.

Franke:2007:FNC

[Fra07] Brice Franke. A functional non–Central Limit Theorem for jump-diffusions with periodic coefficients driven by stable Lévynoise. Journal of Theoretical Probability, 20(4):1087–1100, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-007-0099-5.

Franceschi:2021:GFO

[Fra21] S. Franceschi. Green’s functions with oblique Neumann boundary conditions in the quadrant. Journal of Theoretical Probability, 34(4):1775–1810, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-020-01043-8.

Flandoli:2018:IDC

[FRZ18] Franco Flandoli, Francesco Russo, and Giovanni Zanco. Infinitedimensional calculus under weak spatial regularity of the processes. Journal of Theoretical Probability, 31(2):789–826, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10959-016-0724-2.pdf.

Feinsilver:1992:ASL

[FS92] P. Feinsilver and R. Schott. Appell systems on Lie groups. Journal of Theoretical Probability, 5(2):251–281, April 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01046735.

Fuchs:2013:SRM

[FS13] Florian Fuchs and Robert Stelzer. Spectral representation of multivariate regularly varying Lévy and CARMA processes.

Journal of Theoretical Probability, 26(2):410–436, June 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-011-0369-0.

Forsstrom:2021:AIL

[FS21] Malin P. Forsström and Jeffrey E. Steif. An analysis of the induced linear operators associated to divide and color models. Journal of Theoretical Probability, 34(2):1043–1060, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01001-4.

Feng:2021:SNM

[FSZ21] Xinwei Feng, Qi-Man Shao, and Ofer Zeitouni. Self-normalized moderate deviations for random walk in random scenery. Journal of Theoretical Probability, 34(1):103–124, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-019-
00965-2.

Fageot:2019:SLS

[FU19] Julien Fageot and Michael Unser. Scaling limits of solutions of linear stochastic differential equations driven by Lévy white noises. Journal of Theoretical Probability, 32(3):1166–1189, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Fulman:2008:CRR

[Ful08] Jason Fulman. Convergence rates of random walk on irreducible representations of finite groups. Journal of Theoretical Probability, 21(1):193–211, March 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0102-1.

Fushiya:2010:WCT

[Fus10] Hirotaka Fushiya. Weak convergence theorem of a nonnegative random walk to sticky reflected Brownian motion. Journal of Theoretical Probability, 23(4):1157–1181, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-009-0244-4.

Fageot:2020:TAP

[FUW20] Julien Fageot, Michael Unser, and John Paul Ward. The n-term approximation of periodic generalized Lévy processes. Journal of Theoretical Probability, 33(1):180–200, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Fleischmann:2005:MSC

[FV05] Klaus Fleischmann and Vladimir A. Vatutin. Multi-scale clustering for a non–Markovian spatial branching process. Journal of Theoretical Probability, 18(4):719–755, October 2005. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
005-7524-4.

Falconer:2009:MMO

[FV09] K. J. Falconer and J. Lévy Véhel. Multifractional, multistable, and other processes with prescribed local form. Journal of Theoretical Probability, 22(2):375–401, June 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-008-0147-9.

Falconer:2020:SSP

[FV20] K. J. Falconer and J. Lévy Véhel. Self-stabilizing processes based on random signs. Journal of Theoretical Probability, 33(1):134– 152, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Fu:2020:SPS

[FW20] Zuopeng Fu and Yizao Wang. Stable processes with stationary increments parameterized by metric spaces. Journal of Theoretical Probability, 33(3):1737–1754, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Fang:2018:LMD

[FWS18] Lulu Fang, Min Wu, and Lei Shang. Large and moderate deviation principles for Engel continued fractions. Journal of Theoretical Probability, 31(1):294–318, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Fitzsimmons:2008:TCA

[FY08] Patrick J. Fitzsimmons and Kouji Yano. Time change approach to generalized excursion measures, and its application to limit theorems. Journal of Theoretical Probability, 21(1):246–265, March 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-007-0108-8.

Fang:2001:LDB

[FZ01] Shizan Fang and Tusheng Zhang. Large deviations for the Brownian motion on loop groups. Journal of Theoretical Probability, 14(2):463–483, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1011115831136.

Gabriyelyan:2011:ACS

[Gab11] S. S. Gabriyelyan. Absolute continuity and singularity of two probability measures on a filtered space. Journal of Theoretical Probability, 24(3):595–614, September 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-011-0359-2.

Ganihodzhaev:1991:SPG

[Gan91] N. N. Ganihodzhaev. On stochastic processes generated by quadric operators. Journal of Theoretical Probability, 4(4): 639–653, October 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01259547.

Ganesan:2022:ETS

[Gan22] Ghurumuruhan Ganesan. Euclidean travelling salesman problem with location-dependent and power-weighted edges. Journal of Theoretical Probability, 35(2):819–862, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-021-
01080-x.

Gao:2003:MDL

[Gao03] Fuqing Gao. Moderate deviations and large deviations for kernel density estimators. Journal of Theoretical Probability, 16 (2):401–418, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023574711733.

Garcia:2004:ECP

[Gar04] J. Garcia. An extension of the contraction principle. Journal of Theoretical Probability, 17(2):403–434, April 2004. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000020701.20424.d8.

Garcia:2008:LDP

[Gar08] Jorge Garcia. A large deviation principle for stochastic integrals. Journal of Theoretical Probability, 21(2):476–501, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-007-0136-4.

Gaasemyr:2006:SIM

[Gås06] Jørund Gåsemyr. The spectrum of the independent Metropolis– Hastings algorithm. Journal of Theoretical Probability, 19(1): 152–165, January 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-006-0009-2.

Gaspar:2019:RNO

[Gas19] Pastorel Gaspar. On random normal operators and their spectral measures. Journal of Theoretical Probability, 32(4):2088–2110, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Gaunt:2016:RCN

[Gau16] Robert E. Gaunt. Rates of convergence in normal approximation under moment conditions via new bounds on solutions of the Stein equation. Journal of Theoretical Probability, 29(1): 231–247, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0562-z.

Gaunt:2020:WKE

[Gau20] Robert E. Gaunt. Wasserstein and Kolmogorov error bounds for variance-gamma approximation via Stein’s method I. Journal of Theoretical Probability, 33(1):465–505, March 2020. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-018-0867-4.pdf.

Ghosal:1998:CCM

[GC98] Subhashis Ghosal and Tapas K. Chandra. Complete convergence of martingale arrays. Journal of Theoretical Probability, 11(3):621–631, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022646429754.

Geiss:2001:CPV

[Gei01] Stefan Geiss. Contraction principles for vector valued martingales with respect to random variables having exponential tail with exponent 2 < α < ∞. Journal of Theoretical Probability, 14(1):39–59, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007864813858.

Gerl:1988:RWG

[Ger88] Peter Gerl. Random walks on graphs with a strong isoperimetric property. Journal of Theoretical Probability, 1(2):171–187, April 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01046933.

Gerhard:1992:PSD

[Ger92] W. D. Gerhard. The probabilistic solution of the Dirichlet problem for \frac{1}{2}Δ + ⟨a, ∇⟩ + b with singular coefficients. Journal of Theoretical Probability, 5(3):503–520, July 1992. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01060432.

Gonzalez-Guillen:2018:EDB

[GGPV18] C. E. González-Guillén, C. Palazuelos, and I. Villanueva. Euclidean distance between Haar orthogonal and Gaussian matrices. Journal of Theoretical Probability, 31(1):93–118, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Gao:2024:ABR

[GGX24] Fuqing Gao, Yunshi Gao, and Xianjie Xia. Asymptotic behaviors for random geometric series. Journal of Theoretical Probability, 37(3):2818–2842, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-024-01327-3.

Gregory:1995:GATa

[GH95a] John Gregory and H. R. Hughes. A generalized approximation theory for quadratic forms: Application to randomized spline type Sturm–Liouville problems. Journal of Theoretical Probability, 8(3):703–715, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02218052.

Gregory:1995:GATb

[GH95b] John Gregory and H. R. Hughes. A generalized approximation theory for quadratic forms II: Application to randomized spline type focal/conjugate point problems. Journal of Theoretical Probability, 8(4):963–971, October 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02410119.

Grama:2006:AEP

[GH06] I. G. Grama and E. Haeusler. An asymptotic expansion for probabilities of moderate deviations for multivariate martingales. Journal of Theoretical Probability, 19(1):1–44, January 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-006-0001-x.

Gerencser:2019:IMR

[GH19] Balázs Gerencsér and Julien M. Hendrickx. Improved mixing rates of directed cycles by added connection. Journal of Theoretical Probability, 32(2):684–701, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Giuliano:2021:ELL

[GH21] Rita Giuliano and Milto Hadjikyriakou. On exact laws of large numbers for Oppenheim expansions with infinite mean. Journal of Theoretical Probability, 34(3):1579–1606, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01010-3.

Gong:2018:LBG

[GHL18] Ruoting Gong, Christian Houdré, and Jüri Lember. Lower bounds on the generalized central moments of the optimal alignments score of random sequences. Journal of Theoretical Probability, 31(2):643–683, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Gallegos-Herrada:2024:EGE

[GHLR24] Marco A. Gallegos-Herrada, David Ledvinka, and Jeffrey S. Rosenthal. Equivalences of geometric ergodicity of Markov chains. Journal of Theoretical Probability, 37(2):1230–1256, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01240-1.

Gao:2004:ESB

[GHLT04] F. Gao, J. Hannig, T.-Y. Lee, and F. Torcaso. Exact L^{2} small balls of Gaussian processes. Journal of Theoretical Probability, 17(2):503–520, April 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000020705.28185.4c.

Ghosh:2020:TVC

[Gho20] Subhajit Ghosh. Total variation cutoff for the transpose top-2 with random shuffle. Journal of Theoretical Probability, 33(4): 1832–1854, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-019-00945-6.

Gu:2020:BMI

[GHS20] Yinzheng Gu, Takahiro Hasebe, and Paul Skoufranis. Bimonotonic independence for pairs of algebras. Journal of Theoretical Probability, 33(1):533–566, March 2020. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Guo:2019:HOD

[GHX19] Jingjun Guo, Yaozhong Hu, and Yanping Xiao. Higher-order derivative of intersection local time for two independent fractional Brownian motions. Journal of Theoretical Probability, 32 (3):1190–1201, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Giraudo:2017:HWI

[Gir17] Davide Giraudo. Holderian weak invariance principle for stationary mixing sequences. Journal of Theoretical Probability, 30 (1):196–211, March 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0633-9.

Gairing:2020:HIE

[GIST20] Jan Gairing, Peter Imkeller, Radomyra Shevchenko, and Ciprian Tudor. Hurst index estimation in stochastic differential equations driven by fractional Brownian motion. Journal of Theoretical Probability, 33(3):1691–1714, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Gittenberger:2003:NSS

[Git03] Bernhard Gittenberger. A note on “State Spaces of the Snake and Its Tour-Convergence of the Discrete Snake” by J.-F. Marckert and A. Mokkadem. Journal of Theoretical Probability, 16(4): 1063–1067, October 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000012006.91251.e8. See

[MM03].

Goodman:1988:RCI

[GK88] V. Goodman and J. Kuelbs. Rates of convergence for increments of Brownian motion. Journal of Theoretical Probability, 1(1):27–63, January 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01076287.

Goodman:1991:RCS

[GK91] Victor Goodman and J. Kuelbs. Rates of clustering in Strassen’s LIL for Brownian motion. Journal of Theoretical Probability, 4 (2):285–309, April 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01258738.

Grigorescu:2002:BMF

[GK02] Ilie Grigorescu and Min Kang. Brownian motion on the figure eight. Journal of Theoretical Probability, 15(3):817–844, July 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1016232201962.

Grigorescu:2003:PCI

[GK03] Ilie Grigorescu and Min Kang. Path collapse for an inhomogeneous random walk. Journal of Theoretical Probability, 16(1): 147–159, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022282505543.

Gupta:2023:FPP

[GK23] Neha Gupta and Arun Kumar. Fractional Poisson processes of order k and beyond. Journal of Theoretical Probability, 36(4): 2165–2191, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01268-3.

Ger:2000:CLC

[GKM00] Roman Ger, Michael Keane, and Jolanta K. Misiewicz. On convolutions and linear combinations of pseudo-isotropic distributions. Journal of Theoretical Probability, 13(4):977–995, October 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007809907072.

Gaines:2014:NRP

[GKR14] G. Gaines, K. Kaphle, and F. Ruymgaart. A note on random perturbations of a multiple eigenvalue of a Hermitian operator. Journal of Theoretical Probability, 27(4):1112–1123, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-013-0482-3.

Gun:2015:MAM

[GKS15] Onur Gün, Wolfgang König, and Ozren Sekulović. Moment asymptotics for multitype branching random walks in random environment. Journal of Theoretical Probability, 28(4):1726– 1742, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0551-2.

Gatzouras:1994:SSA

[GL94] Dimitrios Gatzouras and Steven P. Lalley. Statistically selfaffine sets: Hausdorff and box dimensions. Journal of Theoretical Probability, 7(2):437–468, April 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214277.

Gorn:1999:CLC

[GL99] Natalia Gorn and Mikhail Lifshits. Chung’s law and the Csáki function. Journal of Theoretical Probability, 12(2):399–420, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1021678111442.

Gao:2006:LLC

[GL06] Fuchang Gao and Wenbo V. Li. Logarithmic level comparison for small deviation probabilities. Journal of Theoretical Probability, 19(3):535–556, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-006-0026-1.

Gao:2007:LLCa

[GL07a] Fuchang Gao and Wenbo V. Li. Logarithmic level comparison for small deviation probabilities. Journal of Theoretical Probability, 20(1):1–23, March 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-006-0027-0.

Gao:2007:LLCb

[GL07b] Fuchang Gao and Wenbo V. Li. Logarithmic level comparison for small deviation probabilities. Journal of Theoretical Probability, 20(4):1101, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0116-8.

Gehringer:2022:FLT

[GL22] Johann Gehringer and Xue-Mei Li. Functional limit theorems for the fractional Ornstein–Uhlenbeck process. Journal of Theoretical Probability, 35(1):426–456, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-020-01044-7.

Graczyk:2005:HGS

[GLM05] P. Graczyk, G. Letac, and H. Massam. The hyperoctahedral group, symmetric group representations and the moments of the real Wishart distribution. Journal of Theoretical Probability, 18 (1):1–42, January 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-004-0579-9.

Glover:1991:SGM

[Glo91] Joseph Glover. Symmetry groups of Markov processes and the diagonal principle. Journal of Theoretical Probability, 4 (2):417–440, April 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01258746.

Glowacki:2013:CSC

[Glo13] Pawel Glowacki. Convergence of semigroups of complex measures on a Lie group. Journal of Theoretical Probability, 26(1):58–71, March 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10959-011-0385-0.pdf.

Graf:2003:FQS

[GLP03] Siegfried Graf, Harald Luschgy, and Gilles Pagès. Functional quantization and small ball probabilities for Gaussian processes. Journal of Theoretical Probability, 16(4):1047–1062, October 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOTP.0000012005.32667.9d.

Gouet:2007:CMR

[GLS07] Raúl Gouet, F. Javier López, and Gerardo Sanz. A characteristic martingale related to the counting process of records. Journal of Theoretical Probability, 20(3):443–455, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-007-0065-2.

Grahovac:2020:MBI

[GLT20] Danijel Grahovac, Nikolai N. Leonenko, and Murad S. Taqqu. The multifaceted behavior of integrated supOU processes: The infinite variance case. Journal of Theoretical Probability, 33(4): 1801–1831, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-019-00935-8.

Gluck:1999:FHT

[Glu99] David Gluck. First hitting times for some random walks on finite groups. Journal of Theoretical Probability, 12(3):739–755, July 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021679932572.

Gabriel:1991:RPS

[GM91a] J. P. Gabriel and P. Milasevic. A remark on probability spaces. Journal of Theoretical Probability, 4(1):191–195, January 1991.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01047001.

Gong:1991:SPI

[GM91b] J. Gong and A. Mukherjea. Solution of a problem on the identification of parameters by the distribution of the maximum random variable: a multivariate normal case. Journal of Theoretical Probability, 4(4):783–790, October 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01259555.

Gine:1998:LSN

[GM98] Evarist Giné and David M. Mason. On the LIL for selfnormalized sums of IID random variables. Journal of Theoretical Probability, 11(2):351–370, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022675620729.

Gine:2007:LIL

[GM07] Evarist Giné and David M. Mason. Laws of the iterated logarithm for the local U-statistic process. Journal of Theoretical Probability, 20(3):457–485, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-007-0067-0.

Gilch:2011:RWD

[GM11] Lorenz A. Gilch and Sebastian Müller. Random walks on directed covers of graphs. Journal of Theoretical Probability, 24 (1):118–149, March 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0256-0.

Gloter:2018:BSB

[GM18] Arnaud Gloter and Miguel Martinez. Bouncing skew Brownian motions. Journal of Theoretical Probability, 31(1):319–363, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Gut:2016:MSP

[GML16] Allan Gut and Anders Martin-Löf. A maxtrimmed St. Petersburg game. Journal of Theoretical Probability, 29(1):277–291, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-014-0563-y.

Gantert:2010:SBR

[GMPV10] Nina Gantert, Sebastian Müller, Serguei Popov, and Marina Vachkovskaia. Survival of branching random walks in random environment. Journal of Theoretical Probability, 23(4):1002– 1014, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0227-5.

Gilat:2022:NME

[GMS22] David Gilat, Isaac Meilijson, and Laura Sacerdote. A note on the maximal expected local time of L_{2}-bounded martingales. Journal of Theoretical Probability, 35(3):1952–1955, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01118-0.

Gaither:2017:VSD

[GMW17] Jeffrey Gaither, Hosam Mahmoud, and Mark Daniel Ward. On the variety of shapes in digital trees. Journal of Theoretical Probability, 30(4):1225–1254, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Gong:2012:HTD

[GMZ12] Yu Gong, Yong-Hua Mao, and Chi Zhang. Hitting time distributions for denumerable birth and death processes. Journal of Theoretical Probability, 25(4):950–980, December 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-012-0436-1.

Gantert:2024:CTR

[GNS24] Nina Gantert, Evita Nestoridi, and Dominik Schmid. Cutoff on trees is rare. Journal of Theoretical Probability, 37(2): 1417–1444, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01274-5.

Gotze:2020:LSL

[GNT20] F. Götze, A. Naumov, and A. Tikhomirov. Local semicircle law under fourth moment condition. Journal of Theoretical Probability, 33(3):1327–1362, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Gotze:2017:AAS

[GNU17] F. Götze, A. Naumov, and V. Ulyanov. Asymptotic analysis of symmetric functions. Journal of Theoretical Probability, 30(3): 876–897, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Godefroy:2004:MBP

[God04] Laurent Godefroy. Martin boundary of polynomial random walks on the d-dimensional lattice of nonnegative integers. Journal of Theoretical Probability, 17(1):111–129, January 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000020477.86897.82.

Goldstein:1988:MNF

[Gol88] Larry Goldstein. Minimizing noisy functionals in Hilbert space: An extension of the Kiefer–Wolfowitz procedure. Journal of Theoretical Probability, 1(2):189–204, April 1988. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01046934.

Gess:2020:DBS

[GOT20] Benjamin Gess, Cheng Ouyang, and Samy Tindel. Density bounds for solutions to differential equations driven by Gaussian rough paths. Journal of Theoretical Probability, 33(2):611– 648, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Geng:2023:PLE

[GOT23] Xi Geng, Cheng Ouyang, and Samy Tindel. Precise local estimates for differential equations driven by fractional Brownian motion: Elliptic case. Journal of Theoretical Probability, 36(3): 1341–1367, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01208-7.

Graversen:1998:OSM

[GP98] S. E. Graversen and G. Peskir. Optimal stopping and maximal inequalities for linear diffusions. Journal of Theoretical Probability, 11(1):259–277, January 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1021659328029.

Guillotin-Plantard:2001:DRW

[GP01] N. Guillotin-Plantard. Dynamic Z_{d}-random walks in a random scenery: A strong law of large numbers. Journal of Theoretical Probability, 14(1):241–260, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1007885418401.

Guillotin-Plantard:2011:LTR

[GP11] Nadine Guillotin-Plantard. On the local time of random walks associated with Gegenbauer polynomials. Journal of Theoretical Probability, 24(4):1157–1169, December 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0297-4.

Gibson:2015:RDW

[GP15] Lee R. Gibson and Melanie Pivarski. The rate of decay of the Wiener sausage in local Dirichlet space. Journal of Theoretical Probability, 28(4):1253–1270, December 2015. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
014-0553-0.

Guillotin-Plantard:2006:DRW

[GPS06] Nadine Guillotin-Plantard and René Schott. Dynamic random walks on Heisenberg groups. Journal of Theoretical Probability, 19(2):377–395, June 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-006-0021-6.

Gomes:2024:ANO

[GPS24] Pablo A. Gomes, Alan Pereira, and Remy Sanchis. Anisotropic non-oriented bond percolation in high dimensions. Journal of Theoretical Probability, 37(1):121–132, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01254-9.

Guo:2015:MMT

[GPT15] Hongxin Guo, Robert Philipowski, and Anton Thalmaier. Martingales on manifolds with time–dependent connection. Journal of Theoretical Probability, 28(3):1038–1062, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-013-0536-6.

Glover:1988:NMP

[GR88a] Joseph Glover and Murali Rao. Nonsymmetric Markov processes and hypothesis (h). Journal of Theoretical Probability, 1(4): 371–380, October 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01048726.

Glover:1988:SMP

[GR88b] Joseph Glover and Murali Rao. Symmetrizations of Markov processes. Journal of Theoretical Probability, 1(3):305–325, July 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01246632.

Gangopadhyay:1997:SFA

[GR97] Sreela Gangopadhyay and B. V. Rao. Some finitely additive probability: Random walks. Journal of Theoretical Probability, 10(3):643–657, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022649627106.

Gangopadhyay:1998:SPN

[GR98] Sreela Gangopadhyay and B. V. Rao. Strategic purely nonatomic random walks. Journal of Theoretical Probability, 11(2):409–415, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022631805708.

Goldstein:2005:DTO

[GR05] Larry Goldstein and Gesine Reinert. Distributional transformations, orthogonal polynomials, and Stein characterizations. Journal of Theoretical Probability, 18(1):237–260, January 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-004-2602-6.

Goldstein:2007:FBI

[GR07] Larry Goldstein and Yosef Rinott. Functional BKR inequalities, and their duals, with applications. Journal of Theoretical Probability, 20(2):275–293, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0068-z.

Gamboa:2010:CMR

[GR10] F. Gamboa and A. Rouault. Canonical moments and random spectral measures. Journal of Theoretical Probability, 23 (4):1015–1038, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-009-0239-1. See er-

ratum [GR17].

Gamboa:2017:ECM

[GR17] F. Gamboa and A. Rouault. Erratum to: Canonical Moments and Random Spectral Measures. Journal of Theoretical Probability, 30(2):701–702, June 2017. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/content/pdf/10.1007/s10959-
015-0653-5.pdf. See [GR10].

Gerencser:2023:ECM

[GR23] Balázs Gerencsér and Miklós Rásonyi. On the ergodicity of certain Markov chains in random environments. Journal of Theoretical Probability, 36(4):2093–2125, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01256-7.

Graczyk:1994:CTS

[Gra94] Piotr Graczyk. Cramér theorem on symmetric spaces of noncompact type. Journal of Theoretical Probability, 7(3):609–613, July 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02213571.

Graczyk:1999:FTS

[Gra99] Piotr Graczyk. Factorization theorems on symmetric spaces of noncompact type. Journal of Theoretical Probability, 12 (2):375–383, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021674010533.

Graham:2012:SVD

[Gra12] B. T. Graham. Sublinear variance for directed last-passage percolation. Journal of Theoretical Probability, 25(3):687– 702, September 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0315-6.

Grabchak:2015:ILM

[Gra15] Michael Grabchak. Inversions of Lévy measures and the relation between long and short time behavior of Lévy processes. Journal of Theoretical Probability, 28(1):184–197, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-012-0476-6.

Grill:1992:ERC

[Gri92a] Karl Grill. Exact rate of convergence in Strassen’s law of the iterated logarithm. Journal of Theoretical Probability, 5(1):197–204, January 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01046785.

Grimmett:1992:WCU

[Gri92b] Geoffrey Grimmett. Weak convergence using higher-order cumulants. Journal of Theoretical Probability, 5(4):767–773, October 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01058728.

Griffin:1997:NRS

[Gri97] P. S. Griffin. A note on random sums with nonrandom centering. Journal of Theoretical Probability, 10(2):333–348, April 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022608415953.

Griffin:2010:PEB

[Gri10] Philip S. Griffin. Pruitt’s estimates in Banach space. Journal of Theoretical Probability, 23(4):1092–1109, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-009-0250-6.

Gribinski:2024:TSV

[Gri24] Aurelien Gribinski. A theory of singular values for finite free probability. Journal of Theoretical Probability, 37(2):1257– 1298, June 2024. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-023-01295-0.

Grundmann:2014:MFC

[Gru14] Waldemar Grundmann. Moment functions and Central Limit Theorem for Jacobi hypergroups on [0, ∞[. Journal of Theoretical Probability, 27(1):278–300, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-012-0465-9.

Grygierek:2021:MNA

[Gry21] Jens Grygierek. Multivariate normal approximation for functionals of random polytopes. Journal of Theoretical Probability, 34(2):897–922, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-00985-3.

Getoor:1994:LTR

[GS94] R. K. Getoor and M. J. Sharpe. Local times on rays for a class of planar Lévy processes. Journal of Theoretical Probability, 7 (4):799–811, October 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214373.

Goswami:1995:TSP

[GS95] A. Goswami and Arindam Sengupta. Time-space polynomial martingales generated by a discrete-time martingale. Journal of Theoretical Probability, 8(2):417–432, April 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02212886.

Gruet:1996:OTB

[GS96] J.-C. Gruet and Z. Shi. The occupation time of Brownian motion in a ball. Journal of Theoretical Probability, 9(2):429–445, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02214658.

Gut:2010:CSR

[GS10] Allan Gut and Ulrich Stadtmüller. Cesàro summation for random fields. Journal of Theoretical Probability, 23(3):715– 728, September 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0223-9.

Gut:2011:LRF

[GS11] Allan Gut and Ulrich Stadtmüller. On the LSL for random fields. Journal of Theoretical Probability, 24(2):422–449, June 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-009-0265-z.

Gaunt:2024:IBS

[GS24] Robert E. Gaunt and Heather Sutcliffe. Improved bounds in Stein’s method for functions of multivariate normal random vectors. Journal of Theoretical Probability, 37(1):642–670, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01257-6.

Gotze:2021:CIB

[GSS21] Friedrich Götze, Holger Sambale, and Arthur Sinulis. Concentration inequalities for bounded functionals via log-Sobolev-type inequalities. Journal of Theoretical Probability, 34(3):1623– 1652, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01016-x.

Goswami:2024:SRC

[GSY24] Anindya Goswami, Subhamay Saha, and Ravishankar Kapildev Yadav. Semimartingale representation of a class of semi-Markov dynamics. Journal of Theoretical Probability, 37(1):489–510, March 2024. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-023-01259-4.

Giraitis:2001:FNC

[GT01] Liudas Giraitis and Murad S. Taqqu. Functional non-central and Central Limit Theorems for bivariate Appell polynomials. Journal of Theoretical Probability, 14(2):393–426, April 2001.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1011111730227.

Gallardo:2002:RTS

[GT02a] Léonard Gallardo and Khalifa Trimèche. Renewal theorems for singular differential operators. Journal of Theoretical Probability, 15(1):161–205, January 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1013895502747.

Gotze:2002:ADQ

[GT02b] F. Götze and A. Tikhomirov. Asymptotic distribution of quadratic forms and applications. Journal of Theoretical Probability, 15(2):423–475, April 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014867011101.

Gotze:2005:AEN

[GT05] F. Götze and A. N. Tikhomirov. Asymptotic expansions in noncentral limit theorems for quadratic forms. Journal of Theoretical Probability, 18(4):757–811, October 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-005-7525-3.

Geluk:2009:ATP

[GT09] Jaap Geluk and Qihe Tang. Asymptotic tail probabilities of sums of dependent subexponential random variables. Journal of Theoretical Probability, 22(4):871–882, December 2009. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
008-0159-5.

Gim:2018:RCG

[GT18] Minjung Gim and Gerald Trutnau. Recurrence criteria for generalized Dirichlet forms. Journal of Theoretical Probability, 31(4): 2129–2166, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Guerbaz:2009:MCL

[Gue09] Raby Guerbaz. On moduli of continuity for local times of fractional stable processes. Journal of Theoretical Probability, 22(4): 934–954, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0214-x.

Gut:2004:EKF

[Gut04] A. Gut. An extension of the Kolmogorov–Feller weak law of large numbers with an application to the St. Petersburg game. Journal of Theoretical Probability, 17(3):769–779, July 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000040299.15416.0c.

Grothaus:2018:SFP

[GV18] Martin Grothaus and Robert Voßhall. Strong Feller property of sticky reflected distorted Brownian motion. Journal of Theoretical Probability, 31(2):827–852, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Gordin:2002:ASC

[GW02] Mikhail Gordin and Michel Weber. On the almost sure Central Limit Theorem for a class of Z^{d}-actions. Journal of Theoretical Probability, 15(2):477–501, April 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1015014927939.

Garban:2020:CFI

[GW20] Christophe Garban and Hao Wu. On the convergence of FK– Ising percolation to SLE(16/3, (16/3)−6). Journal of Theoretical Probability, 33(2):828–865, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Gine:1996:ILD

[GZ96] Evarist Giné and Cun-Hui Zhang. On integrability in the LIL for degenerate U-statistics. Journal of Theoretical Probability, 9(2):385–412, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214656.

Gotze:1998:ERD

[GZ98] F. Götze and A. Yu. Zaitsev. Estimates for the rapid decay of concentration functions of n-fold convolutions. Journal of Theoretical Probability, 11(3):715–731, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022654631571.

Goldsborough:2024:SPM

[GZ24] Antoine Goldsborough and Stefanie Zbinden. Some properties of Markov chains on the free group F_{2}. Journal of Theoretical Probability, 37(3):2330–2351, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-023-01294-1.

Guo:2024:DMS

[GZM24] Jingjun Guo, Cuiyun Zhang, and Aiqin Ma. Derivative of multiple self-intersection local time for fractional Brownian motion. Journal of Theoretical Probability, 37(1):623–641, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01265-6.

Hajri:2019:CSI

[HA19] Hatem Hajri and Marc Arnaudon. On a coupling of solutions to the interface stochastic differential equation on a star graph. Journal of Theoretical Probability, 32(1):90–105, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Haas:2007:FPI

[Haa07] Bénédicte Haas. Fragmentation processes with an initial mass converging to infinity. Journal of Theoretical Probability, 20(4): 721–758, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0120-z.

Haggstrom:2013:TBB

[Häg13] Olle Häggström. Two badly behaved percolation processes on a nonunimodular graph. Journal of Theoretical Probability, 26(4): 1165–1180, December 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0397-9.

Hamana:2001:AMG

[Ham01] Yuji Hamana. Asymptotics of the moment generating function for the range of random walks. Journal of Theoretical Probability, 14(1):189–197, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1007829300654.

Handjani:1997:STC

[Han97] Shirin J. Handjani. Survival of threshold contact processes. Journal of Theoretical Probability, 10(3):737–746, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022609912993.

Handjani:2003:IVM

[Han03] Shirin J. Handjani. Inhomogeneous voter models in one dimension. Journal of Theoretical Probability, 16(2):325–338, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1023562309007.

Han:2018:EIS

[Han18] Fang Han. An exponential inequality for U-statistics under mixing conditions. Journal of Theoretical Probability, 31(1):556–578, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Harris:2001:NPT

[Har01] Matthew Harris. Nontrivial phase transition in a continuum mirror model. Journal of Theoretical Probability, 14(2):299–317, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1011185511572.

Hariya:2022:IRH

[Har22] Yuu Hariya. Integral representations for the Hartman–Watson density. Journal of Theoretical Probability, 35(1):209–230, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01067-0.

Hassenforder:1988:EKT

[Has88] Claudie Hassenforder. An extension of Knight’s theorem on Cauchy distribution. Journal of Theoretical Probability, 1(2): 205–209, April 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046935.

Hashorva:2005:ABM

[Has05] Enkelejd Hashorva. Asymptotics and bounds for multivariate Gaussian tails. Journal of Theoretical Probability, 18(1):79–97, January 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-004-2577-3.

Hashorva:2010:BNC

[Has10] Enkelejd Hashorva. Boundary non-crossings of Brownian pillow. Journal of Theoretical Probability, 23(1):193–208, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-008-0191-5.

Hasebe:2012:ACF

[Has12] Takahiro Hasebe. Analytic continuations of Fourier and Stieltjes transforms and generalized moments of probability measures. Journal of Theoretical Probability, 25(3):756–770, September 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-011-0344-9.

He:2011:RLV

[He11] Hui He. Rescaled Lotka–Volterra models converge to superstable processes. Journal of Theoretical Probability, 24(3):688– 728, September 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0273-z.

He:2017:CGW

[He17] Xin He. Conditioning Galton–Watson trees on large maximal outdegree. Journal of Theoretical Probability, 30(3):842–851, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

He:2022:LCC

[He22] Xin He. Local convergence of critical random trees and continuous-state branching processes. Journal of Theoretical Probability, 35(2):685–713, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01074-9.

Heck:1996:PRA

[Hec96] Matthias Heck. A perturbation result for the asymptotic behavior of matrix powers. Journal of Theoretical Probability, 9 (3):647–658, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214080.

Heck:1999:PLD

[Hec99] Matthias K. Heck. Principles of large deviations for the empirical processes of the Ornstein–Uhlenbeck process. Journal of Theoretical Probability, 12(1):147–179, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021700811752.

Heinkel:1990:IDL

[Hei90] Bernard Heinkel. An infinite-dimensional law of large numbers in Cesaro’s sense. Journal of Theoretical Probability, 3(4):533–546, October 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01046094.

Heinrich:1996:ZOL

[Hei96] P. Heinrich. Zero–one laws for polynomials in Gaussian random variables: A simple proof. Journal of Theoretical Probability, 9(4):1019–1027, October 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214262.

Heinich:2006:MPB

[Hei06] Henri Heinich. The Monge problem in Banach spaces. Journal of Theoretical Probability, 19(2):509–534, June 2006. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
006-0017-2.

Heilman:2021:LCN

[Hei21] Steven Heilman. Low correlation noise stability of symmetric sets. Journal of Theoretical Probability, 34(4):2192–2240, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01031-y.

Herriger:2013:CRS

[Her13] Philip Herriger. Countable random sets: Uniqueness in law and constructiveness. Journal of Theoretical Probability, 26(3): 781–802, September 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0432-5.

Hermon:2020:SCC

[Her20] Jonathan Hermon. A spectral characterization for concentration of the cover time. Journal of Theoretical Probability, 33(4): 2167–2184, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-019-00946-5.

Hennion:2008:SLP

[HH08] H. Hennion and L. Hervé. Stable laws and products of positive random matrices. Journal of Theoretical Probability, 21(4): 966–981, December 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-008-0153-y.

Harremoes:2009:CMC

[HH09] Peter Harremoës and Klaus Kähler Holst. Convergence of Markov chains in information divergence. Journal of Theoretical Probability, 22(1):186–202, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-007-0133-7.

Hechner:2010:MZL

[HH10] Florian Hechner and Bernard Heinkel. The Marcinkiewicz– Zygmund LLN in Banach spaces: A generalized martingale approach. Journal of Theoretical Probability, 23(2):509–522, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-009-0212-z.

Hahn:1989:OSL

[HHV89] Marjorie G. Hahn, William N. Hudson, and Jerry A. Veeh. Operator stable laws: Series representations and domains of normal attraction. Journal of Theoretical Probability, 2(1):3–35, January 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01048266.

Hildebrand:2001:RLR

[Hil01] Martin Hildebrand. Random lazy random walks on arbitrary finite groups. Journal of Theoretical Probability, 14(4):1019–1034, October 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1012529020690.

Hildebrand:2022:MSV

[Hil22] Martin Hildebrand. A multiplicatively symmetrized version of the Chung–Diaconis–Graham random process. Journal of Theoretical Probability, 35(2):1216–1225, June 2022. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-021-
01088-3.

Hirth:2003:ERO

[Hir03] Ulrich Hirth. Exchangeable random ordered trees by positive definite functions. Journal of Theoretical Probability, 16 (2):339–344, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023514425845.

Howe:1989:AMP

[HJ89] Eric C. Howe and Charles R. Johnson. Aggregation of Markov processes: Axiomatization. Journal of Theoretical Probability, 2(2):201–208, April 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01053410.

Handjani:1996:RCS

[HJ96a] Shirin Handjani and Douglas Jungreis. Rate of convergence for shuffling cards by transpositions. Journal of Theoretical Probability, 9(4):983–993, October 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214260.

Hofmann:1996:SAS

[HJ96b] Karl H. Hofmann and Zbigniew J. Jurek. Some analytical semigroups occurring in probability theory. Journal of Theoretical Probability, 9(3):745–763, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214085.

Hoffmann-Jorgensen:1999:AL

[HJ99] J. Hoffmann-Jørgensen. The arcsine law. Journal of Theoretical Probability, 12(1):131–145, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1021748727681.

Hambly:2002:AOD

[HJ02] B. M. Hambly and O. D. Jones. Asymptotically one-dimensional diffusion on the Sierpiński gasket and multi-type branching processes with varying environment. Journal of Theoretical Probability, 15(2):285–322, April 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014858709284.

Hoffmann-Jorgensen:2007:SDR

[HJKPS07] Jørgen Hoffmann-Jørgensen, Abram M. Kagan, Loren D. Pitt, and Lawrence A. Shepp. Strong decomposition of random variables. Journal of Theoretical Probability, 20(2):211–220, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-007-0061-6.

Hoffmann-Jorgensen:2016:KTL

[HJMLX16] Jørgen Hoffmann-Jørgensen, Yu Miao, Xiao Chun Li, and Shou Fang Xu. Kolmogorov type law of the logarithm for arrays. Journal of Theoretical Probability, 29(1):32–47, March 2016. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
014-0574-8.

Harnett:2019:SSI

[HJN19] Daniel Harnett, Arturo Jaramillo, and David Nualart. Symmetric stochastic integrals with respect to a class of self-similar Gaussian processes. Journal of Theoretical Probability, 32(3): 1105–1144, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Hoffmann-Jorgensen:1997:LLN

[HJST97] J. Hoffmann-Jørgensen, K-L. Su, and R. L. Taylor. The law of large numbers and the Ito–Nisio theorem for vector valued random fields. Journal of Theoretical Probability, 10(1):145–183, January 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022650616646.

Houdre:1995:VIF

[HK95] Christian Houdré and Abram Kagan. Variance inequalities for functions of Gaussian variables. Journal of Theoretical Probability, 8(1):23–30, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02213451.

Hahn:1998:OUL

[HK98] Marjorie G. Hahn and Michael J. Klass. Optimal upper and lower bounds for the upper tails of compound Poisson processes. Journal of Theoretical Probability, 11(2):535–559, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022696125272.

Hafouta:2016:NLL

[HK16] Yeor Hafouta and Yuri Kifer. A nonconventional local limit theorem. Journal of Theoretical Probability, 29(4):1524–1553, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-015-0625-9.

Hoshino:2019:INW

[HK19] Kiyoiki Hoshino and Tetsuya Kazumi. On the identification of noncausal Wiener functionals from the stochastic Fourier coefficients. Journal of Theoretical Probability, 32(4):1973–1989, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Hayashi:2013:LHC

[HKHY13] Masafumi Hayashi, Arturo Kohatsu-Higa, and Gô Yûki. Local Hölder continuity property of the densities of solutions of SDEs with singular coefficients. Journal of Theoretical Probability, 26(4):1117–1134, December 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-012-0430-7.

Hahn:2012:SDT

[HKU12] Marjorie Hahn, Kei Kobayashi, and Sabir Umarov. SDEs driven by a time-changed Lévy process and their associated time-fractional order pseudo-differential equations. Journal of Theoretical Probability, 25(1):262–279, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0289-4.

Hahn:1990:ADM

[HKW90] Marjorie G. Hahn, Jim Kuelbs, and Daniel C. Weiner. The asymptotic distribution of magnitude-Winsorized sums via selfnormalization. Journal of Theoretical Probability, 3(1):137–168, January 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01063331.

Hochstattler:2016:SLM

[HKW16] Winfried Hochstättler, Werner Kirsch, and Simone Warzel. Semicircle law for a matrix ensemble with dependent entries. Journal of Theoretical Probability, 29(3):1047–1068, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-015-0602-3.

Hinrichs:2020:POD

[HKW20] Günter Hinrichs, Martin Kolb, and Vitali Wachtel. Persistence of one-dimensional AR(1)-sequences. Journal of Theoretical Probability, 33(1):65–102, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Hassairi:2001:REF

[HL01] A. Hassairi and S. Lajmi. Riesz exponential families on symmetric cones. Journal of Theoretical Probability, 14(4):927–948, October 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1012592618872.

Heicklen:2003:CIS

[HL03] Deborah Heicklen and Russell Lyons. Change intolerance in spanning forests. Journal of Theoretical Probability, 16(1):47–58, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022222319655.

Hassairi:2004:CRE

[HL04] A. Hassairi and S. Lajmi. Classification of Riesz exponential families on a symmetric cone by invariance properties. Journal of Theoretical Probability, 17(3):521–539, July 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000040287.87081.ba.

Hu:2014:IUE

[HL14] Mingshang Hu and Xiaojuan Li. Independence under the Gexpectation framework. Journal of Theoretical Probability, 27(3): 1011–1020, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0471-y.

Han:2020:MBL

[HL20] Fang Han and Yicheng Li. Moment bounds for large autocovariance matrices under dependence. Journal of Theoretical Probability, 33(3):1445–1492, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Harang:2022:RLT

[HL22] Fabian A. Harang and Chengcheng Ling. Regularity of local times associated with Volterra–Lévy processes and pathwise regularization of stochastic differential equations. Journal of Theoretical Probability, 35(3):1706–1735, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01114-4.

He:2018:CSB

[HLX18] Hui He, Zenghu Li, and Wei Xu. Continuous-state branching processes in Lévy random environments. Journal of Theoretical Probability, 31(4):1952–1974, December 2018. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Hassairi:2008:CRP

[HLZ08] A. Hassairi, S. Lajmi, and R. Zine. A characterization of the Riesz probability distribution. Journal of Theoretical Probability, 21(4):773–790, December 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0142-1.

He:2013:ITT

[HLZ13] Hui He, Zenghu Li, and Xiaowen Zhou. An integral test on time-dependent local extinction for super-coalescing Brownian motion with Lebesgue initial measure. Journal of Theoretical Probability, 26(1):31–45, March 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-011-0372-5.

Hognas:1995:MRW

[HM95] Göran Högnäs and Arunava Mukherjea. A mixed random walk on nonnegative matrices: A law of large numbers. Journal of Theoretical Probability, 8(4):973–990, October 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02410120.

Hognas:2003:MHG

[HM03] Göran Högnäs and Arunava Mukherjea. Maximal homomorphic group image and convergence of convolution sequences on a semigroup. Journal of Theoretical Probability, 16(4):847–854, October 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOTP.0000011996.44195.2b.

Hammond:2005:DEE

[HM05] Christopher Hammond and Steven J. Miller. Distribution of eigenvalues for the ensemble of real symmetric Toeplitz matrices. Journal of Theoretical Probability, 18(3):537–566, July 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-005-3518-5.

Herbin:2006:SIF

[HM06] Erick Herbin and Ely Merzbach. A set-indexed fractional Brownian motion. Journal of Theoretical Probability, 19(2):337–364, June 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-006-0019-0.

Hildebrand:2008:GRV

[HM08] Martin Hildebrand and Joseph McCollum. Generating random vectors in (Z/pZ)^{d} via an affine random process. Journal of Theoretical Probability, 21(4):802–811, December 2008.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-007-0135-5.

Herbin:2009:SSS

[HM09] Erick Herbin and Ely Merzbach. Stationarity and self-similarity characterization of the set-indexed fractional Brownian motion. Journal of Theoretical Probability, 22(4):1010–1029, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-008-0180-8.

Huillet:2016:MDC

[HM16] Thierry Huillet and Servet Martínez. On Möbius duality and coarse–graining. Journal of Theoretical Probability, 29(1): 143–179, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0569-5.

Hirscher:2018:SMC

[HM18] Timo Hirscher and Anders Martinsson. Segregating Markov chains. Journal of Theoretical Probability, 31(3):1512–1538, September 2018. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10959-017-0743-7.pdf.

Hashorva:2021:BNC

[HMS21] Enkelejd Hashorva, Yuliya Mishura, and Georgiy Shevchenko. Boundary non-crossing probabilities of Gaussian processes: Sharp bounds and asymptotics. Journal of Theoretical Probability, 34(2):728–754, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-020-01002-3.

Hu:2024:GMR

[HMW24] Ying Hu, Remi Moreau, and Falei Wang. General mean reflected backward stochastic differential equations. Journal of Theoretical Probability, 37(1):877–904, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-023-01288-z.

Hino:2021:PUN

[HMY21] Masanori Hino, Kouhei Matsuura, and Misaki Yonezawa. Pathwise uniqueness and non-explosion property of Skorohod SDEs with a class of non-Lipschitz coefficients and non-smooth domains. Journal of Theoretical Probability, 34(4):2166–2191, December 2021. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-020-01036-7.

Hu:2004:LDV

[HN04] Y. Hu and H. Nyrhinen. Large deviations view points for heavytailed random walks. Journal of Theoretical Probability, 17 (3):761–768, July 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000040298.43712.e8.

Hu:2005:SPA

[HN05] Y. Hu and D. Nualart. Some processes associated with fractional Bessel processes. Journal of Theoretical Probability, 18 (2):377–397, April 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-005-3508-7.

Harnett:2015:SRF

[HN15] Daniel Harnett and David Nualart. On Simpson’s rule and fractional Brownian motion with H = 1/10. Journal of Theoretical Probability, 28(4):1651–1688, December 2015. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
014-0552-1.

Hansen:2020:NHF

[HN20] Wolfhard Hansen and Ivan Netuka. Nearly hyperharmonic functions are infima of excessive functions. Journal of Theoretical Probability, 33(3):1613–1629, September 2020. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Hu:2014:VDS

[HNS14] Yaozhong Hu, David Nualart, and Jian Song. The \frac{4}{3}-variation of the derivative of the self-intersection Brownian local time and related processes. Journal of Theoretical Probability, 27(3): 789–825, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0469-5.

Ho:1995:LIL

[Ho95] Hwai-Chung Ho. The law of the iterated logarithm for noninstantaneous filters of strongly dependent Gaussian sequences. Journal of Theoretical Probability, 8(2):347–360, April 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02212883.

Hochberg:1996:CSP

[HO96] Kenneth J. Hochberg and Enzo Orsingher. Composition of stochastic processes governed by higher-order parabolic and hyperbolic equations. Journal of Theoretical Probability, 9(2): 511–532, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214661.

Hong:2003:LDS

[Hon03] Wenming Hong. Large deviations for the super–Brownian motion with super–Brownian immigration. Journal of Theoretical Probability, 16(4):899–922, October 2003. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000011999.03918.f7.

Hora:1992:QIM

[Hor92] Akihito Hora. Quasi-invariant measures on the infinite product of SU (2). Journal of Theoretical Probability, 5(1):71–100, January 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01046779.

Hormann:2007:CBA

[Hör07] Siegfried Hörmann. Critical behavior in almost sure central limit theory. Journal of Theoretical Probability, 20(3):613– 636, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0080-3.

Harel:1994:LIL

[HP94] Michel Harel and Madan L. Puri. Law of the iterated logarithm for perturbed empirical distribution functions evaluated at a random point for nonstationary random variables. Journal of Theoretical Probability, 7(4):831–855, October 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02214375.

Heyer:1997:CNT

[HP97] H. Heyer and G. Pap. Convergence of noncommutative triangular arrays of probability measures on a Lie group. Journal of Theoretical Probability, 10(4):1003–1052, October 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022670818425.

Herve:2013:RSP

[HP13] Loïc Hervé and Françoise Pène. On the recurrence set of planar Markov random walks. Journal of Theoretical Probability, 26 (1):169–197, March 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0414-7.

Hu:1995:LIL

[HPLVS95] Y. Hu, D. Pierre-Loti-Viaud, and Z. Shi. Laws of the iterated logarithm for iterated Wiener processes. Journal of Theoretical Probability, 8(2):303–319, April 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02212881.

Hao:2021:HHT

[HPZ21] Zimo Hao, Xuhui Peng, and Xicheng Zhang. Hörmander’s hypoelliptic theorem for nonlocal operators. Journal of Theoretical Probability, 34(4):1870–1916, December 2021. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
01020-1.

Hirth:2000:ERO

[HR00] Ulrich Hirth and Paul Ressel. Exchangeable random orders and almost uniform distributions. Journal of Theoretical Probability, 13(3):609–634, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007805925957.

Heida:2018:LDP

[HR18] Martin Heida and Matthias Röger. Large deviation principle for a stochastic Allen–Cahn equation. Journal of Theoretical Probability, 31(1):364–401, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Hildebrandt:2020:PDM

[HR20] Florian Hildebrandt and Sylvie Roelly. Pinned diffusions and Markov bridges. Journal of Theoretical Probability, 33(2):906– 917, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Hazod:1988:ALG

[HS88] W. Hazod and E. Siebert. Automorphisms on a Lie group contracting modulo a compact subgroup and applications to semistable convolution semigroups. Journal of Theoretical Probability, 1(2):211–225, April 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01046936.

Ho:1991:MTL

[HS91] Hwai-Chung Ho and Tze-Chien Sun. A mixture-type limit theorem for nonlinear functions of Gaussian sequences. Journal of Theoretical Probability, 4(2):407–415, April 1991. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01258745.

Harenbrock:1992:OSS

[HS92a] Marion Harenbrock and Norbert Schmitz. Optional sampling of submartingales with scanned index sets. Journal of Theoretical Probability, 5(2):309–326, April 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01046737.

Hwang:1992:SPM

[HS92b] Chii-Ruey Hwang and Shuenn-Jyi Sheu. Singular perturbed Markov chains and exact behaviors of simulated annealing processes. Journal of Theoretical Probability, 5(2):223–249, April 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01046734.

Hazod:1993:DPA

[HS93] W. Hazod and H. P. Scheffler. The domains of partial attraction of probabilities on groups and on vector spaces. Journal of Theoretical Probability, 6(1):175–186, January 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01046774.

Hanin:1998:DSN

[HS98a] Leonid G. Hanin and Bertram M. Schreiber. Discrete spectrum of nonstationary stochastic processes on groups. Journal of Theoretical Probability, 11(4):1111–1133, October 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022625117616.

Hu:1998:LTS

[HS98b] Yueyun Hu and Zhan Shi. The local time of simple random walk in random environment. Journal of Theoretical Probability, 11(3):765–793, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022658732480.

Hotta:2021:LRM

[HS21] Ikkei Hotta and Sebastian Schleißinger. Limits of radial multiple SLE and a Burgers–Loewner differential equation. Journal of Theoretical Probability, 34(2):755–783, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
00996-0.

Huss:2020:RSR

[HSH20] Wilfried Huss and Ecaterina Sava-Huss. Range and speed of rotor walks on trees. Journal of Theoretical Probability, 33(3): 1657–1690, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-019-00904-1.pdf.

Hsu:1997:SLG

[Hsu97] Elton P. Hsu. Stochastic local Gauss–Bonnet–Chern theorem. Journal of Theoretical Probability, 10(4):819–834, October 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022691430720.

Hoehner:2018:SAD

[HSW18] Steven D. Hoehner, Carsten Schütt, and Elisabeth M. Werner. The surface area deviation of the Euclidean ball and a polytope. Journal of Theoretical Probability, 31(1):244–267, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Han:2022:LTB

[HSZ22] Xiyue Han, Alexander Schied, and Zhenyuan Zhang. A limit theorem for Bernoulli convolutions and the Φ-variation of functions in the Takagi class. Journal of Theoretical Probability, 35(4): 2853–2878, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01157-1.

Hu:2013:SDS

[HT13] Yaozhong Hu and Samy Tindel. Smooth density for some nilpotent rough differential equations. Journal of Theoretical Probability, 26(3):722–749, September 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-011-0388-x.

Hasebe:2016:UFS

[HT16] Takahiro Hasebe and Steen Thorbjørnsen. Unimodality of the freely selfdecomposable probability laws. Journal of Theoretical Probability, 29(3):922–940, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-015-0595-y.

Harang:2024:VED

[HTW24] Fabian Harang, Samy Tindel, and Xiaohua Wang. Volterra equations driven by rough signals 3: Probabilistic construction of the Volterra rough path for fractional Brownian motions. Journal of Theoretical Probability, 37(1):307–351, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01251-y.

Hu:1997:IWC

[Hu97] Yaozhong Hu. Itô-Wiener chaos expansion with exact residual and correlation, variance inequalities. Journal of Theoretical Probability, 10(4):835–848, October 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022654314791.

Hu:1999:HPM

[Hu99] Yueyun Hu. Hausdorff and packing measures of the level sets of iterated Brownian motion. Journal of Theoretical Probability, 12(2):313–346, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021669809625.

Hu:2015:ASL

[Hu15] Yueyun Hu. The almost sure limits of the minimal position and the additive martingale in a branching random walk. Journal of Theoretical Probability, 28(2):467–487, June 2015. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
013-0494-z.

Hu:2017:LTS

[Hu17] Yueyun Hu. Local times of subdiffusive biased walks on trees. Journal of Theoretical Probability, 30(2):529–550, June 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Hu:2020:LTB

[Hu20] Wenqing Hu. On the long-time behavior of a perturbed conservative system with degeneracy. Journal of Theoretical Probability, 33(3):1266–1295, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Huang:2014:ERM

[Hua14] Xueping Huang. Escape rate of Markov chains on infinite graphs. Journal of Theoretical Probability, 27(2):634–682, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-012-0456-x.

Huang:2019:AES

[Hua19] Jiaoyang Huang. Asymptotic expansion of spherical integral. Journal of Theoretical Probability, 32(2):1051–1075, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Hueter:2000:RFT

[Hue00] Irene Hueter. Recovering a family of two-dimensional Gaussian variables from the minimum process. Journal of Theoretical Probability, 13(4):939–950, October 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1007822806163.

Hujo:2006:ARE

[Huj06] Mika Hujo. Is the approximation rate for European pay-offs in the Black–Scholes model always 1/\sqrt{n}. Journal of Theoretical Probability, 19(1):190–203, January 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-006-0008-3.

Hahn:1992:ABS

[HW92] Marjorie G. Hahn and Daniel C. Weiner. Asymptotic behavior of self-normalized trimmed sums: Nonnormal limits II. Journal of Theoretical Probability, 5(1):169–196, January 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01046784.

Hong:2007:QCS

[HZ07] Wenming Hong and Ofer Zeitouni. A quenched CLT for super– Brownian motion with random immigration. Journal of Theoretical Probability, 20(4):807–820, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0079-9.

Huang:2019:MSH

[HZ19] Xing Huang and Shao-Qin Zhang. Mild solutions and Harnack inequality for functional stochastic partial differential equations with Dini drift. Journal of Theoretical Probability, 32(1):303– 329, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Iltis:1995:SAL

[Ilt95] Michael Iltis. Sharp asymptotics of large deviations in R^{d}. Journal of Theoretical Probability, 8(3):501–522, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02218041.

Ivanoff:1996:MCS

[IM96] B. Gail Ivanoff and Ely Merzbach. A martingale characterization of the set-indexed Brownian motion. Journal of Theoretical Probability, 9(4):903–913, October 1996. CODEN JTPREO.

ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02214256.

Imomov:2024:RLT

[IM24] Azam A. Imomov and Misliddin Murtazaev. Renewed limit theorems for noncritical Galton–Watson branching systems. Journal of Theoretical Probability, 37(3):2843–2858, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-024-01330-8.

Ipsen:2020:LDG

[IMS20] Yuguang Ipsen, Ross Maller, and Soudabeh Shemehsavar. Limiting distributions of generalised Poisson–Dirichlet distributions based on negative binomial processes. Journal of Theoretical Probability, 33(4):1974–2000, December 2020. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-019-
00928-7.

Inahama:2010:STL

[Ina10] Yuzuru Inahama. A stochastic Taylor-like expansion in the rough path theory. Journal of Theoretical Probability, 23(3): 671–714, September 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0287-6.

Iafrate:2020:SRB

[IO20] F. Iafrate and E. Orsingher. Some results on the Brownian meander with drift. Journal of Theoretical Probability, 33(2):1034– 1060, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Inahama:2022:PDR

[IP22] Yuzuru Inahama and Bin Pei. Positivity of the density for rough differential equations. Journal of Theoretical Probability, 35(3): 1863–1877, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01116-2.

Ichiba:2022:PPG

[IPT22] Tomoyuki Ichiba, Guodong Pang, and Murad S. Taqqu. Path properties of a generalized fractional Brownian motion. Journal of Theoretical Probability, 35(1):550–574, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01066-1.

Ignatiouk-Robert:2008:MBK

[IR08] Irina Ignatiouk-Robert. Martin boundary of a killed random walk on a half-space. Journal of Theoretical Probability, 21 (1):35–68, March 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0100-3.

Ignatiouk-Robert:2021:HFR

[IR21] Irina Ignatiouk-Robert. Harmonic functions of random walks in a semigroup via ladder heights. Journal of Theoretical Probability, 34(1):34–80, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-019-00974-1.

Issoglio:2024:SDE

[IR24] Elena Issoglio and Francesco Russo. Stochastic differential equations with singular coefficients: The martingale problem view and the stochastic dynamics view. Journal of Theoretical Probability, 37(3):2352–2393, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-024-01325-5.

Sato:2013:IID

[iS13] Ken iti Sato. Inversions of infinitely divisible distributions and conjugates of stochastic integral mappings. Journal of Theoretical Probability, 26(4):901–931, December 2013. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
012-0420-9.

Ibragimov:2001:EEM

[ISC01] R. Ibragimov, Sh. Sharakhmetov, and A. Cecen. Exact estimates for moments of random bilinear forms. Journal of Theoretical Probability, 14(1):21–37, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1007812829787.

Ishak:1992:RTB

[Ish92] S. Ishak. Remarks on a theorem of Burkholder and Gundy. Journal of Theoretical Probability, 5(1):153–158, January 1992.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01046782.

Isozaki:2009:FIA

[Iso09] Yasuki Isozaki. Fluctuation identities applied to the hitting time of a half-line in the plane. Journal of Theoretical Probability, 22 (1):57–81, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-008-0192-4.

Sato:2013:WDI

[iSU13] Ken iti Sato and Yohei Ueda. Weak drifts of infinitely divisible distributions and their applications. Journal of Theoretical Probability, 26(3):885–898, September 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-012-0419-2.

Ito:2023:RSF

[ISY23] Yu Ito, Toru Sera, and Kouji Yano. Resolution of sigma-fields for multiparticle finite-state action evolutions with infinite past. Journal of Theoretical Probability, 36(3):1368–1399, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01219-4.

Imkeller:1999:CDE

[IW99] Peter Imkeller and Ferenc Weisz. Critical dimensions for the existence of self-intersection local times of the n-parameter Brownian motion in R^{d}. Journal of Theoretical Probability, 12(3):721–737, July 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021627815734.

Iksanov:2024:PTB

[IW24] Alexander Iksanov and Vitali Wachtel. Precise tail behaviour of some Dirichlet series. Journal of Theoretical Probability, 37(3): 2704–2737, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-024-01318-4.

Janssen:1990:DBK

[Jan90] Arnold Janssen. The distance between the Kac process and the Wiener process with applications to generalized telegraph equations. Journal of Theoretical Probability, 3(2):349–360, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01045166.

Jankunas:1999:LAN

[Jan99] Andrius Jankunas. Local asymptotic normality for linear homogeneous difference equations with non–Gaussian noise. Journal of Theoretical Probability, 12(3):675–697, July 1999. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021623714825.

Jasiulis:2010:LPR

[Jas10] Barbara H. Jasiulis. Limit property for regular and weak generalized convolution. Journal of Theoretical Probability, 23(1): 315–327, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0238-2.

Jaworski:1997:CAL

[Jaw97] Wojciech Jaworski. Contractive automorphisms of locally compact groups and the concentration function problem. Journal of Theoretical Probability, 10(4):967–989, October 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022666717516.

Jaworski:2004:EMP

[Jaw04] W. Jaworski. Ergodic and mixing probability measures on [SIN] groups. Journal of Theoretical Probability, 17(3):741–759, July 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOTP.0000040297.84097.57.

Jaworski:2007:DCP

[Jaw07] Wojciech Jaworski. Dissipation of convolution powers in a metric group. Journal of Theoretical Probability, 20(3):487– 503, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0072-3.

Jegaraj:2011:STA

[Jeg11] Terence Jegaraj. Small time asymptotics for stochastic evolution equations. Journal of Theoretical Probability, 24(3):756– 788, September 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0336-1.

Jasiulis-Goldyn:2011:UKG

[JGM11] B. H. Jasiulis-Goldyn and J. K. Misiewicz. On the uniqueness of the Kendall generalized convolution. Journal of Theoretical Probability, 24(3):746–755, September 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-010-0279-6.

Jiang:2003:CLT

[JH03] Xinxin Jiang and Marjorie G. Hahn. Central limit theorems for exchangeable random variables when limits are scale mixtures of normals. Journal of Theoretical Probability, 16(3):543–571, July 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1025612330587. See erratum [JH12].

Jiang:2012:ECL

[JH12] Xinxin Jiang and Marjorie G. Hahn. Erratum to: Central Limit Theorems for exchangeable random variables when limits are scale mixtures of normals. Journal of Theoretical Probability, 25(1):310–311, March 2012. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s10959-
012-0405-8.pdf. See [JH03].

Jiang:1999:SLI

[Jia99] Jiming Jiang. Some laws of the iterated logarithm for two parameter martingales. Journal of Theoretical Probability, 12(1):49–74, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021740425864.

Jiang:2010:EHI

[Jia10] Tiefeng Jiang. The entries of Haar-invariant matrices from the classical compact groups. Journal of Theoretical Probability, 23(4):1227–1243, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-009-0241-7.

Jia:2019:SMM

[Jia19] Chen Jia. Sharp moderate maximal inequalities for upward skipfree Markov chains. Journal of Theoretical Probability, 32(3): 1382–1398, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Jin:2017:BMS

[Jin17] Peng Jin. Brownian motion with singular time-dependent drift. Journal of Theoretical Probability, 30(4):1499–1538, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Jin:2019:LEM

[Jin19] Ke Jin. The limit of the empirical measure of the product of two independent Mallows permutations. Journal of Theoretical Probability, 32(4):1688–1728, December 2019. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Jin:2021:ULS

[Jin21] Peng Jin. Uniqueness in law for stable-like processes of variable order. Journal of Theoretical Probability, 34(2):522–552, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-00988-0.

Jirak:2017:WIP

[Jir17] Moritz Jirak. On weak invariance principles for partial sums. Journal of Theoretical Probability, 30(3):703–728, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Jiang:2017:DBD

[JK17] Jianping Jiang and Tom Kennedy. The difference between a discrete and continuous harmonic measure. Journal of Theoretical Probability, 30(4):1424–1444, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Jahnel:2021:GRP

[JK21] Benedikt Jahnel and Christof Külske. Gibbsian representation for point processes via hyperedge potentials. Journal of Theoretical Probability, 34(1):391–417, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-019-00960-7.

Jourdain:2022:SSB

[JK22] Benjamin Jourdain and Ezéchiel Kahn. Strong solutions to a beta-Wishart particle system. Journal of Theoretical Probability, 35(3):1574–1613, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01109-1.

Jing:2008:SLT

[JL08] Bing-Yi Jing and Han-Ying Liang. Strong limit theorems for weighted sums of negatively associated random variables. Journal of Theoretical Probability, 21(4):890–909, December 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-007-0128-4.

Jiang:2015:ARB

[JL15] Tiefeng Jiang and Danning Li. Approximation of rectangular beta–Laguerre ensembles and large deviations. Journal of Theoretical Probability, 28(3):804–847, September 2015. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
013-0519-7.

Jiang:2024:RAS

[JL24] Lianzi Jiang and Gechun Liang. A robust α-stable central limit theorem under sublinear expectation without integrability condition. Journal of Theoretical Probability, 37(3):2394– 2424, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01298-x.

Janson:2005:CDS

[JM05] Svante Janson and Jean-François Marckert. Convergence of discrete snakes. Journal of Theoretical Probability, 18(3):615–645, July 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-005-7252-9.

Jarczyk:2009:WGS

[JM09] W. Jarczyk and J. K. Misiewicz. On weak generalized stability and (c, d)-pseudostable random variables via functional equations. Journal of Theoretical Probability, 22(2):482–505, June 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-008-0195-1.

Jung:2015:HCO

[JM15] Paul Jung and Greg Markowsky. Hölder continuity and occupation-time formulas for fBm self-intersection local time and its derivative. Journal of Theoretical Probability, 28(1):299–312, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-012-0474-8.

Jackson:2012:DEH

[JMP12] Steven Jackson, Steven J. Miller, and Thuy Pham. Distribution of eigenvalues of highly palindromic Toeplitz matrices. Journal of Theoretical Probability, 25(2):464–495, June 2012. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0311-x.

Joffe:2004:SPB

[JMPP04] Anatole Joffe, Éric Marchand, François Perron, and Paul Popadiuk. On sums of products of Bernoulli variables and random permutations. Journal of Theoretical Probability, 17(1):285–292, January 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOTP.0000020485.34082.8c.

Jurek:1999:SLI

[JN99] Zbigniew J. Jurek and Daniel Neuenschwander. s-stable laws in insurance and finance and generalization to nilpotent Lie groups. Journal of Theoretical Probability, 12(4):1089–1107, October 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021653405990.

Jiang:2000:MLD

[JO00] Tiefeng Jiang and George L. O’Brien. The metric of large deviation convergence. Journal of Theoretical Probability, 13 (3):805–824, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007814729591.

Jarni:2021:RTS

[JO21] Imane Jarni and Youssef Ouknine. On reflection with twosided jumps. Journal of Theoretical Probability, 34(4):1811– 1830, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01024-x.

Johnson:1989:SVK

[Joh89] Dudley Paul Johnson. Some variations on Kac’s formula for Brownian motion. Journal of Theoretical Probability, 2(4): 433–436, October 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01051876.

Jolis:2006:MST

[Jol06] Maria Jolis. On a multiple Stratonovich-type integral for some Gaussian processes. Journal of Theoretical Probability, 19(1): 121–133, January 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-006-0006-5.

Jonasson:1998:PRO

[Jon98] Johan Jonasson. On positive random objects. Journal of Theoretical Probability, 11(1):81–125, January 1998. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021694808465.

Jacod:1993:RWC

[JP93] Jean Jacod and Philip Protter. A remark on the weak convergence of processes in the Skorohod topology. Journal of Theoretical Probability, 6(3):463–472, July 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01066712.

Jourdain:2022:COQ

[JP22] Benjamin Jourdain and Gilles Pagès. Convex order, quantization and monotone approximations of ARCH models. Journal of Theoretical Probability, 35(4):2480–2517, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-021-
01141-1.

Jiang:2024:SNC

[JPW24] Hui Jiang, Yajuan Pan, and Xiao Wei. Self-normalized Cramér-Type moderate deviations for explosive Vasicek model. Journal of Theoretical Probability, 37(1):228–250, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01264-7.

Jiang:2017:SRL

[JQ17] Tiefeng Jiang and Yongcheng Qi. Spectral radii of large non– Hermitian random matrices. Journal of Theoretical Probability, 30(1):326–364, March 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0634-8.

Jiang:2019:EDE

[JQ19] Tiefeng Jiang and Yongcheng Qi. Empirical distributions of eigenvalues of product ensembles. Journal of Theoretical Probability, 32(1):353–394, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Jonasson:1999:APT

[JS99] Johan Jonasson and Jeffrey E. Steif. Amenability and phase transition in the Ising model. Journal of Theoretical Probability, 12(2):549–559, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021690414168.

Johnson:2000:ECC

[JS00] Oliver Johnson and Yurii Suhov. Entropy and convergence on compact groups. Journal of Theoretical Probability, 13 (3):843–857, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007818830500.

Jarai:2021:AHD

[JS21] Antal A. Járai and Minwei Sun. Asymptotic height distribution in high-dimensional sandpiles. Journal of Theoretical Probability, 34(1):349–362, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-019-00962-5.

Jin:2023:USM

[JSS23] Xinghu Jin, Tian Shen, and Zhonggen Su. Using Stein’s method to analyze Euler–Maruyama approximations of regime-switching jump diffusion processes. Journal of Theoretical Probability, 36(3):1797–1828, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-022-01221-w.

Jorgensen:2019:RFP

[JT19] Palle Jorgensen and Feng Tian. Realizations and factorizations of positive definite kernels. Journal of Theoretical Probability, 32 (4):1925–1942, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Jockovic:2024:WTS

[JT24] Jelena Jocković and Bojana Todić. Waiting time for a small subcollection in the coupon collector problem with universal coupon. Journal of Theoretical Probability, 37(3):1945–1957, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01312-2.

Juhasz:1990:CVN

[Juh90] Ferenc Juhász. On the characteristic values of non-symmetric block random matrices. Journal of Theoretical Probability, 3 (2):199–205, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01045158.

Jung:2005:CVC

[Jun05] Paul Jung. The critical value of the contact process with added and removed edges. Journal of Theoretical Probability, 18(4): 949–955, October 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-005-7536-0.

Jung:2014:RTI

[Jun14] Paul Jung. Random-time isotropic fractional stable fields. Journal of Theoretical Probability, 27(2):618–633, June 2014. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
012-0433-4.

Jurczak:2015:UEB

[Jur15] Kamil Jurczak. A universal expectation bound on empirical projections of deformed random matrices. Journal of Theoretical Probability, 28(2):650–666, June 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-013-0517-9.

Jurczak:2018:WCE

[Jur18] Kamil Jurczak. Weak convergence of the empirical spectral distribution of high–dimensional band sample covariance matrices. Journal of Theoretical Probability, 31(3):1273–1302, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Jolis:2007:CLR

[JV07] Maria Jolis and Noèlia Viles. Continuity in law with respect to the Hurst parameter of the local time of the fractional Brownian motion. Journal of Theoretical Probability, 20(2):133–152, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-007-0054-5.

Jarai:2014:MCS

[JW14] Antal A. Járai and Nicolás Werning. Minimal configurations and sandpile measures. Journal of Theoretical Probability, 27 (1):153–167, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0446-z.

Jiang:2021:SPE

[JW21] Tiefeng Jiang and Ke Wang. Statistical properties of eigenvalues of Laplace–Beltrami operators. Journal of Theoretical Probability, 34(3):1061–1109, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-020-01061-6.

Jing:2015:CTM

[JWZ15] Bing-Yi Jing, Qiying Wang, and Wang Zhou. Cramér-type moderate deviation for studentized compound Poisson sum. Journal of Theoretical Probability, 28(4):1556–1570, December 2015.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-014-0542-3.

Jiang:2021:MDE

[JWZ21] Hui Jiang, Shaochen Wang, and Wang Zhou. Moderate deviations for extreme eigenvalues of real-valued sample covariance matrices. Journal of Theoretical Probability, 34(2):791–808, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-00999-x.

Jiang:2020:LBL

[JX20] Tiefeng Jiang and Junshan Xie. Limiting behavior of largest entry of random tensor constructed by high-dimensional data. Journal of Theoretical Probability, 33(4):2380–2400, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-019-00958-1.

Jimenez:2002:ASD

[JY02] R. Jiménez and J. E. Yukich. Asymptotics for statistical distances based on Voronoi tessellations. Journal of Theoretical Probability, 15(2):503–541, April 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1014819112010.

Jiang:2022:MDD

[JY22] Hui Jiang and Qingshan Yang. Moderate deviations for drift parameter estimations in reflected Ornstein–Uhlenbeck process. Journal of Theoretical Probability, 35(2):1262–1283, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01096-3.

Jiang:2023:ENB

[JZ23] Hui Jiang and Jingying Zhou. An exponential nonuniform Berry–Esseen bound for the fractional Ornstein–Uhlenbeck process. Journal of Theoretical Probability, 36(2):1037–1058, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01194-w.

Kabluchko:2012:LLS

[Kab12] Zakhar Kabluchko. Limit laws for sums of independent random products: the lattice case. Journal of Theoretical Probability, 25(2):424–437, June 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0296-5.

Kallenberg:1990:ERM

[Kal90] Olav Kallenberg. Exchangeable random measures in the plane. Journal of Theoretical Probability, 3(1):81–136, January 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01063330.

Kalpazidou:1991:CDG

[Kal91] S. Kalpazidou. Circulation distribution on groups. Journal of Theoretical Probability, 4(3):475–483, July 1991. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01210320.

Kallenberg:1992:SRA

[Kal92a] Olav Kallenberg. Symmetries on random arrays and set-indexed processes. Journal of Theoretical Probability, 5(4):727–765, October 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01058727.

Kalpazidou:1992:CPC

[Kal92b] S. Kalpazidou. Circuit processes and the corresponding oneparameter semigroup of weight operators. Journal of Theoretical Probability, 5(1):205–216, January 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01046786.

Kallenberg:1999:MSE

[Kal99] Olav Kallenberg. Multivariate sampling and the estimation problem for exchangeable arrays. Journal of Theoretical Probability, 12(3):859–883, July 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021692202530.

Kallenberg:2002:EPP

[Kal02] Olav Kallenberg. Ergodic properties of Palm and spacing measures. Journal of Theoretical Probability, 15(1):77–96, January 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1013889222507.

Kallenberg:2011:IPC

[Kal11] Olav Kallenberg. Iterated Palm conditioning and some Slivnyaktype theorems for Cox and cluster processes. Journal of Theoretical Probability, 24(3):875–893, September 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0295-6.

Kallenberg:2012:STU

[Kal12] Olav Kallenberg. Schoenberg’s theorem and unitarily invariant random arrays. Journal of Theoretical Probability, 25(4): 1013–1039, December 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0332-5.

Kalsi:2020:EIM

[Kal20] Jasdeep Kalsi. Existence of invariant measures for reflected stochastic partial differential equations. Journal of Theoretical Probability, 33(3):1755–1767, September 2020. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-019-00906-z.pdf.

Kargin:2007:BEF

[Kar07] Vladislav Kargin. Berry–Esséen for free random variables. Journal of Theoretical Probability, 20(2):381–395, June 2007. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0097-7.

Kargin:2011:FSD

[Kar11] Vladislav Kargin. On free stochastic differential equations. Journal of Theoretical Probability, 24(3):821–848, September 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-011-0341-z.

Karouf:2019:RDR

[Kar19] Monia Karouf. Reflected and doubly reflected backward stochastic differential equations with time-delayed generators. Journal of Theoretical Probability, 32(1):216–248, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kargin:2023:SLS

[Kar23] Vladislav Kargin. Scaling limits of slim and fat trees. Journal of Theoretical Probability, 36(4):2192–2228, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01261-w.

Kawai:2013:LAN

[Kaw13] Reiichiro Kawai. Local asymptotic normality property for Ornstein–Uhlenbeck processes with jumps under discrete sampling. Journal of Theoretical Probability, 26(4):932–967, December 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-012-0455-y.

Kazakevicius:2021:SET

[Kaz21] Vytautas Kazakevicius. Subadditive ergodic theorem for double sequences. Journal of Theoretical Probability, 34(1):307– 330, March 2021. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-019-00979-w.

Kuczek:1991:LDR

[KC91] Thomas Kuczek and Keith N. Crank. A large-deviation result for regenerative processes. Journal of Theoretical Probability, 4(3):551–561, July 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01210324.

Kevei:2009:MLC

[KC09] Péter Kevei and Sándor Csörgő. Merging of linear combinations to semistable laws. Journal of Theoretical Probability, 22(3): 772–790, September 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0138-2.

Kasparaviciute:2017:AED

[KD17] Aurelija Kasparaviciute and Dovile Deltuviene. Asymptotic expansion for the distribution density function of the compound Poisson process in large deviations. Journal of Theoretical Probability, 30(4):1655–1676, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kemp:2017:HKE

[Kem17] Todd Kemp. Heat kernel empirical laws on U_{N} and GL_{N}. Journal of Theoretical Probability, 30(2):397–451, June 2017. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kesten:1995:BRW

[Kes95] Harry Kesten. Branching random walk with a critical branching part. Journal of Theoretical Probability, 8(4):921–962, October 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02410118.

Key:1990:LBM

[Key90] Eric S. Key. Lower bounds for the maximal Lyapunov exponent. Journal of Theoretical Probability, 3(3):477–488, July 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01061263.

Key:1992:RN

[Key92] Eric S. Key. Rare numbers. Journal of Theoretical Probability, 5(2):375–389, April 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046742.

Key:1996:DRN

[Key96] Eric S. Key. Divergence rates for the number of rare numbers. Journal of Theoretical Probability, 9(2):413–428, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF02214657.

Key:2005:NRI

[Key05] Eric S. Key. On the number of records in an iid discrete sequence. Journal of Theoretical Probability, 18(1):99–107, January 2005.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-004-2578-2.

Key:2007:NRI

[Key07] Eric S. Key. On the number of records in an iid discrete sequence. Journal of Theoretical Probability, 20(3):695–696, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-007-0053-6.

Khan:2009:VDS

[Kha09] Rasul A. Khan. L_{p}-version of the Dubins–Savage inequality and some exponential inequalities. Journal of Theoretical Probability, 22(2):348–364, June 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-008-0206-2.

Khanchi:2012:AMA

[Kha12] Aziz Khanchi. Asymptotics of Markov additive chains on a halfplane: A ratio limit theorem. Journal of Theoretical Probability, 25(1):62–76, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0384-1.

Kohatsu-Higa:2001:ABD

[KHMCSS01] A. Kohatsu-Higa, D. Márquez-Carreras, and M. Sanz-Solé. Asymptotic behavior of the density in a parabolic SPDE. Journal of Theoretical Probability, 14(2):427–462, April 2001. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1011163714298.

Kohatsu-Higa:2021:LTA

[KHN21] Arturo Kohatsu-Higa and David Nualart. Large time asymptotic properties of the stochastic heat equation. Journal of Theoretical Probability, 34(3):1455–1473, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
01007-y.

Khorunzhiy:2022:APB

[Kho22] O. Khorunzhiy. On asymptotic properties of Bell polynomials and concentration of vertex degree of large random graphs. Journal of Theoretical Probability, 35(1):20–51, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01025-w.

Kiesel:1998:SLS

[Kie98] Rüdiger Kiesel. Strong laws and summability for ϕ-mixing sequences of random variables. Journal of Theoretical Probability, 11(1):209–224, January 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1021655227120.

Kifer:2013:NIP

[Kif13] Yuri Kifer. A nonconventional invariance principle for random fields. Journal of Theoretical Probability, 26(2):489–513, June 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-012-0473-9.

Kim:2008:TPS

[Kim08] Kyeong-Hun Kim. L_{p}-theory of parabolic SPDEs degenerating on the boundary of C^{1} domains. Journal of Theoretical Probability, 21(1):169–192, March 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0098-6.

Kim:2009:TSP

[Kim09] Kyeong-Hun Kim. An L_{p}-theory of stochastic PDEs of divergence form on Lipschitz domains. Journal of Theoretical Probability, 22(1):220–238, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0170-x. See er-

ratum [Kim17].

Kim:2014:WSS

[Kim14] Kyeong-Hun Kim. A weighted Sobolev space theory of parabolic stochastic PDEs on non-smooth domains. Journal of Theoretical Probability, 27(1):107–136, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-012-0459-7.

Kim:2017:EES

[Kim17] Kyeong-Hun Kim. Erratum to: An L_{p}-theory of Stochastic PDEs of Divergence Form on Lipschitz Domains. Journal of Theoretical Probability, 30(1):395–396, March 2017. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-015-0637-5.pdf. See [Kim09].

Kim:2022:LTC

[Kim22] Donghan Kim. Local times for continuous paths of arbitrary regularity. Journal of Theoretical Probability, 35(4):2540–2568, December 2022. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-022-01159-z.

Kinateder:1995:CMP

[Kin95] Kimberly K. J. Kinateder. Corner Markov processes. Journal of Theoretical Probability, 8(3):539–547, July 1995. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02218043.

Kinateder:2000:SMP

[Kin00] Kimberly K. J. Kinateder. Strong Markov properties for Markov random fields. Journal of Theoretical Probability, 13(4):1101– 1114, October 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007822209798.

Kalikow:1994:CVS

[KK94] Steven Kalikow and Jonathan L. King. A countably-valued sleeping stockbroker process. Journal of Theoretical Probability, 7(4): 703–708, October 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214366.

Kim:2008:CCM

[KK08] Tae-Sung Kim and Mi-Hwa Ko. On the complete convergence of moving average process with Banach space valued random elements. Journal of Theoretical Probability, 21(2):431–436, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-007-0118-6.

Konstantinides:2017:DHT

[KK17] Dimitrios G. Konstantinides and Christos E. Kountzakis. Distributions with heavy tails in Orlicz spaces. Journal of Theoretical Probability, 30(4):1726–1762, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kirsch:2018:SLG

[KK18] Werner Kirsch and Thomas Kriecherbauer. Semicircle law for generalized Curie–Weiss matrix ensembles at subcritical temperature. Journal of Theoretical Probability, 31(4):2446–2458, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kataria:2021:LRD

[KK21] K. K. Kataria and M. Khandakar. On the long-range dependence of mixed fractional Poisson process. Journal of Theoretical Probability, 34(3):1607–1622, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-020-01015-y.

Kataria:2022:GFC

[KK22] K. K. Kataria and M. Khandakar. Generalized fractional counting process. Journal of Theoretical Probability, 35(4):2784– 2805, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01160-6.

Kataria:2024:FSP

[KK24] K. K. Kataria and M. Khandakar. Fractional skellam process of order k. Journal of Theoretical Probability, 37(2):1333– 1356, June 2024. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-024-01314-8.

Ko:2009:NAS

[KKH09] Mi-Hwa Ko, Tae-Sung Kim, and Kwang-Hee Han. A note on the almost sure convergence for dependent random variables in a Hilbert space. Journal of Theoretical Probability, 22(2):506–513, June 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0144-z.

Kologlu:2013:LSM

[KKM13] Murat Kologlu, Gene S. Kopp, and Steven J. Miller. The limiting spectral measure for ensembles of symmetric block circulant matrices. Journal of Theoretical Probability, 26(4):1020–1060, December 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-011-0391-2.

Kaimanovich:2004:BHF

[KKR04] V. A. Kaimanovich, Y. Kifer, and B.-Z. Rubshtein. Boundaries and harmonic functions for random walks with random transition probabilities. Journal of Theoretical Probability, 17 (3):605–646, July 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000040291.80182.65.

Katori:2004:DDK

[KKST04] Makoto Katori, Norio Konno, Aidan Sudbury, and Hideki Tanemura. Dualities for the Domany–Kinzel model. Journal of Theoretical Probability, 17(1):131–144, January 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000020478.24536.26.

Kuelbs:1993:SBE

[KL93] James Kuelbs and Wenbo V. Li. Small ball estimates for Brownian motion and the Brownian sheet. Journal of Theoretical Probability, 6(3):547–577, July 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01066717.

Khoshnevisan:1996:UMC

[KL96] Davar Khoshnevisan and Thomas M. Lewis. The uniform modulus of continuity of iterated Brownian motion. Journal of Theoretical Probability, 9(2):317–333, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02214652.

Kuelbs:1998:CLS

[KL98] J. Kuelbs and M. Ledoux. On convex limit sets and Brownian motion. Journal of Theoretical Probability, 11(2):461–492, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022640007525.

Kratz:2001:CLT

[KL01] Marie F. Kratz and José R. León. Central limit theorems for level functionals of stationary Gaussian processes and fields. Journal of Theoretical Probability, 14(3):639–672, July 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1017588905727.

Kuelbs:2002:FLS

[KL02] J. Kuelbs and W. V. Li. A functional LIL and some weighted occupation measure results for fractional Brownian motion. Journal of Theoretical Probability, 15(4):1007–1030, October 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020696922183.

Kuelbs:2005:FLS

[KL05] James Kuelbs and Wenbo Li. A functional LIL for stochastic integrals and the Lévy area process. Journal of Theoretical Probability, 18(2):261–290, April 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-003-2604-9.

Koo:2007:RCM

[KL07] Yooyoung Koo and Sungchul Lee. Rates of convergence of means of Euclidean functionals. Journal of Theoretical Probability, 20(4):821–841, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0089-7.

Kulikov:2008:DGM

[KL08] Vladimir N. Kulikov and Hendrik P. Lopuhaä. Distribution of global measures of deviation between the empirical distribution function and its concave majorant. Journal of Theoretical Probability, 21(2):356–377, June 2008. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-007-0103-0.pdf.

Kharroubi:2014:PEF

[KL14] Idris Kharroubi and Thomas Lim. Progressive enlargement of filtrations and backward stochastic differential equations with jumps. Journal of Theoretical Probability, 27(3):683– 724, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0428-1.

Kabluchko:2017:LEA

[KL17] Zakhar Kabluchko and Mikhail Lifshits. Least energy approximation for processes with stationary increments. Journal of Theoretical Probability, 30(1):268–296, March 2017. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
015-0642-8.

Komisarski:2024:DMR

[KL24] Andrzej Komisarski and Jacques Labuschagne. The distributions of the mean of random vectors with fixed marginal distribution. Journal of Theoretical Probability, 37(3):2121–2129, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01277-2.

Kley:2013:KLI

[Kle13] Oliver Kley. Kuelbs–Li inequalities and metric entropy of convex hulls. Journal of Theoretical Probability, 26(3):649–665, September 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-012-0408-5.

Kleiber:2014:GLD

[Kle14] Christian Kleiber. The generalized lognormal distribution and the Stieltjes moment problem. Journal of Theoretical Probability, 27(4):1167–1177, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-013-0477-0.

Klimsiak:2013:TDF

[Kli13] Tomasz Klimsiak. On time-dependent functionals of diffusions corresponding to divergence form operators. Journal of Theoretical Probability, 26(2):437–473, June 2013. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-011-0381-4.pdf.

Kim:2004:LIL

[KLL04] Younghoon Kim, Sungchul Lee, and Zhengyan Lin. The law of the iterated logarithm for the total length of the nearest neighbor graph. Journal of Theoretical Probability, 17(1):245–260, January 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOTP.0000020483.81884.a2.

Kim:2005:IPT

[KLLW05] Younghoon Kim, Sungchul Lee, Zhengyan Lin, and Wensheng Wang. The invariance principle for the total length of the nearest-neighbor graph. Journal of Theoretical Probability, 18 (3):649–664, July 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-005-7253-8.

Kuelbs:1995:SBP

[KLmS95] J. Kuelbs, W. V. Li, and Qi man Shao. Small ball probabilities for Gaussian processes with stationary increments under Hölder norms. Journal of Theoretical Probability, 8(2):361–386, April 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02212884.

Kahn:1989:CTR

[KLNS89] Jeff D. Kahn, Nathan Linial, Noam Nisan, and Michael E. Saks. On the cover time of random walks on graphs. Journal of Theoretical Probability, 2(1):121–128, January 1989. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01048274.

Klass:2022:DCC

[KLR22] Michael J. Klass, Deli Li, and Andrew Rosalsky. Divergence criterion for a class of random series related to the partial sums of I.I.D. random variables. Journal of Theoretical Probability, 35(3):1556–1573, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01101-9.

Kouritzin:2003:NFD

[KLS03] Michael A. Kouritzin, Hongwei Long, and Wei Sun. Nonlinear filtering for diffusions in random environments. Journal of Theoretical Probability, 16(1):1–20, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022223518746.

Konsowa:1991:TPR

[KM91] Mokhtar Konsowa and Joanna Mitro. The type problem for random walks on trees. Journal of Theoretical Probability, 4 (3):535–550, July 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01210323.

Kasahara:1992:LTT

[KM92] Yuji Kasahara and Makoto Maejima. Limit theorems for trimmed sums. Journal of Theoretical Probability, 5(4):617–628, October 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01058722.

Kesten:1997:DRW

[KM97] Harry Kesten and Ross A. Maller. Divergence of a random walk through deterministic and random subsequences. Journal of Theoretical Probability, 10(2):395–427, April 1997. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022664500932.

Kesten:1998:RWC

[KM98] Harry Kesten and R. A. Maller. Random walks crossing high level curved boundaries. Journal of Theoretical Probability, 11(4):1019–1074, October 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022621016708.

Kumagai:2008:HKE

[KM08] Takashi Kumagai and Jun Misumi. Heat kernel estimates for strongly recurrent random walk on random media. Journal of Theoretical Probability, 21(4):910–935, December 2008. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
008-0183-5.

Kiefer:2010:HDD

[KM10] Richard Kiefer and Peter Mörters. The Hausdorff dimension of the double points on the Brownian frontier. Journal of Theoretical Probability, 23(2):605–623, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-009-0262-2.

Konakov:2011:WES

[KM11] Valentin Konakov and Stéphane Menozzi. Weak error for stable driven stochastic differential equations: Expansion of the densities. Journal of Theoretical Probability, 24(2):454–478, June 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-010-0291-x.

Kerkvliet:2016:UDU

[KM16] Timber Kerkvliet and Ronald Meester. Uniquely determined uniform probability on the natural numbers. Journal of Theoretical Probability, 29(3):797–825, September 2016. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-015-0611-2.pdf.

Kevei:2017:CSA

[KM17] Péter Kevei and David M. Mason. Couplings and strong approximations to time–dependent empirical processes based on I.I.D. fractional Brownian motions. Journal of Theoretical Probability, 30(3):729–770, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kerkvliet:2018:BIB

[KM18] Timber Kerkvliet and Ronald Meester. A behavioral interpretation of belief functions. Journal of Theoretical Probability, 31(4): 2112–2128, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-017-0776-y.pdf.

Kalinkin:2019:SVM

[KM19] Aleksandr V. Kalinkin and Anton V. Mastikhin. On the separating variables method for Markov death-process equations. Journal of Theoretical Probability, 32(1):163–182, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Konarovskyi:2023:CBP

[KM23] Vitalii Konarovskyi and Victor Marx. On conditioning Brownian particles to coalesce. Journal of Theoretical Probability, 36(4): 2126–2164, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01267-4.

Knopova:2024:EPS

[KM24] Victoria Knopova and Yana Mokanu. On ergodic properties of some Lévy-type processes. Journal of Theoretical Probability, 37 (1):582–602, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01252-x.

Kalbasi:2018:APF

[KMV18] Kamran Kalbasi, Thomas S. Mountford, and Frederi G. Viens. Anderson polymer in a fractional Brownian environment: Asymptotic behavior of the partition function. Journal of Theoretical Probability, 31(3):1429–1468, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kern:2018:ABS

[KMX18] Peter Kern, Mark M. Meerschaert, and Yimin Xiao. Asymptotic behavior of semistable Lévy exponents and applications to fractal path properties. Journal of Theoretical Probability, 31 (1):598–617, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Klass:2010:UAQ

[KN10] Michael J. Klass and Krzysztof Nowicki. Uniformly accurate quantile bounds for sums of arbitrary independent random variables. Journal of Theoretical Probability, 23(4):1068–1091, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-009-0254-2.

Kozma:2011:NAC

[KN11] Gady Kozma and Asaf Nachmias. A note about critical percolation on finite graphs. Journal of Theoretical Probability, 24(4): 1087–1096, December 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0283-x.

Klass:2016:UBR

[KN16] Michael J. Klass and Krzysztof Nowicki. Uniform bounds on the relative error in the approximation of upper quantiles for sums of arbitrary independent random variables. Journal of Theoretical Probability, 29(4):1485–1509, December 2016. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
015-0615-y.

Kumagai:2018:LRW

[KN18] Takashi Kumagai and Chikara Nakamura. Lamplighter random walks on fractals. Journal of Theoretical Probability, 31(1):68– 92, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kostler:2021:CLT

[KN21] Claus Köstler and Alexandru Nica. A central limit theorem for star-generators of S_{∞}, which relates to the law of a GUE matrix. Journal of Theoretical Probability, 34(3):1248–1278, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01029-6.

Khachatryan:2023:CFR

[KN23] Linda Khachatryan and Boris S. Nahapetian. On the characterization of a finite random field by conditional distribution and its Gibbs form. Journal of Theoretical Probability, 36(3): 1743–1761, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01209-6.

Knobloch:2018:OSF

[Kno18] Robert Knobloch. One-sided FKPP travelling waves for homogeneous fragmentation processes. Journal of Theoretical Probability, 31(2):895–931, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kasahara:1999:NLT

[KO99] Yuji Kasahara and Namiko Ogawa. A note on the local time of fractional Brownian motion. Journal of Theoretical Probability, 12(1):207–216, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1021756929498.

Kim:2010:SIR

[KO10] Daehong Kim and Yoichi Oshima. Some inequalities related to transience and recurrence of Markov processes and their applications. Journal of Theoretical Probability, 23(1):148–168, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-008-0196-0.

Kawamoto:2019:CDB

[KO19a] Yosuke Kawamoto and Hirofumi Osada. Correction to: Dynamical Bulk Scaling Limit of Gaussian Unitary Ensembles and Stochastic Differential Equation Gaps. Journal of Theoretical Probability, 32(3):1613, September 2019. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-019-00913-0.pdf. See [KO19b].

Kawamoto:2019:DBS

[KO19b] Yosuke Kawamoto and Hirofumi Osada. Dynamical bulk scaling limit of Gaussian unitary ensembles and stochastic differential equation gaps. Journal of Theoretical Probability, 32 (2):907–933, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-018-0816-2.pdf. See cor-

rection [KO19a].

Kobayashi:2011:SCT

[Kob11] Kei Kobayashi. Stochastic calculus for a time-changed semimartingale and the associated stochastic differential equations. Journal of Theoretical Probability, 24(3):789–820, September 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-010-0320-9.

Kochubei:2002:HMS

[Koc02] Anatoly N. Kochubei. Hausdorff measure for a stable-like process over an infinite extension of a local field. Journal of Theoretical Probability, 15(4):951–972, October 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1020789821275.

Kolsrud:1989:THF

[Kol89] Torbjörn Kolsrud. Traces of harmonic functions, capacities, and traces of symmetric Markov processes. Journal of Theoretical Probability, 2(4):399–418, October 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01051874.

Koldobsky:1994:CMP

[Kol94a] Alexander Koldobsky. Characterization of measures by potentials. Journal of Theoretical Probability, 7(1):135–145, January 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF02213364.

Koltchinskii:1994:KMT

[Kol94b] Vladimir I. Koltchinskii. Komlos–Major–Tusnady approximation for the general empirical process and Haar expansions of classes of functions. Journal of Theoretical Probability, 7(1): 73–118, January 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213361.

Koltchinskii:1998:DIO

[Kol98] Vladimir I. Koltchinskii. Differentiability of inverse operators and limit theorems for inverse functions. Journal of Theoretical Probability, 11(3):645–699, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022650530662.

Kolesnik:2001:WCP

[Kol01] Alexander D. Kolesnik. Weak convergence of a planar random evolution to the Wiener process. Journal of Theoretical Probability, 14(2):485–494, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1011167815206.

Kolodziejek:2016:LOR

[Kol16] Bartosz Kolodziejek. The Lukacs–Olkin–Rubin theorem on symmetric cones without invariance of the ”quotient”. Journal of Theoretical Probability, 29(2):550–568, June 2016. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-014-0587-3.pdf.

Kolodziejek:2017:MYP

[Kol17] Bartosz Kolodziejek. The Matsumoto–Yor property and its converse on symmetric cones. Journal of Theoretical Probability, 30(2):624–638, June 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/content/pdf/10.1007/s10959-015-0648-2.pdf.

Komisarski:2008:DBF

[Kom08] Andrzej Komisarski. Distances between σ-fields on a probability space. Journal of Theoretical Probability, 21(4):812–823, December 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0149-7.

Konno:1995:ABB

[Kon95] Norio Konno. Asymptotic behavior of basic contact process with rapid stirring. Journal of Theoretical Probability, 8(4):833–876, October 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02410115.

Kontoyiannis:1998:ARW

[Kon98] Ioannis Kontoyiannis. Asymptotic recurrence and waiting times for stationary processes. Journal of Theoretical Probability, 11 (3):795–811, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022610816550.

Korshunov:2008:KRT

[Kor08] Dmitry Korshunov. The key renewal theorem for a transient Markov chain. Journal of Theoretical Probability, 21(1):234–245, March 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-007-0132-8.

Kortchemski:2016:PPD

[Kor16] Igor Kortchemski. Predator–prey dynamics on infinite trees: A branching random walk approach. Journal of Theoretical Probability, 29(3):1027–1046, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-015-0603-2.

Kosters:2011:CPS

[Kös11] H. Kösters. Characteristic polynomials of sample covariance matrices. Journal of Theoretical Probability, 24(2):545–576, June 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-010-0290-y.

Kouritzin:1996:IAS

[Kou96] Michael A. Kouritzin. On the interrelation of almost sure invariance principles for certain stochastic adaptive algorithms and for partial sums of random variables. Journal of Theoretical Probability, 9(4):811–840, October 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214252.

Koval:2002:NLI

[Kov02] Valery Koval. A new law of the iterated logarithm in R^{d} with application to matrix-normalized sums of random vectors. Journal of Theoretical Probability, 15(1):249–257, January 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013851720494.

Kovchegov:2008:MPP

[Kov08] Yevgeniy Kovchegov. Multi-particle processes with reinforcements. Journal of Theoretical Probability, 21(2):437–448, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-007-0141-7.

Kovchegov:2022:NLP

[Kov22] Yevgeniy Kovchegov. A new life of Pearson’s skewness. Journal of Theoretical Probability, 35(4):2896–2915, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01149-7.

Kowalska:1992:NCL

[Kow92] Anna Kowalska. A new class of libraries. Journal of Theoretical Probability, 5(3):431–445, July 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01060428.

Klusik:2014:NWH

[KP14] Przemysław Klusik and Zbigniew Palmowski. A note on Wiener– Hopf factorization for Markov additive processes. Journal of Theoretical Probability, 27(1):202–219, March 2014. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-012-0425-4.pdf.

Kobayashi:2023:SHC

[KP23] Kei Kobayashi and Hyunchul Park. Spectral heat content for time-changed killed Brownian motions. Journal of Theoretical Probability, 36(2):1148–1180, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-022-01188-8.

Kim:2024:SST

[KP24] Kyeong-Hun Kim and Daehan Park. A Sobolev space theory for time-fractional stochastic partial differential equations driven by Lévy processes. Journal of Theoretical Probability, 37 (1):671–720, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01263-8.

Kyprianou:2014:OTR

[KPP14] A. E. Kyprianou, J. C. Pardo, and J. L. Pérez. Occupation times of refracted Lévy processes. Journal of Theoretical Probability, 27(4):1292–1315, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-013-0501-4.

Khabou:2024:NAC

[KPR24] Mahmoud Khabou, Nicolas Privault, and Anthony Réveillac. Normal approximation of compound Hawkes functionals. Journal of Theoretical Probability, 37(1):549–581, March 2024. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s10959-022-01233-6.

Kolodynski:2003:GSS

[KR03] Slawomir Kolodyński and Jan Rosiński. Group self-similar stable processes in R^{d}. Journal of Theoretical Probability, 16(4): 855–876, October 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000011997.14357.fd.

Kosmala:2021:SIR

[KR21] Tomasz Kosmala and Markus Riedle. Stochastic integration with respect to cylindrical Lévy processes by p-summing operators. Journal of Theoretical Probability, 34(1):477–497, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-019-00978-x.

Krajka:2001:LBR

[Kra01] Andrzej Krajka. The limiting behaviour of randomly indexed sequences. Journal of Theoretical Probability, 14(3):797–811, July 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017549224382.

Kroger:1990:CTD

[Krö90] Pawel Kröger. Comparison theorems for diffusion processes. Journal of Theoretical Probability, 3(4):515–531, October 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01046093.

Kroll:2022:ABE

[Kro22] Marius Kroll. Asymptotic behaviour of the empirical distance covariance for dependent data. Journal of Theoretical Probability, 35(2):1226–1246, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01073-w.

Kyprianou:2010:CSS

[KRS10] Andreas E. Kyprianou, Víctor Rivero, and Renming Song. Convexity and smoothness of scale functions and de Finetti’s control problem. Journal of Theoretical Probability, 23(2):547–564, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-009-0220-z.

Kurien:1993:SGR

[KS93] T. V. Kurien and Jayaram Sethuraman. Singularities in Gaussian random fields. Journal of Theoretical Probability, 6(1):

89–99, January 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046770.

Kiesel:1996:ERS

[KS96] Rüdiger Kiesel and Ulrich Stadtmüller. Erdős–Rényi–Shepp laws and weighted sums of independent identically distributed random variables. Journal of Theoretical Probability, 9(4):961–982, October 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02214259.

Kaj:1998:LPA

[KS98] I. Kaj and S. Sagitov. Limit processes for age-dependent branching particle systems. Journal of Theoretical Probability, 11(1): 225–257, January 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021607311191.

Kerns:2006:DTA

[KS06] G. Jay. Kerns and Gábor J. Székely. Definetti’s theorem for abstract finite exchangeable sequences. Journal of Theoretical Probability, 19(3):589–608, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-006-0028-z.

Kim:2008:BBH

[KS08] Panki Kim and Renming Song. Boundary behavior of harmonic functions for truncated stable processes. Journal of Theoretical Probability, 21(2):287–321, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0145-y.

Knopova:2012:TDE

[KS12] Viktorya Knopova and René L. Schilling. Transition density estimates for a class of Lévy and Lévy-Type processes. Journal of Theoretical Probability, 25(1):144–170, March 2012. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0300-0.

Kalenda:2016:TMD

[KS16a] Ondrej F. K. Kalenda and Jirí Spurný. Typical martingale diverges at a typical point. Journal of Theoretical Probability, 29 (1):180–205, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0567-7.

Kuhn:2016:MDS

[KS16b] Franziska Kühn and René L. Schilling. Moderate deviations and Strassen’s law for additive processes. Journal of Theoretical Probability, 29(2):632–652, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-014-0584-6.

Kessebohmer:2019:SLL

[KS19] Marc Kesseböhmer and Tanja Schindler. Strong laws of large numbers for intermediately trimmed sums of i.i.d. random variables with infinite mean. Journal of Theoretical Probability, 32 (2):702–720, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kagan:2021:CDB

[KS21] Abram M. Kagan and Gábor J. Székely. Calibrating dependence between random elements. Journal of Theoretical Probability, 34 (2):784–790, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-00995-1.

Kremsner:2022:BSC

[KS22] Stefan Kremsner and Alexander Steinicke. L^{p}-solutions and comparison results for Lévy-driven backward stochastic differential equations in a monotonic, general growth setting. Journal of Theoretical Probability, 35(1):231–281, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
01056-3.

Kokoszka:1994:NCS

[KT94] Piotr S. Kokoszka and Murad S. Taqqu. New classes of selfsimilar symmetric stable random fields. Journal of Theoretical Probability, 7(3):527–549, July 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02213567.

Kokoszka:1996:CMP

[KT96] Piotr S. Kokoszka and Murad S. Taqqu. A characterization of mixing processes of type G^{1}. Journal of Theoretical Probability, 9(1):3–17, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213732.

Khan:2000:RDS

[KT00] Rasul A. Khan and R. J. Tomkins. Refinements of the Dubins– Savage inequality. Journal of Theoretical Probability, 13(3): 659–672, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007802326865.

Koralov:2016:QLE

[KT16] Leonid Koralov and Lucas Tcheuko. Quasi-linear equations with a small diffusion term and the evolution of hierarchies of cycles. Journal of Theoretical Probability, 29(3):867–895, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-015-0601-4.

Kevei:2020:DKT

[KT20a] Péter Kevei and Dalia Terhesiu. Darling–Kac theorem for renewal shifts in the absence of regular variation. Journal of Theoretical Probability, 33(4):2027–2060, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-019-
00930-z.

Kirsch:2020:TGC

[KT20b] Werner Kirsch and Gabor Toth. Two groups in a Curie–Weiss model with heterogeneous coupling. Journal of Theoretical Probability, 33(4):2001–2026, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-019-00933-w.

Kevei:2022:SRT

[KT22] Péter Kevei and Dalia Terhesiu. Strong renewal theorem and local limit theorem in the absence of regular variation. Journal of Theoretical Probability, 35(2):1013–1048, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01081-w.

Kratschmer:2023:KCT

[KU23] Volker Krätschmer and Mikhail Urusov. A Kolmogorov– Chentsov type theorem on general metric spaces with applications to limit theorems for Banach-valued processes. Journal of Theoretical Probability, 36(3):1454–1486, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01207-8.

Kuan:2018:TCM

[Kua18] Jeffrey Kuan. Two constructions of Markov chains on the dual of U(n). Journal of Theoretical Probability, 31(3):1411–1428, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kubota:2020:CRF

[Kub20] Naoki Kubota. Continuity for the rate function of the simple random walk on supercritical percolation clusters. Journal of Theoretical Probability, 33(4):1948–1973, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-019-
00936-7.

Kuhn:1991:FAR

[Kuh91] Gabriella Kuhn. Finitely additive random walks on infinitely generated free groups. Journal of Theoretical Probability, 4 (2):311–320, April 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01258739.

Kuhn:2021:SEP

[Küh21] Franziska Kühn. Schauder estimates for Poisson equations associated with non-local Feller generators. Journal of Theoretical Probability, 34(3):1506–1578, September 2021. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
01008-x.

Kulik:2011:ACC

[Kul11] Alexey M. Kulik. Absolute continuity and convergence in variation for distributions of functionals of Poisson point measure. Journal of Theoretical Probability, 24(1):1–38, March 2011. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0325-4.

Kulyk:2023:STL

[Kul23] Oleksii Kulyk. Support theorem for Lévy-driven stochastic differential equations. Journal of Theoretical Probability, 36(3): 1720–1742, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01223-8.

Kumar:2011:CFI

[Kum11] Rohini Kumar. Current fluctuations for independent random walks in multiple dimensions. Journal of Theoretical Probability, 24(4):1170–1195, December 2011. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s10959-
010-0317-4.pdf.

Kunita:1994:GSS

[Kun94a] H. Kunita. Generalized solutions of a stochastic partial differential equation. Journal of Theoretical Probability, 7(2):279–308, April 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02214271.

Kunita:1994:SFA

[Kun94b] H. Kunita. Stochastic flows acting on Schwartz distributions. Journal of Theoretical Probability, 7(2):247–278, April 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02214270.

Kunisada:2022:APF

[Kun22] Ryoichi Kunisada. On the additive property of finitely additive measures. Journal of Theoretical Probability, 35(3):1782– 1794, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01115-3.

Kurenok:2007:SET

[Kur07] V. P. Kurenok. Stochastic equations with time-dependent drift driven by Levy processes. Journal of Theoretical Probability, 20(4):859–869, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0086-x.

Kuwada:2007:UMC

[Kuw07] Kazumasa Kuwada. On uniqueness of maximal coupling for diffusion processes with a reflection. Journal of Theoretical Probability, 20(4):935–957, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0087-9.

Kuznetsov:2020:FRB

[Kuz20] A. Kuznetsov. On free regular and bondesson convolution semigroups. Journal of Theoretical Probability, 33(3):1493–1505, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kingman:2003:SOC

[KV03] J. F. C. Kingman and S. E. Volkov. Solution to the OK corral model via decoupling of Friedman’s urn. Journal of Theoretical Probability, 16(1):267–276, January 2003. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/

A%3A1022294908268. See correction [KV19b].

Kadankova:2007:STB

[KV07] T. Kadankova and N. Veraverbeke. On several two-boundary problems for a particular class of Lévy processes. Journal of Theoretical Probability, 20(4):1073–1085, December 2007. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0088-8.

Kuelbs:2011:WCR

[KV11] James Kuelbs and Anand N. Vidyashankar. Weak convergence results for multiple generations of a branching process. Journal of Theoretical Probability, 24(2):376–396, June 2011. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
009-0266-y.

Kratz:2018:CLT

[KV18] Marie Kratz and Sreekar Vadlamani. Central limit theorem for Lipschitz–Killing curvatures of excursion sets of Gaussian random fields. Journal of Theoretical Probability, 31(3):1729–1758, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kataria:2019:DCS

[KV19a] K. K. Kataria and P. Vellaisamy. On distributions of certain state-dependent fractional point processes. Journal of Theoretical Probability, 32(3):1554–1580, September 2019. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kingman:2019:CSO

[KV19b] J. F. C. Kingman and S. E. Volkov. Correction to: Solution to the OK Corral Model via Decoupling of Friedman’s Urn. Journal of Theoretical Probability, 32(3):1614–1615, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10959-019-00921-0.pdf. See [KV03].

Klass:1993:WDS

[KW93] Michael J. Klass and Rainer Wittmann. Which I.I.D. sums are recurrently dominated by their maximal terms? Journal of Theoretical Probability, 6(2):195–207, April 1993. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01047571.

Kern:2014:HDO

[KW14] Peter Kern and Lina Wedrich. The Hausdorff dimension of operator semistable Lévy processes. Journal of Theoretical Probability, 27(2):383–403, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-012-0422-7.

Kang:2015:QSI

[KW15] Yuanbao Kang and Caishi Wang. Quantum stochastic integral representations on interacting Fock space. Journal of Theoretical Probability, 28(3):1007–1027, September 2015. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
013-0537-5.

Klass:2012:MIA

[KY12] Michael J. Klass and Ming Yang. Maximal inequalities for additive processes. Journal of Theoretical Probability, 25(4):981– 1012, December 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0357-4.

Kaneko:2006:TSG

[KZ06] Hiroshi Kaneko and Xuelei Zhao. Transition semi-groups on a local field induced by Galois group and their representation. Journal of Theoretical Probability, 19(1):221–234, January 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-006-0004-7.

Kuelbs:2008:AVC

[KZ08] Jim Kuelbs and Joel Zinn. Another view of the CLT in Banach spaces. Journal of Theoretical Probability, 21(4):982–1029, December 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0166-6.

Kuelbs:2015:EQC

[KZ15] James Kuelbs and Joel Zinn. Empirical quantile central limit theorems for some self-similar processes. Journal of Theoretical Probability, 28(1):313–336, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-013-0511-2.

Kuelbs:2020:LPM

[KZ20] James Kuelbs and Joel Zinn. Limits for partial maxima of Gaussian random vectors. Journal of Theoretical Probability, 33(2): 788–827, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Kim:2012:SSP

[KZS12] Kunwoo Kim, Zhi Zheng, and Richard B. Sowers. A stochastic Stefan problem. Journal of Theoretical Probability, 25(4): 1040–1080, December 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0392-1.

Lacey:1989:LIL

[Lac89] Michael Lacey. Laws of the iterated logarithm for partial sum processes indexed by functions. Journal of Theoretical Probability, 2(3):377–398, July 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01054022.

Lacey:1991:ASN

[Lac91] Michael Lacey. On almost sure noncentral limit theorems. Journal of Theoretical Probability, 4(4):767–781, October 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01259554.

Lachal:1998:SMR

[Lac98] Aimé Lachal. Some martingales related to the integral of Brownian motion. Applications to the passage times and transience. Journal of Theoretical Probability, 11(1):127–156, January 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021646925303.

Lachal:2000:FET

[Lac00] Aimé Lachal. First exit time from a bounded interval for a certain class of additive functionals of Brownian motion. Journal of Theoretical Probability, 13(3):733–775, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007810528683.

Lacaux:2005:FET

[Lac05] Céline Lacaux. Fields with exceptional tangent fields. Journal of Theoretical Probability, 18(2):481–497, April 2005. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
005-3516-7.

Lachal:2006:SED

[Lac06] Aimé Lachal. Some explicit distributions related to the first exit time from a bounded interval for certain functionals of Brownian motion. Journal of Theoretical Probability, 19(4):757–771, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-006-0039-9.

Lalley:1991:SPA

[Lal91] Steven P. Lalley. Saddle-point approximations and space-time Martin boundary for nearest-neighbor random walk on a homogeneous tree. Journal of Theoretical Probability, 4(4):701–723, October 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01259551.

Lalley:1995:RPR

[Lal95] Steven P. Lalley. Return probabilities for random walk on a halfline. Journal of Theoretical Probability, 8(3):571–599, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02218045.

Lalley:1999:RST

[Lal99] Steven P. Lalley. Rifle shuffles and their associated dynamical systems. Journal of Theoretical Probability, 12(4):903–932, October 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021636902356.

Lasserre:1996:EUI

[Las96] Jean B. Lasserre. Existence and uniqueness of an invariant probability for a class of Feller Markov chains. Journal of Theoretical Probability, 9(3):595–612, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214077.

Last:2010:SRM

[Las10] Günter Last. Stationary random measures on homogeneous spaces. Journal of Theoretical Probability, 23(2):478–497, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-009-0231-9.

Last:2020:ICD

[Las20] Günter Last. An integral characterization of the Dirichlet process. Journal of Theoretical Probability, 33(2):918–930, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Lawi:2008:TCM

[Law08] Stephan Lawi. Towards a characterization of Markov processes enjoying the time-inversion property. Journal of Theoretical Probability, 21(1):144–168, March 2008. CODEN JTPREO.

ISSN 0894-9840 (print), 1572-9230 (electronic). URL http:/

/link.springer.com/article/10.1007/s10959-007-0104-z.

Lam:2011:CCL

[LBBB11] Henry Lam, Jose Blanchet, Damian Burch, and Martin Z. Bazant. Corrections to the Central Limit Theorem for heavytailed probability densities. Journal of Theoretical Probability, 24(4):895–927, December 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-011-0379-y.

Li:2012:APM

[LCP12] Junping Li, Anyue Chen, and Anthony G. Pakes. Asymptotic properties of the Markov branching process with immigration. Journal of Theoretical Probability, 25(1):122–143, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-010-0301-z.

LeGuevel:2019:HDR

[Le 19] R. Le Guével. The Hausdorff dimension of the range of the Lévy multistable processes. Journal of Theoretical Probability, 32(2): 765–780, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Ledoux:2018:LIL

[Led18] Michel Ledoux. A law of the iterated logarithm for directed last passage percolation. Journal of Theoretical Probability, 31(4): 2366–2375, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Lee:1997:CGL

[Lee97] Sungchul Lee. The continuity of m and n in greedy lattice animals. Journal of Theoretical Probability, 10(1):87–100, January 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1022642314829.

Lee:2018:GMI

[Lee18] Kai Lee. Generation and motion of interfaces in one-dimensional stochastic Allen–Cahn equation. Journal of Theoretical Probability, 31(1):268–293, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Lehtomaa:2015:ABR

[Leh15] Jaakko Lehtomaa. Asymptotic behaviour of ruin probabilities in a general discrete risk model using moment indices. Journal of Theoretical Probability, 28(4):1380–1405, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-014-0547-y.

Leitner:1991:SPI

[Lei91] M. Leitner. Stochastic processes indexed by hypergroups II. Journal of Theoretical Probability, 4(2):321–332, April 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01258740.

Lemanczyk:2021:GBL

[Lem21] Michał Lemańczyk. General Bernstein-like inequality for additive functionals of Markov chains. Journal of Theoretical Probability, 34(3):1426–1454, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-020-01006-z.

Leskela:2010:SRR

[Les10] Lasse Leskelä. Stochastic relations of random variables and processes. Journal of Theoretical Probability, 23(2):523–546, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-009-0216-8.

Letac:1989:CWE

[Let89] Gérard Letac. A characterization of the Wishart exponential families by an invariance property. Journal of Theoretical Probability, 2(1):71–86, January 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01048270.

Letac:1993:ECM

[Let93] Gérard Letac. An explicit calculation of the mean of the perimeter of the convex hull of a plane random walk. Journal of Theoretical Probability, 6(2):385–387, April 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01047580.

Levental:1989:UCU

[Lev89] Shlomo Levental. A uniform CLT for uniformly bounded families of martingale differences. Journal of Theoretical Probability, 2 (3):271–287, July 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01054016.

Lewis:1992:SNL

[Lew92] Thomas M. Lewis. A self normalized law of the iterated logarithm for random walk in random scenery. Journal of Theoretical Probability, 5(4):629–659, October 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01058723.

Lewis:1993:LIL

[Lew93] Thomas M. Lewis. A law of the iterated logarithm for random walk in random scenery with deterministic normalizers. Journal of Theoretical Probability, 6(2):209–230, April 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01047572.

Lewis:2001:LLH

[Lew01] Thomas M. Lewis. The length of the longest head-run in a model with long range dependence. Journal of Theoretical Probability, 14(2):357–378, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1011107629319.

Lin:2019:SGG

[LG19] Lin Lin and Hongjun Gao. A stochastic generalized Ginzburg– Landau equation driven by jump noise. Journal of Theoretical Probability, 32(1):460–483, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Lv:2021:HSH

[LH21] Wujun Lv and Xing Huang. Harnack and shift Harnack inequalities for degenerate (functional) stochastic partial differential equations with singular drifts. Journal of Theoretical Probability, 34(2):827–851, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-00989-z.

Lan:2021:PCE

[LHS21] Guolie Lan, Ze-Chun Hu, and Wei Sun. Products of conditional expectation operators: Convergence and divergence. Journal of Theoretical Probability, 34(2):1012–1028, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
01000-5.

Li:1992:CRL

[Li92] Wenbo V. Li. Comparison results for the lower tail of Gaussian seminorms. Journal of Theoretical Probability, 5(1):1–31, January 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01046776.

Li:1999:SDG

[Li99] Wenbo V. Li. Small deviations for Gaussian Markov processes under the sup–norm. Journal of Theoretical Probability, 12(4): 971–984, October 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021771503265.

Li:2016:DFH

[Li16] Huaiqian Li. Dimension-free Harnack inequalities on RCD(K, ∞) spaces. Journal of Theoretical Probability, 29(4):1280–1297, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-015-0621-0.

Li:2018:PTA

[Li18] Li Li. Phase transition for accessibility percolation on hypercubes. Journal of Theoretical Probability, 31(4):2072–2111, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Li:2020:PPB

[Li20] Xinyi Li. Percolative properties of Brownian interlacements and its vacant set. Journal of Theoretical Probability, 33(4):1855– 1893, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-019-00944-7.

Li:2022:CHT

[Li22] Libo Li. Characterisation of honest times and optional semimartingales of class-Σ. Journal of Theoretical Probability, 35(4): 2145–2175, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01154-w.

Liao:1997:RMR

[Lia97] Ming Liao. Random motion of a rigid body. Journal of Theoretical Probability, 10(1):201–211, January 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022654717555.

Liao:1999:ISF

[Lia99] Ming Liao. Isometric stochastic flows on spheres. Journal of Theoretical Probability, 12(2):475–488, April 1999. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021634329189.

Liao:2009:DMP

[Lia09] Ming Liao. A decomposition of Markov processes via group actions. Journal of Theoretical Probability, 22(1):164–185, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-008-0162-x.

Liao:2014:ILP

[Lia14] Ming Liao. Inhomogeneous Lévy processes in Lie groups and homogeneous spaces. Journal of Theoretical Probability, 27 (2):315–357, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0415-6.

Lin:2012:SOS

[Lin12] Jianxi Lin. Second order subexponential distributions with finite mean and their applications to subordinated distributions. Journal of Theoretical Probability, 25(3):834–853, September 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-010-0330-7.

Lin:2018:TBH

[Lin18] Shen Lin. Typical behavior of the harmonic measure in critical Galton–Watson trees with infinite variance offspring distribution. Journal of Theoretical Probability, 31(3):1469–1511, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Liu:2012:RPR

[Liu12] Kai Liu. On regularity property of retarded Ornstein–Uhlenbeck processes in Hilbert spaces. Journal of Theoretical Probability, 25(2):565–593, June 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0374-3.

Liu:2018:SSD

[Liu18] Kai Liu. Sensitivity to small delays of pathwise stability for stochastic retarded evolution equations. Journal of Theoretical Probability, 31(3):1625–1646, September 2018. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-017-0750-8.pdf.

Liu:2021:GTB

[Liu21] Guomin Liu. Girsanov theorem for G-Brownian motion: The degenerate case. Journal of Theoretical Probability, 34(1):125– 140, March 2021. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-020-00997-z.

Lasser:1989:SPI

[LL89] R. Lasser and M. Leitner. Stochastic processes indexed by hypergroups. I. Journal of Theoretical Probability, 2(3):301–311, July 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01054018.

Larsen:1999:CPI

[LL99] Michael Larsen and Russell Lyons. Coalescing particles on an interval. Journal of Theoretical Probability, 12(1):201–205, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021704912660.

Louani:2003:FDS

[LL03] Djamal Louani and Alain Lucas. Fractal dimensions for some increments of the uniform empirical process. Journal of Theoretical Probability, 16(1):59–86, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022274303725.

Li:2004:SDS

[LL04] Wenbo V. Li and Werner Linde. Small deviations of stable processes via metric entropy. Journal of Theoretical Probability, 17 (1):261–284, January 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000020484.26184.c4.

Locherbach:2012:DIC

[LL12] Eva Löcherbach and Dasha Loukianova. Deviation inequalities for centered additive functionals of recurrent Harris processes having general state space. Journal of Theoretical Probability, 25(1):231–261, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0310-y.

Li:2017:GWD

[LL17] Xinpeng Li and Yiqing Lin. Generalized Wasserstein distance and weak convergence of sublinear expectations. Journal of Theoretical Probability, 30(2):581–593, June 2017. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Li:2021:LDR

[LL21] Liuyan Li and Junping Li. Large deviation rates for supercritical branching processes with immigration. Journal of Theoretical Probability, 34(1):162–172, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-019-00968-z.

Li:2023:OII

[LL23] Liping Li and Hanlai Lin. On order isomorphisms intertwining semigroups for Dirichlet forms. Journal of Theoretical Probability, 36(2):1304–1320, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-022-01200-1.

Li:2024:MDR

[LL24a] Hanwu Li and Guomin Liu. Multi-dimensional reflected backward stochastic differential equations driven by G-Brownian motion with diagonal generators. Journal of Theoretical Probability, 37(3):2615–2645, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-024-01334-4.

Li:2024:HML

[LL24b] Liuyan Li and Junping Li. Harmonic moments and large deviations for the Markov branching process with immigration. Journal of Theoretical Probability, 37(2):1397–1416, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01280-7.

LeGuevel:2015:TME

[LLL15] Ronan Le Guével, Jacques Lévy Véhel, and Lining Liu. On two multistable extensions of stable Lévy motion and their semimartingale representations. Journal of Theoretical Probability, 28(3):1125–1144, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-013-0528-6.

Li:2011:LDE

[LLW11] Yi-Ting Li, Dang-Zheng Liu, and Zheng-Dong Wang. Limit distributions of eigenvalues for random block Toeplitz and Hankel matrices. Journal of Theoretical Probability, 24(4):1063– 1086, December 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0326-3.

Li:2008:CDS

[LM08] Zenghu Li and Chunhua Ma. Catalytic discrete state branching models and related limit theorems. Journal of Theoretical Probability, 21(4):936–965, December 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-008-0161-y.

Louani:2012:SFL

[LM12] Djamal Louani and Sidi Mohamed Ould Maouloud. Some functional large deviations principles in nonparametric function estimation. Journal of Theoretical Probability, 25(1):280–309, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-011-0342-y.

Lan:2016:CMT

[LM16a] Kathleen Lan and Kevin McGoff. Coalescence and meeting times on n-block Markov chains. Journal of Theoretical Probability, 29 (2):527–549, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0579-3.

Lanchier:2016:FRT

[LM16b] Nicolas Lanchier and Paul-Henri Moisson. Fixation results for the two-feature Axelrod model with a variable number of opinions. Journal of Theoretical Probability, 29(4):1554–1580, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-015-0623-y.

Liu:2019:CFL

[LM19] Yonghong Liu and Yongxiang Mo. Chung’s functional law of the iterated logarithm for increments of a fractional Brownian motion. Journal of Theoretical Probability, 32(2):721–736, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Lu:2020:ICM

[LM20] Tianshi Lu and Chunsheng Ma. Isotropic covariance matrix functions on compact two-point homogeneous spaces. Journal of Theoretical Probability, 33(3):1630–1656, September 2020. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Ludvik:2022:SFH

[LM22] Pavel Ludvík and Václav Mácha. Stochastic forcing in hydrodynamic models with non-local interactions. Journal of Theoretical Probability, 35(4):2806–2852, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-021-
01137-x.

Lowe:2002:MDL

[LMR02] Matthias Löwe, Franz Merkl, and Silke Rolles. Moderate deviations for longest increasing subsequences: The lower tail. Journal of Theoretical Probability, 15(4):1031–1047, October 2002.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1020649006254.

Ledoan:2012:UPG

[LMS12] Andrew Ledoan, Marco Merkli, and Shannon Starr. A universality property of Gaussian analytic functions. Journal of Theoretical Probability, 25(2):496–504, June 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-011-0356-5.

Last:2024:NAK

[LMS24a] G. Last, I. Molchanov, and M. Schulte. Normal approximation of Kabanov–Skorohod integrals on Poisson spaces. Journal of Theoretical Probability, 37(2):1124–1167, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01287-0.

Li:2024:EAI

[LMS24b] Xiaoyue Li, Xuerong Mao, and Guoting Song. Explicit approximation of invariant measure for stochastic delay differential equations with the nonlinear diffusion term. Journal of Theoretical Probability, 37(2):1850–1881, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-023-01290-5.

Liao:2022:SFP

[LMX22] Zhong-Wei Liao, Yutao Ma, and Aihua Xia. On Stein’s factors for Poisson approximation in Wasserstein distance with nonlinear transportation costs. Journal of Theoretical Probability, 35(4):2383–2412, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01129-x.

Lu:2023:SLN

[LMX23] Tianshi Lu, Chunsheng Ma, and Yimin Xiao. Strong local nondeterminism and exact modulus of continuity for isotropic Gaussian random fields on compact two-point homogeneous spaces. Journal of Theoretical Probability, 36(4):2403–2425, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01231-8.

Locherbach:2019:LDC

[Löc19] E. Löcherbach. Large deviations for cascades of diffusions arising in oscillating systems of interacting Hawkes processes. Journal of Theoretical Probability, 32(1):131–162, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Locherbach:2020:CET

[Löc20] E. Löcherbach. Convergence to equilibrium for time-inhomogeneous jump diffusions with state-dependent jump intensity. Journal of Theoretical Probability, 33(4):2280–2314, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-019-00947-4.

Loper:2020:UEB

[Lop20] Jackson Loper. Uniform ergodicity for Brownian motion in a bounded convex set. Journal of Theoretical Probability, 33 (1):22–35, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-018-0848-7.pdf.

Loukissas:2012:PLD

[Lou12] Fotis Loukissas. Precise large deviations for long-tailed distributions. Journal of Theoretical Probability, 25(4):913–924, December 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-011-0367-2.

Loubaton:2016:ASL

[Lou16] Philippe Loubaton. On the almost sure location of the singular values of certain Gaussian block-Hankel large random matrices. Journal of Theoretical Probability, 29(4):1339–1443, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-015-0614-z.

Lawler:1993:HID

[LP93] Gregory F. Lawler and Thomas W. Polaski. Harnack inequalities and difference estimates for random walks with infinite range. Journal of Theoretical Probability, 6(4):781–802, October 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049175.

Lewis:2003:TPC

[LP03] Thomas M. Lewis and Geoffrey Pritchard. Tail properties of correlation measures. Journal of Theoretical Probability, 16 (3):771–788, July 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1025680718292.

Letac:2018:RWH

[LP18] Gérard Letac and Mauro Piccioni. Random walks in the hyperbolic plane and the Minkowski question mark function. Journal of Theoretical Probability, 31(4):2376–2389, December 2018. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Lechner:2017:EAO

[LPP17] Richard Lechner, Markus Passenbrunner, and Joscha Prochno. Estimating averages of order statistics of bivariate functions. Journal of Theoretical Probability, 30(4):1445–1470, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Li:2022:HPC

[LPW22] Zenghu Li, Etienne Pardoux, and Anton Wakolbinger. The height process of a continuous-state branching process with interaction. Journal of Theoretical Probability, 35(1):142–185, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01054-5.

Li:2011:RKM

[LQR11] Deli Li, Yongcheng Qi, and Andrew Rosalsky. A refinement of the Kolmogorov–Marcinkiewicz–Zygmund strong law of large numbers. Journal of Theoretical Probability, 24(4):1130– 1156, December 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0308-5.

Lii:1988:EDW

[LR88] Keh-Shin Lii and Murray Rosenblatt. Estimation and deconvolution when the transfer function has zeros. Journal of Theoretical Probability, 1(1):93–113, January 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01076289.

Lii:1990:ANC

[LR90] K. S. Lii and M. Rosenblatt. Asymptotic normality of cumulant spectral estimates. Journal of Theoretical Probability, 3 (2):367–385, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01045168.

Lii:1993:BPN

[LR93] Keh-Shin Lii and Murray Rosenblatt. Bispectra and phase of non-Gaussian linear processes. Journal of Theoretical Probability, 6(3):579–593, July 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01066718.

Lachieze-Rey:2012:CMS

[LR12] Raphaël Lachièze-Rey. Concave majorant of stochastic processes and Burgers turbulence. Journal of Theoretical Probability, 25 (2):313–332, June 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0354-7.

Li:1995:CCA

[LRJW95] Deli Li, M. Bhaskara Rao, Tiefeng Jiang, and Xiangchen Wang. Complete convergence and almost sure convergence of weighted sums of random variables. Journal of Theoretical Probability, 8 (1):49–76, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213454.

Liu:2003:LTM

[LRR03] Quansheng Liu, Emmanuel Rio, and Alain Rouault. Limit theorems for multiplicative processes. Journal of Theoretical Probability, 16(4):971–1014, October 2003. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000012003.49768.f6.

Liu:2011:CSB

[LRS11] Rong-Li Liu, Yan-Xia Ren, and Renming Song. L log L condition for supercritical branching hunt processes. Journal of Theoretical Probability, 24(1):170–193, March 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-010-0322-7.

Li:1999:SBE

[LS99] Wenbo V. Li and Qi-Man Shao. Small ball estimates for Gaussian processes under Sobolev type norms. Journal of Theoretical Probability, 12(3):699–720, July 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1021675731663.

Li:2005:RCR

[LS05] Deli Li and Aurel Spataru. Refinement of convergence rates for tail probabilities. Journal of Theoretical Probability, 18(4): 933–947, October 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-005-7534-2.

Liang:2006:SNL

[LS06] Han-Ying Liang and Josef G. Steinebach. Self-normalized LIL for Hanson–Russo type increments. Journal of Theoretical Probability, 19(1):70–88, January 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-006-0011-8.

Lu:2011:FET

[LS11] Dawei Lu and Lixin Song. The first exit time of a Brownian motion from the minimum and maximum parabolic domains. Journal of Theoretical Probability, 24(4):1028–1043, December 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-010-0306-7.

Lambert:2015:ABL

[LS15a] Amaury Lambert and Florian Simatos. Asymptotic behavior of local times of compound Poisson processes with drift in the infinite variance case. Journal of Theoretical Probability, 28 (1):41–91, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0492-1.

Lin:2015:MDP

[LS15b] Gwo Dong Lin and Jordan Stoyanov. Moment determinacy of powers and products of nonnegative random variables. Journal of Theoretical Probability, 28(4):1337–1353, December 2015.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-014-0546-z.

Lember:2021:RVP

[LS21a] Jüri Lember and Joonas Sova. Regenerativity of Viterbi process for pairwise Markov models. Journal of Theoretical Probability, 34(1):1–33, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01022-z.

Li:2021:BSD

[LS21b] Hanwu Li and Yongsheng Song. Backward stochastic differential equations driven by G-Brownian motion with double reflections. Journal of Theoretical Probability, 34(4):2285–2314, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01038-5.

Latz:2023:CMD

[LS23] Jan Niklas Latz and Jan M. Swart. Commutative monoid duality. Journal of Theoretical Probability, 36(2):1088–1115, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01197-7.

Li:1991:CSL

[LT91] Z. F. Li and R. J. Tomkins. Complete stability of large order statistics. Journal of Theoretical Probability, 4(1):213–221, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01047003.

Li:1998:CLI

[LT98] Deli Li and R. J. Tomkins. Compact laws of the iterated logarithm for B-valued random variables with two-dimensional indices. Journal of Theoretical Probability, 11(2):443–459, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1022687923455.

Li:2003:LLW

[LT03] D. Li and R. J. Tomkins. The law of the logarithm for weighted sums of independent random variables. Journal of Theoretical Probability, 16(3):519–542, July 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025684513749.

Leon:2012:MCF

[LT12] Jorge A. León and Samy Tindel. Malliavin calculus for fractional delay equations. Journal of Theoretical Probability, 25(3): 854–889, September 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0349-4.

Lowe:2023:CLT

[LT23] Matthias Löwe and Sara Terveer. A central limit theorem for the mean starting hitting time for a random walk on a random graph. Journal of Theoretical Probability, 36(2):779–810, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01195-9.

Liu:2017:SPC

[LTF17] Wei Liu, Kuanhou Tian, and Mohammud Foondun. On some properties of a class of fractional stochastic heat equations. Journal of Theoretical Probability, 30(4):1310–1333, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-016-0684-6.pdf.

Lefevre:1998:OPP

[LU98] Claude Lefèvre and Sergey Utev. On order-preserving properties of probability metrics. Journal of Theoretical Probability, 11(4): 907–920, October 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022608613982.

Luczak:1997:ESO

[Luc97] Andrzej Luczak. Exponents and symmetry of operator Lévy’s probability measures on finite dimensional vector spaces. Journal of Theoretical Probability, 10(1):117–129, January 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022646415737.

Luczak:2010:CPP

[Luc10] Andrzej Luczak. Centering problems for probability measures on finite-dimensional vector spaces. Journal of Theoretical Probability, 23(3):770–791, September 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-010-0294-7.

Luczak:2024:NMS

[Luc24] Malwina Luczak. A note on the Markovian SIR epidemic on a random graph with given degrees. Journal of Theoretical Probability, 37(2):1039–1051, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-024-01320-w.

Luo:1998:NDW

[Luo98] Shunlong Luo. A nonlinear deformation of Wiener space. Journal of Theoretical Probability, 11(2):331–350, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022623603891.

Luo:2015:QIS

[Luo15] Dejun Luo. Quasi-invariance of the stochastic flow associated to Itô’s SDE with singular time-dependent drift. Journal of Theoretical Probability, 28(4):1743–1762, December 2015. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
014-0554-z.

Luo:2022:ATI

[Luo22] Sijie Luo. On Azuma-type inequalities for Banach spacevalued martingales. Journal of Theoretical Probability, 35(2): 772–800, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01086-5.

LeGall:2012:EPB

[LV12] Jean-François Le Gall and Amandine Véber. Escape probabilities for branching Brownian motion among soft obstacles. Journal of Theoretical Probability, 25(2):505–535, June 2012. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
011-0343-x.

Liu:2021:LDR

[LV21] Jingjia Liu and Quirin Vogel. Large deviations of the range of the planar random walk on the scale of the mean. Journal of Theoretical Probability, 34(4):2315–2345, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01039-4.

Li:2021:SBP

[LVZ21] Doudou Li, Vladimir Vatutin, and Mei Zhang. Subcritical branching processes in random environment with immigration stopped at zero. Journal of Theoretical Probability, 34(2):874– 896, June 2021. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-020-00991-5.

Lifshits:1997:SLI

[LW97] M. A. Lifshits and M. Weber. Strassen laws of iterated logarithm for partially observed processes. Journal of Theoretical Probability, 10(1):101–115, January 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022694331667.

Lin:2004:FLT

[LW04] Z. Lin and X. Wang. A functional limit theorem for observations that change with time. Journal of Theoretical Probability, 17(4): 887–903, October 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-004-0580-3.

Liu:2011:LDE

[LW11] Dang-Zheng Liu and Zheng-Dong Wang. Limit distribution of eigenvalues for random Hankel and Toeplitz band matrices. Journal of Theoretical Probability, 24(4):988–1001, December 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-009-0260-4.

Li:2012:GIE

[LW12] Wenbo V. Li and Ang Wei. A Gaussian inequality for expected absolute products. Journal of Theoretical Probability, 25(1):92–99, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0329-0.

Lin:2016:CSI

[LW16] Zhengyan Lin and Hanchao Wang. On convergence to stochastic integrals. Journal of Theoretical Probability, 29(3):717– 736, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0598-8.

Luo:2019:VSD

[LW19] Peng Luo and Falei Wang. Viability for stochastic differential equations driven by G-Brownian motion. Journal of Theoretical Probability, 32(1):395–416, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Li:2023:WCR

[LW23] Huaiqian Li and Bingyao Wu. Wasserstein convergence rates for empirical measures of subordinated processes on noncompact manifolds. Journal of Theoretical Probability, 36(2):1243– 1268, June 2023. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-022-01196-8.

Lv:2024:CMV

[LWW24] Guangying Lv, Wei Wang, and Jinlong Wei. Coupled McKean– Vlasov equations over convex domains. Journal of Theoretical Probability, 37(2):1824–1849, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-023-01303-3.

Li:2024:LDP

[LWZ24] Ruinan Li, Ran Wang, and Beibei Zhang. A large deviation principle for the stochastic heat equation with general rough noise. Journal of Theoretical Probability, 37(1):251–306, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01228-3.

Li:2012:OTF

[LX12] Yuqiang Li and Yimin Xiao. Occupation time fluctuations of weakly degenerate branching systems. Journal of Theoretical Probability, 25(4):1119–1152, December 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
011-0358-3.

Luks:2017:DPO

[LX17] Tomasz Luks and Yimin Xiao. On the double points of operator stable Lévy processes. Journal of Theoretical Probability, 30 (1):297–325, March 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0638-4.

Luks:2020:MPO

[LX20] Tomasz Luks and Yimin Xiao. Multiple points of operator semistable Lévy processes. Journal of Theoretical Probability, 33(1):153–179, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Li:2019:FPS

[LXY19] Bo Li, Yimin Xiao, and Xiaochuan Yang. On the favorite points of symmetric Lévy processes. Journal of Theoretical Probability, 32(4):1943–1972, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Liu:2016:SNF

[LY16] Junfeng Liu and Litan Yan. Solving a nonlinear fractional stochastic partial differential equation with fractional noise. Journal of Theoretical Probability, 29(1):307–347, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-014-0578-4.

Li:2019:ESS

[LY19] Zhi Li and Litan Yan. Ergodicity and stationary solution for stochastic neutral retarded partial differential equations driven by fractional Brownian motion. Journal of Theoretical Probability, 32(3):1399–1419, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Lyons:2000:SSR

[Lyo00] Russell Lyons. Singularity of some random continued fractions. Journal of Theoretical Probability, 13(2):535–545, April 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007837306508.

Lytova:2018:CLT

[Lyt18] A. Lytova. Central limit theorem for linear eigenvalue statistics for a tensor product version of sample covariance matrices. Journal of Theoretical Probability, 31(2):1024–1057, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10959-017-0741-9.pdf.

Liu:2016:TRG

[LZ16] Yong Liu and Jianliang Zhai. Time regularity of generalized Ornstein–Uhlenbeck processes with Lévy noises in Hilbert spaces. Journal of Theoretical Probability, 29(3):843–866, September 2016. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL http://link.springer.com/
article/10.1007/s10959-015-0594-z.

Lan:2019:RCW

[LZ19] Shi-Yi Lan and Wang Zhou. Random conformal welding for finitely connected regions. Journal of Theoretical Probability, 32 (2):659–683, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Liu:2024:EMC

[LZ24] Huili Liu and Xiaowen Zhou. Exact modulus of continuities for Λ-Fleming–Viot processes with Brownian spatial motion. Journal of Theoretical Probability, 37(2):1710–1744, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-024-01326-4.

Ma:2017:TVI

[Ma17] Chunsheng Ma. Time varying isotropic vector random fields on spheres. Journal of Theoretical Probability, 30(4):1763–1785, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Ma:2023:VRF

[Ma23] Chunsheng Ma. Vector random fields on the probability simplex with metric-dependent covariance matrix functions. Journal of Theoretical Probability, 36(3):1922–1938, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01217-6.

Ma:2024:BBM

[Ma24] Chunsheng Ma. Bifractional Brownian motions on metric spaces. Journal of Theoretical Probability, 37(2):1299–1332, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01284-3.

Madaule:2017:CLB

[Mad17] Thomas Madaule. Convergence in law for the branching random walk seen from its tip. Journal of Theoretical Probability, 30 (1):27–63, March 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0636-6.

Maejima:1993:TSD

[Mae93] Makoto Maejima. Trimmed sums of i.i.d. Banach space valued random variables. Journal of Theoretical Probability, 6(1):57–69, January 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01046768.

Majumdar:1999:LTB

[Maj99] Suman Majumdar. On a Laplace transform based metric for probabilities on a Hilbert space. Journal of Theoretical Probability, 12(1):1–11, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1021784208117.

Maltz:1999:CLT

[Mal99] Alberto L. Maltz. On the Central Limit Theorem for nonuniform ϕ-mixing random fields. Journal of Theoretical Probability, 12 (3):643–660, July 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021619613916.

Malyarenko:2006:FLT

[Mal06] Anatoliy Malyarenko. Functional limit theorems for multiparameter fractional Brownian motion. Journal of Theoretical Probability, 19(2):263–288, June 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-006-0014-5.

Malyarenko:2008:OSE

[Mal08] Anatoliy Malyarenko. An optimal series expansion of the multiparameter fractional Brownian motion. Journal of Theoretical Probability, 21(2):459–475, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0122-x.

Maleki:2019:OSI

[Mal19] Yasaman Maleki. Optimal scale invariant Wigner spectrum estimation of Gaussian locally self-similar processes using Hermite functions. Journal of Theoretical Probability, 32(1):202– 215, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Mansuy:2004:OPG

[Man04] R. Mansuy. On a one-parameter generalization of the Brownian bridge and associated quadratic functionals. Journal of Theoretical Probability, 17(4):1021–1029, October 2004. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
004-0588-8.

Mao:1991:EUA

[Mao91a] Xuerong Mao. Eventual uniform asymptotic stability for stochastic differential equation based on semimartingale with spatial parameter. Journal of Theoretical Probability, 4(2): 441–456, April 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01258747.

Mao:1991:NGS

[Mao91b] Xuerong Mao. A note on global solution to stochastic differential equation based on a semimartingale with spatial parameters. Journal of Theoretical Probability, 4(1):161–167, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01046999.

Mao:1992:SSD

[Mao92] Xuerong Mao. Solutions of stochastic differential-functional equations via bounded stochastic integral contractors. Journal of Theoretical Probability, 5(3):487–502, July 1992. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01060431.

Marraffa:1988:ASC

[Mar88] Valeria Marraffa. On almost sure convergence of amarts and martingales without the Radon–Nikodym property. Journal of Theoretical Probability, 1(3):255–261, July 1988. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01246628.

Marcus:2001:MVS

[Mar01] Michael B. Marcus. The most visited sites of certain Lévy processes. Journal of Theoretical Probability, 14(3):867–885, July 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1012295810270.

Martin:2004:SBA

[Mar04] A. Martin. Small ball asymptotics for the stochastic wave equation. Journal of Theoretical Probability, 17(3):693–703, July 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOTP.0000040294.12188.cd.

Markering:2023:LDP

[Mar23] Maarten Markering. The large deviation principle for inhomogeneous Erdős–Rényi random graphs. Journal of Theoretical Probability, 36(2):711–727, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-022-01181-1.

Mason:2006:CSS

[Mas06] David M. Mason. Cluster sets of self-normalized sums. Journal of Theoretical Probability, 19(4):911–930, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-006-0042-1.

Mason:2013:EAD

[Mas13] David M. Mason. Erratum to: The Asymptotic Distribution of Self-Normalized Triangular Arrays. Journal of Theoretical Probability, 26(2):589–593, June 2013. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s10959-
013-0478-z.pdf. See [Dav05].

Mason:2021:SSC

[Mas21] David M. Mason. Self-standardized central limit theorems for trimmed Lévy processes. Journal of Theoretical Probability, 34(4):2117–2144, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-020-01021-0.

Matthews:1988:SUT

[Mat88] Peter Matthews. A strong uniform time for random transpositions. Journal of Theoretical Probability, 1(4):411–423, October 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01048728.

Matthews:1989:SSP

[Mat89] Peter Matthews. Some sample path properties of a random walk on the cube. Journal of Theoretical Probability, 2(1):129–146, January 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01048275.

Matsumoto:2012:GMI

[Mat12] Sho Matsumoto. General moments of the inverse real Wishart distribution and orthogonal Weingarten functions. Journal of Theoretical Probability, 25(3):798–822, September 2012. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
011-0340-0.

Matysiak:2017:HQB

[Mat17] Wojciech Matysiak. Hypergroups and quantum Bessel processes of non-integer dimensions. Journal of Theoretical Probability, 30 (4):1677–1691, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Meyer-Brandis:2010:ERS

[MBP10] Thilo Meyer-Brandis and Frank Proske. Explicit representation of strong solutions of SDEs driven by infinite-dimensional Lévy processes. Journal of Theoretical Probability, 23(1):301–314, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-009-0226-6.

McCollum:2011:RWD

[McC11] Joseph McCollum. Random walks on dihedral groups. Journal of Theoretical Probability, 24(2):397–408, June 2011. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0307-6.

McGuinness:1991:RNT

[McG91] Sean McGuinness. Recurrent networks and a theorem of Nash– Williams. Journal of Theoretical Probability, 4(1):87–100, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01046995.

Marquez-Carreras:2008:ILL

[MCR08] David Márquez-Carreras and Carles Rovira. Iterated logarithm law for anticipating stochastic differential equations. Journal of Theoretical Probability, 21(3):650–659, September 2008. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0114-x.

Miranda:2007:HMP

[MdCQ07] Enrique Miranda, Gert de Cooman, and Erik Quaeghebeur. The Hausdorff moment problem under finite additivity. Journal of Theoretical Probability, 20(3):663–693, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-007-0055-4.

Meckes:2012:APR

[Mec12] Elizabeth Meckes. Approximation of projections of random vectors. Journal of Theoretical Probability, 25(2):333–352, June 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-010-0299-2.

Meester:1992:CFP

[Mee92] Ronald W. J. Meester. Connectivity in fractal percolation. Journal of Theoretical Probability, 5(4):775–789, October 1992.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01058729.

Meerschaert:1994:NOG

[Mee94] Mark M. Meerschaert. Norming operators for generalized domains of attraction. Journal of Theoretical Probability, 7(4): 793–798, October 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214372.

Megyesi:2002:DGP

[Meg02] Zoltán Megyesi. Domains of geometric partial attraction of maxsemistable laws: Structure, merge and almost sure limit theorems. Journal of Theoretical Probability, 15(4):973–1005, October 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020692805345.

Mendelson:2016:DTT

[Men16] Shahar Mendelson. Dvoretzky type theorems for subgaussian coordinate projections. Journal of Theoretical Probability, 29(4): 1644–1660, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0624-x.

Menguturk:2023:IMF

[Men23] Levent Ali Mengütürk. From irrevocably modulated filtrations to dynamical equations over random networks. Journal of Theoretical Probability, 36(2):845–875, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-022-01201-0.

Merlevede:2003:CLT

[Mer03] Florence Merlevède. On the Central Limit Theorem and its weak invariance principle for strongly mixing sequences with values in a Hilbert space via martingale approximation. Journal of Theoretical Probability, 16(3):625–653, July 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1025668415566.

Mikosch:1994:WIP

[Mik94] Thomas Mikosch. Weak invariance principles for weighted Ustatistics. Journal of Theoretical Probability, 7(1):147–173, January 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02213365.

Miller:1994:TTR

[Mil94] Curtis Miller. Three theorems on ρ^{∗} random fields. Journal of Theoretical Probability, 7(4):867–882, October 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02214377.

Mili:2000:SLT

[Mil00] Maher Mili. Selective limit theorems for random walks on parabolic biangle and triangle hypergroups. Journal of Theoretical Probability, 13(3):717–731, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1007858411844.

Milman:2009:GMU

[Mil09] Emanuel Milman. On Gaussian marginals of uniformly convex bodies. Journal of Theoretical Probability, 22(1):256–278, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0160-z.

Milos:2018:SCL

[Mil18] Piotr Miloś. Spatial central limit theorem for supercritical superprocesses. Journal of Theoretical Probability, 31(1):1–40, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10959-016-0704-6.pdf.

Mimica:2013:HIH

[Mim13] Ante Mimica. Harnack inequality and Hölder regularity estimates for a Lévy process with small jumps of high intensity. Journal of Theoretical Probability, 26(2):329–348, June 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-011-0361-8.

Mindlin:1990:CRM

[Min90] D. S. Mindlin. Convolutions of random measures on a compact topological semigroup. Journal of Theoretical Probability, 3(2):181–197, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01045157.

Minh:2021:RDB

[Min21] Hà Quang Minh. Regularized divergences between covariance operators and Gaussian measures on Hilbert spaces. Journal of Theoretical Probability, 34(2):580–643, June 2021. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
01003-2.

Maejima:1999:SSP

[MiS99] Makoto Maejima and Ken iti Sato. Semi-selfsimilar processes. Journal of Theoretical Probability, 12(2):347–373, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021621926463.

Mirakhmedov:2014:EEG

[MJM14] S. M. Mirakhmedov, S. Rao Jammalamadaka, and Ibrahim B. Mohamed. On Edgeworth expansions in generalized urn models. Journal of Theoretical Probability, 27(3):725–753, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-012-0454-z.

Meng:2017:DPT

[ML17] Wei Wei Meng and Jun Ping Li. Decay properties of n-type Markov branching processes with disasters. Journal of Theoretical Probability, 30(4):1605–1623, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Mayato:2022:GIP

[MLC22] Rafael Sala Mayato, Patrick Loughlin, and Leon Cohen. Generating M-indeterminate probability densities by way of quantum mechanics. Journal of Theoretical Probability, 35(3):1537– 1555, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01121-5.

Marckert:2003:SSS

[MM03] Jean-François Marckert and Abdelkader Mokkadem. States [sic] spaces of the snake and its tour-convergence of the discrete snake. Journal of Theoretical Probability, 16(4):1015–1046, October 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOTP.0000012004.73051.f3. See note [Git03].

Misiewicz:2005:PRV

[MM05] J. K. Misiewicz and G. Mazurkiewicz. On (c, p)-pseudostable random variables. Journal of Theoretical Probability, 18(4): 837–852, October 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-005-7528-0.

Meckes:2007:CLP

[MM07] Elizabeth S. Meckes and Mark W. Meckes. The central limit problem for random vectors with symmetries. Journal of Theoretical Probability, 20(4):697–720, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0119-5.

Mirakhmedov:2010:AER

[MM10] Saidbek S. Mirakhmedov and Sherzod M. Mirakhmedov. On asymptotic expansion in the random allocation of particles by sets. Journal of Theoretical Probability, 23(3):904–919, September 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-009-0225-7.

Ma:2020:TVI

[MM20] Chunsheng Ma and Anatoliy Malyarenko. Time-varying isotropic vector random fields on compact two-point homogeneous spaces. Journal of Theoretical Probability, 33(1):319–339, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10959-018-0872-7.pdf.

Michta:2022:SVF

[MM22] Mariusz Michta and Jerzy Motyl. Set-valued functions of bounded generalized variation and set-valued Young integrals. Journal of Theoretical Probability, 35(1):528–549, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01059-0.

Massey:2007:DER

[MMS07] Adam Massey, Steven J. Miller, and John Sinsheimer. Distribution of eigenvalues of real symmetric palindromic Toeplitz matrices and circulant matrices. Journal of Theoretical Probability, 20(3):637–662, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0078-x.

MacPhee:2013:MET

[MMW13] Iain M. MacPhee, Mikhail V. Menshikov, and Andrew R. Wade. Moments of exit times from wedges for non-homogeneous random walks with asymptotically zero drifts. Journal of Theoretical Probability, 26(1):1–30, March 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-012-0411-x.

Merzbach:1989:GHP

[MN89] Ely Merzbach and David Nualart. Generalized holomorphic processes and differentiability. Journal of Theoretical Probability, 2 (4):419–432, October 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01051875.

Massam:1997:TDW

[MN97] Hélène Massam and Erhard Neher. On transformations and determinants of Wishart variables on symmetric cones. Journal of Theoretical Probability, 10(4):867–902, October 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022658415699.

Moret:2000:QCI

[MN00] S. Moret and D. Nualart. Quadratic covariation and Itô’s formula for smooth nondegenerate martingales. Journal of Theoretical Probability, 13(1):193–224, January 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1007791027791.

Mukherjea:2002:CSL

[MN02] A. Mukherjea and A. Nakassis. On the continuous singularity of the limit distribution of products of I.I.D. d × d stochastic matrices. Journal of Theoretical Probability, 15(4):903–918, October 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1020636603528.

Maticiuc:2015:FBS

[MN15] Lucian Maticiuc and Tianyang Nie. Fractional backward stochastic differential equations and fractional backward variational inequalities. Journal of Theoretical Probability, 28(1): 337–395, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0509-9.

Ma:2020:RCW

[MN20] Nicholas Ma and David Nualart. Rate of convergence for the weighted Hermite variations of the fractional Brownian motion. Journal of Theoretical Probability, 33(4):1919–1947, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-019-00940-x.

Mukherjea:1999:ADL

[MNR99a] A. Mukherjea, A. Nakassis, and J. S. Ratti. Addendum to “On the Distribution of the Limit of Products of i.i.d. 2 × 2 Random Stochastic Matrices”. Journal of Theoretical Probability, 12(4): 1109–1112, October 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021605522829.

Mukherjea:1999:DLP

[MNR99b] A. Mukherjea, A. Nakassis, and J. S. Ratti. On the distribution of the limit of products of I.I.D. 2 × 2 random stochastic matrices. Journal of Theoretical Probability, 12(2):571–583, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021694515077.

Mohan:2022:MSS

[Moh22] Manil T. Mohan. Mild solutions for the stochastic generalized Burgers–Huxley equation. Journal of Theoretical Probability, 35(3):1511–1536, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01100-w.

Montegut:2019:DAR

[Mon19] Fabien Montégut. Double asymptotic for random walks on hypercubes. Journal of Theoretical Probability, 32(4):2044–2065, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Monch:2022:UPE

[Mön22] Christian Mönch. Universality for persistence exponents of local times of self-similar processes with stationary increments. Journal of Theoretical Probability, 35(3):1842–1862, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01102-8.

Mountford:1997:LGR

[MOP97] T. S. Mountford, P. O’hara, and S. C. Port. Large gaps in the range of stable processes. Journal of Theoretical Probability, 10 (1):25–50, January 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022686129850.

Mori:1995:CBN

[Mór95] Tamás F. Móri. Covering with blocks in the non-symmetric case. Journal of Theoretical Probability, 8(1):139–164, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02213458.

Moreno:2010:CLE

[Mor10] Gregorio Moreno. Convergence of the law of the environment seen by the particle for directed polymers in random media in the L^{2} region. Journal of Theoretical Probability, 23(2):466–477, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0203-5.

Morrow:2021:MCT

[Mor21] Steven T. Morrow. An L^{p} multiplicative coboundary theorem for sequences of unitriangular random matrices. Journal of Theoretical Probability, 34(1):206–213, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-019-00981-2.

Mountford:2007:LSM

[Mou07] Thomas S. Mountford. Level sets of multiparameter stable processes. Journal of Theoretical Probability, 20(1):25–46, March 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-006-0034-1.

Merlevede:2006:WIP

[MP06] Florence Merlevède and Magda Peligrad. On the weak invariance principle for stationary sequences under projective criteria. Journal of Theoretical Probability, 19(3):647–689, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-006-0029-y.

Merlevede:2010:MDL

[MP10] Florence Merlevède and Magda Peligrad. Moderate deviations for linear processes generated by martingale-like random variables. Journal of Theoretical Probability, 23(1):277–300, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-009-0218-6.

Menshikov:2014:RLT

[MP14] Mikhail Menshikov and Serguei Popov. On range and local time of many-dimensional submartingales. Journal of Theoretical Probability, 27(2):601–617, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-012-0431-6.

Malyarenko:2024:MRF

[MP24a] Anatoliy Malyarenko and Emilio Porcu. Multivariate random fields evolving temporally over hyperbolic spaces. Journal of Theoretical Probability, 37(2):975–1000, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-024-
01316-6.

Masiero:2024:PSS

[MP24b] Federica Masiero and Enrico Priola. Partial smoothing of the stochastic wave equation and regularization by noise phenomena. Journal of Theoretical Probability, 37(3):2738–2774, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-024-01337-1.

Menoukeu-Pamen:2019:PUN

[MPOT19] Olivier Menoukeu-Pamen, Youssef Ouknine, and Ludovic Tangpi. Pathwise uniqueness of non-uniformly elliptic SDEs with rough coefficients. Journal of Theoretical Probability, 32 (4):1892–1908, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Menshikov:2002:CBO

[MPS02] M. V. Menshikov, S. Yu. Popov, and V. V. Sisko. On the connection between oriented percolation and contact process. Journal of Theoretical Probability, 15(1):207–221, January 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013847619585.

Merlevede:1997:SCC

[MPU97] Florence Merlevède, Magda Peligrad, and Sergey Utev. Sharp conditions for the CLT of linear processes in a Hilbert space. Journal of Theoretical Probability, 10(3):681–693, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022653728014.

Menshikov:2018:HTR

[MPW18] Mikhail V. Menshikov, Dimitri Petritis, and Andrew R. Wade. Heavy-tailed random walks on complexes of half–lines. Journal of Theoretical Probability, 31(3):1819–1859, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-017-0753-5.pdf.

Miao:2021:LSR

[MQ21] Yu Miao and Yongcheng Qi. Limiting spectral radii of circular unitary matrices under light truncation. Journal of Theoretical Probability, 34(4):2145–2165, December 2021. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-020-
01037-6.

Marcus:1992:MCL

[MR92] Michael B. Marcus and Jay Rosen. Moduli of continuity of local times of strongly symmetric Markov processes via Gaussian processes. Journal of Theoretical Probability, 5(4):791–825, October 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01058730.

Maejima:1996:RCO

[MR96] Makoto Maejima and Svetlozar T. Rachev. Rates of convergence in the operator-stable limit theorem. Journal of Theoretical Probability, 9(1):37–85, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02213734.

Mukherjea:1997:CSW

[MR97] A. Mukherjea and J. S. Ratti. Continuous singularity of the weak limit of convolution powers of a discrete probability measure on 2 × 2 stochastic matrices. Journal of Theoretical Probability, 10 (2):499–506, April 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022672802749.

Marcus:1999:MWP

[MR99] Michael B. Marcus and Jay Rosen. Multiple Wick product chaos processes. Journal of Theoretical Probability, 12(2):489–522, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021686313259.

Maejima:2002:TD

[MR02] Makoto Maejima and Jan Rosiński. Type G distributions on R^{d}. Journal of Theoretical Probability, 15(2):323–341, April 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1015044726122.

Marcus:2005:CBI

[MR05] Michael B. Marcus and Jan Rosiński. Continuity and boundedness of infinitely divisible processes: A Poisson point process approach. Journal of Theoretical Probability, 18(1):109–160, January 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-004-2579-1.

Marin-Rubio:2004:SRS

[MRR04] P. Marín-Rubio and J. Real. Some results on stochastic differential equations with reflecting boundary conditions. Journal of Theoretical Probability, 17(3):705–716, July 2004. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000040295.09922.85.

Martinez:2007:LWA

[MRS07] Servet Martínez, Daniel Remenik, and Jaime San Martín. Levelwise approximation of a Markov process associated to the boundary of an infinite tree. Journal of Theoretical Probability, 20(3): 561–579, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0073-2.

Maejima:2015:SIS

[MRU15] Makoto Maejima, Jan Rosiński, and Yohei Ueda. Stochastic integral and series representations for strictly stable distributions. Journal of Theoretical Probability, 28(3):989–1006, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-013-0518-8.

Meerschaert:1996:SRS

[MS96] Mark M. Meerschaert and Hans-Peter Scheffler. Series representation for semistable laws and their domains of semistable attraction. Journal of Theoretical Probability, 9(4):931–959, October 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02214258.

March:1997:LDA

[MS97a] Peter March and Timo Seppäläinen. Large deviations from the almost everywhere Central Limit Theorem. Journal of Theoretical Probability, 10(4):935–965, October 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022614700678.

Meerschaert:1997:SDG

[MS97b] Mark M. Meerschaert and Hans-Peter Scheffler. Spectral decomposition for generalized domains of semistable attraction. Journal of Theoretical Probability, 10(1):51–71, January 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022638213920.

Meerschaert:1999:SCM

[MS99] Mark M. Meerschaert and Hans-Peter Scheffler. Sample covariance matrix for random vectors with heavy tails. Journal of Theoretical Probability, 12(3):821–838, July 1999. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021688101621.

Mountford:2000:EKA

[MS00] Thomas S. Mountford and Ted D. Sweet. An extension of Kuczek’s argument to nonnearest neighbor contact processes. Journal of Theoretical Probability, 13(4):1061–1081, October 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1007818108889.

Martinez:2001:RDP

[MS01a] Servet Martínez and Jaime San Martín. Rates of decay and h-processes for one dimensional diffusions conditioned on nonabsorption. Journal of Theoretical Probability, 14(1):199–212, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007881317492.

Mason:2001:SDS

[MS01b] David M. Mason and Zhan Shi. Small deviations for some multiparameter Gaussian processes. Journal of Theoretical Probability, 14(1):213–239, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1007833401562.

Matoussi:2002:SPD

[MS02] Anis Matoussi and Michael Scheutzow. Stochastic PDEs driven by nonlinear noise and backward doubly SDEs. Journal of Theoretical Probability, 15(1):1–39, January 2002. CODEN JTPREO.

ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1013803104760.

Mueller:2005:HIS

[MS05] C. Mueller and A. Stan. A Heisenberg inequality for stochastic integrals. Journal of Theoretical Probability, 18(2):291–315, April 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-004-2605-3.

Maejima:2006:MPS

[MS06] Makoto Maejima and Riddhi Shah. Moments and projections of semistable probability measures on p-adic vector spaces. Journal of Theoretical Probability, 19(2):447–459, June 2006. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
006-0018-1.

Ma:2012:GFK

[MS12] Li Ma and Wei Sun. On the generalized Feynman–Kac transformation for nearly symmetric Markov processes. Journal of Theoretical Probability, 25(3):733–755, September 2012. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0318-3.

Mueller:2013:LLI

[MS13] Carl Mueller and Shannon Starr. The length of the longest increasing subsequence of a random Mallows permutation. Journal of Theoretical Probability, 26(2):514–540, June 2013. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
011-0364-5.

Meerschaert:2016:TFS

[MS16] Mark M. Meerschaert and Farzad Sabzikar. Tempered fractional stable motion. Journal of Theoretical Probability, 29(2):681–706, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-014-0585-5.

Mimica:2019:HIS

[MS19] Ante Mimica and Stjepan Sebek. Harnack inequality for subordinate random walks. Journal of Theoretical Probability, 32(2):

737–764, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Matsui:2021:TIR

[MŚ21] Muneya Matsui and Witold Światkowski. Tail indices for AX+B recursion with triangular matrices. Journal of Theoretical Probability, 34(4):1831–1869, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-020-01019-8.

Mao:2022:CGA

[MS22] Yong-Hua Mao and Yan-Hong Song. Criteria for geometric and algebraic transience for discrete-time Markov chains. Journal of Theoretical Probability, 35(3):1974–2008, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01105-5.

Maejima:2023:SFA

[MS23] Makoto Maejima and Noriyoshi Sakuma. Selfsimilar free additive processes and freely selfdecomposable distributions. Journal of Theoretical Probability, 36(3):1667–1697, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01227-4.

Musial:2007:FTP

[MSM07] K. Musial, W. Strauss, and N. D. Macheras. Fubini type products for densities and liftings. Journal of Theoretical Probability, 20(3):545–560, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0075-0.

Mai:2017:TNC

[MSS17] Jan-Frederik Mai, Steffen Schenk, and Matthias Scherer. Two novel characterizations of self-decomposability on the half-line. Journal of Theoretical Probability, 30(1):365–383, March 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-015-0644-6.

Matthes:1993:SZO

[MSSW93] K. Matthes, R. Siegmund-Schultze, and A. Wakolbinger. A spatial zero-one law for equilibria of branching particle systems.

Journal of Theoretical Probability, 6(4):699–711, October 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049172.

Moricz:2008:SLB

[MST08] Ferenc Móricz, Ulrich Stadtmüller, and Monika Thalmaier. Strong laws for blockwise m-dependent random fields. Journal of Theoretical Probability, 21(3):660–671, September 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-007-0127-5.

Misiats:2016:EUI

[MSY16] Oleksandr Misiats, Oleksandr Stanzhytskyi, and Nung Kwan Yip. Existence and uniqueness of invariant measures for stochastic reaction–diffusion equations in unbounded domains. Journal of Theoretical Probability, 29(3):996–1026, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-015-0606-z.

Marcus:1988:CTS

[MT88] Michael B. Marcus and Michel Talagrand. Chevet’s theorem for stable processes II. Journal of Theoretical Probability, 1 (1):65–92, January 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01076288.

Marcus:1998:CCC

[MT98] Michael B. Marcus and Michel Talagrand. Continuity conditions for a class of Gaussian chaos processes related to continuous additive functionals of Lévy processes. Journal of Theoretical Probability, 11(1):157–179, January 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1021699009373.

Morayne:2008:FWA

[MT08] Michał Morayne and Krzysztof Tabisz. Filtrations for which all H^{2} martingales are of integrable variation; distances between σalgebras. Journal of Theoretical Probability, 21(1):1–13, March 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-007-0131-9.

Maejima:2014:DRT

[MTU14] Makoto Maejima, Taisuke Takamune, and Yohei Ueda. The dichotomy of recurrence and transience of semi-Lévy processes. Journal of Theoretical Probability, 27(3):982–996, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-012-0452-1.

Mijatovic:2015:LSP

[MU15] Aleksandar Mijatović and Mikhail Urusov. On the loss of the semimartingale property at the hitting time of a level. Journal of Theoretical Probability, 28(3):892–922, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-013-0527-7.

Mukherjea:1997:P

[Muk97] A. Mukherjea. Preface. Journal of Theoretical Probability, 10(2):277, April 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022648130066.

Mukeru:2019:ANR

[Muk19] Safari Mukeru. Average number of real zeros of random algebraic polynomials defined by the increments of fractional Brownian motion. Journal of Theoretical Probability, 32(3):1502–1524, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Mukeru:2021:CSD

[Muk21] Safari Mukeru. On the convergence of series of dependent random variables. Journal of Theoretical Probability, 34(3):1299– 1320, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01018-9.

Murphy:2019:ECR

[Mur19] James T. Murphy III. Exact coupling of random walks on Polish groups. Journal of Theoretical Probability, 32(4):1729–1745, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Mustapha:2006:GER

[Mus06] Sami Mustapha. Gaussian estimates for random walks on some unimodular p-adic groups. Journal of Theoretical Probability, 19(4):773–787, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-006-0044-z.

Meerschaert:1993:SES

[MV93] Mark M. Meerschaert and Jerry Alan Veeh. The structure of the exponents and symmetries of an operator stable measure. Journal of Theoretical Probability, 6(4):713–726, October 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01049173.

Martinez:1998:PRR

[MV98] F. Martínez and A. R. Villena. Path regularity and regularity in mean of 1st order processes. Journal of Theoretical Probability, 11(2):435–442, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022635906617.

Mishura:2000:IAG

[MV00] Yulia Mishura and Esko Valkeila. An isometric approach to generalized stochastic integrals. Journal of Theoretical Probability, 13(3):673–693, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007854310936.

Maheshwari:2019:FPP

[MV19a] A. Maheshwari and P. Vellaisamy. Fractional Poisson process time-changed by Lévy subordinator and its inverse. Journal of Theoretical Probability, 32(3):1278–1305, September 2019. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Mountford:2019:ECF

[MV19b] Thomas Mountford and Glauco Valle. Exponential convergence for the Fredrickson–Andersen one-spin facilitated model. Journal of Theoretical Probability, 32(1):282–302, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Mao:2022:CRU

[MW22] Yong-Hua Mao and Tao Wang. Convergence rates in uniform ergodicity by hitting times and L^{2}-exponential convergence rates. Journal of Theoretical Probability, 35(4):2690–2711, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01155-9.

Mao:2015:LDB

[MWW15] Yong-Hua Mao, Feng Wang, and Xian-Yuan Wu. Large deviation behavior for the longest head run in an IID Bernoulli sequence. Journal of Theoretical Probability, 28(1):259–268, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-013-0498-8.

Mytnik:2004:TFS

[MX04] L. Mytnik and K.-N. Xiang. Tanaka formulae for (α, d, β)- superprocesses. Journal of Theoretical Probability, 17(2):483– 502, April 2004. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL http://link.springer.com/
article/10.1023/B%3AJOTP.0000020704.68569.25.

Merzbach:2009:SIB

[MY09] Ely Merzbach and Arthur Yosef. Set-indexed Brownian motion on increasing paths. Journal of Theoretical Probability, 22(4): 883–890, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-008-0188-0.

Myronyuk:2020:ILF

[Myr20] Margaryta Myronyuk. Independent linear forms on the group Ω_{p}. Journal of Theoretical Probability, 33(1):1–21, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Myronyuk:2024:AKT

[Myr24] Margaryta Myronyuk. An analogue of the Klebanov Theorem for locally compact Abelian groups. Journal of Theoretical Probability, 37(3):2646–2664, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-024-01339-z.

Mytnik:1998:CMC

[Myt98] Leonid Mytnik. Collision measure and collision local time for (α, d, β) superprocesses. Journal of Theoretical Probability, 11 (3):733–763, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022606715641.

Maciuca:2006:FHT

[MZ06] Romeo Maciuca and Song-Chun Zhu. First hitting time analysis of the independence Metropolis sampler. Journal of Theoretical Probability, 19(1):235–261, January 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-006-0002-9.

Mikulevicius:2024:CWE

[MZ24] Remigijus Mikulevicius and Changyong Zhang. Convergence of weak Euler approximation for nondegenerate stochastic differential equations driven by point and martingale measures. Journal of Theoretical Probability, 37(1):43–80, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01260-x.

Nahla:2011:RPC

[NA11] Ben Salah Nahla and Masmoudi Afif. The real powers of the convolution of a gamma distribution and a Bernoulli distribution. Journal of Theoretical Probability, 24(2):450–453, June 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10959-009-0253-3.pdf.

Naganuma:2015:AED

[Nag15] Nobuaki Naganuma. Asymptotic error distributions of the Crank–Nicolson scheme for SDEs driven by fractional Brownian motion. Journal of Theoretical Probability, 28(3):1082– 1124, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0539-y.

Nagel:2021:FCP

[Nag21] Jan Nagel. A functional CLT for partial traces of random matrices. Journal of Theoretical Probability, 34(2):953–974, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-019-00982-1. See correction [Nag22].

Nagel:2022:CFC

[Nag22] Jan Nagel. Correction to: A functional CLT for partial traces of random matrices. Journal of Theoretical Probability, 35(3): 2080–2081, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01132-2. See [Nag21].

Nakashima:2013:MPB

[Nak13] Makoto Nakashima. Minimal position of branching random walks in random environment. Journal of Theoretical Probability, 26(4):1181–1217, December 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-012-0410-y.

Nakatsu:2019:SPD

[Nak19] Tomonori Nakatsu. Some properties of density functions on maxima of solutions to one–dimensional stochastic differential equations. Journal of Theoretical Probability, 32(4):1746–1779, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Namba:2022:EEC

[Nam22] Ryuya Namba. Edgeworth expansions for centered random walks on covering graphs of polynomial volume growth. Journal of Theoretical Probability, 35(3):1898–1938, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01111-7.

Nass:2022:LPS

[Nas22] Aminu Ma’aruf Nass. Lie point symmetries of autonomous scalar first-order Itô stochastic delay ordinary differential equations. Journal of Theoretical Probability, 35(3):1939–1951, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01103-7.

Nazarov:2009:ESB

[Naz09] Alexander I. Nazarov. Exact L_{2}-small ball asymptotics of Gaussian processes and the spectrum of boundary-value problems. Journal of Theoretical Probability, 22(3):640–665, September 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-008-0173-7.

Ni:2021:HMR

[NC21] Wenqing Ni and Zhenlong Chen. Hausdorff measure of the range of space-time anisotropic Gaussian random fields. Journal of Theoretical Probability, 34(1):264–282, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-019-
00964-3.

Neal:1992:CSP

[Nea92] David Neal. Characterization of O-summable processes. Journal of Theoretical Probability, 5(3):585–596, July 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01060438.

Nejib:1994:CSC

[Nej94] Ben Salem Nejib. Convolution semigroups and central limit theorem associated with a dual convolution structure. Journal of Theoretical Probability, 7(2):417–436, April 1994. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02214276.

Nemish:2018:NOS

[Nem18] Yuriy Nemish. No outliers in the spectrum of the product of independent non–Hermitian random matrices with independent entries. Journal of Theoretical Probability, 31(1):402–444, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Neuenschwander:1992:LCT

[Neu92] Daniel Neuenschwander. Limits of commutative triangular systems on simply connected step 2-nilpotent Lie groups. Journal of Theoretical Probability, 5(1):217–222, January 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01046787.

Neuenschwander:1995:GCL

[Neu95] Daniel Neuenschwander. A Gaussian central limit theorem for trimmed products on simply connected step 2-nilpotent Lie groups. Journal of Theoretical Probability, 8(1):165–174, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02213459.

Neuenschwander:2008:UEG

[Neu08] Daniel Neuenschwander. Uniqueness of embedding into a Gaussian semigroup and a Poisson semigroup with determinate jump law on a simply connected nilpotent Lie group. Journal of Theoretical Probability, 21(4):791–801, December 2008. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-

008-0175-5. See erratum [Neu09].

Neuenschwander:2009:EUE

[Neu09] Daniel Neuenschwander. Erratum to: Uniqueness of Embedding into a Gaussian Semigroup and a Poisson Semigroup with Determinate Jump Law on a Simply Connected Nilpotent Lie Group. Journal of Theoretical Probability, 22(4):1058–1060, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10959-009-0243-5.pdf. See [Neu08].

Neyman:2013:MVM

[Ney13] Abraham Neyman. The maximal variation of martingales of probabilities and repeated games with incomplete information. Journal of Theoretical Probability, 26(2):557–567, June 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-012-0447-y.

Nickl:2009:CCR

[Nic09] Richard Nickl. On convergence and convolutions of random signed measures. Journal of Theoretical Probability, 22(1):38–56, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0177-3.

Nieznaj:2022:COR

[Nie22] Ernest Nieznaj. On a certain operator related to Blackwell’s Markov chain. Journal of Theoretical Probability, 35(3):1501– 1510, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01110-8.

Nikeghbali:2006:MDF

[Nik06] Ashkan Nikeghbali. Multiplicative decompositions and frequency of vanishing of nonnegative submartingales. Journal of Theoretical Probability, 19(4):931–949, December 2006. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
006-0035-0.

Neuenkirch:2007:ERC

[NN07] Andreas Neuenkirch and Ivan Nourdin. Exact rate of convergence of some approximation schemes associated to SDEs driven by a fractional Brownian motion. Journal of Theoretical Probability, 20(4):871–899, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0083-0.

Nourdin:2010:CLT

[NN10] Ivan Nourdin and David Nualart. Central limit theorems for multiple Skorokhod integrals. Journal of Theoretical Probability, 23(1):39–64, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0258-y.

Najnudel:2016:LTO

[NNR16] Joseph Najnudel, Ashkan Nikeghbali, and Alain Rouault. Limit theorems for orthogonal polynomials related to circular ensembles. Journal of Theoretical Probability, 29(4):1199–1239, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-015-0632-x.

Nikitin:2000:SDP

[NO00] Y. Nikitin and E. Orsingher. On sojourn distributions of processes related to some higher-order heat-type equations. Journal of Theoretical Probability, 13(4):997–1012, October 2000.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007861923910.

Nualart:2003:BRS

[NO03] David Nualart and Youssef Ouknine. Besov regularity of stochastic integrals with respect to the fractional Brownian motion with parameter H > 1/2. Journal of Theoretical Probability, 16 (2):451–470, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023530929480.

Nobel:1991:LTP

[Nob91] Siegfried Nobel. Limit theorems for probability measures on simply connected nilpotent lie groups. Journal of Theoretical Probability, 4(2):261–284, April 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01258737.

Noda:2017:LDB

[Nod17] Hideaki Noda. Large deviations for Brownian motion on scale irregular Sierpiński gaskets. Journal of Theoretical Probability, 30(3):852–875, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Noiry:2021:SMS

[Noi21] Nathan Noiry. Spectral measures of spiked random matrices. Journal of Theoretical Probability, 34(2):923–952, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-00987-1.

Nualart:2007:ILT

[NOL07] David Nualart and Salvador Ortiz-Latorre. Intersection local time for two independent fractional Brownian motions. Journal of Theoretical Probability, 20(4):759–767, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-007-0106-x.

Novak:2024:PA

[Nov24] S. Y. Novak. On Poisson approximation. Journal of Theoretical Probability, 37(3):2277–2303, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01310-4.

Norvaisa:1991:RCC

[NP91] R. Norvaisa and V. Paulauskas. Rate of convergence in the central limit theorem for empirical processes. Journal of Theoretical Probability, 4(3):511–534, July 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01210322.

Nualart:1996:SIP

[NP96] David Nualart and Philip Protter. Skorohod integral of a product of two stochastic processes. Journal of Theoretical Probability, 9(4):1029–1037, October 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214263.

Nickl:2007:BME

[NP07] Richard Nickl and Benedikt M. Pötscher. Bracketing metric entropy rates and empirical Central Limit Theorems for function classes of Besov- and Sobolev-type. Journal of Theoretical Probability, 20(2):177–199, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0058-1.

Neuenkirch:2018:OAS

[NP18] Andreas Neuenkirch and Peter Parczewski. Optimal approximation of Skorohod integrals. Journal of Theoretical Probability, 31 (1):206–231, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Nevo:2024:SMB

[NP24] Amos Nevo and Felix Pogorzelski. Shannon–McMillan–Breiman theorem along almost geodesics in negatively curved groups. Journal of Theoretical Probability, 37(1):814–859, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s10959-023-01291-4.

Nourdin:2016:CFF

[NPPS16] Ivan Nourdin, Giovanni Peccati, Guillaume Poly, and Rosaria Simone. Classical and free fourth moment theorems: Universality and thresholds. Journal of Theoretical Probability, 29 (2):653–680, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0590-8.

Nourdin:2022:MNA

[NPY22] Ivan Nourdin, Giovanni Peccati, and Xiaochuan Yang. Multivariate normal approximation on the Wiener space: New bounds in the convex distance. Journal of Theoretical Probability, 35(3): 2020–2037, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01112-6.

Neuenschwander:1995:LAB

[NS95] Daniel Neuenschwander and René Schott. On the local and asymptotic behavior of Brownian motion on simply connected nilpotent Lie groups. Journal of Theoretical Probability, 8(4): 795–806, October 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02410112.

Noda:2023:ENZ

[NS23a] Kohei Noda and Tomoyuki Shirai. Expected number of zeros of random power series with finitely dependent Gaussian coefficients. Journal of Theoretical Probability, 36(3):1534–1554, September 2023. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-022-01203-y.

Nualart:2023:EDE

[NS23b] David Nualart and Bhargobjyoti Saikia. Error distribution of the Euler approximation scheme for stochastic Volterra equations. Journal of Theoretical Probability, 36(3):1829–1876, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01222-9.

Nicol:2002:FSS

[NSB02] Matthew Nicol, Nikita Sidorov, and David Broomhead. On the fine structure of stationary measures in systems which contracton-average. Journal of Theoretical Probability, 15(3):715–730, July 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1016224000145.

Nourdin:2014:CNC

[NT14] Ivan Nourdin and Murad S. Taqqu. Central and non-central limit theorems in a free probability setting. Journal of Theoretical Probability, 27(1):220–248, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-012-0443-2.

Nummelin:1994:DRP

[Num94] Esa Nummelin. On distributionally regenerative processes. Journal of Theoretical Probability, 7(4):739–756, October 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02214369.

Nagnibeda:2002:RWT

[NW02] Tatiana Nagnibeda and Wolfgang Woess. Random walks on trees with finitely many cone types. Journal of Theoretical Probability, 15(2):383–422, April 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014810827031.

Nestoridi:2019:SLD

[NW19] Evita Nestoridi and Graham White. Shuffling large decks of cards and the Bernoulli–Laplace urn model. Journal of Theoretical Probability, 32(1):417–446, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Nyrhinen:2009:LDM

[Nyr09] Harri Nyrhinen. On large deviations of multivariate heavy-tailed random walks. Journal of Theoretical Probability, 22(1):1–17, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0194-2.

Nishimura:2022:HOE

[NYY22] Tsubasa Nishimura, Kenji Yasutomi, and Tomooki Yuasa. Higher-order error estimates of the discrete-time Clark–Ocone formula. Journal of Theoretical Probability, 35(4):2518–2539, December 2022. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-021-01134-0.

Nourdin:2022:ABL

[NZ22] Ivan Nourdin and Guangqu Zheng. Asymptotic behavior of large Gaussian correlated Wishart matrices. Journal of Theoretical Probability, 35(4):2239–2268, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-021-
01133-1.

Owo:2023:GBD

[OA23] Jean-Marc Owo and Auguste Aman. Generalized backward doubly stochastic differential equations driven by Lévy processes with discontinuous and linear growth coefficients. Journal of Theoretical Probability, 36(4):2311–2338, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01270-9.

OBrien:1996:SCC

[O’B96] George L. O’Brien. Sequences of capacities, with connections to large-deviation theory. Journal of Theoretical Probability, 9 (1):19–35, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213733.

Oualaid:2023:RBS

[OBO23] Abdelkarim Oualaid, Khaled Bahlali, and Youssef Ouknine. Reflected backward stochastic differential equations associated to jump Markov processes and application to partial differential equations. Journal of Theoretical Probability, 36(3):1400– 1436, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01212-x.

Orsingher:2007:RFH

[OD07] E. Orsingher and A. De Gregorio. Random flights in higher spaces. Journal of Theoretical Probability, 20(4):769–806, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-007-0093-y.

Orsingher:2015:HOL

[OD15] Enzo Orsingher and Mirko D’Ovidio. Higher-order Laplace equations and hyper–Cauchy distributions. Journal of Theoretical Probability, 28(1):92–118, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-013-0480-5.

Okamura:2014:SRF

[Oka14] Kazuki Okamura. Singularity results for functional equations driven by linear fractional transformations. Journal of Theoretical Probability, 27(4):1316–1328, December 2014. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
013-0516-x.

Okamura:2021:SRR

[Oka21] Kazuki Okamura. Some results for range of random walk on graph with spectral dimension two. Journal of Theoretical Probability, 34(3):1653–1688, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-020-01013-0.

Okazaki:2024:CMJ

[Oka24] Fumiya Okazaki. Convergence of martingales with jumps on submanifolds of Euclidean spaces and its applications to harmonic maps. Journal of Theoretical Probability, 37(2):1168–1198, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01273-6.

Oksendal:1990:WSI

[Øks90] Bernt Øksendal. When is a stochastic integral a time change of a diffusion? Journal of Theoretical Probability, 3(2):207–226, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01045159.

Obiang:2022:ICC

[OMO22] Fulgence Eyi Obiang, Octave Moutsinga, and Youssef Ouknine. An ideal class to construct solutions for skew Brownian motion equations. Journal of Theoretical Probability, 35 (2):894–916, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01078-5.

Obloj:2009:ESE

[OP09] Jan Oblój and Martijn Pistorius. On an explicit Skorokhod embedding for spectrally negative Lévy processes. Journal of Theoretical Probability, 22(2):418–440, June 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-008-0157-7.

Oh:2024:GFG

[OPST24] Sewoong Oh, Soumik Pal, Raghav Somani, and Raghavendra Tripathi. Gradient flows on graphons: Existence, convergence, continuity equations. Journal of Theoretical Probability, 37(2): 1469–1522, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01271-8.

ORourke:2016:CLT

[OR16] Sean O’Rourke and David Renfrew. Central limit theorem for linear eigenvalue statistics of elliptic random matrices. Journal of Theoretical Probability, 29(3):1121–1191, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-015-0609-9.

Oraby:2007:LSG

[Ora07] Tamer Oraby. The limiting spectra of Girko’s block-matrix. Journal of Theoretical Probability, 20(4):959–970, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-007-0084-z.

ORourke:2013:FME

[ORS13] Sean O’Rourke, David Renfrew, and Alexander Soshnikov. On fluctuations of matrix entries of regular functions of Wigner matrices with non-identically distributed entries. Journal of Theoretical Probability, 26(3):750–780, September 2013. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
011-0396-x.

ORourke:2015:PLE

[OS15] Sean O’Rourke and Alexander Soshnikov. Partial linear eigenvalue statistics for Wigner and sample covariance random matrices. Journal of Theoretical Probability, 28(2):726–744, June 2015.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-013-0491-2.

Osekowski:2009:SDS

[Ose09] Adam Osekowski. Strong differential subordination and sharp inequalities for orthogonal processes. Journal of Theoretical Probability, 22(4):837–855, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-009-0236-4.

Osekowski:2011:LES

[Ose11] Adam Osekowski. Logarithmic estimates for submartingales and their differential subordinates. Journal of Theoretical Probability, 24(3):849–874, September 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-009-0268-9.

Osekowski:2014:MIM

[Ose14] Adam Oseķowski. Maximal inequalities for martingales and their differential subordinates. Journal of Theoretical Probability, 27 (1):1–21, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-012-0458-8.pdf.

Osswald:2009:AGT

[Oss09a] Horst Osswald. On anticipative Girsanov transformations for Lévy processes. Journal of Theoretical Probability, 22(2): 474–481, June 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-008-0197-z.

Osswald:2009:SAM

[Oss09b] Horst Osswald. A smooth approach to Malliavin calculus for Lévy processes. Journal of Theoretical Probability, 22(2): 441–473, June 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-008-0148-8.

Orsingher:2017:STF

[OT17] Enzo Orsingher and Bruno Toaldo. Space-time fractional equations and the related stable processes at random time. Journal of Theoretical Probability, 30(1):1–26, March 2017. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
015-0641-9.

Ogawa:2014:SFC

[OU14] Shigeyoshi Ogawa and Hideaki Uemura. On a stochastic Fourier coefficient: Case of noncausal functions. Journal of Theoretical Probability, 27(2):370–382, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-012-0464-x.

Okamura:2021:SST

[OU21] Haruna Okamura and Toshihiro Uemura. On symmetric stable-type processes with degenerate/singular Lévy densities. Journal of Theoretical Probability, 34(2):809–826, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-00990-6.

Ouchti:2008:CCW

[OV08] L. Ouchti and D. Volný. A conditional CLT which fails for ergodic components. Journal of Theoretical Probability, 21(3): 687–703, September 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0126-6.

Ondrejat:2014:WCS

[OV14] Martin Ondreját and Mark Veraar. Weak characterizations of stochastic integrability and Dudley’s theorem in infinite dimensions. Journal of Theoretical Probability, 27(4):1350–1374, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-013-0479-y.

OBrien:1998:RCC

[OW98] George L. O’Brien and Stephen Watson. Relative compactness for capacities, measures, upper semicontinuous functions and closed sets. Journal of Theoretical Probability, 11(3):577–588, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022659912007.

ORourke:2019:LDF

[OW19] Sean O’Rourke and Philip Matchett Wood. Low-degree factors of random polynomials. Journal of Theoretical Probability, 32 (2):1076–1104, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Owada:2016:LTS

[Owa16] Takashi Owada. Limit theory for the sample autocovariance for heavy-tailed stationary infinitely divisible processes generated by conservative flows. Journal of Theoretical Probability, 29(1):63–95, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0565-9.

Ozdemir:2019:RCT

[Özd19] Alperen Y. Özdemir. The random (n−k)-cycle to transpositions walk on the symmetric group. Journal of Theoretical Probability, 32(3):1438–1460, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Pachon:2012:NLD

[Pac12] Daniel Andrés Díaz Pachón. A note on large deviations for the stable marriage of Poisson and Lebesgue with random appetites. Journal of Theoretical Probability, 25(1):77–91, March 2012. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0304-9.

Pacheco:2016:PID

[Pac16] Carlos G. Pacheco. Picard iterations for diffusions on symmetric matrices. Journal of Theoretical Probability, 29(4):1444– 1457, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0618-8.

Pahlajani:2015:SAH

[Pah15] Chetan D. Pahlajani. Stochastic averaging for a Hamiltonian system with skew random perturbations. Journal of Theoretical Probability, 28(3):1165–1226, September 2015. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
013-0533-9.

Pak:2000:TRW

[Pak00] Igor Pak. Two random walks on upper triangular matrices. Journal of Theoretical Probability, 13(4):1083–1100, October 2000.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007870125727.

Pakes:2020:SDL

[Pak20a] Anthony G. Pakes. Self-decomposable laws from continuous branching processes. Journal of Theoretical Probability, 33(1): 361–395, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Pakzad:2020:LDP

[Pak20b] Cambyse Pakzad. Large deviations principle for the largest eigenvalue of the Gaussian β-ensemble at high temperature. Journal of Theoretical Probability, 33(1):428–443, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Palacios:1992:ECT

[Pal92] José Luis Palacios. Expected cover times of random walks on symmetric graphs. Journal of Theoretical Probability, 5 (3):597–600, July 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01060439.

Pang:2019:LAM

[Pan19] C. Y. Amy Pang. Lumpings of algebraic Markov chains arise from subquotients. Journal of Theoretical Probability, 32(4): 1804–1844, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Pap:1990:AHE

[Pap90] Gyula Pap. Analog of heat equation for Gaussian measure of a ball in Hilbert space. Journal of Theoretical Probability, 3(4): 563–577, October 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046096.

Pap:2011:AMA

[Pap11] Gyula Pap. The accuracy of merging approximation in generalized St. Petersburg games. Journal of Theoretical Probability, 24 (1):240–270, March 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0247-1.

Perez-Abreu:2007:RID

[PAR07] Victor Pérez-Abreu and Jan Rosiński. Representation of infinitely divisible distributions on cones. Journal of Theoretical Probability, 20(3):535–544, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-007-0076-z.

Pardo:2009:UEP

[Par09] J. C. Pardo. The upper envelope of positive self-similar Markov processes. Journal of Theoretical Probability, 22(2):514–542, June 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0152-z.

Pardon:2012:CLT

[Par12] John Pardon. Central limit theorems for uniform model random polygons. Journal of Theoretical Probability, 25(3):823– 833, September 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0335-2.

Perez-Abreu:2012:FID

[PAS12] Victor Pérez-Abreu and Noriyoshi Sakuma. Free infinite divisibility of free multiplicative mixtures of the Wigner distribution. Journal of Theoretical Probability, 25(1):100–121, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-010-0288-5.

Perez-Abreu:1994:LDC

[PAT94] Víctor Pérez-Abreu and Constantin Tudor. Large deviations for a class of chaos expansions. Journal of Theoretical Probability, 7 (4):757–765, October 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214370.

Pakes:1991:FTS

[PD91] Anthony G. Pakes and F. M. Dekking. On family trees and subtrees of simple branching processes. Journal of Theoretical Probability, 4(2):353–369, April 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01258742.

Penda:2023:CLT

[PD23] S. Valère Bitseki Penda and Jean-François Delmas. Central limit theorem for kernel estimator of invariant density in bifurcating Markov chains models. Journal of Theoretical Probability, 36(3): 1591–1625, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01205-w.

Pechtl:1993:AMI

[Pec93] Andreas Pechtl. Arithmetic means and invariance principles in stochastic approximation. Journal of Theoretical Probability, 6 (1):153–173, January 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046773.

Peligrad:1996:ANS

[Pel96] Magda Peligrad. On the asymptotic normality of sequences of weak dependent random variables. Journal of Theoretical Probability, 9(3):703–715, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214083.

Peligrad:2001:NUL

[Pel01] Magda Peligrad. A note on the uniform laws for dependent processes via coupling. Journal of Theoretical Probability, 14 (4):979–988, October 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012524819781.

Peligrad:2024:QCS

[Pel24] Magda Peligrad. On the quenched CLT for stationary Markov chains. Journal of Theoretical Probability, 37(1):603–622, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01241-0.

Pemantle:1989:RTO

[Pem89] Robin Pemantle. Randomization time for the overhand shuffle. Journal of Theoretical Probability, 2(1):37–49, January 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01048267.

Pemantle:1990:TDV

[Pem90] Robin Pemantle. A time-dependent version of Pólya’s urn. Journal of Theoretical Probability, 3(4):627–637, October 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01046101.

Perera:1997:GCL

[Per97] Gonzalo Perera. Geometry of Z^{d} and the Central Limit Theorem for weakly dependent random fields. Journal of Theoretical Probability, 10(3):581–603, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022693309359.

Pereira:2018:ALL

[Per18] Luísa Pereira. On the asymptotic locations of the largest and smallest extremes of a stationary sequence. Journal of Theoretical Probability, 31(2):853–866, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Peskir:1995:UCR

[Pes95] Goran Peskir. Uniform convergence of reversed martingales. Journal of Theoretical Probability, 8(2):387–415, April 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02212885.

Peskir:2001:BMH

[Pes01] Goran Peskir. Bounding the maximal height of a diffusion by the time elapsed. Journal of Theoretical Probability, 14(3):845–855, July 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017505509361.

Peskir:2005:CVF

[Pes05] Goran Peskir. A change-of-variable formula with local time on curves. Journal of Theoretical Probability, 18(3):499–535, July 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-005-3517-6.

Papamarcou:1991:SAS

[PF91] Adrian Papamarcou and Terrence L. Fine. Stationarity and almost sure divergence of time averages in interval-valued probability. Journal of Theoretical Probability, 4(2):239–260, April 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01258736.

Prosdocimi:2016:CPE

[PF16] Cecilia Prosdocimi and Lorenzo Finesso. Countable partially exchangeable mixtures. Journal of Theoretical Probability, 29 (2):423–442, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0573-9.

Pfiefer:1990:MMS

[Pfi90] Richard E. Pfiefer. Maximum and minimum sets for some geometric mean values. Journal of Theoretical Probability, 3 (2):169–179, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01045156.

Peligrad:1999:ASR

[PG99] Magda Peligrad and Allan Gut. Almost-sure results for a class of dependent random variables. Journal of Theoretical Probability, 12(1):87–104, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021744626773.

Pajor-Gyulai:2016:DSP

[PGS16] Zsolt Pajor-Gyulai and Michael Salins. On dynamical systems perturbed by a null-recurrent fast motion: The continuous coefficient case with independent driving noises. Journal of Theoretical Probability, 29(3):1083–1099, September 2016. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
015-0600-5.

Pinsky:1993:FIM

[Pin93] Mark A. Pinsky. Fourier inversion for multidimensional characteristic functions. Journal of Theoretical Probability, 6(1): 187–193, January 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046775.

Pincus:1994:SSM

[Pin94] Steve Pincus. Singular stationary measures are not always fractal. Journal of Theoretical Probability, 7(1):199–208, January 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF02213368.

Pinsky:2001:BCM

[Pin01] Mark A. Pinsky. Brownian continuity modulus via series expansions. Journal of Theoretical Probability, 14(1):261–266, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007837502471.

Pinelis:2011:PPM

[Pin11] Iosif Pinelis. Positive-part moments via the Fourier–Laplace transform. Journal of Theoretical Probability, 24(2):409–421, June 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-010-0276-9.

Pinelis:2018:PPM

[Pin18] Iosif Pinelis. Positive-part moments via characteristic functions, and more general expressions. Journal of Theoretical Probability, 31(1):527–555, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Pistorius:2004:EES

[Pis04] M. R. Pistorius. On exit and ergodicity of the spectrally one-sided Lévy process reflected at its infimum. Journal of Theoretical Probability, 17(1):183–220, January 2004. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000020481.14371.37.

Planchet:1989:CC

[Pla89] Bernard Planchet. Credibility and conditioning. Journal of Theoretical Probability, 2(3):289–299, July 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01054017.

Plachky:1992:SCD

[Pla92] D. Plachky. On semiregular conditional distributions. Journal of Theoretical Probability, 5(3):577–584, July 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01060437.

Popov:2021:CTD

[Pop21] Serguei Popov. Conditioned two-dimensional simple random walk: Green’s function and harmonic measure. Journal of Theoretical Probability, 34(1):418–437, March 2021. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-019-
00963-4.

Proske:2002:SLL

[PP02] Frank N. Proske and Madan L. Puri. Strong law of large numbers for Banach space valued fuzzy random variables. Journal of Theoretical Probability, 15(2):543–551, April 2002. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1014823228848.

Pascu:2018:CBM

[PP18] Mihai N. Pascu and Ionel Popescu. Couplings of Brownian motions of deterministic distance in model spaces of constant curvature. Journal of Theoretical Probability, 31(4):2005–2031, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Peigne:2024:CLL

[PP24] Marc Peigné and Da Cam Pham. A conditioned local limit theorem for nonnegative random matrices. Journal of Theoretical Probability, 37(2):1882–1901, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-024-01336-2.

Pardo:2016:RMA

[PPPA16] Juan Carlos Pardo, José-Luis Pérez, and Victor Pérez-Abreu. A random matrix approximation for the non-commutative fractional Brownian motion. Journal of Theoretical Probability, 29(4):1581–1598, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-015-0627-7.

Pacheco:2022:EBD

[PPR22] Carlos G. Pacheco and Mariana Pérez-Rojas. Excursions of the Brox diffusion. Journal of Theoretical Probability, 35(3):1479– 1500, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01113-5.

Pages:2023:DEFa

[PR23a] Gilles Pagès and Clément Rey. Discretization of the ergodic functional Central Limit Theorem. Journal of Theoretical Probability, 36(1):1–44, March 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01237-w.

Pages:2023:DEFb

[PR23b] Gilles Pagès and Clément Rey. Discretization of the ergodic functional Central Limit Theorem. Journal of Theoretical Probability, 36(4):2359–2402, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-023-01237-w.

Pritchard:1996:SBS

[Pri96] Geoffrey Pritchard. Sample behavior of some random series. Journal of Theoretical Probability, 9(4):1039–1056, October 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02214264.

Privault:2015:COL

[Pri15] Nicolas Privault. Cumulant operators for Lie–Wiener–Itô– Poisson stochastic integrals. Journal of Theoretical Probability, 28(1):269–298, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0532-x.

Privault:2019:TCS

[Pri19] Nicolas Privault. Third cumulant Stein approximation for Poisson stochastic integrals. Journal of Theoretical Probability, 32 (3):1461–1481, September 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Profeta:2021:PET

[Pro21] Christophe Profeta. Persistence and exit times for some additive functionals of skew Bessel processes. Journal of Theoretical Probability, 34(1):363–390, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-019-00966-1.

Palmowski:2024:ETD

[PRP24] Zbigniew Palmowski, Lewis Ramsden, and Apostolos D. Papaioannou. Exit times for a discrete Markov additive process. Journal of Theoretical Probability, 37(2):1052–1078, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-024-01322-8.

Puccetti:2019:CPM

[PRWW19] Giovanni Puccetti, Pietro Rigo, Bin Wang, and Ruodu Wang. Centers of probability measures without the mean. Journal of Theoretical Probability, 32(3):1482–1501, September 2019. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Pop-Stojanovic:1989:EHF

[PS89] Z. R. Pop-Stojanović. Excessiveness of harmonic functions for certain diffusions. Journal of Theoretical Probability, 2(4): 503–508, October 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01051881.

Peligrad:1994:SNC

[PS94] Magda Peligrad and Qi-Man Shao. Self-normalized central limit theorem for sums of weakly dependent random variables. Journal of Theoretical Probability, 7(2):309–338, April 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02214272.

Pedersen:2005:CDP

[PS05] Jan Pedersen and Ken-Iti Sato. The class of distributions of periodic Ornstein–Uhlenbeck processes driven by Lévy processes. Journal of Theoretical Probability, 18(1):209–235, January 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-004-2595-1.

Peligrad:2013:CLT

[PS13] Magda Peligrad and Hailin Sang. Central limit theorem for linear processes with infinite variance. Journal of Theoretical Probability, 26(1):222–239, March 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-011-0393-0.

Peres:2015:MTH

[PS15] Yuval Peres and Perla Sousi. Mixing times are hitting times of large sets. Journal of Theoretical Probability, 28(2):488–519, June 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-013-0497-9.

Popivoda:2024:STG

[PS24] Goran Popivoda and Sinisa Stamatović. Sojourn times of Gaussian processes with random parameters. Journal of Theoretical Probability, 37(3):2023–2053, September 2024. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01305-1.

Pipiras:2004:DFS

[PT04] Vladas Pipiras and Murad S. Taqqu. Dilated fractional stable motions. Journal of Theoretical Probability, 17(1):51–84, January 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOTP.0000020475.95139.37.

Peccati:2008:SCM

[PT08] Giovanni Peccati and Murad S. Taqqu. Stable convergence of multiple Wiener–Itô integrals. Journal of Theoretical Probability, 21(3):527–570, September 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0154-x.

Pereira:2017:ASC

[PT17] Luísa Pereira and Zhongquan Tan. Almost sure convergence for the maximum of nonstationary random fields. Journal of Theoretical Probability, 30(3):996–1013, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Puha:1999:RNP

[Puh99] Amber L. Puha. A reversible nearest particle system on the homogeneous tree. Journal of Theoretical Probability, 12(1): 217–254, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021709013569.

Puhalskii:2003:LDC

[Puh03] Anatolii A. Puhalskii. On large deviation convergence of invariant measures. Journal of Theoretical Probability, 16(3):689–724, July 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1025624600545.

Picco:1994:LIL

[PV94] Pierre Picco and Maria Eulalia Vares. A law of the iterated logarithm for geometrically weighted martingale difference sequences. Journal of Theoretical Probability, 7(2):375–415, April 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF02214275.

Pemantle:1998:MCF

[PV98] Robin Pemantle and Stanislav Volkov. Markov chains in a field of traps. Journal of Theoretical Probability, 11(2):561–569, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1022648209343.

Passeggeri:2019:MPM

[PV19] Riccardo Passeggeri and Almut E. D. Veraart. Mixing properties of multivariate infinitely divisible random fields. Journal of Theoretical Probability, 32(4):1845–1879, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-018-0864-7.pdf.

Peligrad:2020:QIP

[PV20] Magda Peligrad and Dalibor Volný. Quenched invariance principles for orthomartingale-like sequences. Journal of Theoretical Probability, 33(3):1238–1265, September 2020. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Pagacz:2022:RPM

[PW22] Patryk Pagacz and Michał Wojtylak. Random perturbations of matrix polynomials. Journal of Theoretical Probability, 35 (1):52–88, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01048-3.

Park:2020:UDR

[PXY20] Hyunchul Park, Yimin Xiao, and Xiaochuan Yang. Uniform dimension results for the inverse images of symmetric Lévy processes. Journal of Theoretical Probability, 33(4):2213–2232, December 2020. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-019-00956-3.

Peng:2022:WZA

[PZ22] Shige Peng and Huilin Zhang. Wong–Zakai approximation for stochastic differential equations driven by G-Brownian1 motion. Journal of Theoretical Probability, 35(1):410–425, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01058-1.

Qiao:2014:EES

[Qia14] Huijie Qiao. Exponential ergodicity for SDEs with jumps and non–Lipschitz coefficients. Journal of Theoretical Probability, 27 (1):137–152, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0440-5.

Qin:2024:ASM

[Qin24] Yifeng Qin. Approximation schemes for McKean–Vlasov and Boltzmann-type equations (error analysis in total variation distance). Journal of Theoretical Probability, 37(2):1523–1596, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-024-01324-6.

Qiu:2023:IWA

[QK23] Qinjing Qiu and Reiichiro Kawai. Iterative weak approximation and hard bounds for switching diffusion. Journal of Theoretical Probability, 36(2):1003–1036, June 2023. CODEN JTPREO.

ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-022-01185-x.

Quang:2018:SKU

[QSS18] Nguyen Van Quang, Do The Son, and Le Hong Son. Some kinds of uniform integrability and laws of large numbers in noncommutative probability. Journal of Theoretical Probability, 31 (2):1212–1234, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Queffelec:2012:MAA

[QV12] Hervé Queffélec and Dalibor Volný. On martingale approximation of adapted processes. Journal of Theoretical Probability, 25 (2):438–449, June 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0386-z.

Qi:2020:SRP

[QX20] Yongcheng Qi and Mengzi Xie. Spectral radii of products of random rectangular matrices. Journal of Theoretical Probability, 33(4):2185–2212, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-019-00942-9.

Qian:2011:EPR

[QZ11] Min Qian and Fuxi Zhang. Entropy production rate of the coupled diffusion process. Journal of Theoretical Probability, 24(3): 729–745, September 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0352-9.

Radulovic:1998:CWB

[Rad98] Dragan Radulović. Can we bootstrap even if CLT fails? Journal of Theoretical Probability, 11(3):813–830, July 1998. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022662800620.

Rafler:2013:PSP

[Raf13] Mathias Rafler. The Pólya sum process: Limit theorems for conditioned random fields. Journal of Theoretical Probability, 26(4): 1097–1116, December 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0404-9.

Rahmani:2022:MTC

[Rah22] John Rahmani. Mixing times for the commuting chain on CA groups. Journal of Theoretical Probability, 35(1):457–483, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01049-2.

Rains:1997:CPB

[Rai97] E. M. Rains. Combinatorial properties of Brownian motion on the compact classical groups. Journal of Theoretical Probability, 10(3):659–679, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022601711176.

Raic:2004:MCD

[Rai04] M. Raic. A multivariate CLT for decomposable random vectors with finite second moments. Journal of Theoretical Probability, 17(3):573–603, July 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000040290.44087.68.

Rai:2007:SRG

[Rai07] Sanatan Rai. The spectrum of a random geometric graph is concentrated. Journal of Theoretical Probability, 20(2):119–132, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-006-0049-7.

Raja:1999:CHS

[Raj99] C. Robinson Edward Raja. On a class of Hungarian semigroups and the factorization theorem of Khinchin. Journal of Theoretical Probability, 12(2):561–569, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1021642531006.

Raja:2015:ODM

[Raj15] C. R. E. Raja. Operator decomposable measures and stochastic difference equations. Journal of Theoretical Probability, 28(3): 785–803, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0534-8.

Randjiou:2006:JIT

[Ran06] Youssef Randjiou. A joint integral test for the locations of extrema for Brownian motion. Journal of Theoretical Probability, 19(3):701–720, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-006-0030-5.

Rao:2005:MNC

[Rao05] Murali Rao. More on a new concept of entropy and information. Journal of Theoretical Probability, 18(4):967–981, October 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-005-7541-3.

Ratanov:2020:KLP

[Rat20] Nikita Ratanov. Kac–Lévy processes. Journal of Theoretical Probability, 33(1):239–267, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Rubin:1988:SSP

[RC88] Herman Rubin and Jeesen Chen. Some stochastic processes related to random density functions. Journal of Theoretical Probability, 1(2):227–237, April 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01046937.

Rebrova:2020:CRR

[Reb20] Elizaveta Rebrova. Constructive regularization of the random matrix norm. Journal of Theoretical Probability, 33(3):1768– 1790, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Redelmeier:2011:GER

[Red11] C. Emily I. Redelmeier. Genus expansion for real Wishart matrices. Journal of Theoretical Probability, 24(4):1044–1062, December 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-010-0278-7.

Ressel:1994:NHF

[Res94] Paul Ressel. Non-homogeneous de Finetti-type theorems. Journal of Theoretical Probability, 7(2):469–482, April 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02214278.

Ressel:2013:FEL

[Res13] Paul Ressel. Finite exchangeability, Lévy–Frailty copulas and higher-order monotonic sequences. Journal of Theoretical Probability, 26(3):666–675, September 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-011-0389-9.

Retzlaff:2003:RDC

[Ret03] Todd Retzlaff. Rate of decay of concentration functions on discrete groups. Journal of Theoretical Probability, 16(2):391–399, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1023522727662.

Retzlaff:2004:DCF

[Ret04] T. M. Retzlaff. Decay of concentration functions for adapted probabilities on discrete groups. Journal of Theoretical Probability, 17(4):1031–1040, October 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-004-0589-7.

Revesz:1993:PPI

[Rév93] P. Révész. Path properties of an infinite system of Wiener processes. Journal of Theoretical Probability, 6(2):353–383, April 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01047579.

Revesz:1998:MCB

[Rév98] P. Révész. The maximum of a critical branching Wiener process. Journal of Theoretical Probability, 11(4):953–977, October 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022664831729.

Revelle:2001:BRW

[Rev01] David Revelle. Biased random walk on lamplighter groups and graphs. Journal of Theoretical Probability, 14(2):379–391, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1011159613389.

Richard:2017:SSS

[Ric17] Alexandre Richard. Some singular sample path properties of a multiparameter fractional Brownian motion. Journal of Theoretical Probability, 30(4):1285–1309, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Rigo:2020:NDT

[Rig20] Pietro Rigo. A note on duality theorems in mass transportation. Journal of Theoretical Probability, 33(4):2337–2350, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-019-00932-x.

Rio:2009:MIS

[Rio09] Emmanuel Rio. Moment inequalities for sums of dependent random variables under projective conditions. Journal of Theoretical Probability, 22(1):146–163, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-008-0155-9.

Rivero:2003:RSC

[Riv03] Víctor Manuel Rivero. On random sets connected to the partial records of Poisson point process. Journal of Theoretical Probability, 16(1):277–307, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022247025107.

Rozora:2024:CUN

[RMVP24] Iryna Rozora, Yurii Mlavets, Olga Vasylyk, and Volodymyr Polishchuk. On convergence of the uniform norm and approximation for stochastic processes from the space F_{ψ}(Ω). Journal of Theoretical Probability, 37(2):1627–1653, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01309-x.

Rochet:2017:COH

[Roc17] Jean Rochet. Complex outliers of Hermitian random matrices. Journal of Theoretical Probability, 30(4):1624–1654, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Rosinski:1994:USR

[Ros94] Jan Rosiński. On uniqueness of the spectral representation of stable processes. Journal of Theoretical Probability, 7(3):615–634, July 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02213572.

Rosenblatt:1997:SSR

[Ros97] Murray Rosenblatt. Some simple remarks on an autoregressive scheme and an implied problem. Journal of Theoretical Probability, 10(2):295–305, April 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022604315045.

Rosenthal:2003:GCR

[Ros03] Jeffrey S. Rosenthal. Geometric convergence rates for timesampled Markov chains. Journal of Theoretical Probability, 16 (3):671–688, July 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1025672516474.

Rossi:2019:DRH

[Ros19] Maurizia Rossi. The defect of random hyperspherical harmonics. Journal of Theoretical Probability, 32(4):2135–2165, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Rozkosz:2013:SRW

[Roz13] Andrzej Rozkosz. Stochastic representation of weak solutions of viscous conservation laws: A BSDE approach. Journal of Theoretical Probability, 26(4):1061–1083, December 2013. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-011-0395-y.pdf.

Rozovsky:2019:SBP

[Roz19] Leonid V. Rozovsky. Small ball probabilities for certain Gaussian random fields. Journal of Theoretical Probability, 32(2):934–949, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Rothmann:1992:BCS

[RR92] Mark D. Rothmann and Ralph P. Russo. On boundary crossings of the sample p th quantile. Journal of Theoretical Probability, 5(2):407–416, April 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046744.

Rothmann:1994:SCM

[RR94] Mark D. Rothmann and Ralph P. Russo. Strong consistency of the mean on randomly deleted sets. Journal of Theoretical Probability, 7(1):119–128, January 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02213362.

Rothmann:1997:LPT

[RR97] Mark D. Rothmann and Ralph P. Russo. On the limiting proportion of types. Journal of Theoretical Probability, 10(1):131–143, January 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022698532576.

Rothmann:2000:LLN

[RR00] Mark D. Rothmann and Ralph P. Russo. Laws of large numbers for observations that change with time. Journal of Theoretical Probability, 13(4):1013–1025, October 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1007814007981.

Ramirez:2008:SGS

[RR08] José A. Ramírez and Brian Rider. Shape of the ground state energy density of Hill’s equation with nice Gaussian potential. Journal of Theoretical Probability, 21(2):413–430, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-007-0105-y.

Racz:2019:STW

[RR19] Miklós Z. Rácz and Jacob Richey. A smooth transition from Wishart to GOE. Journal of Theoretical Probability, 32(2):898– 906, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Rajput:2007:UCT

[RRM07] Balram S. Rajput and Kavi Rama-Murthy. Uniform comparison of tails of (non-symmetric) probability measures and their symmetrized counterparts with applications. Journal of Theoretical Probability, 20(1):87–105, March 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-006-0050-1.

Rajput:1994:SSS

[RRMZ94] Balram S. Rajput, Kavi Rama-Murthy, and Tomasz Zak. Supports of semi-stable probability measures on locally convex spaces. Journal of Theoretical Probability, 7(4):931–942, October 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02214379.

Rachev:1992:UCL

[RRS92] S. T. Rachev, L. Rüschendorf, and A. Schief. Uniformities for the convergence in law and in probability. Journal of Theoretical Probability, 5(1):33–44, January 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01046777.

Rukhin:1991:BEW

[RS91] Andrew L. Rukhin and Morris Skibinsky. Bayes estimators whose range has convex hull of zero prior probability. Journal of Theoretical Probability, 4(2):465–474, April 1991. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01258749.

Rachev:1994:GSD

[RS94] Svetlozar T. Rachev and Gennady Samorodnitsky. Geometric stable distributions in Banach spaces. Journal of Theoretical Probability, 7(2):351–373, April 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214274.

Rao:1995:LDC

[RS95] B. L. S. Prakasa Rao and Arusharka Sen. Limit distributions of conditional U-statistics. Journal of Theoretical Probability, 8 (2):261–301, April 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02212880.

Rackauskas:2004:NSC

[RS04] Alfredas Rackauskas and Charles Suquet. Necessary and sufficient condition for the functional Central Limit Theorem in Hölder spaces. Journal of Theoretical Probability, 17(1):221–243, January 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOTP.0000020482.66224.6c.

Roy:2008:SSS

[RS08] Parthanil Roy and Gennady Samorodnitsky. Stationary symmetric α-stable discrete parameter random fields. Journal of Theoretical Probability, 21(1):212–233, March 2008. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0107-9.

Ramanan:2018:BLC

[RS18] Kavita Ramanan and Aaron Smith. Bounds on lifting continuous-state Markov chains to speed up mixing. Journal of Theoretical Probability, 31(3):1647–1678, September 2018. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Rickelhoff:2024:TDS

[RS24] Sebastian Rickelhoff and Alexander Schnurr. The time– dependent symbol of a non-homogeneous Itô process and corresponding maximal inequalities. Journal of Theoretical Probability, 37(3):2508–2533, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-023-01308-y.

Rovira:1996:LSN

[RSS96] Carles Rovira and Marta Sanz-Solé. The law of the solution to a nonlinear hyperbolic SPDE. Journal of Theoretical Probability, 9 (4):863–901, October 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214255.

Ren:2024:LDS

[RSZ24] Yan-Xia Ren, Renming Song, and Rui Zhang. Lower deviation for the supremum of the support of super-Brownian motion. Journal of Theoretical Probability, 37(2):1079–1123, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01292-3.

Rifi:2000:ABW

[RU00] Mohamed Rifi and Frederic Utzet. On the asymptotic behavior of weighted U-statistics. Journal of Theoretical Probability, 13 (1):141–167, January 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007786926882.

Rubshtein:1995:CRM

[Rub95] Ben-Zion Rubshtein. Convolutions of random measures on compact groups. Journal of Theoretical Probability, 8(3):523–538, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02218042.

Rukhin:1993:IPD

[Ruk93] Andrew L. Rukhin. Influence of the prior distribution on the risk of the Bayes rule. Journal of Theoretical Probability, 6(1):71–87, January 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01046769.

Rushton:2007:FLD

[Rus07] Joshua Rushton. A functional LIL for d-dimensional stable processes; invariance for Lévy- and other weakly convergent processes. Journal of Theoretical Probability, 20(3):397– 427, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0059-0.

Ruzsa:1988:IDI

[Ruz88] Imre Z. Ruzsa. Infinite divisibility II. Journal of Theoretical Probability, 1(4):327–339, October 1988. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01048723.

Rosalsky:2004:WLB

[RV04] A. Rosalsky and A. I. Volodin. On the weak limiting behavior of almost surely convergent row sums from infinite arrays of rowwise independent random elements in Banach spaces. Journal of Theoretical Probability, 17(2):327–346, April 2004.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000020697.69668.e7.

Rost:1996:RKF

[RW96] Daniel Rost and Christian Wieckenberg. A representation of the Kantorovich-functional. Journal of Theoretical Probability, 9(1):87–103, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213735.

Rempala:2002:CLT

[RW02] Grzegorz A. Rempala and Jacek Wesolowski. Central limit theorems for random permanents with correlation structure. Journal of Theoretical Probability, 15(1):63–76, January 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1013837205669.

Robinson:2005:SNC

[RW05] John Robinson and Qiying Wang. On the self-normalized Cramér-type large deviation. Journal of Theoretical Probability, 18(4):891–909, October 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-005-7531-5.

Ren:2015:GLD

[RWZ15] Jiagang Ren, Jing Wu, and Hua Zhang. General large deviations and functional iterated logarithm law for multivalued stochastic differential equations. Journal of Theoretical Probability, 28 (2):550–586, June 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0531-y.

Ross:1993:CTC

[RX93] Kenneth A. Ross and Daming Xu. A comparison theorem on convergence rates of random walks on groups. Journal of Theoretical Probability, 6(2):323–343, April 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01047577.

Ross:1994:HDM

[RX94] Kenneth A. Ross and Daming Xu. Hypergroup deformations and Markov chains. Journal of Theoretical Probability, 7(4):813–830, October 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02214374.

Ren:2010:LDM

[RXZ10] Jiagang Ren, Siyan Xu, and Xicheng Zhang. Large deviations for multivalued stochastic differential equations. Journal of Theoretical Probability, 23(4):1142–1156, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-009-0274-y.

Rachev:1991:RCS

[RY91] S. T. Rachev and J. E. Yukich. Rates of convergence of αstable random motions. Journal of Theoretical Probability, 4 (2):333–352, April 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01258741.

Ray:2023:QRS

[RY23] Gourab Ray and Tingzhou Yu. Quantitative Russo–Seymour– Welsh for random walk on random graphs and decorrelation of uniform spanning trees. Journal of Theoretical Probability, 36(4): 2284–2310, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01248-7.

Ren:2024:LDC

[RY24] Yan-Xia Ren and Ting Yang. Limiting distributions for a class of super-Brownian motions with spatially dependent branching mechanisms. Journal of Theoretical Probability, 37(3):2457– 2507, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01304-2.

Ryabko:2010:DBP

[Rya10] Daniil Ryabko. Discrimination between B-processes is impossible. Journal of Theoretical Probability, 23(2):565–575, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-009-0263-1.

Ryvkina:2015:FBM

[Ryv15] Jelena Ryvkina. Fractional Brownian motion with variable Hurst parameter: Definition and properties. Journal of Theoretical Probability, 28(3):866–891, September 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-013-0502-3.

Rouault:2002:LDP

[RYZ02] Alain Rouault, Marc Yor, and Marguerite Zani. A large deviations principle related to the strong arc-sine law. Journal of Theoretical Probability, 15(3):793–815, July 2002. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1016280117892.

Ryznar:1990:MTB

[RZ90] Michał Ryznar and Tomasz Zak. The measure of a translated ball in uniformly convex spaces. Journal of Theoretical Probability, 3 (4):547–562, October 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046095.

Rosinski:1997:EEW

[RZ97] Jan Rosiński and Tomasz Zak. The equivalence of ergodicity and weak mixing for infinitely divisible processes. Journal of Theoretical Probability, 10(1):73–86, January 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022690230759.

Ran:2010:EUB

[RZ10] Qikang Ran and Tusheng Zhang. Existence and uniqueness of bounded weak solutions of a semilinear parabolic PDE. Journal of Theoretical Probability, 23(4):951–971, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-009-0252-4.

Rafler:2015:LPB

[RZ15] Mathias Rafler and Hans Zessin. The logical postulates of Böge, Carnap and Johnson in the context of Papangelou processes. Journal of Theoretical Probability, 28(4):1431–1446, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-014-0543-2.

Ren:2019:RLS

[RZ19] Jiagang Ren and Hua Zhang. Regularity of the law of stochastic differential equations with jumps under Hörmander’s conditions: The lent particle method. Journal of Theoretical Probability, 32 (2):1023–1050, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Saada:2001:ABC

[Saa01] Diane Saada. Asymptotic behavior of continuous set-indexed martingales. Journal of Theoretical Probability, 14(2):319–332, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1011103428410.

Sadasue:2008:QII

[Sad08] Gaku Sadasue. On quasi-invariance of infinite product measures. Journal of Theoretical Probability, 21(3):571–585, September 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0171-9.

Sahi:2008:FIP

[Sah08] Siddhartha Sahi. The FKG inequality for partially ordered algebras. Journal of Theoretical Probability, 21(2):449–458, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-007-0117-7.

Salvatori:1992:NGI

[Sal92] Maura Salvatori. On the norms of group-invariant transition operators on graphs. Journal of Theoretical Probability, 5 (3):563–576, July 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01060436.

Samia:2021:TCL

[Sam21] Sakrani Samia. Time-changed local martingales under signed measures. Journal of Theoretical Probability, 34(2):644–659, June 2021. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-020-00994-2.

Sampedro:2024:SPL

[Sam24] Juan Carlos Sampedro. On the L^{p}-spaces of projective limits of probability measures. Journal of Theoretical Probability, 37(3): 2665–2703, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-024-01329-1.

Sancetta:2007:WCL

[San07] Alessio Sancetta. Weak convergence of laws on R^{K} with common marginals. Journal of Theoretical Probability, 20(2):371–380, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-007-0096-8.

Sandric:2014:RTC

[San14] Nikola Sandrić. Recurrence and transience criteria for two cases of stable-like Markov chains. Journal of Theoretical Probability, 27(3):754–788, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-012-0445-0.

Sandric:2016:EPS

[San16] Nikola Sandrić. Ergodic property of stable-like Markov chains. Journal of Theoretical Probability, 29(2):459–490, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-014-0586-4.

Sarantsev:2017:RBM

[Sar17] Andrey Sarantsev. Reflected Brownian motion in a convex polyhedral cone: Tail estimates for the stationary distribution. Journal of Theoretical Probability, 30(3):1200–1223, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Sarantsev:2019:CTC

[Sar19] Andrey Sarantsev. Comparison techniques for competing Brownian particles. Journal of Theoretical Probability, 32(2):545–585, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Sato:1992:NND

[Sat92] Hiroshi Sato. Nuclearity of a nonnegative definite integral kernel on a separable metric space. Journal of Theoretical Probability, 5(2):349–353, April 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046740.

Sauer:2016:UKO

[Sau16] Martin Sauer. L^{1}-uniqueness of Kolmogorov operators associated with two-dimensional stochastic Navier–Stokes Coriolis equations with space–time white noise. Journal of Theoretical Probability, 29(2):569–589, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-014-0582-8.

Savov:2010:STO

[Sav10] Mladen Savov. Small time one-sided LIL behavior for Lévy processes at zero. Journal of Theoretical Probability, 23(1):209–236, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0202-6.

Sawyer:2001:CLT

[Saw01] P. Sawyer. The Central Limit Theorem on SO_{0}(p, q)/SO(p) × SO(q). Journal of Theoretical Probability, 14(3):857–866, July 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1017557526199.

Stan:2021:MRV

[SC21] Aurel I. Stan and Florin Catrina. 1-Meixner random vectors. Journal of Theoretical Probability, 34(4):2033–2080, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01023-y.

Scheffler:1994:DAS

[Sch94a] Hans-Peter Scheffler. D-domains of attraction of stable measures on stratified Lie groups of attraction of stable measures on stratified Lie groups. Journal of Theoretical Probability, 7 (4):767–792, October 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214371.

Schuette:1994:LDT

[Sch94b] Paul H. Schuette. Large deviations for trajectories of sums of independent random variables. Journal of Theoretical Probability, 7(1):3–45, January 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02213359.

Schilling:1998:FPG

[Sch98] René L. Schilling. Feller processes generated by pseudodifferential operators: On the Hausdorff dimension of their sample paths. Journal of Theoretical Probability, 11(2):303–330, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022678219821.

Schied:1999:ERC

[Sch99] Alexander Schied. Existence and regularity for a class of infinitemeasure (ξ, ψ, K)-superprocesses. Journal of Theoretical Probability, 12(4):1011–1035, October 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1021645204173.

Schoolfield:2002:RWW

[Sch02a] Clyde H. Schoolfield, Jr. Random walks on wreath products of groups. Journal of Theoretical Probability, 15(3):667–693, July 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1016219932004.

Schoolfield:2002:SGB

[Sch02b] Clyde H. Schoolfield, Jr. A signed generalization of the Bernoulli–Laplace diffusion model. Journal of Theoretical Probability, 15(1):97–127, January 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1013841306577.

Schoolfield:2005:GRS

[Sch05] Clyde H. Schoolfield, Jr. Generating a random signed permutation with random reversals. Journal of Theoretical Probability, 18 (4):911–931, October 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-005-7532-4.

Schweinsberg:2008:LER

[Sch08] Jason Schweinsberg. Loop-erased random walk on finite graphs and the Rayleigh process. Journal of Theoretical Probability, 21 (2):378–396, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0125-7.

Schack:2009:OWS

[Sch09] Helga Schack. An optimal wavelet series expansion of the Riemann–Liouville process. Journal of Theoretical Probability, 22(4):1030–1057, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0187-1.

Schlemm:2016:KSI

[Sch16a] Eckhard Schlemm. The Kearns–Saul inequality for Bernoulli and Poisson–binomial distributions. Journal of Theoretical Probability, 29(1):48–62, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-014-0564-x.

Schulte:2016:NAP

[Sch16b] Matthias Schulte. Normal approximation of Poisson functionals in Kolmogorov distance. Journal of Theoretical Probability, 29 (1):96–117, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0576-6.

Schaeffer:2024:MSL

[Sch24] Nicolas Schaeffer. Multilinear smoothing and local wellposedness of a stochastic quadratic nonlinear Schrödinger equation. Journal of Theoretical Probability, 37(1):160–208, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01282-5.

Scott:1990:NID

[Sco90] Damon Scott. A non-integral-dimensional random walk. Journal of Theoretical Probability, 3(1):1–7, January 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01063325.

Son:2022:CLT

[SD22] Ta Cong Son and Le Van Dung. Central limit theorems for weighted sums of dependent random vectors in Hilbert spaces via the theory of the regular variation. Journal of Theoretical Probability, 35(2):988–1012, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01079-4.

Seneta:1993:EPI

[Sen93] E. Seneta. Ergodicity for products of infinite stochastic matrices. Journal of Theoretical Probability, 6(2):345–352, April 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01047578.

Sengupta:2000:TSH

[Sen00] Arindam Sengupta. Time-space harmonic polynomials for continuous-time processes and an extension. Journal of Theoretical Probability, 13(4):951–976, October 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1007857823002.

Sencimen:2013:SCP

[Sen13] Celaleddin Sençimen. Statistical convergence in probability for a sequence of random functions. Journal of Theoretical Probability, 26(1):94–106, March 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0346-7.

Sepanski:1994:PCG

[Sep94] Steven J. Sepanski. Probabilistic characterizations of the generalized domain of attraction of the multivariate normal law. Journal of Theoretical Probability, 7(4):857–866, October 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02214376.

Sepanski:1997:SIP

[Sep97] Steven J. Sepanski. Some invariance principles for random vectors in the generalized domain of attraction of the multivariate normal law. Journal of Theoretical Probability, 10(4):1053–1063, October 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022622902495.

Sepanski:1999:CLI

[Sep99] Steven J. Sepanski. A compact law of the iterated logarithm for random vectors in the generalized domain of attraction of the multivariate Gaussian law. Journal of Theoretical Probability, 12(3):757–778, July 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021632016642.

Sepanski:2001:EVM

[Sep01] Steven J. Sepanski. Extreme values and the multivariate compact law of the iterated logarithm. Journal of Theoretical Probability, 14(4):989–1018, October 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017576903851.

Serlet:1996:OMS

[Ser96] Laurent Serlet. The occupation measure of super-Brownian motion conditioned to nonextinction. Journal of Theoretical Probability, 9(3):561–578, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214075.

Sarymsakov:1990:AMT

[SG90] T. A. Sarymsakov and N. N. Ganihodzhaev. Analytic methods in the theory of quadric stochastic operators. Journal of Theoretical Probability, 3(1):51–70, January 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01063328.

Sgibnev:1998:ABH

[Sgi98] M. S. Sgibnev. On the asymptotic behavior of the harmonic renewal measure. Journal of Theoretical Probability, 11(2): 371–382, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022627704800.

Son:2022:CTS

[SGQ22] Do The Son, Duong Xuan Giap, and Nguyen Van Quang. Chungtype strong laws and almost complete convergence for arrays of measurable operators. Journal of Theoretical Probability, 35(3): 1391–1411, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01108-2.

Shah:1991:IDM

[Sha91] Riddhi Shah. Infinitely divisible measures on p-adic groups. Journal of Theoretical Probability, 4(2):391–405, April 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01258744.

Shao:1993:NSB

[Sha93] Qi-Man Shao. A note on small ball probability of a Gaussian process with stationary increments. Journal of Theoretical Probability, 6(3):595–602, July 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01066719.

Shao:1999:CTL

[Sha99] Qi-Man Shao. A Cramér type large deviation result for Student’s t-statistic. Journal of Theoretical Probability, 12(2):385–398, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021626127372.

Shah:2000:SMS

[Sha00a] Riddhi Shah. Selfdecomposable measures on simply connected nilpotent groups. Journal of Theoretical Probability, 13(1):65–83, January 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007778725065.

Shao:2000:CTM

[Sha00b] Qi-Man Shao. A comparison theorem on moment inequalities between negatively associated and independent random variables. Journal of Theoretical Probability, 13(2):343–356, April 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007849609234.

Shields:1993:TDR

[Shi93a] Paul C. Shields. Two divergence-rate counterexamples. Journal of Theoretical Probability, 6(3):521–545, July 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01066716.

Shields:1993:WTP

[Shi93b] Paul C. Shields. Waiting times: Positive and negative results on the Wyner–Ziv problem. Journal of Theoretical Probability, 6(3):499–519, July 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01066715.

Shi:1996:SBP

[Shi96] Z. Shi. Small ball probabilities for a Wiener process under weighted sup-norms, with an application to the supremum of Bessel local times. Journal of Theoretical Probability, 9(4): 915–929, October 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214257.

Shi:1997:TDB

[Shi97] Z. Shi. Time difference on the Brownian curve. Journal of Theoretical Probability, 10(3):625–642, July 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1022697510267.

Shiu:2013:CCC

[Shi13] Shang-Yuan Shiu. Covering a connected curve on the torus with squares. Journal of Theoretical Probability, 26(2):480–488, June 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-011-0387-y.

Siegel:1990:POS

[Sie90] G. Siegel. Properties of operator stable distributions in infinitedimensional Banach spaces. Journal of Theoretical Probability, 3(2):227–244, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01045160.

Siebert:1992:ODG

[Sie92] Eberhard Siebert. Operator-decomposability of Gaussian measures on separable Banach spaces. Journal of Theoretical Probability, 5(2):333–347, April 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01046739.

Siegrist:1994:EDS

[Sie94] Kyle Siegrist. Exponential distributions on semigroups. Journal of Theoretical Probability, 7(4):725–737, October 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02214368.

Siegrist:2006:DED

[Sie06] Kyle Siegrist. Decomposition of exponential distributions on positive semigroups. Journal of Theoretical Probability, 19(1): 204–220, January 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-006-0005-6.

Sifi:1995:CLT

[Sif95] Mohamed Sifi. Central limit theorem and infinitely divisible probabilities associated with partial differential operators. Journal of Theoretical Probability, 8(3):475–499, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF02218040.

Sikic:2000:ZOL

[Sik00] Hrvoje Sikić. Zero–one law for some Brownian functionals. Journal of Theoretical Probability, 13(2):571–574, April 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1007833205600.

Simon:2004:SBE

[Sim04] T. Simon. Small ball estimates in p-variation for stable processes. Journal of Theoretical Probability, 17(4):979–1002, October 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-004-0586-x.

Simper:2024:RTC

[Sim24] Mackenzie Simper. Random transpositions on contingency tables. Journal of Theoretical Probability, 37(3):1988–2022, September 2024. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-023-01286-1.

Singh:2007:STD

[Sin07] Arvind Singh. A slow transient diffusion in a drifted stable potential. Journal of Theoretical Probability, 20(2):153–166, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-007-0056-3.

Shao:1998:ERP

[SJ98] Yongzhao Shao and Raúl Jiménez. Entropy for random partitions and its applications. Journal of Theoretical Probability, 11 (2):417–433, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022683822547.

Schirokauer:2007:UDN

[SK07] Oliver Schirokauer and Joseph B. Kadane. Uniform distributions on the natural numbers. Journal of Theoretical Probability, 20(3):429–441, September 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0066-1.

Stern:2015:CCM

[SK15] Rafael B. Stern and Joseph B. Kadane. Coherence of countably many bets. Journal of Theoretical Probability, 28(2):520–538, June 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-013-0489-9.

Spece:2020:PRC

[SK20] Michael Spece and Joseph B. Kadane. Prime residue class of uniform charges on the integers. Journal of Theoretical Probability, 33(1):340–360, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Slaby:1989:LIL

[Sla89] Marek Slaby. The law of the iterated logarithm for subsequences and characterization of the cluster set of S_{nk}/(2n_{k} log log n_{k})^{1/2} in Banach spaces. Journal of Theoretical Probability, 2(3): 343–376, July 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01054021.

Slavik:2022:LMD

[Sla22] Jakub Slavík. Large and moderate deviations principles and central limit theorem for the stochastic 3D primitive equations with gradient-dependent noise. Journal of Theoretical Probability, 35(3):1736–1781, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01125-1.

Slonowsky:2001:SMT

[Slo01] Dean Slonowsky. Strong martingales: Their decompositions and quadratic variation. Journal of Theoretical Probability, 14 (3):609–638, July 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017536921656.

Smith:2015:CTM

[Smi15] Aaron Smith. Comparison theory for Markov chains on different state spaces and application to random walk on derangements. Journal of Theoretical Probability, 28(4):1406–1430, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-014-0559-7.

Sobieczky:2010:ITS

[Sob10] Florian Sobieczky. An interlacing technique for spectra of random walks and its application to finite percolation clusters. Journal of Theoretical Probability, 23(3):639–670, September 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-010-0298-3.

Solomyak:1999:ESF

[Sol99] Rita Solomyak. Essential spanning forests and electrical networks on groups. Journal of Theoretical Probability, 12(2): 523–548, April 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021638430098.

Solan:2003:CVC

[Sol03] Eilon Solan. Continuity of the value of competitive Markov decision processes. Journal of Theoretical Probability, 16(4):831–845, October 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/B%3AJOTP.0000011995.28536.ef.

Soltanifar:2014:OVS

[Sol14] Mohsen Soltanifar. On operator-valued semicircular random variables. Journal of Theoretical Probability, 27(1):41–56, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-012-0406-7.

Song:1995:FKS

[Son95] Renming Song. Feynman–Kac semigroup with discontinuous additive functionals. Journal of Theoretical Probability, 8(4): 727–762, October 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02410109.

Song:2006:ETD

[Son06] Renming Song. Estimates on the transition densities of Girsanov transforms of symmetric stable processes. Journal of Theoretical Probability, 19(2):487–507, June 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-006-0023-4.

Song:2020:GEE

[Son20] Yulin Song. Gradient estimates and exponential ergodicity for mean-field SDEs with jumps. Journal of Theoretical Probability, 33(1):201–238, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Song:2021:SER

[Son21a] Shiyu Song. Some explicit results on first exit times for a jump diffusion process involving semimartingale local time. Journal of Theoretical Probability, 34(4):2346–2367, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01040-x.

Sonmez:2021:SPP

[Sön21b] Ercan Sönmez. Sample path properties of generalized random sheets with operator scaling. Journal of Theoretical Probability, 34(3):1279–1298, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-020-01045-6.

Sottinen:2004:GPE

[Sot04] T. Sottinen. On Gaussian processes equivalent in law to fractional Brownian motion. Journal of Theoretical Probability, 17 (2):309–325, April 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000020696.99064.5d.

Sowers:2008:RPC

[Sow08] Richard B. Sowers. Random perturbations of canards. Journal of Theoretical Probability, 21(4):824–889, December 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-008-0150-1.

Spataru:1999:PAS

[Spa99] Aurel Spataru. Precise asymptotics in Spitzer’s law of large numbers. Journal of Theoretical Probability, 12(3):811–819, July 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/A%3A1021636117551.

Spataru:2004:EAL

[Spa04] A. Spataru. Exact asymptotics in log log laws for random fields. Journal of Theoretical Probability, 17(4):943–965, October 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-004-0584-z.

Spataru:2010:LIL

[Spa10] Aurel Spataru. The law of the iterated logarithm for finitely inhomogeneous random walks. Journal of Theoretical Probability, 23(2):417–427, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0213-y.

Spataru:2016:CPA

[Spa16] Aurel Spataru. Convergence and precise asymptotics for series involving self-normalized sums. Journal of Theoretical Probability, 29(1):267–276, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-014-0560-1.

Spiliopoulos:2012:LDP

[Spi12] Konstantinos Spiliopoulos. Large deviations principle for a large class of one-dimensional Markov processes. Journal of Theoretical Probability, 25(4):925–949, December 2012. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
011-0345-8.

Shyamalkumar:2007:OPO

[SRL07] Nariankadu D. Shyamalkumar, Ralph P. Russo, and Ramon Lawrence. Optimal policies to obtain the most join results. Journal of Theoretical Probability, 20(2):237–256, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-007-0063-4.

Silvestrov:1998:ETI

[SS98] Dmitrii S. Silvestrov and Örjan Stenflo. Ergodic theorems for iterated function systems controlled by regenerative sequences. Journal of Theoretical Probability, 11(3):589–608, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022642328845.

Saada:2006:NSL

[SS06] Diane Saada and Dean Slonowsky. A notion of stopping line for set-indexed processes. Journal of Theoretical Probability, 19 (2):397–410, June 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-006-0015-4.

Szabados:2009:SIB

[SS09] Tamás Szabados and Balázs Székely. Stochastic integration based on simple, symmetric random walks. Journal of Theoretical Probability, 22(1):203–219, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-007-0140-8.

Sendov:2015:NRT

[SS15] Hristo Sendov and Shen Shan. New representation theorems for completely monotone and Bernstein functions with convexity properties on their measures. Journal of Theoretical Probability, 28(4):1689–1725, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-014-0557-9.

Sturm:2018:PDM

[SS18] Anja Sturm and Jan M. Swart. Pathwise duals of monotone and additive Markov processes. Journal of Theoretical Probability, 31(2):932–983, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Salminen:2021:OTO

[SS21] Paavo Salminen and David Stenlund. On occupation times of one-dimensional diffusions. Journal of Theoretical Probability, 34(2):975–1011, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-00993-3.

Sambale:2022:CIM

[SS22] Holger Sambale and Arthur Sinulis. Concentration inequalities on the multislice and for sampling without replacement. Journal of Theoretical Probability, 35(4):2712–2737, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01139-9.

Samorodnitsky:1990:MSI

[ST90] Gennady Samorodnitsky and Murad S. Taqqu. Multiple stable integrals of Banach-valued functions. Journal of Theoretical Probability, 3(2):267–287, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01045162.

Sottinen:2006:EMG

[ST06] Tommi Sottinen and Ciprian A. Tudor. On the equivalence of multiparameter Gaussian processes. Journal of Theoretical Probability, 19(2):461–485, June 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-006-0022-5.

Sanchez:2015:CBM

[ST15] Juan Fernández Sánchez and Wolfgang Trutschnig. Conditioningbased metrics on the space of multivariate copulas and their interrelation with uniform and levelwise convergence and Iterated Function Systems. Journal of Theoretical Probability, 28(4):

1311–1336, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0541-4.

Stadje:1998:TES

[Sta98] Wolfgang Stadje. Two ergodic sample-path properties of the Poisson process. Journal of Theoretical Probability, 11(1): 197–208, January 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021603210282.

Staples:2007:GTA

[Sta07] George Stacey Staples. Graph-theoretic approach to stochastic integrals with Clifford algebras. Journal of Theoretical Probability, 20(2):257–274, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-007-0064-3.

Stan:2011:TDM

[Sta11] Aurel I. Stan. Two-dimensional Meixner random vectors of class M_{L}. Journal of Theoretical Probability, 24(1):39–65, March 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-010-0309-4.

Stankov:2022:CTE

[Sta22] Bogdan Stankov. Convergence towards the end space for random walks on Schreier graphs. Journal of Theoretical Probability, 35(3):1412–1422, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01104-6.

Steinicke:2016:FLP

[Ste16] Alexander Steinicke. Functionals of a Lévy process on canonical and generic probability spaces. Journal of Theoretical Probability, 29(2):443–458, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-014-0583-7.

Stephenson:2018:LCL

[Ste18] Robin Stephenson. Local convergence of large critical multi-type Galton–Watson trees and applications to random maps. Journal of Theoretical Probability, 31(1):159–205, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Stewart:2020:TVA

[Ste20] Kathryn Stewart. Total variation approximation of random orthogonal matrices by Gaussian matrices. Journal of Theoretical Probability, 33(2):1111–1143, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Stong:1991:CRS

[Sto91] Richard Stong. Choosing a random spanning subtree: a case study. Journal of Theoretical Probability, 4(4):753–766, October 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01259553.

Stolz:1996:SSB

[Sto96] Wolfgang Stolz. Some small ball probabilities for Gaussian processes under nonuniform norms. Journal of Theoretical Probability, 9(3):613–630, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214078.

Strobl:1995:RSM

[Str95] Franz Strobl. On the reversed sub-martingale property of empirical discrepancies in arbitrary sample spaces. Journal of Theoretical Probability, 8(4):825–831, October 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02410114.

Stufler:2022:QLC

[Stu22a] Benedikt Stufler. Quenched local convergence of Boltzmann planar maps. Journal of Theoretical Probability, 35(2):1324– 1342, June 2022. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-021-01089-2.

Stufler:2022:RMT

[Stu22b] Benedikt Stufler. Rerooting multi-type branching trees: The infinite spine case. Journal of Theoretical Probability, 35(2): 653–684, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-01069-y.

Su:1996:LIL

[Su96] Zhonggen Su. The law of the iterated logarithm and Marcinkiewicz law of large numbers for B-valued U-statistics. Journal of Theoretical Probability, 9(3):679–701, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02214082.

Shiozawa:2014:EJT

[SU14] Yuichi Shiozawa and Toshihiro Uemura. Explosion of jump-type symmetric Dirichlet forms on R^{d}. Journal of Theoretical Probability, 27(2):404–432, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-012-0424-5.

Schneider:2019:LSC

[SU19] Jan Schneider and Roman Urban. Lévy subordinators in cones of fuzzy sets. Journal of Theoretical Probability, 32(4):1909–1924, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Sudbury:2000:DFI

[Sud00] Aidan Sudbury. Dual families of interacting particle systems on graphs. Journal of Theoretical Probability, 13(3):695–716, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007806427774.

Sun:1995:PED

[Sun95] Shan Sun. Perturbed empirical distribution functions and quantiles under dependence. Journal of Theoretical Probability, 8(4): 763–777, October 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02410110.

Sun:1997:MRH

[Sun97] T. C. Sun. Murray Rosenblatt: His contributions to probability and statistics. Journal of Theoretical Probability, 10(2):279–286, April 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022600214136.

Sung:2014:MZT

[Sun14] Soo Hak Sung. Marcinkiewicz–Zygmund type strong law of large numbers for pairwise i.i.d. random variables. Journal of Theoretical Probability, 27(1):96–106, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-012-0417-4.

Sun:2022:BSD

[Sun22] Shengqiu Sun. Backward stochastic differential equations driven by G-Brownian motion with uniformly continuous coefficients in (y, z). Journal of Theoretical Probability, 35(1):370–409, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01057-2.

Salminen:2005:FRT

[SV05] Paavo Salminen and Pierre Vallois. On first range times of linear diffusions. Journal of Theoretical Probability, 18(3):567–593, July 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-005-3519-4.

Song:2006:PTS

[SV06] Renming Song and Zoran Vondracek. Potential theory of special subordinators and subordinate killed stable processes. Journal of Theoretical Probability, 19(4):817–847, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-006-0045-y.

Sottinen:2016:PII

[SV16] Tommi Sottinen and Lauri Viitasaari. Pathwise integrals and Itô–Tanaka formula for Gaussian processes. Journal of Theoretical Probability, 29(2):590–616, June 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-014-0588-2.

Szabo:2020:IPL

[SV20] Réka Szabó and Daniel Valesin. Inhomogeneous percolation on ladder graphs. Journal of Theoretical Probability, 33(2): 992–1010, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-019-00896-y.pdf.

Shafer:2012:LZO

[SVT12] Glenn Shafer, Vladimir Vovk, and Akimichi Takemura. Lévy’s zero–one law in game-theoretic probability. Journal of Theoretical Probability, 25(1):1–24, March 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-011-0390-3.

Sundberg:1992:CMM

[SW92] Carl Sundberg and Carl Wagner. Characterizations of monotone and 2-monotone capacities. Journal of Theoretical Probability, 5 (1):159–167, January 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01046783.

Stan:2009:SPM

[SW09] Aurel I. Stan and John J. Whitaker. A study of probability measures through commutators. Journal of Theoretical Probability, 22(1):123–145, March 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-008-0167-5.

Su:2020:DTT

[SW20] Zhonggen Su and Hanchao Wang. A Donsker-type theorem for log-likelihood processes. Journal of Theoretical Probability, 33 (3):1401–1425, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Shiozawa:2021:MNL

[SW21] Yuichi Shiozawa and Jian Wang. Martingale nature and laws of the iterated logarithm for Markov processes of pure-jump type. Journal of Theoretical Probability, 34(4):2005–2032, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01035-8.

Swart:2009:CPS

[Swa09] Jan M. Swart. The contact process seen from a typical infected site. Journal of Theoretical Probability, 22(3):711–740, September 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-008-0184-4.

Sweeting:1989:CWC

[Swe89] T. J. Sweeting. On conditional weak convergence. Journal of Theoretical Probability, 2(4):461–474, October 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01051878.

Shi:2022:GEE

[SWF22] Zhiyan Shi, Zhongzhi Wang, and Yan Fan. The generalized entropy ergodic theorem for nonhomogeneous bifurcating Markov chains indexed by a binary tree. Journal of Theoretical Probability, 35(3):1367–1390, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01117-1.

Sun:2022:LLN

[SYG22] Xichao Sun, Litan Yan, and Yong Ge. The laws of large numbers associated with the linear self-attracting diffusion driven by fractional Brownian motion and applications. Journal of Theoretical Probability, 35(3):1423–1478, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01126-0.

Szekely:1990:CDC

[SZ90] Gábor J. Székely and Wei-Bin Zeng. The Choquet–Deny convolution equation μ = μ ∗ σ for probability measures on Abelian semigroups. Journal of Theoretical Probability, 3(2):361–365, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01045167.

Szabados:2012:SIL

[Sza12] Tamás Szabados. Self-intersection local time of planar Brownian motion based on a strong approximation by random walks. Journal of Theoretical Probability, 25(4):1081–1118, December 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-011-0351-x.

Szewczak:2009:LTC

[Sze09] Zbigniew S. Szewczak. On limit theorems for continued fractions. Journal of Theoretical Probability, 22(1):239–255, March 2009.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-007-0139-1.

Sztonyk:2010:ETS

[Szt10] Pawel Sztonyk. Estimates of tempered stable densities. Journal of Theoretical Probability, 23(1):127–147, March 2010. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
009-0208-8.

Szulga:1992:SEM

[Szu92] Jerzy Szulga. Series expansions of multiple Lévy integrals. Journal of Theoretical Probability, 5(4):601–615, October 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01058721.

Takacs:2001:MRE

[Tak01] Christiane Takacs. More randomness of environment does not always slow down a random walk. Journal of Theoretical Probability, 14(3):699–715, July 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017593006635.

Takeda:2008:LDA

[Tak08] Masayoshi Takeda. Large deviations for additive functionals of symmetric stable processes. Journal of Theoretical Probability, 21(2):336–355, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0111-0.

Takeuchi:2010:BEL

[Tak10] Atsushi Takeuchi. Bismut–Elworthy–Li-type formulae for stochastic differential equations with jumps. Journal of Theoretical Probability, 23(2):576–604, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-010-0280-0.

Takeda:2011:LDP

[Tak11] Masayoshi Takeda. A large deviation principle for symmetric Markov processes with Feynman–Kac functional. Journal of Theoretical Probability, 24(4):1097–1129, December 2011. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0324-5.

Takeda:2018:BST

[Tak18] Masayoshi Takeda. The bottom of the spectrum of time-changed processes and the maximum principle of Schrödinger operators. Journal of Theoretical Probability, 31(2):741–756, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Takeda:2019:EUQ

[Tak19] Masayoshi Takeda. Existence and uniqueness of quasi-stationary distributions for symmetric Markov processes with tightness property. Journal of Theoretical Probability, 32(4):2006–2019, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Talagrand:1992:VVR

[Tal92] Michel Talagrand. On vector-valued random Fourier series. Journal of Theoretical Probability, 5(2):327–331, April 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01046738.

Talagrand:1996:GCP

[Tal96a] Michel Talagrand. The Glivenko–Cantelli problem, ten years later. Journal of Theoretical Probability, 9(2):371–384, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF02214655.

Talagrand:1996:LCF

[Tal96b] Michel Talagrand. Lower classes for fractional Brownian motion. Journal of Theoretical Probability, 9(1):191–213, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02213740.

Talagrand:1998:GLB

[Tal98] Michel Talagrand. A general lower bound for Gaussian chaos. Journal of Theoretical Probability, 11(1):289–301, January 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021663428937.

Tang:2007:HTI

[Tan07] Yongqiang Tang. A Hoeffding-type inequality for ergodic time series. Journal of Theoretical Probability, 20(2):167–176, June 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-007-0057-2.

Tanguy:2020:TIS

[Tan20] Kevin Tanguy. Talagrand inequality at second order and application to Boolean analysis. Journal of Theoretical Probability, 33(2):692–714, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Teicher:1992:CSN

[Tei92] Henry Teicher. Convergence of self-normalized generalized Ustatistics. Journal of Theoretical Probability, 5(2):391–405, April 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01046743.

Teicher:1995:MRS

[Tei95] Henry Teicher. Moments of randomly stopped sums-revisited. Journal of Theoretical Probability, 8(4):779–793, October 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02410111.

Teicher:1998:MZS

[Tei98a] Henry Teicher. On the Marcinkiewicz–Zygmund strong law for U-statistics. Journal of Theoretical Probability, 11(1):279–288, January 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1021611412099.

Teicher:1998:SLM

[Tei98b] Henry Teicher. Strong laws for martingale differences and independent random variables. Journal of Theoretical Probability, 11 (4):979–995, October 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022616915799.

Telcs:1995:SGF

[Tel95] András Telcs. Spectra of graphs and fractal dimensions II. Journal of Theoretical Probability, 8(1):77–96, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02213455.

Telcs:2006:SGS

[Tel06] András Telcs. Sub-Gaussian short time asymptotics for measure metric Dirichlet spaces. Journal of Theoretical Probability, 19(3): 631–645, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-006-0031-4.

Temmel:2014:SMS

[Tem14] Christoph Temmel. Shearer’s measure and stochastic domination of product measures. Journal of Theoretical Probability, 27(1):22–40, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0423-6.

Tempelman:2023:ETR

[Tem23] Arkady Tempelman. Ergodic theorems with random weights for stationary random processes and fields. Journal of Theoretical Probability, 36(3):1877–1901, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-022-
01226-5.

Teran:2016:MSL

[Ter16] Pedro Terán. A multivalued strong law of large numbers. Journal of Theoretical Probability, 29(2):349–358, June 2016. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
014-0572-x.

Tetali:1991:RWE

[Tet91] Prasad Tetali. Random walks and the effective resistance of networks. Journal of Theoretical Probability, 4(1):101–109, January 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01046996.

Thompson:2019:DFK

[Tho19] James Thompson. Derivatives of Feynman–Kac semigroups. Journal of Theoretical Probability, 32(2):950–973, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Thompson:2020:FIF

[Tho20] James Thompson. Functional inequalities for Feynman–Kac semigroups. Journal of Theoretical Probability, 33(3):1523–1540, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Turki-Moalla:1998:RCL

[TM98] Kaoukeb Turki-Moalla. Rates of convergence and law of the iterated logarithm for U-processes. Journal of Theoretical Probability, 11(4):869–906, October 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022656529912.

Teran:2006:LLN

[TM06] Pedro Terán and Ilya Molchanov. The law of large numbers in a metric space with a convex combination operation. Journal of Theoretical Probability, 19(4):875–898, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-006-0043-0.

Toby:1990:BDS

[Tob90] E. H. Toby. Birth and death of a stationary Markov process. Journal of Theoretical Probability, 3(2):289–309, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01045163.

Tomkins:1996:RCS

[Tom96] R. J. Tomkins. Rates of convergence for the stability of large order statistics. Journal of Theoretical Probability, 9(4):841–851, October 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02214253.

Trashorras:2008:LDS

[Tra08] José Trashorras. Large deviations for symmetrised empirical measures. Journal of Theoretical Probability, 21(2):397–412, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-007-0121-y.

Trinh:2019:GSF

[Tri19] Khanh Duy Trinh. Global spectrum fluctuations for Gaussian beta ensembles: A martingale approach. Journal of Theoretical Probability, 32(3):1420–1437, September 2019. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Trivellato:2024:SES

[Tri24] Barbara Trivellato. Sub-exponentiality in statistical exponential models. Journal of Theoretical Probability, 37(3):2076– 2096, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01281-6.

Tsoi:1993:IPL

[Tso93a] Allanus H. Tsoi. Integration by parts for a Lie group valued Brownian motion. Journal of Theoretical Probability, 6(4): 693–698, October 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01049171.

Tsoi:1993:TRI

[Tso93b] Allanus H. Tsoi. Time reversal of infinite-dimensional point processes. Journal of Theoretical Probability, 6(3):451–461, July 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01066711.

Tsou:2020:DPM

[Tso20] Benjamin Tsou. The distribution of permutation matrix entries under randomized basis. Journal of Theoretical Probability, 33 (3):1506–1522, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Terrin:1990:NLT

[TT90] Norma Terrin and Murad S. Taqqu. A noncentral limit theorem for quadratic forms of Gaussian stationary sequences. Journal of Theoretical Probability, 3(3):449–475, July 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01061262.

Takacs:1998:RWT

[TT98] Christiane Takacs and Roland Takacs. Random walks on trees and an inequality of means. Journal of Theoretical Probability, 11(3):701–714, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022602614733.

Tudor:2002:SPS

[TT02] C. Tudor and M. Tudor. Some properties of solutions of double stochastic differential equations. Journal of Theoretical Probability, 15(1):129–151, January 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1013893301839.

Torres:2018:MRW

[TT18] Soledad Torres and Ciprian A. Tudor. The multifractal random walk as pathwise stochastic integral: Construction and simulation. Journal of Theoretical Probability, 31(1):445–465, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Thang:2014:GRS

[TTQ14] Dang Hung Thang, Nguyen Thinh, and Tran Xuan Quy. Generalized random spectral measures. Journal of Theoretical Probability, 27(2):576–600, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-012-0461-0.

Tugaut:2014:SSP

[Tug14] Julian Tugaut. Self-stabilizing processes in multi-wells landscape in R^{d}-invariant probabilities. Journal of Theoretical Probability, 27(1):57–79, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0435-2.

Tugaut:2018:EPM

[Tug18] Julian Tugaut. Exit-problem of McKean–Vlasov diffusions in double-well landscape. Journal of Theoretical Probability, 31 (2):1013–1023, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Tanre:2006:RBM

[TV06] Etienne Tanré and Pierre Vallois. Range of Brownian motion with drift. Journal of Theoretical Probability, 19(1):45–69, January 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-006-0012-7.

Tuden:2016:DES

[TV16] Ivana Gecek Tuden and Zoran Vondracek. A distributional equality for suprema of spectrally positive Lévy processes. Journal of Theoretical Probability, 29(3):826–842, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-014-0593-5.

Talyigas:2020:BPF

[TV20] Zsófia Talyigás and Bálint Vető. Borodin–Péché fluctuations of the free energy in directed random polymer models. Journal of Theoretical Probability, 33(3):1426–1444, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-019-00919-8.pdf.

Tucci:2014:ABM

[TW14] Gabriel H. Tucci and Philip A. Whiting. Asymptotic behavior of the maximum and minimum singular value of random Vandermonde matrices. Journal of Theoretical Probability, 27(3): 826–862, September 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0466-8.

Tang:2024:CRW

[TW24] Lanlan Tang and Hua-Ming Wang. Cutpoints of (1,2) and (2,1) random walks on the lattice of positive half line. Journal of Theoretical Probability, 37(1):409–445, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01293-2.

Talagrand:1996:FBM

[TX96] Michel Talagrand and Yimin Xiao. Fractional Brownian motion and packing dimension. Journal of Theoretical Probability, 9 (3):579–593, July 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214076.

Tikhomirov:2019:NRJ

[TY19] Konstantin Tikhomirov and Pierre Youssef. On the norm of a random jointly exchangeable matrix. Journal of Theoretical Probability, 32(4):1990–2005, December 2019. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Tykesson:2009:CPA

[Tyk09] Johan H. Tykesson. Continuum percolation at and above the uniqueness threshold on homogeneous spaces. Journal of Theoretical Probability, 22(2):402–417, June 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-008-0179-1.

Teicher:1996:DST

[TZ96] Henry Teicher and Cun-Hui Zhang. A decomposition for some Utype statistics. Journal of Theoretical Probability, 9(1):161–170, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02213738.

Tan:2020:ETL

[TZ20] Zhongquan Tan and Shengchao Zheng. Extremes of a type of locally stationary Gaussian random fields with applications to Shepp statistics. Journal of Theoretical Probability, 33(4): 2258–2279, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-019-00953-6.

Uludag:2019:SSL

[UA19] Muhammed Uludag and Hakan Ayral. A subtle symmetry of Lebesgue’s measure. Journal of Theoretical Probability, 32(1): 527–540, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Utev:1998:LRS

[UB98] Sergey Utev and Niels G. Becker. A limit result for U-statistics of binary variables. Journal of Theoretical Probability, 11(3): 853–856, July 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1022666901529.

Uboe:1995:CVM

[Ubø95] Jan Ubøe. Complex valued multiparameter stochastic integrals. Journal of Theoretical Probability, 8(3):601–624, July 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF02218046.

Uchiyama:2012:AED

[Uch12a] Kôhei Uchiyama. Asymptotic estimates of the distribution of Brownian hitting time of a disc. Journal of Theoretical Probability, 25(2):450–463, June 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-010-0305-8. See er-

ratum [Uch12b].

Uchiyama:2012:EAE

[Uch12b] Kôhei Uchiyama. Erratum to: Asymptotic estimates of the distribution of Brownian hitting time of a disc. Journal of Theoretical Probability, 25(3):910–911, September 2012. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10959-012-0439-y.pdf. See [Uch12a].

Uchiyama:2016:HDL

[Uch16] Kôhei Uchiyama. The hitting distribution of a line segment for two-dimensional random walks. Journal of Theoretical Probability, 29(4):1661–1684, December 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-015-0629-5.

Uchiyama:2018:TDB

[Uch18] Kôhei Uchiyama. The transition density of Brownian motion killed on a bounded set. Journal of Theoretical Probability, 31 (3):1380–1410, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Uchiyama:2020:SLR

[Uch20] Kôhei Uchiyama. Scaling limits of random walk bridges conditioned to avoid a finite set. Journal of Theoretical Probability, 33 (3):1296–1326, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Ueda:2019:MLP

[Ued19] Yoshimichi Ueda. Matrix liberation process I: Large deviation upper bound and almost sure convergence. Journal of Theoretical Probability, 32(2):806–847, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Ueda:2022:LTC

[Ued22] Yuki Ueda. Limit theorems for classical, freely and Boolean maxinfinitely divisible distributions. Journal of Theoretical Probability, 35(1):89–114, March 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-020-01060-7.

Uemura:2004:TFM

[Uem04a] H. Uemura. Tanaka formula for multidimensional Brownian motions. Journal of Theoretical Probability, 17(2):347–366, April 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOTP.0000020698.51262.24.

Uemura:2004:SSL

[Uem04b] T. Uemura. On symmetric stable-like processes: Some path properties and generators. Journal of Theoretical Probability, 17(3):541–555, July 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000040288.26024.50.

Utev:2014:RIA

[UM14] Sergey Utev and Syeda Rabab Mudakkar. Rademacher inequalities with applications. Journal of Theoretical Probability, 27 (1):301–314, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0442-3.

Utev:2003:MII

[UP03] Sergey Utev and Magda Peligrad. Maximal inequalities and an invariance principle for a class of weakly dependent random variables. Journal of Theoretical Probability, 16(1):101–115, January 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022278404634.

Vuolle-Apiala:1996:SBM

[VA96] Juha Vuolle-Apiala. Skew Brownian motion-type of extensions. Journal of Theoretical Probability, 9(4):853–861, October 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02214254.

Vuolle-Apiala:2001:WBM

[VA01] Juha Vuolle-Apiala. Walsh’s Brownian motion-type of extensions. Journal of Theoretical Probability, 14(1):115–124, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1007873115675.

Vuolle-Apiala:2003:CER

[VA03] Juha Vuolle-Apiala. On certain extensions of a rotation invariant Markov process. Journal of Theoretical Probability, 16(4): 957–969, October 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/B%3AJOTP.0000012002.08107.34.

Vardar-Acar:2021:MDD

[VAÇA21] Ceren Vardar-Acar, Mine Çaglar, and Florin Avram. Maximum drawdown and drawdown duration of spectrally negative Lévy processes decomposed at extremes. Journal of Theoretical Probability, 34(3):1486–1505, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-020-01014-z.

Vadlamani:2010:FBF

[Vad10] Sreekar Vadlamani. Fractional Brownian flows. Journal of Theoretical Probability, 23(1):257–276, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
008-0185-3.

Vazquez-Becerra:2023:FMI

[VB23] Josue Vazquez-Becerra. Fluctuation moments induced by conjugation with asymptotically liberating random matrix ensembles.

Journal of Theoretical Probability, 36(4):1972–2039, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01246-9.

Vigelis:2013:FPD

[VC13] Rui F. Vigelis and Charles C. Cavalcante. On ϕ-families of probability distributions. Journal of Theoretical Probability, 26(3): 870–884, September 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0400-5.

Vellaisamy:2018:IDA

[VC18] P. Vellaisamy and V. Cekanavicius. Infinitely divisible approximations for sums of m-dependent random variables. Journal of Theoretical Probability, 31(4):2432–2445, December 2018. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

vanDoorn:2017:OPA

[vD17] Erik A. van Doorn. An orthogonal-polynomial approach to first-hitting times of birth–death processes. Journal of Theoretical Probability, 30(2):594–607, June 2017. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/content/pdf/10.1007/s10959-
015-0659-z.pdf.

Vechambre:2022:GSS

[Véc22] Grégoire Véchambre. General self-similarity properties for Markov processes and exponential functionals of Lévy processes. Journal of Theoretical Probability, 35(4):2083–2144, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01097-2.

Vechambre:2023:ASB

[Véc23] Grégoire Véchambre. Almost sure behavior for the local time of a diffusion in a spectrally negative Lévy environment. Journal of Theoretical Probability, 36(2):876–925, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01191-z.

Vershynin:2012:HCS

[Ver12] Roman Vershynin. How close is the sample covariance matrix to the actual covariance matrix? Journal of Theoretical Probability, 25(3):655–686, September 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-010-0338-z.

Vernic:2016:DSS

[Ver16] Raluca Vernic. On the distribution of a sum of Sarmanov distributed random variables. Journal of Theoretical Probability, 29 (1):118–142, March 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-014-0571-y.

Vidotto:2020:ISO

[Vid20] Anna Vidotto. An improved second-order Poincaré inequality for functionals of Gaussian fields. Journal of Theoretical Probability, 33(1):396–427, March 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Villemonais:2020:LBC

[Vil20] Denis Villemonais. Lower bound for the coarse Ricci curvature of continuous-time pure-jump processes. Journal of Theoretical Probability, 33(2):954–991, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Virag:1998:RWF

[Vir98] Balint Virag. Random walks on finite convex sets of lattice points. Journal of Theoretical Probability, 11(4):935–951, October 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1022612814891.

Vo:2023:ORW

[Vo23] Tran Duy Vo. The oscillating random walk on Z. Journal of Theoretical Probability, 36(4):2426–2447, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01250-z.

Voit:1990:LLN

[Voi90] Michael Voit. Laws of large numbers for polynomial hypergroups and some applications. Journal of Theoretical Probability, 3 (2):245–266, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01045161.

Voit:1993:LIL

[Voi93] Michael Voit. A law of the iterated logarithm for Markov chains on N_{0} associated with orthogonal polynomials. Journal of Theoretical Probability, 6(4):653–669, October 1993. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01049169.

Voit:1996:LTC

[Voi96] Michael Voit. Limit theorems for compact two-point homogeneous spaces of large dimensions. Journal of Theoretical Probability, 9(2):353–370, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214654.

Voit:2009:BCM

[Voi09] Michael Voit. Bessel convolutions on matrix cones: Algebraic properties and random walks. Journal of Theoretical Probability, 22(3):741–771, September 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0186-2.

Voiculescu:2017:FPP

[Voi17a] Dan-Virgil Voiculescu. Free probability for pairs of faces IV: Bifree extremes in the plane. Journal of Theoretical Probability, 30 (1):384–394, March 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-015-0635-7.

Voit:2017:DLT

[Voi17b] Michael Voit. Dispersion and limit theorems for random walks associated with hypergeometric functions of type BC. Journal of Theoretical Probability, 30(3):1130–1169, September 2017. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Voigtlaender:2021:GVP

[Voi21] Felix Voigtlaender. A general version of Price’s Theorem. Journal of Theoretical Probability, 34(3):1474–1485, September 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-020-01017-w.

Volkovich:1992:SSC

[Vol92] V. Vol’kovich. On symmetric stochastic convolutions. Journal of Theoretical Probability, 5(3):417–430, July 1992. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01060427.

Volkov:2006:PTV

[Vol06] Stanislav Volkov. Phase transition in vertex-reinforced random walks on Z with non-linear reinforcement. Journal of Theoretical Probability, 19(3):691–700, December 2006. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-006-0033-2.

Volny:2010:MAO

[Vol10] Dalibor Volný. Martingale approximation and optimality of some conditions for the Central Limit Theorem. Journal of Theoretical Probability, 23(3):888–903, September 2010. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
010-0275-x.

Vondracek:1991:CBM

[Von91] Zoran Vondracek. A characterization of Brownian motion in a Lipschitz domain by its killing distributions. Journal of Theoretical Probability, 4(2):457–464, April 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF01258748.

Vondracek:1993:MFT

[Von93] Zoran Vondracek. Markov functions of a time-changed recurrent diffusion. Journal of Theoretical Probability, 6(3):485–497, July 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01066714.

Vondracek:1996:CCD

[Von96] Zoran Vondracek. A characterization of certain diffusions on the Sierpiński gasket by the exit distributions. Journal of Theoretical Probability, 9(2):335–352, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02214653.

Vondracek:2000:AFP

[Von00] Zoran Vondracek. Asymptotics of first-passage time over a onesided stochastic boundary. Journal of Theoretical Probability, 13 (1):279–309, January 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007747312770.

Videla:2022:ESS

[VR22] Leonardo Videla and Rolando Rebolledo. Evolving systems of stochastic differential equations. Journal of Theoretical Probability, 35(3):1662–1705, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01098-1.

Vuillermot:2014:BDC

[VZ14] Pierre A. Vuillermot and Jean C. Zambrini. Bernstein diffusions for a class of linear parabolic partial differential equations. Journal of Theoretical Probability, 27(2):449–492, June 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-012-0426-3.

vanZuijlen:2018:LDC

[vZ18] W. van Zuijlen. Large deviations of continuous regular conditional probabilities. Journal of Theoretical Probability, 31(2): 1058–1096, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10959-016-0733-1.pdf.

Valle:2023:VRD

[VZ23] Glauco Valle and Leonel Zuaznábar. A version of the random directed forest and its convergence to the Brownian web. Journal of Theoretical Probability, 36(2):948–1002, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01202-z.

Wada:2019:ETL

[Wad19] Masaki Wada. Ergodic-type limit theorem for fundamental solutions of critical Schrödinger operators. Journal of Theoretical Probability, 32(1):447–459, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Wada:2021:ABF

[Wad21] Masaki Wada. Asymptotic behavior of the fundamental solutions for critical Schrödinger forms. Journal of Theoretical Probability, 34(4):1951–1958, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-020-01027-8.

Wang:2005:FLT

[Wan05] Wensheng Wang. Functional limit theorems for the increments of Gaussian samples. Journal of Theoretical Probability, 18 (2):327–343, April 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-005-3505-x.

Wang:2010:ASL

[Wan10] Li Wang. An almost sure limit theorem for super–Brownian motion. Journal of Theoretical Probability, 23(2):401–416, June 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-008-0200-8.

Wang:2011:RSN

[Wan11] Qiying Wang. Refined self-normalized large deviations for independent random variables. Journal of Theoretical Probability, 24(2):307–329, June 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0347-6.

Wang:2013:SCL

[Wan13] Jie-Ming Wang. Stochastic comparison for Lévy-Type processes. Journal of Theoretical Probability, 26(4):997–1019, December 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-011-0394-z.

Wang:2014:DFG

[Wan14a] Feng-Yu Wang. Derivative formula and gradient estimates for Gruschin type semigroups. Journal of Theoretical Probability, 27(1):80–95, March 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0427-2.

Wang:2014:EEE

[Wan14b] Jian Wang. On the existence and explicit estimates for the coupling property of Lévy processes with drift. Journal of Theoretical Probability, 27(3):1021–1044, September 2014. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
012-0463-y.

Wang:2014:IPF

[Wan14c] Yizao Wang. An invariance principle for fractional Brownian sheets. Journal of Theoretical Probability, 27(4):1124–1139, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-013-0483-2.

Wang:2017:LDS

[Wan17] Wensheng Wang. Large deviations for sums of random vectors attracted to operator semi-stable laws. Journal of Theoretical Probability, 30(1):64–84, March 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-015-0645-5.

Wang:2018:LLN

[Wan18a] Hua-Ming Wang. Law of large numbers for random walk with unbounded jumps and birth and death process with bounded jumps in random environment. Journal of Theoretical Probability, 31 (2):619–642, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Wang:2018:CLT

[Wan18b] Li Wang. Central limit theorems for supercritical superprocesses with immigration. Journal of Theoretical Probability, 31(2):984– 1012, June 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Wang:2019:CDE

[Wan19] Jian Wang. Compactness and density estimates for weighted fractional heat semigroups. Journal of Theoretical Probability, 32 (4):2066–2087, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Wang:2024:MCO

[Wan24] Wensheng Wang. The moduli of continuity for operator fractional Brownian motion. Journal of Theoretical Probability, 37(3):2097–2120, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-023-01307-z.

Watanabe:2000:CPD

[Wat00] Toshiro Watanabe. Continuity properties of distributions with some decomposability. Journal of Theoretical Probability, 13(1): 169–191, January 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007739010953.

Watanabe:2002:SSS

[Wat02] Toshiro Watanabe. Shift self-similar additive random sequences associated with supercritical branching processes. Journal of Theoretical Probability, 15(3):631–665, July 2002. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1016267815166.

Watanabe:2016:ERM

[Wat16] Toshiro Watanabe. Escape rates for multidimensional shift self-similar additive sequences. Journal of Theoretical Probability, 29(3):896–921, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-015-0599-7.

Watanabe:2021:THE

[Wat21] Toshiro Watanabe. Two hypotheses on the exponential class in the class of O-subexponential infinitely divisible distributions. Journal of Theoretical Probability, 34(2):852–873, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-019-00976-z.

Watanabe:2022:EGP

[Wat22a] Toshiro Watanabe. Embrechts–Goldie’s problem on the class of lattice convolution equivalent distributions. Journal of Theoretical Probability, 35(4):2622–2642, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-021-
01130-4.

Watanabe:2022:SOB

[Wat22b] Toshiro Watanabe. Second-order behaviour for self-decomposable distributions with two-sided regularly varying densities. Journal of Theoretical Probability, 35(2):1343–1366, June 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-021-
01083-8.

Weihrauch:2023:PRF

[WB23] Tobias Weihrauch and Stefan Bachmann. On the probabilistic representation of the free effective resistance of infinite graphs. Journal of Theoretical Probability, 36(4):1956–1971, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01218-5.

Wu:2024:EUD

[WCL24] Xiaohan Wu, Anyue Chen, and Junping Li. Existence and uniqueness of denumerable Markov processes with instantaneous states. Journal of Theoretical Probability, 37(1):511–532, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01299-w.

Weber:1996:CGS

[Web96] Michel Weber. Coupling of the GB set property for ergodic averages. Journal of Theoretical Probability, 9(1):105–112, January 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF02213736.

Weisz:1996:DRI

[Wei96] Ferenc Weisz. Duality results and inequalities with respect to Hardy spaces containing function sequences. Journal of Theoretical Probability, 9(2):301–316, April 1996. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/BF02214651.

Wei:2014:RAV

[Wei14] Ang Wei. Representations of the absolute value function and applications in Gaussian estimates. Journal of Theoretical Probability, 27(4):1059–1070, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-013-0486-z.

Weisbart:2022:ECE

[Wei22] David Weisbart. Estimates of certain exit probabilities for padic Brownian bridges. Journal of Theoretical Probability, 35(3): 1878–1897, September 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01099-0.

Werner:2005:CMS

[Wer05] Ivan Werner. Contractive Markov system with constant probabilities. Journal of Theoretical Probability, 18(2):469–479, April 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-005-3514-9.

Wang:2024:RBS

[WGYX24] Bingjun Wang, Hongjun Gao, Mingxia Yuan, and Qingkun Xiao. Reflected backward stochastic differential equations driven by G-Brownian motion under monotonicity condition. Journal of Theoretical Probability, 37(2):1902–1926, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-023-
01279-0.

Wang:2017:CMM

[WH17] Xiaoqiang Wang and Chunmao Huang. Convergence of martingale and moderate deviations for a branching random walk with a random environment in time. Journal of Theoretical Probability, 30(3):961–995, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Whalen:1992:ADM

[Wha92] Edward Whalen. The asymptotic distribution of magnitude trimmed sums for distributions in the Feller class. Journal of Theoretical Probability, 5(3):447–463, July 1992. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

http://link.springer.com/article/10.1007/BF01060429.

Wieand:2003:PMW

[Wie03] Kelly Wieand. Permutation matrices, wreath products, and the distribution of eigenvalues. Journal of Theoretical Probability, 16(3):599–623, July 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1025616431496.

Wilmer:2003:LLT

[Wil03] Elizabeth L. Wilmer. A local limit theorem for a family of nonreversible Markov chains. Journal of Theoretical Probability, 16 (3):751–770, July 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1025628701453.

Wittje:2000:SSL

[Wit00] Berthold Wittje. Spitzer’s strong law of large numbers in nonseparable Banach spaces. Journal of Theoretical Probability, 13 (1):85–92, January 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007730809136.

Wang:2003:SAL

[WLG03] Qiying Wang, Yan-Xia Lin, and Chandra M. Gulati. Strong approximation for long memory processes with applications. Journal of Theoretical Probability, 16(2):377–389, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023570510824.

Wolf:2010:RFD

[Wol10] Jamison Wolf. Random fractals determined by Lévy processes. Journal of Theoretical Probability, 23(4):1182–1203, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-009-0224-8.

Wu:1999:ECP

[Wu99] Liming Wu. Exponential convergence in probability for empirical means of Brownian motion and of random walks. Journal of Theoretical Probability, 12(3):661–673, July 1999. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021671630755.

Wu:2004:LDP

[Wu04] L. Wu. Large deviation principle for exchangeable sequences: Necessary and sufficient condition. Journal of Theoretical Probability, 17(4):967–978, October 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-004-0585-y.

Wu:2009:ECI

[Wu09] Liming Wu. A Φ-entropy contraction inequality for Gaussian vectors. Journal of Theoretical Probability, 22(4):983–991, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-009-0211-0.

Wu:2023:APR

[Wu23] Da Wu. Asymptotic properties of random contingency tables with uniform margin. Journal of Theoretical Probability, 36(4): 2066–2092, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01234-5.

Wubker:2013:AOI

[Wüb13] Achim Wübker. Asymptotic optimality of isoperimetric constants. Journal of Theoretical Probability, 26(1):198–221, March 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-011-0366-3.

Wang:2009:ECA

[WW09] Yuebao Wang and Kaiyong Wang. Equivalent conditions of asymptotics for the density of the supremum of a random walk in the intermediate case. Journal of Theoretical Probability, 22 (2):281–293, June 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0217-7.

Wang:2015:FIS

[WW15] Feng-Yu Wang and Jian Wang. Functional inequalities for stable-like Dirichlet forms. Journal of Theoretical Probability, 28(2):423–448, June 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-013-0500-5.

Wang:2014:EBC

[WWLH14] Xuejun Wang, Xinghui Wang, Xiaoqin Li, and Shuhe Hu. Extensions of the Borel–Cantelli lemma in general measure spaces.

Journal of Theoretical Probability, 27(4):1229–1248, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-013-0526-8.

Wu:2010:RIL

[WX10] Dongsheng Wu and Yimin Xiao. Regularity of intersection local times of fractional Brownian motions. Journal of Theoretical Probability, 23(4):972–1001, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-009-0221-y.

Wang:2022:EUM

[WX22] Ran Wang and Yimin Xiao. Exact uniform modulus of continuity and Chung’s LIL for the generalized fractional Brownian motion. Journal of Theoretical Probability, 35(4):2442–2479, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-021-01148-8.

Wang:2010:PIP

[WY10a] Feng-Yu Wang and Chenggui Yuan. Poincaré inequality on the path space of Poisson point processes. Journal of Theoretical Probability, 23(3):824–833, September 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-009-0232-8.

Watanabe:2010:LSS

[WY10b] Toshiro Watanabe and Kouji Yamamuro. Local subexponentiality and self-decomposability. Journal of Theoretical Probability, 23(4):1039–1067, December 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-009-0240-8.

Wang:2016:GEE

[WY16] Zhongzhi Wang and Weiguo Yang. The generalized entropy ergodic theorem for nonhomogeneous Markov chains. Journal of Theoretical Probability, 29(3):761–775, September 2016. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
015-0597-9.

Watanabe:2017:TNC

[WY17] Toshiro Watanabe and Kouji Yamamuro. Two non-closure properties on the class of subexponential densities. Journal of Theoretical Probability, 30(3):1059–1075, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Wu:2023:GAR

[WY23] Yue Wu and Chenggui Yuan. The Galerkin analysis for the random periodic solution of semilinear stochastic evolution equations. Journal of Theoretical Probability, 36(1):??, March 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01236-x.

Wu:2024:GAR

[WY24] Yue Wu and Chenggui Yuan. The Galerkin analysis for the random periodic solution of semilinear stochastic evolution equations. Journal of Theoretical Probability, 37(1):133–159, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01236-x.

Wang:2021:BSD

[WZ21a] Falei Wang and Guoqiang Zheng. Backward stochastic differential equations driven by G-brownian motion with uniformly continuous generators. Journal of Theoretical Probability, 34 (2):660–681, June 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-020-00998-y.

Wang:2021:DRS

[WZ21b] Wenyuan Wang and Xiaowen Zhou. A drawdown reflected spectrally negative Lévy process. Journal of Theoretical Probability, 34(1):283–306, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-019-00971-4.

Wollgast:2023:WDI

[WZK23] Dennis Wollgast and Pavel Zorin-Kranich. Weighted Davis inequalities for martingale square functions. Journal of Theoretical Probability, 36(3):1520–1533, September 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-022-
01204-x.

Wu:2017:OTG

[WZY17] Lan Wu, Jiang Zhou, and Shuang Yu. Occupation times of general Lévy processes. Journal of Theoretical Probability, 30(4): 1565–1604, December 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Wang:2022:STD

[WZZ22] Wei Wang, Jianliang Zhai, and Tusheng Zhang. Stochastic two-dimensional Navier–Stokes equations on time-dependent domains. Journal of Theoretical Probability, 35(4):2916–2939, December 2022. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-021-01150-0.

Xiao:1997:WVG

[Xia97] Yimin Xiao. Weak variation of Gaussian processes. Journal of Theoretical Probability, 10(4):849–866, October 1997. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1022606431629.

Xiao:1998:LTR

[Xia98] Yimin Xiao. Local times and related properties of multidimensional iterated Brownian motion. Journal of Theoretical Probability, 11(2):383–408, April 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1022679721638.

Xu:2023:SCR

[XL23] Jie Xu and Qiqi Lian. A strong convergence rate of the averaging principle for two-time-scale forward–backward stochastic differential equations. Journal of Theoretical Probability, 36(4): 2590–2610, December 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-023-01278-1.

Xu:2023:STS

[XSR23] Jie Xu, Yanhua Sun, and Jie Ren. A support theorem for stochastic differential equations driven by a fractional Brownian motion. Journal of Theoretical Probability, 36(2):728–761, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-022-01186-w.

Xu:1998:EPZ

[Xu98] Shaogang Xu. An ergodic process of zero divergence-distance from the class of all stationary processes. Journal of Theoretical Probability, 11(1):181–195, January 1998. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1021651126212.

Xu:2007:RBS

[Xu07] Mingyu Xu. Reflected backward SDEs with two barriers under monotonicity and general increasing conditions. Journal of Theoretical Probability, 20(4):1005–1039, December 2007. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
007-0109-7.

Xu:2013:RHF

[Xu13] Fangjun Xu. Regularity of harmonic functions for some Markov chains with unbounded range. Journal of Theoretical Probability, 26(2):541–556, June 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-012-0412-9.

Xu:2018:CLT

[Xu18] Lu Xu. A central limit theorem for stochastic heat equations in random environment. Journal of Theoretical Probability, 31(3): 1356–1379, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Xue:1991:CCS

[Xue91] Xing-Hong Xue. Convergence of classes of subpramarts and games which become better with time. Journal of Theoretical Probability, 4(3):605–623, July 1991. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01210326.

Xue:2015:ABC

[Xue15] Xiaofeng Xue. Asymptotic behavior of critical infection rates for threshold–one contact processes on lattices and regular trees. Journal of Theoretical Probability, 28(4):1447–1467, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/s10959-014-0544-1.

Xue:2018:ACV

[Xue18] Xiaofeng Xue. Asymptotic of the critical value of the large– dimensional SIR epidemic on clusters. Journal of Theoretical Probability, 31(4):2343–2365, December 2018. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Xing:2010:SEI

[XY10] Guodong Xing and Shanchao Yang. Some exponential inequalities for positively associated random variables and rates of convergence of the strong law of large numbers. Journal of Theoretical Probability, 23(1):169–192, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-008-0205-3.

Xia:2009:PBD

[XZ09] Aihua Xia and Fuxi Zhang. Polynomial birth–death distribution approximation in the Wasserstein distance. Journal of Theoretical Probability, 22(2):294–310, June 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-008-0207-1.

Xiao:2010:ASL

[XZ10] Han Xiao and Wang Zhou. Almost sure limit of the smallest eigenvalue of some sample correlation matrices. Journal of Theoretical Probability, 23(1):1–20, March 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-009-0270-2.

Xu:2018:NFL

[XZZ18] Lifeng Xu, Shaoyi Zhang, and Ren Zhang. A note on Foster– Lyapunov drift condition for recurrence of Markov chains on general state spaces. Journal of Theoretical Probability, 31(4): 1923–1928, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Yamato:2024:RDA

[Yam24] Kosuke Yamato. Renewal dynamical approach for nonminimal quasi-stationary distributions of one-dimensional diffusions. Journal of Theoretical Probability, 37(1):533–548, March 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01239-8.

Yan:2009:CIO

[Yan09] Liqing Yan. Comparison inequalities for one sided normal probabilities. Journal of Theoretical Probability, 22(4):827–836, December 2009. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-009-0248-0.

Yano:2013:RWF

[Yan13] Kouji Yano. Random walk in a finite directed graph subject to a road coloring. Journal of Theoretical Probability, 26(1):259–283, March 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-011-0398-8.

Yano:2017:JLO

[Yan17] Yuko Yano. On the joint law of the occupation times for a diffusion process on multiray. Journal of Theoretical Probability, 30(2):490–509, June 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Yang:2018:HDR

[Yan18] Xiaochuan Yang. Hausdorff dimension of the range and the graph of stable-like processes. Journal of Theoretical Probability, 31(4):2412–2431, December 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Yasuda:2000:IDD

[Yas00] Kumi Yasuda. On infinitely divisible distributions on locally compact Abelian groups. Journal of Theoretical Probability, 13 (3):635–657, July 2000. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007850210027.

Yasuda:2010:MPA

[Yas10] Kumi Yasuda. Markov processes on the Adeles and representations of Euler products. Journal of Theoretical Probability, 23(3): 748–769, September 2010. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0222-x.

Yaskov:2018:LQF

[Yas18] Pavel Yaskov. LLN for quadratic forms of long memory time series and its applications in random matrix theory. Journal of Theoretical Probability, 31(4):2032–2055, December 2018. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Yasuda:2020:LDS

[Yas20] Kumi Yasuda. Large deviations for scaled sums of p-adic-valued rotation-symmetric independent and identically distributed random variables. Journal of Theoretical Probability, 33(2):1196– 1210, June 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Yin:2016:SLL

[YBH16] Yanqing Yin, Zhidong Bai, and Jiang Hu. On the semicircular law of large–dimensional random quaternion matrices. Journal of Theoretical Probability, 29(3):1100–1120, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-015-0610-3.

Yuan:2024:HMU

[YC24] Weijie Yuan and Zhenlong Chen. Hausdorff measure and uniform dimension for space–time anisotropic Gaussian random fields. Journal of Theoretical Probability, 37(3):2304–2329, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-024-01323-7.

Ye:2024:SDE

[Ye24] Wenjie Ye. Stochastic differential equations with local growth singular drifts. Journal of Theoretical Probability, 37(3):2576– 2614, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-024-01333-5.

Yengibarian:2004:FMC

[Yen04] N. B. Yengibarian. Factorization of Markov chains. Journal of Theoretical Probability, 17(2):459–481, April 2004. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOTP.0000020703.46248.19.

Ye:2008:ERN

[YGHL08] Liuer Ye, Xianping Guo, and Onésimo Hernández-Lerma. Existence and regularity of a nonhomogeneous transition matrix under measurability conditions. Journal of Theoretical Probability, 21(3):604–627, September 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-008-0163-9.

Yi:2023:MMF

[Yi23] Jaeyun Yi. Macroscopic multi-fractality of Gaussian random fields and linear stochastic partial differential equations with colored noise. Journal of Theoretical Probability, 36(2):926– 947, June 2023. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-022-01198-6.

Yin:2001:LRC

[Yin01] G. Yin. On limit results for a class of singularly perturbed switching diffusions. Journal of Theoretical Probability, 14(3):673–697, July 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1017541022565.

Yang:2022:SOT

[YLY22] Yang Yang, Shuang Liu, and Kam Chuen Yuen. Second-order tail behavior for stochastic discounted value of aggregate net losses in a discrete-time risk model. Journal of Theoretical Probability, 35(4):2600–2621, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://

link.springer.com/article/10.1007/s10959-021-01143-z.

Yokoyama:2021:SPM

[Yok21] Satoshi Yokoyama. A stochastically perturbed mean curvature flow by colored noise. Journal of Theoretical Probability, 34 (1):214–240, March 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-019-00983-0.

Yoshida:2020:RFA

[Yos20] Hiroaki Yoshida. Remarks on a free analogue of the beta prime distribution. Journal of Theoretical Probability, 33(3):1363–1400, September 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Yosef:2021:STG

[Yos21] Arthur Yosef. Selected topics in the generalized mixed setindexed fractional Brownian motion. Journal of Theoretical Probability, 34(3):1366–1381, September 2021. CODEN JT-PREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL

https://link.springer.com/article/10.1007/s10959-021-
01077-6.

Yuan:2003:CLT

[YSH03] Ming Yuan, Chun Su, and Taizhong Hu. A Central Limit Theorem for random fields of negatively associated processes. Journal of Theoretical Probability, 16(2):309–323, April 2003. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1023538824937.

Yan:2008:LFS

[YSL08] Litan Yan, Yu Sun, and Yunsheng Lu. On the linear fractional self-attracting diffusion. Journal of Theoretical Probability, 21 (2):502–516, June 2008. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0113-y.

Yu:1989:FWC

[Yu89] Kai Fun Yu. On fixed-width confidence intervals associated with maximum likelihood estimation. Journal of Theoretical Probability, 2(2):193–199, April 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01053409.

Yu:1990:CCW

[Yu90] Kai Fun Yu. Complete convergence of weighted sums of martingale differences. Journal of Theoretical Probability, 3(2):339–347, April 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/BF01045165.

Yu:2021:HOD

[Yu21] Qian Yu. Higher-order derivative of self-intersection local time for fractional Brownian motion. Journal of Theoretical Probability, 34(4):1749–1774, December 2021. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.

springer.com/article/10.1007/s10959-021-01093-6.

Yu:2018:LMD

[YWC18] Shihang Yu, Dehui Wang, and Xia Chen. Large and moderate deviations for the total population arising from a sub-critical Galton–Watson process with immigration. Journal of Theoretical Probability, 31(1):41–67, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Yin:2022:WPS

[YWS22] Xiuwei Yin, Jiang-Lun Wu, and Guangjun Shen. Well-posedness for stochastic fractional Navier–Stokes equation in the critical Fourier–Besov space. Journal of Theoretical Probability, 35(4): 2940–2959, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01152-y.

Yen:2015:VPT

[YY15] Ju-Yi Yen and Marc Yor. A variant of Pitman’s theorem on (2J_{s} − R_{s}, s ≥ 0) for a general transient Bessel process R_{(+)} and its implications for the corresponding Itô’s measure n_{(−)}. Journal of Theoretical Probability, 28(1):223–230, March 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-013-0505-0.

Yan:2019:ABH

[YY19] Litan Yan and Xianye Yu. Asymptotic behavior for high moments of the fractional heat equation with fractional noise. Journal of Theoretical Probability, 32(4):1617–1646, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Yu:2024:LTS

[YY24] Qian Yu and Xianye Yu. Limit theorem for self-intersection local time derivative of multidimensional fractional Brownian motion. Journal of Theoretical Probability, 37(3):2054–2075, September 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01300-6.

Yang:2014:EUI

[YZ14] Juan Yang and Tusheng Zhang. Existence and uniqueness of invariant measures for SPDEs with two reflecting walls. Journal of Theoretical Probability, 27(3):863–877, September 2014.

CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-012-0448-x.

Yue:2015:ACL

[YZ15] Wen Yue and Tusheng Zhang. Absolute continuity of the laws of perturbed diffusion processes and perturbed reflected diffusion processes. Journal of Theoretical Probability, 28(2):587–618, June 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-013-0499-7.

Zabell:1993:LTE

[Zab93] S. L. Zabell. A limit theorem for expectations conditional on a sum. Journal of Theoretical Probability, 6(2):267–283, April 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01047574.

Zabell:1995:CME

[Zab95] S. L. Zabell. Characterizing Markov exchangeable sequences. Journal of Theoretical Probability, 8(1):175–178, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02213460.

Zhou:2023:LSD

[ZBH23] Huanchao Zhou, Zhidong Bai, and Jiang Hu. The limiting spectral distribution of large-dimensional general information-plus-noise-type matrices. Journal of Theoretical Probability, 36(2): 1203–1226, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01193-x.

Zeindler:2013:CLT

[Zei13] Dirk Zeindler. Central limit theorem for multiplicative class functions on the symmetric group. Journal of Theoretical Probability, 26(4):968–996, December 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/s10959-011-0382-3.

Zeineddine:2018:ABW

[Zei18] Raghid Zeineddine. Asymptotic behavior of weighted power variations of fractional Brownian motion in Brownian time. Journal of Theoretical Probability, 31(3):1539–1589, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Zempleni:1990:HHH

[Zem90] András Zempléni. On the heredity of Hun and Hungarian property. Journal of Theoretical Probability, 3(4):599–609, October 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1007/BF01046099.

Zeng:1995:CMS

[Zen95] Wei-Bin Zeng. On characterization of multivariate stable distributions via random linear statistics. Journal of Theoretical Probability, 8(1):1–15, January 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF02213449.

Zeuner:1989:CLT

[Zeu89] Hansmartin Zeuner. The central limit theorem for Chébli– Trimèche hypergroups. Journal of Theoretical Probability, 2 (1):51–63, January 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01048268.

Zeuner:1994:IPR

[Zeu94] Hansmartin Zeuner. Invariance principles for random walks on hypergroups on R_{2} and n. Journal of Theoretical Probability, 7(2):225–245, April 1994. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02214269.

Zhang:2012:HBM

[ZH12] Wei Zhang and Marjorie G. Hahn. Haar-based multiresolution stochastic processes. Journal of Theoretical Probability, 25(3): 890–909, September 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-010-0333-4.

Zhang:2001:SME

[Zha01a] Cun-Hui Zhang. Some moment and exponential inequalities for V -statistics with bounded kernels. Journal of Theoretical Probability, 14(2):511–525, April 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1011171916115.

Zhang:2001:MAS

[Zha01b] Yu Zhang. A martingale approach in the study of percolation clusters on the Z^{d} lattice. Journal of Theoretical Probability, 14 (1):165–187, January 2001. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1007877216583.

Zhang:2005:FCL

[Zha05] Mei Zhang. Functional Central Limit Theorem for the super– Brownian motion with super–Brownian immigration. Journal of Theoretical Probability, 18(3):665–685, July 2005. CO-DEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

URL http://link.springer.com/article/10.1007/s10959-
005-7254-7.

Zhang:2011:CLT

[Zha11] Mei Zhang. Central limit theorems for a super-diffusion over a stochastic flow. Journal of Theoretical Probability, 24(1): 294–306, March 2011. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-009-0261-3.

Zhang:2017:TRM

[Zha17] Deng Zhang. Tridiagonal random matrix: Gaussian fluctuations and deviations. Journal of Theoretical Probability, 30(3):1076– 1103, September 2017. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Zhang:2018:NEN

[Zha18] Yi Zhang. On the nonexplosion and explosion for nonhomogeneous Markov pure jump processes. Journal of Theoretical Probability, 31(3):1322–1355, September 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Zhang:2019:GFM

[Zha19] Deng Zhang. Gaussian fluctuations and moderate deviations of eigenvalues in unitary invariant ensembles. Journal of Theoretical Probability, 32(4):1647–1687, December 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Zhang:2023:LDP

[Zha23] Shuxiong Zhang. Lower deviation probabilities for level sets of the branching random walk. Journal of Theoretical Probability, 36(2):811–844, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01183-z.

Zhou:2024:ALS

[ZHBS24] Huanchao Zhou, Jiang Hu, Zhidong Bai, and Jack W. Silverstein. Analysis of the limiting spectral distribution of large-dimensional general information–plus-noise-type matrices. Journal of Theoretical Probability, 37(2):1199–1229, June 2024. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL https://link.springer.com/article/10.
1007/s10959-023-01276-3.

Zhou:1993:RDR

[Zho93] Xian Yin Zhou. Resistance dimension, random walk dimension and fractal dimension. Journal of Theoretical Probability, 6(4): 635–652, October 1993. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01049168.

Zhou:1995:RRR

[Zho95] Xian Yin Zhou. On the range of reversible random walks on Z^{2} in a random environment. Journal of Theoretical Probability, 8 (2):453–473, April 1995. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF02212888.

Zhong:2015:FCF

[Zho15] Ping Zhong. On the free convolution with a free multiplicative analogue of the normal distribution. Journal of Theoretical Probability, 28(4):1354–1379, December 2015. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1007/s10959-014-0556-x.

Zhu:2023:SER

[Zhu23] Yizhe Zhu. On the second eigenvalue of random bipartite biregular graphs. Journal of Theoretical Probability, 36(2):1269– 1303, June 2023. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-022-01190-0.

Zitkovic:2002:FVB

[Zit02] Gordan Zitković. A filtered version of the bipolar theorem of Brannath and Schachermayer. Journal of Theoretical Probability, 15(1):41–61, January 2002. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1013885121598.

Zitelli:2022:AFL

[Zit22] G. L. Zitelli. Amalgamated free Lévy processes as limits of sample covariance matrices. Journal of Theoretical Probability, 35(4): 2176–2193, December 2022. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-021-01153-x.

Zheng:2018:RDC

[ZL18] Shiqiu Zheng and Shoumei Li. On the representation for dynamically consistent nonlinear evaluations: Uniformly continuous case. Journal of Theoretical Probability, 31(1):119–158, March 2018. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Zhang:2023:EEN

[ZM23] Wei Zhang and Hui Min. L^{p}-error estimates for numerical schemes for solving certain kinds of mean-field backward stochastic differential equations. Journal of Theoretical Probability, 36 (2):762–778, June 2023. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-022-01184-y.

Zhang:2020:QCL

[ZRP20] Na Zhang, Lucas Reding, and Magda Peligrad. On the quenched central limit theorem for stationary random fields under projective criteria. Journal of Theoretical Probability, 33(4):2351– 2379, December 2020. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL https://link.springer.

com/article/10.1007/s10959-019-00943-8.

Zuckerman:1989:CTR

[Zuc89] David Zuckerman. Covering times of random walks on bounded degree trees and other graphs. Journal of Theoretical Probability, 2(1):147–157, January 1989. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/BF01048276.

Zweimuller:2007:MLT

[Zwe07] Roland Zweimüller. Mixing limit theorems for ergodic transformations. Journal of Theoretical Probability, 20(4):1059– 1071, December 2007. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-007-0085-y.

Zhao:2004:BEB

[ZWW04] L. C. Zhao, C. Q. Wu, and Q. Wang. Berry–Esséen bound for a sample sum from a finite set of independent random variables. Journal of Theoretical Probability, 17(3):557–572, July 2004. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230

(electronic). URL http://link.springer.com/article/10.
1023/B%3AJOTP.0000040289.99786.ca.

Zhao:2024:VMS

[ZX24] Linjie Zhao and Xiaofeng Xue. The voter model with a slow membrane. Journal of Theoretical Probability, 37(3):2775–2817, September 2024. CODEN JTPREO. ISSN 0894-9840 (print),

1572-9230 (electronic). URL https://link.springer.com/
article/10.1007/s10959-024-01321-9.

Zaharopol:1999:DCE

[ZZ99] Radu Zaharopol and Gheorghita Zbaganu. Dobrushin coefficients of ergodicity and asymptotically stable L^{1}-contractions. Journal of Theoretical Probability, 12(4):885–902, October 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1023/
A%3A1021684818286.

Zhang:2012:PRW

[ZZ12] Qi Zhang and Huaizhong Zhao. Probabilistic representation of weak solutions of partial differential equations with polynomial growth coefficients. Journal of Theoretical Probability, 25 (2):396–423, June 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0350-y.

Zang:2019:ABT

[ZZ19] Qingpei Zang and Lixin Zhang. Asymptotic behaviour of the trajectory fitting estimator for reflected Ornstein–Uhlenbeck processes. Journal of Theoretical Probability, 32(1):183–201, March REFERENCES 416

2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).